A Bibliography of Computer Graphics Forum: 1982–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 3.01

Title word cross-reference

2 [BW09, BG89, BCF94, CLME09, Day90, DGF98, DHH02, DF90, EFGS96, FG88, GD96, GL94a, GL94b, HLM97, IEGC08, MSM^{+}08, MB08, The02a, TRS03a, TRS03b, TSH01, WG93, WTHS06, WO92, WOBT09, ZLK05, van90]. 2\frac{1}{2}D [FW99]. 3 [Ais85, AS96b, AA09, AD01, ANF97a, ANF97b, AMSF08, BG95, BP98, BCG^{+}96, BB00a, Bel87, BR02, BGK^{+}96, BMG99, BTG95, BG08, BYB09, BH93, BCD01, BCF94, BMWM01, BNRSV01, BB08, CCFM08, CTSO03a, CTSO03b, CLB^{+}09, CMPS93, CMS94, CD94, CYJ02, CKSW08, DCPS08, DH93a, EWHS08, EFGS96, FFD93, GD96, GL94a, GL94b, GE98, GTS86, Han97, HE94, Hub93, IIS08, IP99, IK01a, Jac85, JBTS08, Joo86, Kil82, Kil85, KK07, KQWM08, LGMT00, LJN02, LVT08, LCM^{+}09, LCWK07, MFT02, MK99, MGAF95, NK99, NMOT01, O’H02, OMT02, OPC96, PP89, PGG^{+}09, Rob93, RCM^{+}01, RWP88, RL09, Ros97, Sab82, Sam93a, SW09, Šár07, SSB08, Sco02, SLSK07, SW92b, Sni95, SB99b, SP03a, SFWS03a, SP03b, SFWS03b, SS95, SSB05, SOM04, TL01, TTW90, TW97b, TVD09, TGS96, VPLL08]. 3 [VCDF95, Vel93, VW95, VB99, WG93, WL08, WWG07, Wat96, WTHS04, WLSG03, YD88, ZHC^{+}00, ZPS03, ZLK05, van90, vJB85]. 4 [MDC93, NCKG00, VG00, WMWG09]. 7^{th} [Kun04]. 9^{th} [Kun04]. ^{st} [Hec01]. th [HJL07]. C^{1} [THN93]. C^{2} [KFK94]. Δ [Mil90]. E^{3} [Ska96]. ϵ [Pic86b]. G^{1} [BV99]. G^{2} [LS08b]. k [VG00]. L [DKY98]. λ_{2} [SVG^{+}08]. N [Déc05, NSY09, ILRS03a, ILRS03b]. Q [Hal99]. r [AAK^{+}09, FR92]. Σ [Mil90]. √3 [LG00]. z [SDD^{+}92].

-based [SVG^{+}08]. -buffer [SDD^{+}92]. -Buffers [Déc05]. -D [BW09, Bel87, BG89, DF90, HLM97, MGAF95, YD88, ZHC^{+}00]. -Dimensional [BMG99, ILRS03a, ILRS03b]. -mapping [Hal99]. -Sampled [AAK^{+}09]. -sets [FR92]. -Sided [NSY09]. -Subdivision [LG00]. -surface [THN93]. -systems [DKY98].

//www.eg.org/sbm/= [Che06].

10th [Ano95a]. 11th [Ano96a]. 12 [HJL07]. 13th [Ano92]. 14th [Ano93, Ano95p, Ano96p, LMD04]. 15th [Ano96b, Ano94]. 16th [Ano97q, Ano97z, Kon06]. 17th [Ano96l, Ano96s]. 18th [AR06]. 1979 [ACM80]. 1980 [Van80]. 1982 [BP82]. 1983 [tH83a]. 1984 [Kil85]. 1985 [Duc86]. 1987 [Cre88]. 1988 [DJ88]. 1989 [Bro90, HHS89, tHS90]. 1991 [HB91, Jan91]. 1992 [Arn91]. 1994 [HK94]. 1995 [Gob95, HP95, PB95, SvZ95, TT95b]. 1996 [Ano97-29, BH96, Gob96, PS96b]. 1997 [Ano97w, Ano97s, Ano97z, Ano97-28, Ano97-32]. 1998 [Ano97t, Duc98, MJ98]. 1a [Dam91]. 1st [Ano95b, Arn84, Kon06].

2003 [CA05, DT04, FMK04, KSH04, Kun04, LMD04, SCA04]. 2004 [Ano04b, BPB^{+}04, DL04, Kei04, LLR^{+}04, Oll04, Raf05, SZAB04]. 2005 [Des06, Kon06, LLD05, NSGP06, Wil06a]. 2006 [Ano07j, BH06, Duc06b, LLRD07, San06, SP06]. 2007 [Bot07, Duc07, Kau07]. 2008 [Ano08e, MTPS08, Wei08]. 2009 [Ano08f, CDD09, DS09, SJ09b]. 2010 [Enc98]. 25th [Ano04b]. 26th [Ano05b]. 27th [Ano06b, Kil85]. 28th [Ano07g]. 29th [Kon06, Ano08d]. 2nd [Kid84, Ano98b, Ano98c, Che06, CG07a].

3-D [WHL^{+}04]. 3-Manifolds [MDP08]. 30th [LLRD07, Ano09c]. 31st [LLD05]. 3D [DPT^{+}08, Fre90, PNR89, TW97a]. 3D-audio [DPT^{+}08]. 3rd [Ano97-32, Dun91, LLRD07].

4th [Ano97d, LLD05, Ano06g].

5th [Ano96c, Arn84, CG07a, Ano97e, Ano98d].

6th [Ano95c, ZY04, tH83b, Ano88a, Ano98e].

7th [Ano95d, Gal84, Kil85, Ano89].

’81 [Enc81]. ’82 [WG82, Wat82, Kid82, Kuh82]. ’83 [HP84, SW83, tH83a, End83b, HEtH^{+}83, Joo86, tH84]. ’84 [Arn84, Hop84, TB84]. ’85 [Van85, Gre85]. ’86 [HD86, Req86, Rob87, Mor86]. ’87 [How87]. ’88 [DJ88, Wat88]. ’89 [HHS89, Arn89, Jan89]. 8th [Ano97f, Ano97-32, Ano98f, Ano98g, DS98, TvdP98].

’90 [Duc90b]. ’91 [HB91]. ’92 [Arn91, Ano91c, Ano92]. ’93 [Ano93]. ’94 [Ano94]. ’95 [Ano95j, Ano95k, Ano95o, Ano95z, Gob95, HP95, PB95, SvZ95, TT95b, Ano95i]. ’96 [Ano95l, Ano95m, Ano96g, Ano96h, BH96, Gob96, PS96b, Ano96l]. ’97 [Ano96j, Ano96k, Ano97v, Ano97w, Ano97o, Ano97r, Ano97-28, Ano97-32, Ano96i, Ano97u]. ’98 [Ano95n, Ano98d, Ano97y, Ano97x, Ano98u, Ano98v, Ano98w]. 9th [Aus91, Ano91a, Ano97g, Ano98h, Ano98i, Ano98j]. 9x [DDR93, ND94].

= [Che06].

A-buffer [LW92]. Aachen [KSH04]. Abingdon [MJ98]. Abstract [DH93b, Han05, MG87, TC93]. Abstraction [Gna82, IIS09, JC08, KL08, KKD09, WT09]. Abstractions [BGB^{+}08a]. Abstractive [CZCE08]. Accelerated [AGG^{+}08, BKES00, BS02, Gar09, KZ08, LMS04, MMFE08, McC96, RPZ02, WBP98, DSSD99]. Accelerating [HLL07, LD08c, LCD09b]. Acceleration [LLHY09, LCD09b, SPH^{+}09, YN00]. Access [KCB97]. Accessibility [HD02]. Accessible [CH09, KCL06]. Accumulation [BG09, LG95]. Accurate [BPMG08, DF93, EL01, FBP08, HK09c, KHM09, KGL^{+}98, LW92, RH06, SLTM08, TCRS00, ZCP07]. achieved [RPP93]. Acid [DTA94]. ACM [Des06, BPB^{+}04, Rob87]. ACM/EG [BPB^{+}04]. Acoustic [BMD^{+}08, MF00]. Acquisition [BPM06, BR02, DCG87, MMS^{+}05, RA94, SHS99, WHL^{+}04]. Actions [HKG06, KMTT92]. Active [Gla99, HH98]. Activities [Ano84a, Mud83]. Activity [EBMT00, JNM^{+}09, RL84]. Actors [NT95, ST94]. ADA [Mac84, Mil87]. Adaptable [WW99, KSH92]. Adaptation [CC06, PWP08, WW09b]. Adapted [KLD^{+}09]. Adaptive [AG01, BW09, BT95, BPWG07, BS08, BP01, CG02, CD08, CB09, CGG^{+}03a, CGG^{+}03b, CAE08, CZGF05, CG07b, CPK09, DTG96, DMAC03a, DMAC03b, DDC09, DRS09, FP04, FCGW02, GGW98, GMC^{+}06, GVV05, GBP07, HWC^{+}05, HH98, HJ99, IP00, KJ92, KMS07, MTF03a, MTF03b, PJ94, PCK09, PWH98, SHS99, SS09, SP01, SD94b, ST08, Ure00, WDGT01, WN09, YWY08, ZBP99, dS96, LBT92, Rap92]. Adaptively [DRS08]. Adding [Bak88, LAA08]. addressing [Ger92]. Adjustments [LAA08]. Advanced [BDS^{+}03, DDPL00, HSS^{+}05, HE01, HP02, Kil85, XLTP03a, XLTP03b]. Advances [Ale02, BW01]. Advection [CKSW08, LTH08, SWPL08, Wei04]. aerodynamicist [PW93]. Aesthetics [Ano08e, Ano08f, NSGP06]. Affect [DBHM03a, DBHM03b]. affecting [Kin95]. Affective [LC09]. Affine [KHK^{+}09, NG92, NR95, dFSV03, dS96]. Africa [GS06]. AFRIGRAPH [GS06]. Against [HL02]. Age [Fra83]. Agenda [Arn08]. Agent [EBMT00, SGS05]. Agents [BBT99, NG03a, NG03b, Bad93]. Agile [MRT08]. Aging [BPMG04, DGA04, GMW04, HCG08, KB04, SSSB07]. AGM [Ano95p, Ano96b]. AGRELs [RK94]. Aided [BF84, Owe88, Owe89b, Owe94, Owe90b, Owe92a, Owe93, Owe95]. Aiming [DKYN96]. Alain [Fiu01]. Albedo [WL08]. Albuquerque [Osl82]. Algae [DTA94]. Algebraic [EBV01, Gna83, GGG08, RS08]. ALGOL [MA85]. Algorithm [And89, AS95, AHT04, BS98, CFFP84, CY89, COF95, Day90, DBG99, FP04, FP94, FBW01, Fis98, FA87, Hew84a, HS98, HB94, HL02, KSKAC02, Kuz90, Kuz95, LLA06, Lin85, Liu94, LŽY04, LSŽ08, Mil84, Moh87, MGAF95, NW01, O’H02, PJ94, RHv95, RR94, Ska87, Sug94, SN84, TT95a, VW90, WC05, ŽCO97, vKB94, AP92, BP93, BLS93, BBA08, Day92, FND92, Kra92, Rap92, Ska96]. Algorithmic [BR97, BP83a]. Algorithms [AR94, BCS96, Day88, JEO00, JFS09, LS89, Mil83, MBW^{+}05, SM86, SG96b, SN86, TIS^{+}95, UWP06, Vel99, VCC98, SDD^{+}92, VW91]. Alias [SEA08]. Alias-free [SEA08]. Aliased [DS05a, Liu94, TNF89]. Aliasing [BK01, Chr86, Pil85, GM06]. Alignment [AGP08, COF95]. All-Integer [Liu94]. Alone [WMTG05, ZHC^{+}00]. Along [LSŽ08]. Alpha [GMW97, WS09a, Wil06b]. Alphabetic [LO95]. Alternative [MKV09, Gui92]. Alvey [Cre88]. Ambient [PB07, RMSD^{+}08]. Ambiguous [AHM09]. Analysis [Ano99m, BBT^{+}06, Ber09, BCWR08, CCSG^{+}09, DG97, FdABS99, GBAL09, GD96, GL94b, HDSD99, HWAG09, JNM^{+}09, hKLS00, KHIK01, LZ07, LZSCO09, MTR08, MKO^{+}08, Mun83, NGB^{+}09, NMOT01, PB90, Ric87a, SW04a, SOG09, SKFNC97, TWC^{+}09, WMS^{+}08, YMMS06, BG93, GL94a, Kra92, MK08]. Analytic [TT97]. Analytical [PP09a]. Analytics [BCWR08, HKD^{+}08, IF09, JZF^{+}09, MK08, WMS^{+}08]. Analyzer [LAD02]. Anatomical [CLME09, JBB^{+}08]. Anatomy [ZPS03]. Anatomy-Based [ZPS03]. Anchor [AR95]. Andreas [Ano07b]. Aneurysm [NGB^{+}09]. Angeles [Ano97-32, LLR^{+}04, LLD05, Ano95z]. Angle [LAD02]. Angle-Analyzer [LAD02]. Anglia [Ano97z, Ano97-28, AEL^{+}82]. Angular [KB89]. ANIMA [MM93]. Animacy [RAP08]. Animal [GJL^{+}09, KOB^{+}08]. Animals [Ter02, WF97]. Animatable [LGMT00]. Animated [BT95, GFW^{+}06, PCR89, SS96b, SWG08, TGS96, WMG^{+}09, WS02, van89a, Bad93, SSK93]. Animating [BCN03b, BCN03a, BY08, BSVS04, CCI^{+}07, HK03b, JL00, KRFC09, NJ04, OAIS09, QNTN03b, SRH^{+}09, SOH99, WF97, WHL^{+}04, DTKT93]. Animation [ATBG08, Ano95-27, Ano98f, Ary86, AB97, BCN03b, BCH^{+}95, BPB^{+}04, CKB04, CYJ02, Col05, CMT02, DAP08, DTTS08, DN08, DO00, DF90, DKY98, Dur01, GLHB09, GGK06, GFW^{+}06, GW05, HKW09, HE01, HKG06, Heg90, Hég91, HKMS08, HE94, HK03a, HK03b, ITYI09, JL08, KŽ05, KHIK01, KMT03a, KMT03b, KMA05, KMB94, LL00, LL05, LCL07, MTT89, MCH94, MB97, MH07, MBCN09, MAA^{+}09, Moh87, MCT01, NB88, NAS07, OIST91, OTH02, Osh08, PG95, PPD07, Pat95, PG08, PTB^{+}03a, PH94, PH96, PT88, RAP08, RLN06, RHC08, RSKN08, SCA04, SSSK04a, SL09, SNI06, SLTM08, SiK– DM05, Sni95, SKC01, TFK^{+}03a, TFK^{+}03b, TMO04, TMT86, TvdP98, TG98, VSC01, WSG05, WH96, Ye08, YM06, YKH^{+}09, YYPZ07, ZISS04, ZLK05, van97, vdP97, BH96, DTG96, DRBR09, ESP92, HGW92, HPT89]. animation [JW89, KB92, LG92, MM93, Rob93, TT95b, TD00, WO92, YNBH09]. Animations [AM00, BG95, DDM03, DTTS08, HVAPB08, OZS08, PC94, VS09, VT94]. Anisotropic [AA06, DGF98, HMS09, HP04, KMG96, KKD09]. Anisotropy [GCP^{+}09, PWP08]. Anno [Enc98]. Annotated [How89]. Annotation [BCD01]. Announcement [Ano03a, Ano98u]. Announcements [Ano98k, Ano98l, Ano99a, Ano99b, Ano00a]. Annual [Ano87a, Ano88a, Ano91a, Ano92, Ano93, Ano94, Ano96l, Ano96s, Ano97o, Ano97q, Ano97z, Ano97-28, Ano98a, SCA04]. Anomalous [WW09a]. ANSC [Arn84]. Anti [BK01, Chr86, DS05a, GM06, Liu94, Pil85, TNF89]. Anti-Aliased [DS05a, Liu94, TNF89]. Anti-Aliasing [BK01, Chr86, Pil85, GM06]. Antialiased [PWC^{+}09]. Antialiasing [CW99, LA06, Sch01, NG92]. Any [Pas02, ML91]. Aperiodic [PGGM09b]. APEX [GS85]. Apparent [SLTM08]. Appearance [BPV^{+}09, DWL^{+}09, DDKL09, EKM01, GCP^{+}09, PWC^{+}09, RK09a, SSCO09, YMMS06, YIC^{+}09, VSD09]. Application [ABC^{+}04, AM95, BB08, Büh01, CL99, DMCP94, End84b, Haw85, LDB07, Mac84, ML91, PA06, PS95, SMTG07, BM93, CFM93, HS92, HK92]. application/graphics [HK92]. Applications [AKB^{+}95, BR07, BMD^{+}08, BF84, BGB^{+}08a, BBC^{+}05, Bry96, DDM03, DWL08, DDV^{+}02, Dia84, ED07b, FR00, HSC^{+}05, HTSFP09, JNX^{+}08, KOB^{+}08, LD03, Mik82, MMHL08, NG03a, NG03b, NMOT01, PT03, PP89, Sal96, SRH^{+}09, SHPS08, SG96b, TLG99, VG00, VCC98, VBB^{+}06, WBCG09a, WVKR08, WVV09, WKM^{+}09, XYL09, ZHC^{+}00, ZK08, vvT84, vL90, Bar92, Sas92, SGM^{+}93, TT94, Bar06, Kje92]. Applied [Des04, SKFNC97, CRW83]. Applying [Rok97, YR97]. Appreciation [Fiu01]. Approach [AE86, AM95, Ary84, Ary86, BWS03a, BWS03b, BB91, BPMG08, BBL^{+}09, Buc98, CN05, Chr86, CMT05, CAH00, DJW^{+}06, Dvt90, DDtR94, Enc82, Fau96, FM04, Gna83, GCL^{+}06, HBK02, IK00, JC94, KVLS99, KB00, LD04, LL00, LZX^{+}08, ML03, Men95, MKO^{+}08, ND94, NMOT01, OMT02, PPD07, PL94, Pic86b, PB07, RPZ02, RGTC98, Sch00, SLCZ09, TWJ06, TC94, WW87a, Wei04, WE97, XYL09, YD88, ZSP98, ZQK04, ZQQS08, ZCHM09, ZCG98, BS02, BPZ96, EZK08, ESP92, GMDW09, HCL93, NN93, SW92a].

Approximate [AHT04, CW99, SKALP05, TPBC09, VG96]. Approximated [VC04]. Approximating [KZ08, LSW09, dFSV03]. Approximation [AMSF08, BGAM04, Bou88, BB08, EBV05, Kob05, LSJK09, Mai00, RR96]. Approximations [GKD07, WB02]. April [Ano97s, BP82, Duc98, Gob96, Kje91a]. Arbitrary [And89, AO87, BV99, FS91, GM06, GP09, GE04, HS08, KHK^{+}09, Kob96, Lee99, NN89, NKF09, PSF04, RSS96, SMAB02, The02a, WD09, YF85, FS08, MLP92]. arc [GS85]. Arches [PGGM09a]. Architectural [Ais85, CLDD09, DKYN96, LFGG08, LCZ99, SC08a]. Architecture [Joo86, KH96, PLT^{+}97, Jac85, KP87, San92]. Architectures [AO86, AO87, DKS01, FG88, HMK^{+}95]. Archive [WH91]. arcs [Liu93a]. Area [HSC^{+}05, HR87, SM86, Sun92, TT97, Wil87a, YF85, ZBP99, ten82a, KJ92, NN93]. Areas [CG07b]. ARGOSI [ADS90, SGM^{+}93]. ARGOST [Ano91b]. Arithmetic [KHK^{+}09, dFSV03, NG92, dS96]. Arrangement [GS09, IMIM08]. Arrangements [EDM^{+}08, ELS08, IMIM08, YM09]. Array [TP88]. Art [Bar05, BIWG08, BDG^{+}04, BBC^{+}05, CON08, Coo05, DDM03, EWHS08, Elb99, Fra83, Hec01, xHMC09, Kau04, LHD^{+}04, LW99, MFS08, ME98, MMS^{+}05, NMK^{+}06, OLG^{+}07, RD05, Sal96, Smi85, Sor06, SKU08, SKUP^{+}09, Szy91b, TKH^{+}05, Tru84, UWP06, VBB^{+}06, WMG^{+}09, YM09]. Articulated [CZ08, DGV08, GOT^{+}07, LS09, OM01, PG08, RT08a, YLD07, BT92]. Artificial [Ter02, TMT86, DKW94a]. Artist [MAA^{+}09, RSC01]. Artist-directable [MAA^{+}09]. Artist-Directed [RSC01]. Artistic [EWHS08, KSL^{+}08, OKP^{+}08, SE02a, SE02b]. arts [CFM93]. Aspects [Ric87c]. Assembly [Ano95r, Ano97-30, Ano04b, Ano05b, Ano06b, Ano07g, Ano08d, Ano09c]. Assessment [GMSK09, Cla92]. Assisted [CSI09]. Association [Ano84a, Ano92, Ano93, Ano96l, Ano96s, Ano97o, Ano04g, Ano05e, Ano06e, Ano07l]. Astrolabes [Zot08]. atlas [GK03a, GK03b]. Atmosphere [JW97]. Atmospheric [BN08b, Wil87b]. Attribute [Gos89, SJH08]. Attribute-Specific [SJH08]. audio [DPT^{+}08]. Audition [NT95]. Auditor [Ano95e, Ano97h, Ano08a]. Auditors [Ano04g, Ano05e, Ano06e, Ano07l]. Augmented [AKB^{+}95, BES00, BWRT96, HVM^{+}08, SG97, WCB^{+}95, MG96, MC02]. Augmenting [KMN^{+}08]. August [Ano97-29, Ano97-32, BH96, Che06, DL04, Kil85, LLR^{+}04, LLD05, LLRD07, Req86, tH83a, Rob87]. August-2 [tH83a]. Ausgraph [HD86]. AUSGRAPH’90 [Mae90]. Australian [Mae90]. Austria [CG07a, Des06, HK94, HB91, Pur07]. Authentication [XYL09]. Author [Ano98m, Ano04h, Ano05f, Ano06f]. Auto [GBAL09]. Automatic [AAB^{+}96, AGM^{+}06, AS00, BTG95, BWRT96, CZ08, DTTS08, FCOL00, GK03a, GK03b, JBTS08, Koc93, LJN02, LL05, MK05, MB97, MZT09, ML03, McC83, MKO^{+}00, MSF00, ST94, SL01, SAA09, Váz07, WT09, WS09b, XJJ^{+}08, PFC^{+}05, SKSK07]. Autonomous [BBT99, PO02, RL84]. Autonomy [Col05]. Autostereoscopic [EBA^{+}09].

Avatars [VSC01]. Avoidance [Neb00]. Award [Ano95s, Bot07, Dre07, Duc06b, Duc07, Kau07, JW01]. Awards [Ano05g, Ano97v]. Aware [CC00, CWM09, DGEG09, HK09c, LAA08, MDWK08, WSLG07, YSL08, ZHM08, CLE07, LZSCO09, YCLE09]. Axes [KOB^{+}08]. Axis [BTG95].

B [DF90, FS91, GM06, GBS99, KS92, ND06, PS95, PS96a, Vas97, VMG09, YP95]. B-Spline [DF90, FS91, GM06, ND06, YP95]. B-Splines [GBS99, PS95, PS96a, Vas97, KS92, VMG09]. B2 [KFK94]. B2-splines [KFK94]. B2-splines/S-splines [KFK94]. Back [BES00]. Back-Projection [BES00]. Background [LL05]. Balance [TySK00]. Balancing [MB99, WGS04]. Ballistic [RAP08]. Balls [LAM09b]. Ballyhoo [Mur85]. Barcelona [Gob95, Jan91]. Barnsley [Jon90]. Barycentric [LS08a, RLF09, WBCG09a]. Base [DSC09b]. Based [ABC^{+}04, ABB^{+}07, Ara94, BSJ08, BB09, BB00a, BPMG04, BK03c, BK03d, BPWG07, CS00, CTSO03a, CTSO03b, Che06, CJ90, CDPS09, DS02a, Dav07, DMKP07, DO00, DG95, DJZ^{+}09, DKY98, EP09, EKM01, ESK03a, ESK03b, FCH^{+}06, FP04, FL06, Fau96, FCOL00, FML06, GMY97, GHH01, Gol85, GCL^{+}06, GBP05, HK09a, HS04, HH02, HE94, HMB08, HL03c, HL02, HCG08, HM86, HJ99, IP99, KMN^{+}05, KZ08, KFG09, Kei04, KB04, KMT03a, KMT03b, LW94, LB06, LG96, LHD^{+}04, LLSL98, LJN02, LFGG08, LLA06, LG95, MCH94, MPT98, ML03, ME98, Men95, Mil88b, Mil88a, MSK06, NAB86, NMK^{+}06, Neb00, NMP98, NLED08, OS08, OW91, PA06, PSF04, PL94, PJ94, PTB^{+}03a, PTB^{+}03b, PBP96, Rey86, RA94, RGTC98, RMS^{+}08, SWPL08, SSS02, SSO08, Sch94b, Sch00, SLS04, SSCO09, SL08]. Based [SEA08, SS96b, SLCZ09, SOG09, SKCA01, TYK^{+}09, TRSKK08, UWP06, UGB^{+}04, VPLL08, VSD09, VW08, WBP98, WM09, WWG07, Wei04, WR05, WT09, XTLP02, XGL^{+}07, XYL09, XWT^{+}08, YWB03, YN00, YSY94, ŽCO97, ZPS03, ZV09, ZCG08, ARC05, BJCO03a, BJCO03b, BWPP04, BS03a, BN08a, CLH^{+}08, CTW92, CLHL08, CLT^{+}08, CYJ02, DCOM00, DTTS08, Den03a, DSW09, DM92, Dut04, EBSC99, FFD93, FLW00, GA96, GJ02, GKS00, Got03, GGK06, GBKG04, GBP04, HKMS08, HBK02, IMIM08, IK01b, Jac85, JZF^{+}09, KMTT92, KSN08, KMG96, KJC^{+}09, hKLS00, KSK97a, KSK97b, LL00, LNS05, LLY09, LMS04, LTH08, LYP^{+}08, MG96, MFS08, MMS07, MRS08, Mum86, OOI05, ÖGG09, RZLG08, RCM^{+}01, Sbe93, SVG^{+}08, SSB08, SS08, SS09, TZF04, TO97, TPS09, TWJ06, VB00, WSE04, WHCO08, WS09a, WBCG09b, WK04]. based [YK92, YFGL09, YK06, YBS07, YWY08, ZY02, ZQQS08, hZCK98]. Bases [FS08, HS08, MKB^{+}08]. Basic [Arn08, St88, Bar92]. Basis [BK05b, DKN^{+}95, JBL^{+}06, LWP^{+}04, MKB^{+}08, RSC01, VS09, WSC06, DRBR09]. Bat [SS95]. Batched [CGG^{+}03b, GMC^{+}06, CGG^{+}03a]. Bayesian [BBL^{+}09, JWB^{+}06]. BDAM [CGG^{+}03a, CGG^{+}03b, GMC^{+}06]. Be [Hec01]. beacon [MG96]. Beam [JZJ08b]. Beauty [Pic91a]. Behavior [PP05, UT02]. behaviors [Bad93, SGC04]. Behaviour [AVF04, LD04, SEASM09, SGC04]. Behavioural [MCH94, MCT01, ZV09, TD00]. Beijing [van89b]. Belgium [LMD04, BP83b]. Beltrami [CLB^{+}09]. Bending [BW07]. Bernstein [RGTC98]. Better [DF85, Gro01, Gue82, SJ09a]. Between [BMWM01, EL01, FG04, HHS01, DMYN08, HMRSK92, KGL^{+}98, NHH97, PB90, RKC02, RSS96, ZHC^{+}00]. Beyond [Gos86, HHD03a, HHD03b, KBKŠ09, LD09, MMS09, Saw07, SD94a, SKMS06, YNBH09]. Bézier [BV96, BV99, DRS08, Hei95, NN94, RK94, RGTC98, RDG01]. Bibliography [How89, Owe86, Owe88, Owe89b, Owe94, Col93, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. Bidirectional [MMS^{+}05, RMN05, SBLD03, SvLD03, WÅ09, XWT^{+}09, ZBA^{+}07]. Bifurcation [Pic87]. Billboard [WWG07, BCF^{+}05]. Billboards [DN09]. Binary [CWA^{+}08, JD00, JD01, LLHY09]. Binding [MA85, Ric87c, SD00, Bak91a]. bindings [Spa85]. Biography [Owe87]. Biological [DKY98, Pic86a]. Biologically [SH02]. Biologically-Parameterized [SH02]. Biology [Dia84, MMHL08]. Biomechanically [hZCK98]. Biomechanically-based [hZCK98]. Biomechanics [Col05]. Biomedical [JNM^{+}09, NGB^{+}09, RGM85, YCLE09]. Biomorphs [Pic86a]. Biophysically [KB04]. Biophysically-Based [KB04]. Biorthogonal [WM09, ZQQS08]. BIPs [RK94]. Bit [DF90]. Bitmask [SS07]. Bivariate [SCM^{+}09]. Blaubeuren [Duc98]. Bleeding [RMSD^{+}08]. Blend [BK01, GW07]. Blended [DRS08]. Blending [HL03a, HL03b, HL03c, LM07, Li07, WGG99, LM93]. Blends [FG04]. Blessing [Mur85]. Blind [BB00b]. Blob [GA96]. Blobs [RTN03a, RTN03b]. BlobTree [FJW^{+}05, GLA00]. Block [Szy91a]. BlockMaps [CDG^{+}07]. Blocks [KO88]. Blue [Ano04c, SJ09a]. Blur [Wei04]. Bodies [BT95, DGV08, RKC02]. Body [EMK09, HS04, HSS^{+}09, HKG06, KSK97a, WDGT01, ZV09, KSK97b]. Bologna [ACM80]. Bombay [Mud83]. Bonsai [BPF^{+}03a, BPF^{+}03b]. Book [Ano97i, Ano97j, Ano97k, Ano98n, Ano98o, Ano98p, Ano99c, Ano99d, Ano00b, Ano00c, Ano02a, Ano08b, Cal07, How91a, How91b, How91c, How97, Lan07, Owe89a, Owe90a]. Boolean [KP87, KGMM97, PG93, WGG99]. Booleans [BF09]. Boston [LLRD07]. Botanical [OOI05, Hol94]. Boulogne [Ano97s]. Boulogne-sur-Mer [Ano97s]. Bound [Dan96]. Boundaries [DLRW09]. Boundary [AAB^{+}96, ADF85, KGMM97, PG94, YYPZ07]. Bounded [SSS97]. Bounding [DHK08, HS98, LAM09b, LCWK07, MW06, YM06]. Bounds [ED07a]. Box [McC96]. Boxing [FB94]. BOXTREE [BCG^{+}96]. Brain [JBB^{+}08, JNM^{+}09]. Branched [KNP07]. Branching [FCGA97, Wai88]. Branislav [Ano05c]. Brazilian [AR06]. BRDF [LF97, MTR08, PCF05, RK09b, Sch94b]. BRDFs [GK03b, LLSS03a, LLSS03b, MG09, NP00, NKF09, SvLD03, XS06]. Bresenham [Kuz95, Ska87]. Bright [DMHS08]. Brightness [WBEF97]. Bringing [Edm83]. Brittle [SWB01]. Broadening [Saw07]. Browser [Bov90, WH91]. Browsing [WM85]. Brush [XTLP02, XLTP03a, XLTP03b]. Brushing [IF09]. Brushstrokes [Pud94]. BSP [AMT02, CS92]. Bspline [KE97]. BTF [GMSK09, MG09, RK09a]. BTF-CIELab [GMSK09]. BTFs [GMSK09]. Bubble [Dur01, MMS09]. Bubbles [HK03a, HK03b]. Budapest [Ano97w]. Buffer [LW95, LW92, SDD^{+}92, Sun92, TT94]. Buffers [Déc05]. Building [BPM06, Fuc97, LJN02, vL90]. Buildings [HMDO05, MGAF95, Koc93]. Built [Kuz95]. Built-in [Kuz95]. Bulge [Blo97]. Bump [TCRS00]. Bundles [MSSK08]. Bundling [HV09]. Bus [ZHC^{+}00]. BVH [LGS^{+}09]. Bye [VD90]. Bye-laws [VD90].

C [Bak91a, GMC^{+}06, Ric87c]. C-BDAM [GMC^{+}06]. Cable [SL07]. Cache [BGB08b, HMS09, SS96a, YM06]. Cache-Efficient [YM06]. Caching [BG02, CLF^{+}03a, CLF^{+}03b, DDB^{+}09, GKB09, SiK– DM05]. CAD [Ais85, Arb90, BK05a, Enc82, Gol85, KH96, Kwi89, Mil90, NC99, OYSO92, Owe86, Owe87, RGM85, RL84, Sei94, van89b]. CAD-System [Enc82]. CAD/CAM [KH96, RGM85]. CAD/CIM [Sei94]. Cagliari [SP06]. Calculation [DKYN96, PA06, Gui92, MMAG93]. Calculations [Pic86b]. Calculus [Des04, van87]. Calendar [Ano97l, Ano97m, Ano98q, Ano98r, Ano98s, Ano99e, Gre97, Gre98]. Calender [Ano97n]. Calibrated [CZGF05]. Calibration [BD04, RvBWR04, WH04, WCB^{+}95]. California [LLR^{+}04, LLD05, SCA04]. Call [Ano95n, Ano95l, Ano95m, Ano95k, Ano96g, Ano96h, Ano98v, Ano98b, Ano98c, Ano98h, Ano98i, Ano07i, Ano08f, Ano97y, Ano97e, Ano97g]. Calligraphic [XJJ^{+}08]. CAM [KH96, RGM85]. Cambridge [Arn91, Cre88]. Camera [BCS96, BIWG08, CN05, CON08, DJZ^{+}09, FCOL00, HHS01, JL98, MPS05, WH04, MG96]. Cameraman [JL98]. Cameras [MRT08]. CAMP [Joo86]. Can [Ros97, Šár07]. Canal [KL03]. Canonical [EGKT08]. Capabilities [Sco02, SD00]. Capability [Bel87, Ben94]. Capture [BK03a, GBU00, GPD09, JL08, MMS07, PKG03a, PG08, PH03b, PH03a, RBB01, SKSK07, SSK^{+}05, YLD07, PCF05]. Captured [AA09, CZ09, PZB^{+}09]. Capturing [CLHL08, DWL^{+}09, HS99]. Car [GC09, RMS^{+}08, OYSO92]. Caricature [LCM^{+}09]. Carla [Duc00]. Carlo [Gob95, Gob96, BBL^{+}09, PWP08, Pic86b, Sbe97, SKFNC97]. Carmel [Arn84]. Cartographic [CJ90, VW91]. Cartography [CG02]. Cartoon [JCJ09, SDC09, WBCG09a]. Cartoons [SDC09]. Case [DCPS08, LP95]. Casted [CDG^{+}07]. Casteljau [AP92]. Casting [FQK08, HSS^{+}05, KZ08, KSN08, LA05, RS08]. Categorization [RTN03a, RTN03b]. Catmull [CLT^{+}08]. Causal [BY08]. Caustic [BAJ08, WN09]. Caustics [IDN03a, IDN03b, Jen97, PP09a, SJ09a, WS03a, WS03b, WN09]. Cave [LW99, LW99]. Cavities [BGB^{+}08a]. CCD [MG96]. CCD-camera [MG96]. ceiling [Lam09a]. Cell [BNRSV01, CKM^{+}99, NRJS03a, NRJS03b, St88, FR92, HDS03b, HDS03a, LCD09a, MMFE08]. cell-and-portal [HDS03b, HDS03a]. cell-decomposition [FR92]. Cell-to-Cell [NRJS03a, NRJS03b]. Cell-Visibility [BNRSV01]. Cells [BPWG07, St88]. Centered [Enc98, EMK09]. Centrality [VW08]. Centroidal [VC04]. Century [Hec01]. Cerebral [NGB^{+}09]. Certification [BP82, BP83b, Gna83]. CFD [SCD05]. CG [MH07, van89b]. CGF [Ano07h]. CGForum [Ano06c, Ano04c, Ano05c, Ano07b, MTPS08, SJ09b]. CGI [Ano88b]. CGM [Duc89, BHM87, CP88, MN87, Mum89]. Chairs [BWS05]. Challenge [tHS90, Gue82]. Challenges [BR07, BBC^{+}05, Bry96, Enc98]. chance [Bak90]. Change [BCWR08]. changes [CS92]. Changing [AMS09]. Channel [MZT09]. Channelling [Kin95]. Chaos [Jon90, PGR96, Pic87]. Chapter [Ano95p, Ano97z, Duc89, Pin85]. Character [DAP08, DTTS08, HK09b, ITYI09, JL08, yKL08, MAA^{+}09, OTH02, OZS08, PG08, RAP08, RB03a, RB03b, SNI06, TGS96]. Characteristic [WDR09]. Characteristics [MTKO02]. Characterization [KK08]. Characters [BCH^{+}95, GOT^{+}07, KRFC09, LNS05, MH07, RHC08, TG98]. Charles [Ano06c]. Charts [Coc83, JKS05, TPBC09]. CHC [MBW08]. Chebyshev [FB94]. chemistry [CRW83]. Chia [SvZ95]. Children [Bur95]. China [van89b]. Chinese [AHM09, WS01, XJJ^{+}08]. Chisl [WGK88]. Choosing [Kje83]. Christopher [Ano06c]. Chromatic [WW09b]. Chunks [GW05]. Ciechomski [Ano05c]. CIELab [GMSK09]. CIM [Sei94]. Cinematic [OKP^{+}08]. CIP [KySK08]. Circle [Kuz90]. Circle-Drawing [Kuz90]. Circles [RF96, HCL93]. Circular [NW01, Liu93a]. Cities [CLDD09, PGGM09a, WMWG09, TD00]. Civil [GSHN94]. Clark [CLT^{+}08]. Class [BB07, SNLH09]. Classes [Zar06]. Classification [DKC00, ES94, HW91, JL06, JNX^{+}08, LAA08, SGM^{+}93]. Classroom [Kel86]. Clausal [Mil90]. Clavien [Ano05c]. Climbing [PWH98]. Clinical [Bar05]. Clip [HH90, Sal96, Zac89, HR88, Hub93]. Clipping [And89, Hei95, HL02, Kuz95, Mat85, Nie95, NN94, TP89, Day92, Kra92, Ska96]. Closed [ND06]. closing [MK08]. Cloth [MBCN09, PH03b, PH03a, VSC01, Ye08]. Clothes [CMT05]. ClothesPEG [Cla92]. Clothing [DJW^{+}06]. Cloud [EBV05, HL01, JWB^{+}06, KZ04, KB00, SWK07]. Clouds [LSW09, MMG06, SWG08, WWG07, BCF^{+}05, SSK93]. Clues [HDM98]. Cluster [Dur01, FQK08, FK09, GSA03a, GSA03b]. Clustered [CZCE08, SSS02]. Clustering [Ano99m, CDP95, DHS04, HVAPB08, HDSD99, LJN02, MSF00, WS09b, ZYQ^{+}08, GH96]. Clusters [Gar09, LL09, RL09, SSS97]. CMYK [DDPL00, DDV^{+}02]. Co [BSG^{+}95, IF09]. Co-located [IF09]. Co-operative [BSG^{+}95]. Coarse [BC01, LD06, Lov06, SLS^{+}06]. Coarse-to-fine [BC01]. Coarsening [DSC09a, VC04]. COBRA [VS09]. Code [TA08, ŽCO97]. Code-Based [ŽCO97]. Codec [LAD02]. Coded [BE95, SSK^{+}05]. Coding [Got03, KPRW05, MA91, Sch84, TW96, SM84]. Codings [SC84]. Cognitive [Bur95]. Coherence [HHS01, ZY02, HMRSK92]. Coherence-based [ZY02]. Coherent [BWPP04, FMS01, MW06, MBW08, SiK– DM05, WSBW01]. Coliseum [GBU00]. Collaboration [Coo05, ZHC^{+}00]. Collaborative [AKB^{+}95, BDG^{+}04, CWKS00, IF09, LLP00, cLCtLL98, PLT^{+}97, SNLW01, Wat96]. Collections [IF09]. College [Che06]. Collision [BT95, BGAM04, CJC^{+}09, DO00, Dur01, GGK06, JFSO06, KAAT03a, KAAT03b, KŽ05, KHH^{+}09, KZ04, Neb00, NG03a, PG95, RKC02, SOM04, TKH^{+}05, TPBC09, VCC98, VT94, WLML99, WP04, ZY02]. Collision-Free [KAAT03a, KAAT03b]. Collisions [BW00]. colon [HGW92]. Color [Ano02b, AW07, BFH^{+}98, BE95, Čad08, CWM09, DDPL00, Fis98, GVWD06, IK00, hKAC07, MG05, MMTH09, NMP98, RMSD^{+}08, RP98, RTN03a, RTN03b, SSK^{+}05, WHCO08, WVVV08, XM09, XTJ^{+}07, OYSO92]. Color-Coded [BE95, SSK^{+}05]. Color-to-Grayscale [Čad08]. Coloring [PP09b, RGW05]. Colormap [BS98]. Colors [hKAC07]. Colour [Ano96d, Ano98t, Ano00d, AEL^{+}82, AGM^{+}06, DCPS08, Den86, LR88, MTCT84, Mur85, OW91, Wil84, WM85, WW88, Wil06b, van87]. Colouration [Pat95]. colours [Sch94a]. Combination [SKCA01]. Combinations [KGMM97]. Combined [BT92, SSSK04b]. Combining [BTG95, FHL^{+}08, RSK90, CCFM08]. CoMEdiA [San92]. Comes [Pic91a]. Comic [WB02]. Command [Her82]. Commands [KMB94]. Commodity [YWB03]. Common [HB00]. Communication [van82]. Compact [LD08b, NNSK99b, WSC06, XS06]. Comparative [NMOT01, vKP06]. Comparing [STD09]. Comparison [CRW09, HV08, HTSFP09, LD08a, SC84, SDD^{+}92]. Competing [SLS^{+}06]. Competition [Ano95f, Ano95g, Ano95h, Ano96e, Ano96f, Ano97c, Ano97t, Ano97-41, Ano97a, Ano97b]. Completion [LYP^{+}08, SDK09]. Complex [BG08, CDP95, CATM09, CLF^{+}03a, CLF^{+}03b, DG95, FML06, GSA03a, GSA03b, Kle06, KHM09, LD04, MB99, OPC96, ORDP96, PGGM09a, Ric87a, Ric87b, Ros97, Sad09, SSS02, Szy91a, Szy91c, TL01, WS02, WBCG09a, AVF04, BSV92, GH96]. Complexity [FdABS99, FBW01, FBP08, PG93]. Components [AM00]. Composed [LC99]. Composing [PC92, SSB08]. Composite [CG07b]. Compositing [BCS96, FW99, GVWD06, GCL^{+}06, LW95, SS08, Wil06b]. Composition [CN05, EKM01, Sch00, dBD^{+}92]. Compound [BD08, GBD09]. Comprehensive [VB99]. Compressed [Cot85, GMC^{+}06]. Compression [BJCO03a, CC06, CH09, DDPL00, DDV^{+}02, DGGP05, DSW09, GMC^{+}06, GPD09, GMSK09, HB96, HSS^{+}09, HKMS08, ILRS03a, IP99, Ise01, IK00, KCL06, LS09, LGK97, MGS07, MMP08, MMG06, MCHAM08, NS01, OS08, PD04, RK09a, TRS03a, TRS03b, TWC^{+}09, VCP09, VS09, WD09, WS09a, YN00, ILRS03b, LCL^{+}06]. Compressive [SD09]. Computation [BPZ96, BS98, CLM09, Elb95, FQK08, Fau96, KGMM97, LCWK07, MNR94, NHH97, OLG^{+}07, PK08, SLSK07, SOM04, SSSK04b, TW97b, TPBC09, Ure00, YLL^{+}09, CD94, Sbe93, TW97a]. Computational [AAB09, Bur95, DMHS08, FHL^{+}08, Kol08, LAA08, NSGP06, SD09, SSCO09, Vel99, de 97, GY93, Vel93, Ano08e, Ano08f, NSGP06]. computations [PM93]. Computed [FBW01, FAVM09]. Computer [AHKS94, AR06, Ano82, Ano84a, Ano84b, Ano88b, Ano92, Ano93, Ano95a, Ano96l, Ano96s, Ano97o, Ano98f, Ano03c, BDŠ84, Bar06, BCS96, BF84, BB88, Bou90, BPB^{+}04, Bro90, CB09, CON08, Chr86, Dav07, Den86, Dia84, DJ88, Duc91, Duf88, Dun91, Dut04, Edm83, FS85, Fra83, FR00, GS06, Gal84, GC09, Gos86, GSHN94, HHS01, Han05, HHS89, Her89, Hew84a, Hew90, KH95, Kel86, Kil85, KHIK01, Kje89, Kje90, Kje91b, Kje91c, Kje91d, Kol08, KW05, LL05, Lei94, Mac94, MGJ06, Mar82, MB97, May99, May00, MB99, Mik82, NYTN87, NMK^{+}06, NMNP98, OIST91, Owe86, Owe87, Owe88, Owe89b, Owe90b, Owe92a, Owe92b, Owe93, Owe94, Owe95, Pat95, Pic86a, PNR89, PH96, PT88, RMA88, Req86, RP98, SCA04, Sal96, Šár07, Sch84, Sei94, Sni95, SB99b]. Computer [Sta06, SHPS08, SM84, Suz89, Szy91b, TMT86, TvdP98, Tru84, Tur84, VBB^{+}06, Zar06, tH83b, tHS90, vD98, van82, AS92, BJ94, CRW83, Jac85, KB92, LG92, MMAG93, Sch94a, TC93, Ano03b, Bon85, BH96, DP93, DS04b, TT95b]. computer- [Sch94a]. Computer-Aided [Owe88, Owe89b, Owe94, Owe90b, Owe92a, Owe93, Owe95]. Computer-Generated [Chr86, SB99b]. Computerized [DF90]. Computers [Smi85]. Computing [ABC^{+}04, Ano97s, Ano97f, Arn08, BIWG08, CCI08, CS92, DHvOS00, Duc98, Enc98, GGRZ06, HMS09, HCG08, JEO00, Le 90, PN97, SvZ95, Ano98j, PH91]. Concepts [VG00]. Conceptual [Mac85]. Concerning [Duc85]. Concise [SOG09]. Concurrent [Mil87, SJH08]. Concurrently [FCP^{+}90]. Condensation [BNC96]. Conditions [AMS09]. cone [SAE93]. Cones [Ano04c]. Conference [Ano92, Ano93, Ano94, Ano96l, Ano96j, Ano96s, Ano97o, Ano97r, Ano97-28, Arn89, Aus91, Bar06, Cre88, DJ88, Duc89, HHS89, HP84, Kid84, Mum90, Pin85, Req86, Sei88, Suz89, TB84, Van85, Wat88, ACM80, Enc81, Van80, WG82, tH83a, van89b, Ano85, Ano86, Ano87a, Ano88a, Ano89, Ano91a, Ano96k, Ano96p, Dun91, Mae90, Mat86, Neu06]. Configurable [AHR84, WGK88]. Configurations [BŠ90]. Conflict [JD00, JD01]. Confocal [FHL^{+}08]. Conformal [BCGB08, MTAD08]. Conformance [Ano88b]. congressi [ACM80]. conic [GS85]. Conical [BJCW09]. Conics [Her89, QSW92]. Connected [BCN03b, BCN03a, RR94, SS91]. Connectivity [AD01, Got03, KBS00, MGS07]. Consensus [YWB03]. Consensus-Based [YWB03]. Conservative [BNRSV01, COFHZ98, JEO00]. Conserving [MW06]. considerations [DDR93]. Considered [WGS04]. Considering [IIS09, MTKO02]. consistency [SNLH09]. Consistent [AW00, DMKP07, Ger92, Sug94, DF93]. Constant [Got94]. Constant-Time [Got94]. Constrained [BAU05, Dwy09, LD08c, LBH^{+}01, PBP96, YYPZ07]. Constraint [EP09, HHS01, LM96b, LWP^{+}04, MCH94, FFD93, LM96a, Pin92, VR95]. Constraint-Based [EP09, MCH94, FFD93]. Constraints [ESG01, IIS09, JL98, LVT08, Mil90, Lam09a]. Constructing [CFFP84, LM07, SDB99, TIS^{+}95, TLG99, Wal87, CTW92, Rap92]. Construction [AW07, BV96, CDP95, Duf88, HB94, JC94, LGS^{+}09, MNP08, SSK07, SG08, Str84, TGS96, TG98, WTHS04, BM93, Vol93]. Constructive [CT00, HL03c, VPP^{+}04]. Contact [Fau96, GOT^{+}07, HS04, KGL^{+}98, OTSG09, ST08]. containing [BP93]. Content [BDA^{+}09, CCP09, ZHM08]. Content-Aware [ZHM08]. Context [BJCO03a, BJCO03b, BPBD08, CLME09, DCOM00, DKW94b, FBT99, Her84, MDWK08, MKO^{+}08, SSA^{+}08, WSLG07, WVKR08, WKM^{+}09, YCLE09, ZK08]. Context-Aware [WSLG07, YCLE09]. Context-based [BJCO03a, BJCO03b, DCOM00]. Context-Frames [Her84]. Context-Preserving [WKM^{+}09, ZK08]. Contextual [KMB94]. continuities [Vel92]. continuity [Bar92]. Continuous [BW09, GCL^{+}06, KO88, KHH^{+}09, LO95, RKC02, SG03, SKCA01]. Continuously [CZGF05]. Continuum [HMT01, TPS09]. Continuum-based [TPS09]. Contour [Coc83, GP87, JC94, LZ07, RWP88, TW96, DMP93]. Contouring [McC83, Sab86, SW05]. Contours [CMH^{+}01, OPC96, SPOK95]. Contrast [DMAC03a, DMAC03b, GVWD06, KMS07]. Contributions [Ano98b, Ano98c, Dre07, Ano97y]. Control [ADJ^{+}01, BDŠ84, BK03a, BK03b, CON08, CZGF05, DAP08, EP09, FCH^{+}06, FS91, IP00, LNS05, MS96b, OKP^{+}08, OM01, ST94, SG03, SKC01, TYK^{+}09, ZV09, BT92, HPT89, KS92, MS96a, STM93]. Controllable [KFK94, SWB01]. Controlled [BW07, FL06]. Controller [KSK97a, KSK97b, JW89]. Controlling [HTH96]. Controls [DN08, HL03a, HL03b]. Convergence [Fol98]. Conversation [MH07]. Conversion [CW99, DTTS08, DKS01, LO95, Rok97, SLTM08, YR97, Che97, VG96]. Conversions [Čad08]. Converting [WW87b]. Convex [AMSF08, BG01, Dan96, Day88, Day90, EL01, Sug94, SN86, WBG07, CD94, HMRSK92, Rap92, Ska96]. Convolution [MS98, TZF04]. Convolutional [Blo97]. cooperation [Des06]. Cooperative [DGC^{+}98, GY93, PW93, San92, ST93]. Coordinate [MSK06, SWB98]. Coordinates [FS08, HK09b, HS08, JC08, LS08a, MKB^{+}08, MM08, Nie95, RLF09, WBCG09a, ZYQ^{+}08, ZCQ^{+}09]. Copenhagen [TB84]. COPERNICUS [ZK08]. copper [Lei94]. Core [CY89, FCGW02, FK09, SVG^{+}08, BBS^{+}09, ILRS03a, ILRS03b]. Correct [AHM09, TIS^{+}95]. Correction [KTMN07, MZT09, TRSKK08, MMTH09]. Corrective [HMYS01]. Correlated [SSSK04b]. Correlation [CAM08a, MZT09, NAS07]. Correspondence [LSP08, ZSCO^{+}08]. Corresponding [MH07]. Cosserat [Pai02]. Cost [FBW01, ML03, MS93, RCM^{+}01]. Cost-model [ML03]. Countershading [KMS07]. Course [Kil85, Las84, Zar06]. Cover [Ano95f, Ano95g, Ano95h, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97t, Ano97-41, Ano04c, Ano05c, Ano06c, Ano07b, IIS09, Lut02, MTPS08, SJ09b]. Covering [ND06]. Coverings [JFSO06, KNP07]. Covers [TNF89]. CP3 [BG01]. CPU [MFS08]. CPU-based [MFS08]. CPUs [KHH^{+}09]. Cracks [VPLL08]. Craniofacial [PZY08]. Crayons [RMN05]. Creating [BAHS06, LCZ99, LD98, MS98]. Creation [ML03, MSF00, SG03]. Credit [Ano97-31]. CREST [Kil85]. Critical [CLM09, TIS^{+}95]. Cross [BV09, GVWD06, MZT09, MB08, POB^{+}07, Spe91, TT94]. Cross-Channel [MZT09]. Cross-Indexed [Spe91]. Cross-Sections [BV09, MB08]. Crowd [AVF04, LD04, PPD07, SGC04, UT02]. CrowdBrush [Ano05c]. Crowds [LCL07, NAS07, OCV^{+}02, PPD07, RD05, TLC02]. Csaba [Ano04c]. CSG [BR96, Ger92, JA95, Pat89, Wie96, WGG99, ZC95]. Cubes [DZTS08, MJC01, RW08, RHv95, SW05, The02b]. Cubic [EMK09, ND06, Guo93]. Cubical [BRS01, HWC^{+}05]. CUDA [STK08, SS09]. Cues [FMB^{+}00]. Culling [ASVNB00, BKES00, BWPP04, CATM09, MBW08, SBW06]. Cultural [Arn08, CL99, DCPS08, PSK09, VPP^{+}04, Zot08]. Current [PF90]. Curse [Pas02]. Curvature [ADS06, BCGB08, CCSLT09, DGEG09, ESP08, EP09, HTH96, WB01]. Curvature-Domain [ESP08]. Curve [BYP95, BV96, BF84, CK84, Elb95, GA98, JW95, LBD^{+}08, RK94, SB99a, SLSK07, SL89, ZCG98]. Curve-skeleton [SLSK07]. Curved [RvBWR04, BP93, SAE93]. Curves [BNRSV01, DCOM00, EBV01, Fiu95, FB94, HM83, Haw85, Hew90, KM83, Nas03, ND06, Pic91a, VW95, YR91, dFSV03, AP92, BSV92, KFK94, MNR94, MS93, SW92b, Vel92]. Curvilinear [YM09]. Customising [PCS94]. customizable [TC05]. Cut [FP94, LAM09b]. Cutaway [DWE03a, DWE03b]. Cuts [BMG99, GCMS00]. Cutting [DLC05, JLCW06]. Cycles [DGEG09]. Cyclide [FG04]. Cylinder [RSVP02]. Cyprus [Ano07j]. Czech [Gob96].

D [GL94a, GL94b, Ais85, AS96b, AA09, AD01, ANF97a, ANF97b, AMSF08, BW09, BG95, BP98, BCG^{+}96, BB00a, Bel87, BG89, BR02, BGK^{+}96, BTG95, BG08, BYB09, BH93, BCD01, BCF94, BMWM01, BNRSV01, BB08, CCFM08, CTSO03a, CTSO03b, CLME09, CLB^{+}09, CMPS93, CMS94, CD94, CYJ02, CKSW08, Day90, DCPS08, DGF98, DH93a, DHH02, DF90, EWHS08, EFGS96, FFD93, FG88, GD96, GL94a, GL94b, GE98, GTS86, Han97, HLM97, HE94, Hub93, IIS08, IP99, IEGC08, IK01a, Jac85, JBTS08, Joo86, JKS05, Kil82, Kil85, KK07, KQWM08, LGMT00, LJN02, LVT08, LCM^{+}09, LCWK07, MFT02, MSM^{+}08, MK99, MDC93, MB08, MGAF95, NK99, NCKG00, NMOT01, O’H02, OMT02, OPC96, PP89, PGG^{+}09, Pri85, Rob93, RCM^{+}01, RWP88, RL09, Ros97, Sab82, Sam93a, SW09, Šár07, SSB08, Sco02]. D [SLSK07, SW92b, Sni95, SB99b, SP03a, SFWS03a, SP03b, SFWS03b, SS95, SSB05, SOM04, TL01, The02a, TRS03a, TRS03b, TTW90, TW97b, TVD09, TSH01, TGS96, VPLL08, VCDF95, Vel93, VW95, VB99, WG93, WHL^{+}04, WL08, WWG07, Wat96, WMWG09, WTHS04, WTHS06, WO92, WOBT09, WLSG03, YD88, ZHC^{+}00, ZPS03, ZLK05, van90, vJB85]. D-Charts [JKS05]. D-clip [Hub93]. D-Reconstruction [BB00a]. Dallas [Rob87]. Dancing [SNI06]. Dancing-to-Music [SNI06]. Darmstadt [Enc81, BP82, tHS90]. Dart [CJW^{+}09]. DaScript3D [Sam93a]. Data [AS96b, AGDJ08, BB91, BTG95, BBS^{+}09, CCS95, CS99a, Cal96, CJC^{+}09, CMS94, Col93, CMT05, CPK09, DAP08, DTA94, DMCP94, DKC00, FK09, Frü94, FH09, GBU00, GSGC08, GMDW09, GLW96, GBKG04, GJL^{+}09, GPD09, HKD^{+}08, HWC^{+}05, HV08, IP99, JBB^{+}08, JBL^{+}06, JNM^{+}09, JC08, JC94, JL08, KZ08, KK07, KHH^{+}09, Kle06, KSS97, LPK09, LF97, LL09, MG87, Mar95, McC83, MKO^{+}08, MH00, MSK06, NGB^{+}09, NCKG00, NS01, NKP93, PC94, PS95, PS96a, PD04, PSK09, PZB^{+}09, RSTK08, RRRP08, RNV07, RSS96, SB99a, SHLS02, SNLH09, SW04b, TIS^{+}95, TPBC09, VHB08, Cot85, DKW94a, FT93, Jon96, JR08, MK08, MM93]. Data-Dependent [SW04b]. Data-Driven [CMT05, PZB^{+}09, MSK06, RNV07]. Data-Faithful [KK07]. Databases [BB91, SDB99, FT93]. Dataset [BGK^{+}96]. Datasets [LC99]. David [DCPS08]. Daylight [QNTN03a, QNTN03b, ZBP99]. Dead [COS95]. Dead-Zones [COS95]. Deblurring [AAB09]. December [Duc86, tHS90]. decision [MK08]. Declarative [BB91, DH93a]. Decoding [DSW09]. Decomposing [SN86]. Decomposition [EL01, Got94, GFW^{+}06, HWAG09, KE97, RK09a, STK02, FR92]. Decompression [CC06, GMC^{+}06, KCL06, ILRS03a, ILRS03b]. Decoupled [AW00]. Dedicated [FBT99, NYTN87, TD00]. Deep [YK08]. Deferred [AMT02, GBP04, WN09]. Define [MG87]. Defined [PA06, RHv95, dD85, BR96]. Defining [BP82, RSS96]. Definition [DDtR94, Ste85, WM85, WW87b]. Deflection [MG95]. Defocus [BD07]. Deformable [BCH^{+}95, BNC96, CZ09, GVV05, GGK06, GW05, HE01, KŽ05, NMK^{+}06, OTSG09, SP97, SE04, STK08, TKH^{+}05, TG98, Vas97, WDGT01]. Deformation [AB97, BPWG07, BBP09, CLME09, CWK07, EP09, GOT^{+}07, xHMC09, ITYI09, KFG09, LC09, ON05, PGGM09b, PZB^{+}09, RLN06, SHB07, WSLG07, WBCG09a, WBG07, YBS07, ZRKS05, ZSCO^{+}08, hZCK98]. Deformation-Driven [ZSCO^{+}08]. Deformations [BiA06, BPWG07, EP09, HAWG08, KBS00, Lov06, Mai00, OHBKH09, OTSG09, PBP96, TPS09, WSBZ08, WRS01, YK06, YBS07, KMTT92]. Deforming [AKP^{+}05, ATBG08, CZ08, CCI^{+}07, HAWG08, LLG97, LSP08, SG08, SWG08, ZSCO^{+}08]. Degradation [DO00]. Degree [GM06, QSW92]. Delaunay [BYB09, CMPS93, GKS00, NMOT01, dCTAD09]. Delta [DM92]. Demand [SG96a]. Demand-Driven [SG96a]. Demographic [vvT84]. Demonstration [Duc89]. Dendritic [JNX^{+}08]. Denmark [TB84]. Dense [DCPS08, LHD^{+}04, MMG06, RR00]. Densely [COFHZ98]. Density [EBV05, HKD^{+}08, BLS93]. Dental [SG08]. Dependent [CKB04, DKW94b, ED07a, ESV99, ESC00, ESK03a, ESK03b, FM04, GLW96, MTR08, RBG08, SW04b, TSH01, WTHS06]. Depiction [SKMS06]. Deposition [RMN05, RPP93]. Depth [BG08, FBP08, KTMN07, KS07, LKC08, LSP08, PG08, TRSKK08, ZCP07, Déc05]. Depth-of-Field [KS07, LKC08]. Dequantization [hKAC07]. Derivation [FAT07]. Deriving [Sch94a]. Descattering [FHL^{+}08]. Description [BŠ90, DG97, BSV92, Sam93a]. descriptors [CCFM08]. Design [AR94, AKB^{+}95, Ano98d, BF84, BPF^{+}03a, BPF^{+}03b, Coh95, DJW^{+}06, DKN^{+}94, DKN^{+}95, DKYN96, Gre94, GC96, Haw85, IIS09, KKD09, LLSL98, MJ98, MMP09, ND94, NC99, OOI05, Owe88, Owe89b, Owe94, PCS94, PL94, PP05, SL01, WH04, WOBT09, WT09, XLTP03a, XLTP03b, DH93a, HHB93, KSH92, MS93, OYSO92, Owe90b, Owe92a, Owe93, Owe95, WBCG09b, dBv93, Ano98d, PB95]. Designing [AA09, Ara94, MOT99, The02a, vJB85]. Desktop [SS95]. Detail [ASVNB00, BPBD08, IOI06, MB97, OCV^{+}02, SW08a, SG03, YSL08, BS02]. Detail-In-Context [BPBD08]. Detailed [DCPS08]. Details [CCS95]. Detecting [YM09]. Detection [BT95, BB00b, BBW^{+}09, BGAM04, BCWR08, CJC^{+}09, GGK06, HBK02, JBTS08, JFSO06, KŽ05, KHH^{+}09, KZ04, PG95, RKC02, SVG^{+}08, SWK07, SOM04, TKH^{+}05, TPBC09, VCC98, VBP^{+}09, VT94, WB01, WLML99, WP04, YM09, ZY02]. Determination [JD98, KGL^{+}98, SZG93]. Determining [BBMR88, SN84]. Deterministic [SKS09, BPZ96]. Developable [BW07, JKS05, TPBC09]. Developing [BB07, KJC^{+}09, PCR89, WT93]. Development [KM83, NM91, OIST91, Sab82, VPLL08, BFTL82, CTW92, EFGS96]. Developments [Bou88, Bro90, LD03, de 97, CRW83]. Device [ADJ^{+}01, BP82, HR85, van90]. Devices [BCF94, Dvt90, FR00, LD03, vt87]. Diagnostics [PSK09]. Diagram [OPC96, YLL^{+}09, DZM08]. Diagrams [RL09, VC04, WKM^{+}09]. Dialogue [Gre84, MSWK02, St88]. Diamond [WD09]. Dick [Wil06a]. Diego [SCA04]. Dietrich [Ano07b]. Difference [BG09, GMSK09, NMP98]. Differences [ZK09, Rap92]. Different [BMWM01, OZS08, SC84]. Differential [Des04, Sor06, Gui92]. Differentials [CCSLT09, CLF^{+}03a, CLF^{+}03b, CLB^{+}09, LSW09]. Diffuse [GSA03a, GSA03b, Hei01, KJ92, SK99]. Diffusion [CZCE08, GBAL09, KT97, LDR09, LSW09, SSO08, SOG09, SKS09]. DiFi [SOM04]. DIGIS [dBv93, vL90]. DIGIS-a [dBv93]. Digital [BDFG07, BB88, COS95, DGEG09, DKW94b, KT09, LSJK09, NT95, XTJ^{+}07, XYL09, YGL^{+}09, BM93, CS93, Lev99]. Digitised [AO89]. digitising [VW91]. Digraphs [BD08, GBD09]. Dimension [Pas02, Pic86b, RGW05, WD09]. Dimensional [BMG99, GKS00, KySK08, LD06, MAA^{+}09, SS96a, TMT86, ZR96a, AHKS94, BDS^{+}03, Day92, ILRS03a, ILRS03b, SD94b, SNLH09, ZR96b]. Dimensionality [Pas02]. Dimensions [HS94, Nic85]. Diorama [AW07]. Direct [AGG^{+}08, BBP09, CAM08a, CAM08b, DQ00, FFD93, RSK06, SGS05, SPH^{+}09, SSFS06, Sni95, SSSK04b, WÅ09, ZCG98, van90, vL90, BT92]. directable [MAA^{+}09]. Directed [Bov90, HV09, RSC01]. Directional [BKES00, GKB09, SPH^{+}09, ABB^{+}07]. Directions [LF97, FLBS07]. Disambiguation [FMB^{+}00]. Disc [RNtH03]. Discontinuities [Thü01, TC05]. discontinuous [STM93]. Discovering [Ano98w]. Discrete [AGDJ08, AM02b, BB08, CLM09, Des04, EBV05, GW07, HSS^{+}05, MDP08, RL09, SWPL08, SW08a, SC95, TIS^{+}95, TW97b, Thü01, VPLL08, YGL^{+}09, BLS93, FT93, GGRZ06, TW97a]. Discretization [BT98, LBT92].

Discretized [BKES00, VT94]. Discussed [PNR89]. Disk [LD08a]. Dispersion [WW09a]. Displacement [AM02b, BK03c, BK03d, CC08, DKS01, LYP^{+}08, LP95, PHL91, SKU08]. Displacements [PBP96]. Display [AGG^{+}08, AG06, AMS09, BG01, Hea90, Hei01, PH87, Pat89, Pil85, SD94a, SL89, SF83, TTW90, Jac85]. Displaying [Coq85, DMAC03a, DMAC03b, Hei95, NN94, PMW86, SK86]. Displays [CZGF05, Den86, DMHS08, HKD^{+}08, MG05, MS96b, NGB^{+}09, Pic86a, RvBWR04, WO94, YMMS06, MS96a]. Dissection [SSA^{+}08]. Dissolve [GVWD06]. Distance [CCI08, KZ04, LDR09, MGS07, MRS08, OZ09, RLF09, SOM04, SKALP05, TPBC09, WP04]. Distance-Ranked [MGS07]. Distinctive [JBTS08]. Distortion [Ara94, BGI08, LYP^{+}08, MDP08]. Distortion-Free [BGI08]. Distortions [KLD^{+}09]. Distributed [AKB^{+}95, Ano91b, BDG^{+}04, DGC^{+}98, LGK97, Mil88b, MO08, NNB97, SG96a, HK94, Kin92]. Distributing [HHD03a, HHD03b]. Distribution [CLF^{+}03a, CLF^{+}03b, FK09, IK00, KSS97]. Distributions [LD08a]. Dither [VB00]. Dithered [BFH^{+}98]. dithering [CFM93]. Diversity [PZY08]. Diving [WH96]. division [FND92]. do [VR95]. Doctoral [Kje89, Kje90, Kje91b, Kje91c, Kje91d, Kje95]. DocuBurst [CCP09]. Document [CCP09, IF09]. Documented [KCB97]. DOF [SKSK07]. Domain [ESP08, GP09, JA95, LGK97, OMT02, PP09b]. Domains [BV09, DSC09b, FS08]. Dominant [LDR09, MK06]. Doo [SMAB02]. Doodle [BiA06]. doTS [GBKG04]. Douglas [VW90]. Down [Sla88]. Draping [WOBT09]. Draw [SS91]. Drawing [Ben94, BCD01, Bur95, JCJ09, Kuz90, Liu94, Sla88, SP03a, SFWS03a, SP03b, SFWS03b, SDC09, WBCG09a, FND92]. Drawings [Bij87, Bur95, DHvOS00, LFGG08, PKL88, SC08b, WG93]. drawn [SDC09]. Dressed [CMT02, ZY02]. Driven [AD01, CMT05, GH97, MBGS01, MF00, PZB^{+}09, Red96, SG96a, SWB98, SBW06, SB98, ZSCO^{+}08, ZV09, BGB08b, CAE08, LCL^{+}06, MAA^{+}09, MMHL08, MSK06, RNV07, SKC01, TMO04]. Driver [HR85]. droplet [MMS09]. Drypoint [TMO04]. DSV [Ano98d, Ano97e]. DSV-IS [Ano98d, Ano97e]. Dual [DWL^{+}09, PA06, SW05, SD09, ZQQS08]. Dual-microfacet [DWL^{+}09]. Duality [BV96]. Dublin [Che06, HP95]. Due [IDN03a, IDN03b]. Dunhuang [LLP00, LW99]. Dupin [FG04]. Dürer [Jon90]. DVR [LLHY09]. Dyadic [VGSS04]. Dye [LTH08, Wei04]. Dynamic [AES94, BW00, BDA^{+}09, BRS01, BD08, CGG^{+}03a, CGG^{+}03b, CYJ02, CG07b, Fis98, FK09, Gar09, GMC^{+}06, GGG08, GJW08, HMT01, KK08, KBS00, KMS05, MKV09, NS09, NKF09, ORDP96, OM01, PLPB07, PB07, RIF^{+}09, RMSD^{+}08, SS00, Sha97, SC08b, SSK07, SN08, SG96b, VPLL08, WHL^{+}04, WDGT01, YMMS06, YIC^{+}09, BH93, CS92, Vol93]. Dynamical [RP01, Sta06, GJ02]. Dynamically [AMS09]. Dynamics [Ai98, Col05, Dur01, PGR96, Sta97, WVKR08, CC93, GY93].

East [Ano97z, Ano97-28, AEL^{+}82]. Easter [Bro90]. Easy [GCL^{+}06, JLCW06]. ecological [WY92]. Economic [Bel87]. Edge [ADF85, HV09, LAA08, PG94]. Edge-Aware [LAA08]. Edge-Face [ADF85]. Edges [CLM09, IHS02, MBCN09, PO85, WG09]. Edit [XWT^{+}09]. Editable [MG09]. Editing [AG06, BK05b, BN08a, Čad08, FAT07, GA98, IOI06, KL08, KB98, LW95, Lee99, LYP^{+}08, MZT09, MBW^{+}05, OOI05, Sal96, SEASM09, SG08, Ste84, WHCO08, XM09, XWT^{+}09, BT92, CTW92, DOS93, LW92, San92, ST93, XGL^{+}07]. Edition [End84a]. Editor [cLCtLL98, Mun83, Wal87]. Editorial [AR06, Ano01a, Ano01b, Ano03d, Ano04a, Ano05a, BWS05, CS97a, CS97b, CS97c, CS98b, CS98c, CS98d, CS99b, CS99c, CS99d, CD00a, CD00b, CD00c, DS01, DS02b, DS02c, DS02d, DS03, DS04a, DS04b, DS05b, DS05c, DS06a, DS06b, DS06c, DS07b, DS07c, DS07a, SG09, SC97, CFT86]. Editors [SD94a]. Education [Ano97d, Ano97-29, Kol08, Tur84, BH06, DS09]. EEC [BP82, BP83b, Duc82a]. Effect [RAP08, TTW90]. Effective [FAT07]. Effects [Buc96, GSMA08, IDN02, JZJ08b, KW05, RMSD^{+}08, Sta97, SKUP^{+}09, ZR96a, MMS09, ZR96b]. Efficiency [Sch84, SM84]. Efficient [AEW90, AKP^{+}05, ATBG08, AHL^{+}06, BW09, BPM06, BT98, BBT99, CAH00, DSW09, FP94, FA87, FW99, GH96, GKD07, IDN02, JEO00, KHM09, Kob03a, Kob03b, KK02, Kuz90, LPK09, LeYTM08, MFS08, MK99, MKB^{+}05, Nar95, NB94, RR94, SHLS02, SWK07, SC84, SSFS06, SC08b, SDB97, SSG^{+}00, SBW06, TW96, VT94, WPG02, WC05, WSG05, YM06, ŽCO97, ZPS03, Déc05, Rap92, Ska96]. Efficiently [HMS09]. EG [Ano97e, Ano97g, Ano98f, Ano98b, Ano98e, Ano98h, Ano98g, CG07a, Ano95n, Ano96a, Ano96c, Ano96p, Ano97o, Ano97p, Ano97q, Ano97r, Ano97d, Ano97f, Ano98-30, Ano03e, Ano05a, Arn08, BPB^{+}04, BWS05, DCPS08, Kei04, Zot08]. EG/IEEE [Kei04]. EG2008 [Ano07i]. Egocentric [PIWB98]. Eighth [Ano97s]. Elastically [TG98]. Elber [Ano97-31]. Electrical [tH83a]. Electronic [PR93]. electronics [Pri85]. Element [Bak91b, IMIM08, SKCA01, WBG07]. Elements [BNC96, KB98, MKB^{+}08, SR96, VMA^{+}04, WDGT01]. Elevation [DKW94b, TIS^{+}95, BM93]. Elimination [Blo97, FA87, LS89, HB92]. Ellipsoid [LWP^{+}04]. Ellipsoidal [JBL^{+}06]. Embedded [GWT^{+}08, xHMC09]. Embedding [LZ07]. Emission [NGN^{+}01]. Emotion [KB98, RNtH03]. Emotional [LC09]. Emperor [Ano06c]. Emphasis [TO97]. Empirical [PIWB98]. enabling [FS92]. Enclosed [PN97]. Enclosing [RL09]. Encoded [JBL^{+}06]. Encoding [AD01, GP87, HW91]. Enderle [Ano95s]. Endoscopy [Bar05]. Energy [BV96, CWM09, Fau96, GLW96, KFG09, VW95]. Energy-Based [Fau96, KFG09]. Engine [FMS01, GMDW09, HHS01, ZK08, Kni93]. Engineering [Gol85, GSHN94, Jen07, tH83a, Sas92]. England [Arn91]. English [HR92]. Engraver [Zot08]. enhance [SSK93]. Enhanced [SKMS06, TK98]. Enhancement [DMHS08]. Enhancements [Joo86, SPS94]. Enhancing [JBL^{+}06]. Enrichment [Zar06]. Entertainment [Saw07]. Entropy [HS08]. Entry [Ben94]. enumeration [dS96]. Environment [AHR85, BSG^{+}95, BCH^{+}95, CWKS00, FBT99, GKD07, HK00, LD06, LLP00, Mil88a, MO08, MN87, MFDA86, PRS89, RTN03a, RTN03b, Sch01, SSSK04b, VGSS04, WKG85, YIC^{+}09, CFT86, CTW92, DM92, KSH92, VCDF95, WT93, dBv93]. Environmental [GSHN94, vvT84]. Environments [Ai98, AVF04, BGAM04, CBSF07, Fuc04, GSA03a, GSA03b, Hei01, KCB97, KW05, Kun04, LD04, LL01, LGK97, MCO97, MG95, MSWK02, MF00, PLT^{+}97, PIWB98, SS00, SG96a, Sha97, SDM99, SD00, SDB99, SNLW01, SGC04, WBP98, WS99, GH96, Gob95, Gob96, HK94, VCDF95, HJL07]. eNVyMyCar [GC09]. Equalization [AGM^{+}06]. Equation [YW97, PM93]. Equidistant [MLP92]. Erosion [KBKŠ09, RPP93]. Errata [Ano06a]. Erratum [Ano99f, Ano07a, Ano09a, JD01, SE02b]. Error [AKSA09, CRS98, ED07a, HS98, LA06, SW04a, SKS09]. Error-Bounding [HS98]. Errors [RAP08, VW91]. Escher [SDG99]. Escher-Like [SDG99]. Estimate [JZJ08b, WW09b]. Estimating [CLB^{+}09, GCP^{+}09, WMTG05]. Estimation [BPV^{+}09, BAJ08, HP02, KHIK01, NCKG00]. Estimations [Büh01]. Estimators [FCH^{+}06]. EUROGRAPHICS [Ano88a, Ano89, Ano91a, Ano92, Ano93, Ano94, Ano96l, Ano04b, Ano05b, Ano05d, Ano06b, Ano06d, Ano07g, Ano08d, Ano09c, Kwi89, Ano97o, Ano97s, Ano97-28, Ano97-29, Ano98d, Ano98c, Ano98i, Ano98j, Ano04g, Ano05e, Ano06e, Ano07c, Ano07d, Ano07e, Ano07l, Arb90, AS98, BH96, BH06, BJ94, Che06, CDD09, Des06, DS09, DS98, Duc98, Gob95, Gob96, HP95, Hég91, HK94, HJL07, Jan91, Kon06, Kui91, Kun04, LMD04, MJ98, NSGP06, Ott90, PB95, PH91, PS96b, Raf05, Rei03, San06, SvZ95, TT95b, TvdP98, Wei08, ACM80, Ano84a, Ano85, Ano86, Ano95i, Ano95n, Ano95l, Ano95m, Ano95b, Ano95j, Ano95k, Ano95o, Ano95p, Ano95a, Ano95d, Ano95c, Ano95q, Ano95-27, Ano07k, Ano08a, Ano09d, Arn89, ADS90, Arn91, Bou90, Cla89, Dau90]. EUROGRAPHICS [Duc89, Duc90a, Duc91, Enc81, Gre85, Heg90, HEtH^{+}83, Kid82, Kid84, Le 90, Lis90, Mat86, Mor86, Pin85, Req86, SW83, Van85, VD90, Vel91, WG82, Wat82, Ano87a, Ano88b, Ano91c, Ano96i, Ano96g, Ano96h, Ano96j, Ano96k, Ano97v, Ano97w, Ano97y, Ano97t, Ano97z, Ano97u, Ano97x, Ano98u, Ano98v, Ano98w, Ano98x, Ano99g, Ano99h, Ano99i, Ano00e, Ano00f, Ano00g, Ano01c, Ano01d, Ano01e, Ano02c, Ano02d, Ano02e, Ano04c, Ano04b, Ano04d, Ano04f, Ano04e, Ano04g, Ano05c, Ano05b, Ano05d, Ano05e, Ano05g, Ano06c, Ano06b, Ano06d, Ano06e, Ano07b, Ano07c, Ano07d, Ano07e, Ano07j, Aus91, Bla88, Bro90, Cla88, DJ88, DT04, GP06, HHS89, How87, HB91, Kje91a, LMD04, Mum90, Oll04, Sch98, TB84, Van80, Wat88, tH83a, tH84]. EUROGRAPHICS/ARGOSI [ADS90]. EUROGRAPHICS’2006 [Pur07]. EUROGRAPHICS’90 [KB90]. EUROGRAPHICS’96 [Ano96s]. Eurographics’98 [Ano97-27]. Europe [Ano82]. EUROPE’92 [tHS90]. European [Ano96s, Ano97o, DJ88, HHS89, Req86, TB84, Van85, Ano84a, Ano92, Ano93, Ano96l]. Evacuation [HK00]. Evaluation [BLD^{+}09, Čad08, DDV^{+}02, HMDO05, JR08, Klo87, McC96, Pat89, PIWB98, ZC95, vKP06, BR96, GA93, VW90, VW91]. Evenly [JL00, SLCZ09]. Evenly-Spaced [JL00]. Event [Ano98y, Ano99j, Ano99k, Ano99l, Ano00h, Ano00i, Ano02f, Ano02g, Ano02h, Ano02i]. Events [Ano97n, Ano97l, Ano97m, Ano98q, Ano98r, Ano98s, Ano99e, Gre97, Gre98, KB92]. Everyone [Ros97]. Evolution [Col05, GBD09, TA08]. Evolutionary [HCG08]. evolutions [CL92]. EWA [RPZ02]. Exact [BF09, ED07a, KRG03, MG95, NAB86, SPD07, The02b, YLL^{+}09, BR96]. Exaggerating [yKL08]. examination [Kni93]. Example [BB09, IMIM08, LCL07, OOI05, PCS94, WHCO08, DD92]. Example-Based [BB09, IMIM08, OOI05, WHCO08]. Examples [OAIS09]. EXCELL [Tam82]. Excentric [BRL09]. Exchange [KKK09]. Exchanging [BSVS04]. Exhaustive [HS98]. Exhibition [Ano93, Ano94, Ano96l, Ano96s, DJ88, HHS89, Req86, TB84, Van85, ACM80, Enc81, Van80, WG82, tH83a, Wat82]. Expansion [BDA^{+}09, XS06]. expect [Šár07]. Expeditions [Wat96]. Expeditious [CBSF07]. Experience [BPF^{+}03b, ESK03b, MGJ06, TCH^{+}03a, VGSS04, WGK88]. Experiences [AEL^{+}82, CRW83]. Experiment [Moh87, Kin92]. Experiments [CD94]. Exploiting [CAM08a, PFC^{+}05, CDPS09]. Exploration [BRS01, BCWR08, CZCE08, JBB^{+}08, LL09, SSCO09, SKKS08]. Exploratory [IEGC08, KOB^{+}08]. Explore [RNtH03]. Exploring [BB07, EBSC99, PSK09]. Exposure [MKV09]. Expression [RNtH03, SL09, WVKR08, ZPS03, MMAG93]. Expressions [BIMO04, WHL^{+}04, WR05]. Expressive [DN08, JL08]. Extended [BRL09, RW08, PCF05]. Extending [WGG99]. Extension [Bak88, JA95, KPiAS01]. External [ESC00]. Extracting [DMP93, TIS^{+}95]. Extraction [AGDJ08, GWT^{+}08, GE98, HWC^{+}05, OZ06, PKG03a, PKG03b, PWH98, SWPL08, SVG^{+}08, SR96, WG09]. Extraordinary [ADS06, LS08b, Nas03]. Extremely [DCPS08]. Extruded [WW87b]. Eye [BB09, KTMN07, RIF^{+}09, VGSS04, MH07].

F.R. [Enc81]. F.R.G [HHS89]. Fabrics [ME98]. Facade [LJN02]. Face [ADF85, BAHS06, BL08, DPT^{+}08, GSA03a, GSA03b, IIS08, JD98, OZS08, PZY08, SKC01, WL08, ZPS03, PG93, PG94]. Faces [BBPV03a, BBPV03b, BSVS04, KK07, KMT03a, MH07, NJ04, RB03a, SSSB07, WHL^{+}04]. Facets [MNP08]. Facial [JL08, NJ04, Pat95, RNtH03, SKC01, TSP05, WHL^{+}04, WR05, YD88, ZPS03, KMTT92]. Facilitate [Gna82]. Facilities [Ros83]. FACS [WR05]. Facsimile [XJJ^{+}08]. factor [MMAG93, Sbe93]. Factoring [AEW90]. Factorization [MTR08]. factors [Gui92, SZG93]. Faculty [tH83a]. Fade [GVWD06]. Fair [KPS95, MS93]. Fairing [Gre94, LBH^{+}01]. Faithful [KK07]. Family [BAU05, HAML05]. Fast [AES94, AO86, AO89, AMSF08, BT95, BF09, BB99, BS98, BB08, CSN04, CKM^{+}99, CMH^{+}01, DHK08, DMYN08, DKC00, EL01, GOT^{+}07, HKS09, HHS05, HSC^{+}05, IDN03a, IDN03b, KSN08, KŽ05, KHK^{+}09, KH92, KWF^{+}01, LD08b, LGS^{+}09, LLSL98, LG95, LCD09a, MMP09, MNP08, NNSK99a, NS09, PWC^{+}09, PPD98, PB94, Pat93, QNTN03a, QNTN03b, RKC02, RHv95, SS09, She99, SSK07, SWB01, SLTM08, SN08, SOM04, TT95a, TT97, VSC01, VCC98, WS99, YLL^{+}09, YF85, ZPS03, ZQK04, ZY04, vKP06, Sbe93, SAE93, WBCG09b]. Faster [KB89, WGS04]. FastV [CATM09]. Fat [YR91]. Faulted [McC83]. Feather [SH02, DSC95]. Feature [AGDJ08, BBW^{+}09, BYB09, BK01, BBN02, FGM99, GWT^{+}08, HWC^{+}05, KSL^{+}08, LM96b, LLSL98, ÖGG09, PKG03a, PKG03b, PK08, SWPL08, SVG^{+}08, SKC01, TE99, VRKS01, ZLK05, LM93, LM96a]. Feature-Based [LLSL98]. Feature-guided [KSL^{+}08]. Feature-point-driven [SKC01]. Featured [TCRS00]. Features [DMHS08, HP04, HBK02, LWY08, MKO^{+}08, PMW86, WB01]. February [Gob96]. Feed [KSK97a, KSK97b]. Feed-forward [KSK97a, KSK97b]. Feedback [LNS05, Pic86a, BH93]. Fellow [Ano07c]. Fellows [Ano08a, Ano04f, Ano05d, Ano06d, Ano07k, Ano09d]. Fellowship [Ano07d, Ano07e]. FEM [hZCK98]. Few [AR95, POB^{+}07, SP03a, SP03b]. FFT [McC96]. Fiber [CZCE08]. Fibers [SSA^{+}08]. Fibre [PVtHR09]. Fiction [Fuc97]. Fidelity [DDC09, RP98]. Field [AGG^{+}08, BS02, BB08, DYN04, HL03a, HL03b, KTMN07, KS07, LKC08, LCM^{+}06, MMP08, PLPB07, RSTK08, RR00, SP97, SS96b, SOM04, TRSKK08, ZCP07, ZBA^{+}07]. Fields [AGDJ09, AM02b, BDS^{+}03, BRS01, CCI08, Coq85, EBA^{+}09, FLBS07, HE94, IEGC08, KZ04, MRS08, NS09, OZ09, SEA08, SN08, SOM04, The02a, TRS03a, TRS03b, TSH01, UGLY08, VMA^{+}04, WTHS04, WTHS06, WP04, vP94, AM92, ILRS03a, ILRS03b, JBTS08]. Fifteenth [Ano97-28]. Fifth [Ano95q, Lis90, Wil84, MJ98]. figure [BT92]. Figures [OM01]. File [DHH02, SABG^{+}05]. Fill [AHT04, GP87, HR87, Ric87b, SM86]. Filled [Rok97, YR97]. Filling [DCOM00, NSY09, Ric87b, YF85]. film [SSK93]. Films [SW08b]. Filter [KR05]. Filtered [Kla00, KC08]. Filtering [Baj03a, Baj03b, CDP95, Got94, HP04, KKD09, PO85, TySK00, dCTAD09]. Filters [Bak88, GM06]. Final [HHS05, SSS02]. Finally [Ros97]. Finder [AVF04]. Finding [Szy91c]. Fine [CYJ02, NJ04, SLS^{+}06, BC01]. Fine-Scale [CYJ02]. Finishing [ten82b]. Finite [Bak91b, BNC96, KB98, MKB^{+}08, SKCA01, WBG07, WDGT01]. Finite-Element [SKCA01]. Fins [CJW^{+}06]. FIRD [MMP09]. First [Ano84a, Dav07, End84c, Heg90, Kel86, Le 90, RSD^{+}88, GA93, Ott90]. First-Year [Dav07]. Fit [ADJ^{+}01]. Fitting [LWP^{+}04, LWY08, PS95, PS96a, Pud94, SB99a, SHLS02, LDB07]. Five [Ano84a, SD94b]. Five-dimensional [SD94b]. flash [PCF05]. flatland [Hec92, ORDP96]. Flats [CWA^{+}08]. Flattening [BCGB08, MFT02]. Flexible [BBT99, BS08, GHH01, LW95, PLT^{+}97, RSS96, Sta97, SDC09]. Floating [DHvOS00, EDM^{+}08]. Flocking [O’H02]. Flood [AHT04]. Flood-Fill [AHT04]. floor [Lam09a]. Flow [AGDJ08, CKSW08, CPK09, GWT^{+}08, JL00, LHD^{+}04, LTH08, OHBKH09, PSL98, PKPH09, RSVP02, Sad09, SWPL08, SVG^{+}08, SWS09, VCC98, Wei04, WT09, YGL^{+}09, MMS09].

Flow-Based [WT09]. Flow-Embedded [GWT^{+}08]. Flower [IOI06]. Flows [TA08]. Fluid [DYN04, KySK08, KMN^{+}08, MMS07, OAIS09, SKSK07, TFK^{+}03a, TFK^{+}03b, YKH^{+}09, YLD07, GY93, KWF^{+}01]. Fluids [AM02a, BCN03b, DMYN08, GLHB09, KySK08, KBKŠ09, KMN^{+}08, LD09, MMS09, PTB^{+}03a, PTB^{+}03b, ST08, TFK^{+}03a, YNBH09]. Fly [MH00, SMS01, SLSK07]. Flying [SC08a]. fMRI [JNM^{+}09]. Foam [TFK^{+}03a, TFK^{+}03b]. Focus [MKO^{+}08]. Fonts [Gos86, SR96]. Footprints [van97, vdP97]. Force [BB08, Fau96, HV09]. Force-Directed [HV09]. Forecasting [BAJ08, SSK93]. Forest [FBW01]. Form [FGM99, FM04, Hec01, MMAG93, RS08, UGLY08, ZT96, ZCG98, AE97, GS85, Gui92, KMTT92, KHR02, Sbe93, SZG93]. form-factor [Sbe93]. form-factors [SZG93]. Formal [Dam91, Duc82a, DF85, Duc91, DD92, DDtR94, TC94, DP93, FZP92]. Formation [Dur01, TYK^{+}09]. Formats [DHH02]. Forms [Pic86a]. Forum [Hew84a, Ano03b, DS04b]. forward [KSK97a, KSK97b]. Four [HKS09, HS94, MKO^{+}08, Nic85]. Four-level [MKO^{+}08]. Four-Way [HKS09]. Fourier [NMMK05, XS06]. Fournier [Fiu01]. Fourth [SWS09, Ano97-29, Arb90, Cla89, Rei03]. Foveal [CG07b]. Fractal [Pic86b, ZT96, BM93, SSK93, WY92]. Fractal-Dimension [Pic86b]. Fractals [Gro92, NR95]. Fraction [AGDJ08]. Fractional [NMMK05]. Fragment [TRSKK08]. Frame [CLHL08, FMS01, HHS01, SW08a]. Frame-Coherent [FMS01]. Framebuffer [YSL08]. Frames [Her84, SPR^{+}94]. Framespace [AW00]. Framework [CLM09, FQK08, FdABS99, GD09, GA98, HKMS08, LW94, LS89, MCH94, PGGM09a, RSW^{+}97, ZCG08, DP93, LDB07, MC02]. Frameworks [BDFG07]. France [Ano97s, Ano97-29, Ano04b, BH96, DL04, DJ88, PB95, SZAB04, Van85]. Frayed [MBCN09]. Free [ANF97a, ANF97b, AE97, BGI08, CC08, FGM99, FM04, HK09c, KPRW05, KAAT03a, KAAT03b, KHR02, UGLY08, WB02, WLSG03, ZT96, ZCG98, ITYI09, KMTT92, SEA08]. Free-Form [FM04, ZCG98, AE97, KHR02]. Free-Viewpoint [WLSG03]. Freeform [Elb99, Kob03a, Kob03b, YK06]. Freehand [OOI05]. FreeLence [KPRW05]. Frequency [OMT02]. Frequency-Domain [OMT02]. FRG [BP82]. Friendly [SC04]. From-point [CATM09]. Frontiers [Enc98, vD98]. Frontmatter [Ano08c, Ano09b]. Fronts [SLS^{+}06]. Frusta [BJCW09]. Frustum [RS08]. Full [CLHL08, KH02, RR96]. Full-Frame [CLHL08]. Full-range [RR96]. Fully [DJW^{+}06, SAA09]. Function [AGDJ09, GBAL09, RSC01, SPOK95, Bar93]. Functional [Ary84, Ary86, JBB^{+}08, KO88, LS89, PS96a, SW92a]. Functionalities [PCS94]. Functionally [HL03c]. Functionally-Based [HL03c]. Functionals [KPS95]. Functions [AGDJ09, AR95, BK05b, BG07, CLB^{+}09, DKN^{+}95, FP94, HL03a, HL03b, JBL^{+}06, LWP^{+}04, MKB^{+}08, MMS^{+}05, NP00, PP09b, RSS96, SK86, SCM^{+}09, SBLD03, SvLD03, TC94, WS09b, XWT^{+}09, DF93, Sch94a].

Fundamental [DGEG09]. Furniture [GS09]. Fused [LLHY09]. Fusion [MKV09]. Future [Bak88, Bro90, Enc98, Han05, Joo86].

g [MG09]. g-BRDFs [MG09]. G. [Ano97-31]. Galleries [BBW^{+}09, RSTK08, VBP^{+}09, YM09]. Galliformes [DSC95]. Game [Col05, FMS01, GC09, Jon90]. Games [BB00a, HHS01, Hec01, KW05, Saw07]. Gamuts [RGW05]. gap [MK08]. Garden [WBEF97]. Garment [ML91, SSK^{+}05]. Garments [DJW^{+}06, PZB^{+}09]. Gaseous [GLHB09, SW92a]. Gaskets [Jon90]. Gather [SSS02]. Gathering [HHS05]. Gaussian [ADS06]. Gaze [VGSS04]. GEncode [LCL^{+}06]. General [Ano95r, Ano97q, Ano97z, Ano97-30, Ano04b, Ano04g, Ano05b, Ano05e, Ano06b, Ano06e, Ano07g, Ano07l, Ano08d, Ano09c, CLM09, Coc83, HR85, LCL^{+}06, LSŽ08, LP95, OLG^{+}07, Pil85, SSS97, GMDW09, GGRZ06]. General-Purpose [OLG^{+}07, Pil85]. Generalized [ESV99, OPC96, SMS01, TSYK01, YGL^{+}09]. Generated [Chr86, Pat95, Pic86a, SB99b, Lei94]. Generating [AW00, BDA^{+}09, BBDM85, GC96, LD08a, LF97, LGMT00, Mar95, Par86, WVVV08]. Generation [ABC^{+}91, AAB^{+}96, DG95, GMY97, GD96, HSC^{+}05, HTH96, KMN^{+}05, KK07, Kuz95, LW95, LCM^{+}09, MK05, MG87, MKO^{+}00, Mül86, PJ94, PGGM09b, RW08, RMA88, RHv95, SGS05, TON^{+}02, TTW90, Wil84, WR05, WT09, BYB09, DKW94a, GK03a, GK03b, HGW92, HDS03b, HDS03a, HCL93, KH92, LW92, Liu93a, Liu93b, MK08]. Generator [Her82, YLK08]. generators [ES94]. Generic [MS08, Vel99, XTJ^{+}07]. Geneva [Van80, Ano04g, Ano05e, Ano06e, Ano07l]. Genome [WVKR08]. Genomic [WVKR08]. Genuine [PCK09]. Genuinity [AA09]. Genus [GP09, PSF04]. Geodesic [BW07, HM83, Haw85]. Geodesic-Controlled [BW07]. Geodesics [TWC^{+}09]. GeoFilter [KR05]. Geographic [PCS94, Sam93b]. Geographical [TIS^{+}95, TON^{+}02, HS92]. Geography [Mil83]. Geometric [AS92, AR94, BDŠ84, BŠ90, Bar92, BMWM01, CDPS09, DJW^{+}06, Den03a, Den03b, DG97, HH02, KR05, Las84, LL02, NAB86, NB94, Sei88, Vel99, WMWG09, WB02, BC01, ESP92, YN93]. Geometrical [DG95, LC09, VT94]. Geometrically [CC00]. Geometrically-Aware [CC00]. Geometry [ABC^{+}04, ADS06, BK05a, BK05b, BHGS06, BB08, CT00, CC06, CLF^{+}03a, CLF^{+}03b, CDPS09, CKSW08, DWL08, Des04, Des06, DGGP05, DGEG09, DSW09, ESP08, EMK09, ESK03a, ESK03b, FAVM09, GP09, GMC^{+}06, GE04, GBP05, HP04, HWC^{+}05, HL03c, JWB^{+}06, JLCW06, JKS05, KPRW05, KT09, KR05, KCL06, Kob03a, Kob03b, KSH04, Kob05, Kol08, KQWM08, LJN02, LCL^{+}06, LTH08, LSJK09, MK05, MMHL08, OZ06, PFC^{+}05, RL09, SWPL08, SG96a, STK08, SZAB04, SLS^{+}06, SP06, SSB05, VC04, VL08, Vel99, VMG09, WC05, YSL08, YGL^{+}09, ZRKS05, de 97, PCF05, Sbe93]. Geometry-Aware [YSL08]. Geometry-driven [LCL^{+}06, MMHL08]. Geospatial [CBSF07]. Gerd [Ano07b]. Germany [Duc98, Enc81, KSH04, Kon06]. Gestural [Duk95, KMB94]. Gesture [EBSC99, BH93]. Gestures [JL08, KHIK01, SKSK07]. Getting [SSSS98]. Gino [Duc82b]. Girona [NSGP06]. GIS [Bar93, But94, GL94a, GL94b]. GIS-product [Bar93]. given [THN93]. GIzMOs [PCK09]. GKS [Ano87b, AHR85, AH89, Bak88, Bak91a, BBMR88, BP83a, BHM87, Dam91, DDR93, DD92, ELM^{+}83, End83a, End84a, FCP^{+}90, Fre90, GD85, Her84, HR85, HR87, HM86, Mac84, MA85, Mil87, Mil88b, Mil88a, MN87, Mum86, ND94, PNR89, RGM85, Ric87c, Ros83, SC84, SK86, SM86, Sla84, SAHt91, Ste84, WH89, ten82b]. GKS-3D [Fre90, PNR89]. GKS-9x [DDR93, ND94]. GKS-Based [Mil88a]. GKS-implementations [End83a]. GKS/GKS [PNR89]. GKS/GKS-3D [PNR89]. Glare [KMN^{+}05, RIF^{+}09]. Glasgow [Kil85, Mat86]. Global [BW00, BWS03a, BWS03b, BBL^{+}09, BAJ08, CAE08, DDM03, DGV08, DDB^{+}09, EZK08, FD09, FP04, FLBS07, GSA03a, GSA03b, GD01, GKD07, HVAPB08, Hei01, HHK^{+}07, HMS09, HP02, HLS96, IDN03b, LSP08, LZX^{+}08, NSW09, NS09, NKF09, OSG08, PLPB07, STK08, SW04a, SIP07, SPS95, SSG^{+}00, SW99, SSSK04b, SK99, SKCA01, WGS04, WS03a, WMS^{+}08, WÅ09, YW97, ZSP98, ZBA^{+}07, ITYI09]. Global-Illumination [HLS96]. Globally [WSC06]. Glossy [GD01, SSG^{+}00, YWY08]. Glyph [KJC^{+}09]. Glyph-based [KJC^{+}09]. Goal [ZV09]. Goal-Driven [ZV09]. Gödel [GMDW09]. Good [MB08, SP03a, SP03b, SNLH09, VBP^{+}09]. GosiP [Fre90]. Gouraud [Nar95]. GPU [AGG^{+}08, BJCW09, BWPP04, BS08, BBA08, BBS^{+}09, CSI09, CLH^{+}08, CZGF05, DSW09, DRS08, ED07b, FQK08, GGK06, KZ08, KSN08, KW05, LGS^{+}09, LMS04, LCD09a, Lov06, MMFE08, MO08, MRS08, PGSS07, RS08, RLN06, SS09, SKALP05, SKU08, SKUP^{+}09, TRSKK08, WSE04, WSBZ08, YWY08]. GPU-Accelerated [MMFE08]. GPU-Assisted [CSI09]. GPU-Based [TRSKK08, BWPP04, CLH^{+}08, DSW09, GGK06, KSN08, LMS04, MRS08, SS09, WSE04, YWY08]. GPUs [HKS09, KHH^{+}09, LGS^{+}09]. Graceful [DO00]. Gradation [MTCT84]. Gradient [GCP^{+}09, NCKG00, RZLG08, XM09]. Gradient-based [RZLG08]. Gradient-Preserving [XM09]. Gradients [JZJ08a, LSW09]. GRAFLOG [PKL88]. Grammar [KJC^{+}09]. Granada [DT04]. Granite [SMTG07]. Granular [LD09, ON05, SMTG07]. Graph [BW00, Dwy09, GSGC08, GBD09, HV09, HLH96, Men95, NHH97, PSF04, PV08, SAA09, SKKS08, VW08, ZHC^{+}00, CTW92, PSF04]. Graph-Based [Men95, PSF04, CTW92]. Graphic [DCG87, FG88, TC93, TC94, AS92, DM92]. Graphical [Bad93, BW87, Kin95, NM91, Ric87a, ZHC^{+}00, BH93, HR92, dBv93, BFTL82]. Graphics [AEW90, AHKS94, AR06, Ano82, Ano84b, Ano88b, Ano92, Ano93, Ano95a, Ano95q, Ano96l, Ano96a, Ano96s, Ano97d, Ano97-29, Ano98b, Ano98e, Ano98c, Ano98-30, Ano03b, Ano03c, AS98, AEL^{+}82, AHR84, ADS90, AMS09, BIMO04, BDŠ84, Bar06, BB91, Bla88, BB88, Bou90, BS03a, BS03b, BP82, BP83b, BG85, Bro90, But94, CB09, CON08, CWM09, Cla88, Cla89, CDD09, Dau90, Dav07, DSS90, Den86, Den03a, Den03b, Des04, Dia84, Duc82a, DJ88, Duc91, DHH02, Duf88, DS04b, Dun91, Edm83, ELM^{+}83, End82, Ert02, FS85, Fol98, FG88, Fra83, FR00, FMK04, GS06, Gal84, GC09, GD85, Gna82, Gna83, GVV05, Gos86, GSHN94, Gro01, Han05, HHS89, Her89, HHA82, Hew84a, Hew84b, Hew90, HGG^{+}84, HM86, Hd89, IDN02, Kei04, Kel86, Kil82, Kil85, Kje89]. Graphics [Kje90, Kje91b, Kje91c, Kje91d, Klo87, KH96, Kol08, Kui91, LC99, Lis90, Mac94, Mae90, MMTH09, MGJ06, Mar82, May99, May00, Mik82, Mil88a, MFDA86, Mud83, Mur85, NYTN87, NMK^{+}06, NMNP98, NSGP06, OIST91, Owe86, Owe87, Owe88, Owe89b, Owe94, OLG^{+}07, Pic87, Pil85, PP89, PNR89, Raf05, Rei03, Req86, Rey86, Ric87b, RIF^{+}09, RSK90, Ros82, RP98, Sab82, Sal96, San06, Šár07, SCD05, SD94a, Sch98, Sch84, Sei94, SG03, Spa85, Sta06, SHPS08, Str83, Str82, SOM04, Suz89, Tho86, TB84, Tur84, Van85, Vel91, VBB^{+}06, Wal87, Wat87, Wei08, WKG85, Wil84, YWB03, Zar06, ten82a, tHS90, vJB85, van87, BJ94, CFT86, CRW83, DP93, ESP92, FZP92, HK92, Jac85, Owe90b, Owe92a, Owe92b, Owe93, Owe95, Sam93a, SM84, SGM^{+}93, TC93, UH92, Ano84a]. Graphics [Ano96c, Ano97o, Ano06g, Bon85, DL04, Ott90, tH83b]. Graphics-Based [Dav07]. Graphite [SB99b, SB00]. Graphs [Bov90, EGKT08, Pic86b, PV08, TIS^{+}95, VW08]. GRAPP [Bar06]. Grasping [KAAT03a, KAAT03b, ST94]. gravity [Hol94]. Grayscale [Čad08]. Graz [HK94]. Green [AMS09, CWM09, MMTH09, RIF^{+}09]. Grenoble [Ano04b, DL04]. Grey [WO94]. Greyscale [SLTM08]. Grid [Ara94, OHBKH09, SSS02]. Grid-Distortion [Ara94]. Grid-Less [OHBKH09]. Grids [CWA^{+}08, CKM^{+}99, DMYN08, KRG03, LD08b, SW05, SCM^{+}09, TSM94, Liu93a, Liu93b]. Grimsdale [Wil06a]. Gromov [CCSG^{+}09]. Gromov-Hausdorff [CCSG^{+}09]. Grooved [BPMG08]. Group [Ano84b, Duc82b, End84b, Mac85, Str84, TYK^{+}09, AS92]. group-theoretic [AS92]. Groups [OCV^{+}02]. Growing [GMW04, Pas02, WF97]. Growth [DKY98, SFS05, DGA04]. Guaranteed [TRS03a, TRS03b]. guarantees [BC01]. Guest [AR06]. Guidance [ZRKS05]. Guide [Spe91, TON^{+}02]. Guide-Map [TON^{+}02]. Guided [BRS01, HKW09, HMYS01, AVF04, KSL^{+}08]. Guidelines [BBMR88]. Guiding [NMMK05]. Gunter [Ano95s]. GVE [Ano97-29]. GWB [Man83].

HAGI [HK92]. Hair [BCN03b, BCN03a, BPV^{+}09, HMT01, PTB^{+}03a, TFK^{+}03a, DTKT93]. Hairy [XTLP02]. Half [Buc96]. Half-toning [Buc96]. Halftoning [SB98, VB99]. Haloed [FA87]. Hamburg [HHS89]. Hand [KHIK01, Pud94, ST94, SDC09, HR92, SKSK07]. Hand-drawn [SDC09]. Hand-Sketched [Pud94]. Handed [SG97, TGS96]. Handling [DO00, Edm83, FML06, GOT^{+}07, OTSG09, SM86, WLML99]. Handy [MK99]. Haptic [ADJ^{+}01, FMB^{+}00, HS04, LD03, LD07, MS96b, TCH^{+}03a, TCH^{+}03b, MS96a]. Haptics [DYN04, HS04, NC99, SLS04]. Hard [ESG01]. Hardly [ASVNB00]. Hardly-Visible [ASVNB00]. Hardware [BIMO04, BWPP04, Bla88, BS02, BS03a, BS03b, Cla88, Cla89, CZGF05, DS02a, Den03a, Den03b, DKS01, DL04, Ert02, IDN02, KZ08, Kau04, KW05, LMS04, LLHY09, Lis90, LCD09b, MPT98, OLG^{+}07, RPZ02, SCD05, SPH^{+}09, SG03, SOM04, SKALP05, YWB03, Sun92, Ano95a, Ano96a, Ano98-30, Kui91, Sch98]. Hardware-Accelerated [KZ08, LMS04]. Hardware-Based [DS02a, MPT98, BS03a, Den03a]. Harmful [WGS04]. Harmonic [MKB^{+}08, XS06, ZRKS05]. Harmonics [JCJ09, VL08]. Hashing [CJC^{+}09]. Hatching [ZISS04, Bak90]. Hausdorff [CCSG^{+}09]. having [SMAB02]. Haze [Wil87b]. HDR [BDA^{+}09, DMHS08, MCHAM08, SKMS06]. HDTV [Ste85]. Head [MH07, MG95, SS00, TGS96, WO94, DPT^{+}08]. Head-eye [MH07]. Head-Mounted [WO94]. Head-Slaved [SS00]. Head-Tracked [TGS96]. Hearing [RB03a, RB03b]. Heart [PVtHR09]. Heat [SOG09]. Heavier [Sil97]. Height [NS09, SN08]. Heightfield [RK09b]. Held [tHS90, Kje91a]. Helical [KOB^{+}08]. Helices [Ber09, HWAG09, LZSCO09]. Help [HD95]. Heras [Ano05c]. Heritage [Arn08, CL99, DCPS08, GBU00, PSK09, VPP^{+}04, Zot08]. Hermite [FS08]. Hermitian [GBS99]. Hershey [Gos86]. Heterogeneous [MB99, NG03a, NG03b, SMTG07, ZHC^{+}00]. Heuristic [WGS04]. hexagonal [Liu93a, Liu93b]. Hidden [FA87, Hea89, HB92, JD98, LS89, SD94a, Ska87, Tam82, YLK08, SDD^{+}92]. Hidden-Line [Ska87]. Hidden-picture [YLK08]. Hiding [WC05]. Hierarchal [YFGL09]. Hierarchical [Ano99m, AMSF08, BJCO03a, BJCO03b, BCG^{+}96, BV99, BWPP04, CSN04, CH09, DHS04, DGV08, Dvt90, FML06, GSA03a, GSA03b, GE98, HB96, HDSD99, Hew84b, HS98, KSS97, LA05, LL09, MPT98, MB97, MBW08, MB99, NJ04, NSW09, NG03a, NG03b, NP00, O’H02, PKPH09, Sha97, SSSS98, WGK88, DMP93, GH96]. Hierarchically [CZCE08, HV08, PC92]. Hierarchies [CH09, DHK08, LAM09b, MW06, MSF00, VCP09, WD09, YM06]. Hierarchy [CDP95, yKL08, ML03, PPD98, Ste84]. Hierarchyless [DGGP05]. High [BDA^{+}09, Bel87, BBDM85, CSI09, CGG^{+}03a, CGG^{+}03b, DDM03, Den86, DMAC03a, DMAC03b, DDC09, DZTS08, FAVM09, GLHB09, GBP07, HK09a, HE01, HMS09, KZ08, KK08, Kle06, KMS05, Los97, LD97, MKV09, MRS08, MHG00, NS01, PK08, PGSS07, RPZ02, SiK– DM05, Ste85, WH04, WHL^{+}04, WM85, WW87b, ZISS04, HK92, SNLH09]. high-dimensional [SNLH09]. High-Fidelity [DDC09]. high-level [HK92]. High-Quality [FAVM09, GBP07, HMS09, KZ08, Los97, LD97, MHG00, DDM03]. High-Resolution [PK08]. High-speed [DZTS08]. Higher [CAH00, DHS04, GSA03a, GSA03b, LS08a, WTHS04, BPZ96]. Higher-Order [DHS04]. Highly [DCPS08, DG95, Ros97, SSK07]. HistoPyramids [DZTS08]. history [CFM93]. HL [AE86]. HL/HSR [AE86]. Hochschule [Enc81]. Hologram [ZBA^{+}07]. Holy [Ano06c]. Homogeneous [AEW90, Nie95, PP09a, XGL^{+}07]. Homotopy [Sad09]. Honorary [Ano07d, Ano07e]. Hoops [BNRSV01]. Horizon [DAP08, RBB01]. Host [Ano95n]. Hough [OZ06]. HPCCD [KHH^{+}09]. HSR [AE86]. Huge [SC95]. Hull [Day88, Day90]. Hulls [BG08, LMS04, PT03, CD94]. Human [AMS09, BT98, BGK^{+}96, BCH^{+}95, CWM09, DAP08, Enc98, GPD09, HSS^{+}09, KTMN07, KSK97a, KB04, LB06, LL00, LC09, MTT89, MMTH09, MTKO02, OM01, RIF^{+}09, TO97, WH96, ZY02, ZLK05, Gue82, HGW92, KSK97b, KMA05, Lam09a, SKSK07]. Human-Centered [Enc98]. Human-like [OM01, KMA05]. human-machine [Gue82]. Human-Media [Enc98]. Humanoid [EBMT00, BCH^{+}95]. Humans [BB09, CMT02, FBT99, HMDO05, LD04, Lam09a, LGMT00, PLT^{+}97, PO02, TYK^{+}09, Ter02]. Hungary [Ano97w]. Hurricane [MSM^{+}08]. Huygens [MNR94]. HW [PRS89]. HW/SW [PRS89]. Hybrid [BBS^{+}09, HH02, JBB^{+}08, KHH^{+}09, Kob05, ML03, Sch00, SDM99, WT09, ZQK04]. Hydraulic [KBKŠ09]. Hydrodynamics [KBKŠ09]. Hypermedia [Ben94, HS92, San92, HK94]. Hypertexture [GD09]. Hypertextures [SG04]. Hypothetical [SKC01].

IBRAC [YN00]. iCheat [OKP^{+}08]. Icons [Ais85, SABG^{+}05]. Idea [AHKS94]. Ideas [Fuc97]. IDECAP [vvT84]. IEEE [Kei04]. IFSs [Hor90]. II [ARC05, BPM06, BBT^{+}06, BAHS06, CZ08, Déc05, DGGP05, Gar09, GKB09, HAWG08, KPRW05, MK05, PCF05, SiK– DM05, SWG08]. III [BK05a, HWC^{+}05, PFC^{+}05]. Illuminated [SW09]. Illumination [AHT04, BWS03a, BWS03b, BGB08b, BBL^{+}09, BAJ08, CAM08a, CAM08b, CAE08, CDPS09, DDM03, DDB^{+}09, FD09, FP04, FLBS07, GCP^{+}09, GSA03a, GSA03b, GD01, GKD07, GJW08, HVAPB08, Hei01, HHK^{+}07, HMS09, HP02, HLS96, IDN03b, JL06, KC08, MTR08, MSW04, NSW09, NS09, NKF09, PLPB07, SSSK04a, STK08, SW04a, SIP07, SP01, SPS95, SSS97, SSG^{+}00, SW99, SSSK04b, SK99, SKCA01, WGS04, WS03a, WL08, WDC^{+}08, WÅ09, WW09b, YW97, YWY08, YIC^{+}09, ZSP98, ZBA^{+}07, BPZ96, DF93, PM93]. Illumination-related [CDPS09]. Illuminations [NN89]. Illustration [BCD01, CLE07, JR08]. illustration-inspired [JR08]. Illustrations [Ano98t, CLME09, DWE02, DWE03a, DWE03b, FMS01]. Illustrative [BG07, JBB^{+}08, MM08, RBG08, RSTK08, DD92]. Image [AAB09, AR06, ABB^{+}07, Ano97-31, Ano04c, Ano05c, Ano06c, Ano07b, AR95, AHT04, AW07, BCS96, BD04, Čad08, CYJ02, CDPS09, CL99, DDV^{+}02, DJZ^{+}09, DMAC03b, Dut04, Edm83, ESK03a, ESK03b, FMB^{+}00, FCH^{+}06, FP04, FCOL00, FLW00, GMY97, GHH01, GW07, GVWD06, GCL^{+}06, HK09a, HH02, xHMC09, HCG08, KL08, KMG96, KFG09, hKAC07, KSL^{+}08, KS07, KKD09, LB06, LW95, LD06, LAA08, LYP^{+}08, LWX^{+}09, Lut02, MG87, MA91, MZT09, MG95, McN01, MTPS08, Mül86, MSK06, NMP98, NLED08, NSW09, OS08, PRS89, Par89, PCK09, PB90, RGW05, RvBWR04, RTN03a, RTN03b, RMS^{+}08, Sal96, SS96a, SGS05, SSK^{+}05, SD09, SSCO09, SC08a, SLCZ09, SJ09b, SLAM08, SW04b, TE99, TO97, TTW90, TSP05, TW96, VSD09, WPG02, WH04, WL08, WWG07, WHCO08, WSBZ08, WKG85, WMTG05, WC02]. Image [XM09, XTJ^{+}07, XGL^{+}07, XWT^{+}09, XWT^{+}08, YD88, YN00, ZCHM09, Hor90, Koc93, LW92, VL93]. Image-Based [CDPS09, DJZ^{+}09, FCOL00, GHH01, HK09a, HCG08, MSK06, NLED08, OS08, RMS^{+}08, SLCZ09, VSD09, XWT^{+}08, YN00, CYJ02, Dut04]. Image-Space [AHT04, GW07, NSW09]. Image-Swept [WC02]. Image-to-Geometry [CDPS09]. Imagery [BPBD08, Han05, Hol94]. Images [AO86, AO87, AO89, BBPV03a, BBPV03b, BSVS04, BAHS06, BE95, CLHL08, Chr86, DDPL00, GP09, HLM97, HK09c, xHMC09, JBS^{+}06, KPiAS01, KO88, KLD^{+}09, KMS05, LCZ99, RGW05, RWP88, RM89, SKMS06, SLTM08, SLAM08, WO02, WSBZ08, WW09b, WOBT09, YMMS06, ZR96a, ZHM08, BYB09, CFT86, FLBS07, HGW92, PC92, ZR96b]. Imaging [AGP08, BD07, BDA^{+}09, BPBD08, DMHS08, Fol98, FHL^{+}08, JNX^{+}08, KTMN07, KK08, KS07, KMS07, LAA08, MS08, MRT08, NSGP06, POB^{+}07, SL08, TSY^{+}07, BG93, CS93]. Immersive [Fuc04, Kun04, MSWK02, RvBWR04, VGSS04]. Impact [Saw07, VW91]. Impaired [RB03a, RB03b]. Implementation [Day90, ELM^{+}83, FBW01, Mac84, NB94, PL94, SC84, DDR93, NS93, VW91]. Implementations [End84a, WH89, MMAG93, End83a]. Implementer [MN87]. Implementing [Ara94, AH89, SPS94]. implements [MS93]. Implicit [AG01, Ano95b, BDS^{+}03, BTG95, CBS96, FJW^{+}05, FAT07, GMW04, Har97, HJ99, KHK^{+}09, LWP^{+}04, Li07, NOS09, OTSG09, RRS97b, She99, SS96b, TSYK01, WSC06, WGG99, DTG96, Guo93, KHR02, RRS97a, VG96, dS96]. Importance [CLH^{+}08, CAM08b, CAE08, JCJ09, KC08, MMP08, NNSK99a, SWB98, SB98, SSSK04b, SKS09, WÅ09, PM93]. Important [CJW^{+}09, LLY09, WDR09]. Impossible [SDG99]. Impostor [ABC^{+}04, HMDO05, SDB97]. Impostor-Based [ABC^{+}04]. Impostors [ABB^{+}07, O’H02, PMDS06, SKALP05, DSSD99]. Impressions [Ano97-32, Kid82, RSD^{+}88, Wat82]. Improved [DTA94, LLA06, VMM99, WO02]. Improving [HR87, MSSK08]. Impulse [BW00]. Impulses [Ye08]. In-Class [BB07]. In-Out [MTR08]. Incident [NMNP98, UGLY08]. Including [Sch00, KJ92]. Inclusion [JFS09]. Incoherent [DHK08]. Incremental [COF95, GD01, GM96, HB96, KQWM08, LM96a, LM96b, SL89, TP88, VCP09]. Independent [BPMG08, Chr86, ME04, PGG^{+}09, KMA05]. Index [Ano98m, Ano98a, Ano04h, Ano05f, Ano06f, Ano07f]. Indexed [Owe86, Owe87, Owe88, Owe89b, Owe94, Spe91, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. Indexing [GPD09]. Indirect [BBP08, GJW08, OKP^{+}08, SP01]. Individual [SSSB07, ZK09]. individualized [DPT^{+}08]. Indoor [FML06]. Industry [ML91]. Inexpensive [HAML05, Sch94b, WH04]. Inextensible [Ye08]. Influence [Fra83, SPR^{+}94, SF92]. Information [BSG^{+}95, CDPS09, DCPS08, DCG87, FdABS99, HDM98, JEO00, KCB97, PCS94, Rob93, SD94a, TK98, Váz07, WC05, vvT84, HS92, Sam93b]. Informative [SOG09]. Informed [FBT99]. Infrared [WDC^{+}08]. Inhomogeneous [RZLG08, PCF05]. Initial [DDtR94, IOI06]. initialization [SKSK07]. Ink [FJW^{+}05, FMS01, SFWS03a, SFWS03b]. Inlay [SG08].

Innovation [Kin95]. Input [ADJ^{+}01, Dvt90, GM96, Ros82, Ros83, Rus01, SAHt91, van90, vt87, FZP92, FS92]. inquiry [End83a]. Insight [CMS94]. Inspection [NW91, PV08]. Inspired [Arn08, JR08]. Instant [BG09, CZGF05, DDB^{+}09, ME04, SIP07, WWS01]. Instantaneous [KOB^{+}08]. Institute [WG82, tHS90]. Integer [AP92, Liu94, LŽY04, NG92]. Integral [Rok97, YR97, Sbe93]. Integrated [GBU00, MFDA86, WE97, vKB94, DTKT93]. Integrating [ASVNB00, HM91, HKMS08, MCT01, Mum88, PH96, Sam93a]. Integration [FR00, Fuc04, IOI06, Joo86, LJN02, PO85, SDB99, SKFNC97, YW97]. Intelligence [Ter02, TMT86]. Intelligent [Arb90, BBW^{+}09, Kwi89, MS96a, MS96b, OIST91, RL84, VBP^{+}09, YM09]. Intensity [BG09, CS98a, MHG00]. Intention [CLHL08]. Intention-based [CLHL08]. Inter [MTKO02]. Inter-reflection [MTKO02]. Interacting [RSW^{+}97, YLD07]. Interaction [AHM09, BW87, BRL09, BCF94, BG07, CLE07, CN05, CG07b, DMKP07, Duk95, FJW^{+}05, FR00, Han97, Hd89, IEGC08, JR08, KB04, LL01, LNS05, LF00, MSM^{+}08, PL94, PIWB98, RBG08, Ros82, RSKN08, SS95, STD09, SG97, TGS96, VGSS04, WKG85, ZK09, vD98, vt87, BH93, DH93b, EFGS96, FS92, Gue82, Kje92]. Interaction-Dependent [RBG08]. Interactions [DMYN08, HS04]. Interactive [AGG^{+}08, AGDJ09, Ano98d, AG06, BP98, BWS03a, BWS03b, BGK^{+}96, BMG99, BS02, BPF^{+}03a, BPF^{+}03b, BCH^{+}95, BWRT96, BN08a, BG93, CLDD09, CB09, CC00, CPP08, CKM^{+}99, CRW83, CG07b, CH09, CKSW08, CPK09, DDM03, DDB^{+}09, DWE02, DWE03a, DWE03b, DKN^{+}95, DKYN96, DYN04, DQ00, DDC09, Duf88, Elb99, Ert02, ESK03a, ESK03b, FD09, FG88, GCMS00, GD09, GLW96, HK09a, Hei01, Her82, Hew84b, HE94, HK09c, IIS09, IP99, IHS02, IEGC08, JBL^{+}06, JZF^{+}09, KTMN07, KAAT03a, KAAT03b, KH95, KHIK01, KB00, KGMM97, KS92, LG96, LNS05, LeYTM08, Lee99, LGB^{+}03, LLG97, LL09, Los97, LD97, MJ98, MAA^{+}09, MOT99, MHG00, MKO^{+}08, Mun83, MBW^{+}05, NC99, NSW09, OOI05, ON05, OTH02, Pai02, PVtHR09, PGGM09a, PMDS06, RP01, RRRP08, RRS97a, RMSD^{+}08, RRS97b, SSS00, SSB08]. Interactive [SS08, STK08, SLS04, SPH^{+}09, SC04, SSK07, Sla88, SSG^{+}00, SBLD03, SvLD03, TG98, UT02, Vas97, WSBW01, Wat96, WMWG09, Wie96, WLML99, WKM^{+}09, YWY08, YIC^{+}09, ZK08, vvT84, vL90, AM92, DTG96, FFD93, GMDW09, Hor90, KMA05, LW92, MM93, PB95, Rob93, VSD09]. Interconnected [SKKS08]. Interface [AGDJ08, Ano88b, DR87, Duc90a, End84b, End84c, HK09a, HM91, KCB97, Kin95, LFGG08, LD98, Mac85, PSK09, PF90, RSS96, Sni95, SG97, WGK88, BH93, HK92, Req86, Sas92, ZK92, dBv93]. Interfaces [Che06, CG07b, Duc85, Hd89, NM91, PB07, STD09, vJB85, KSH92]. Interference [FS85, SW08b, WP04]. InterHyper [Cla92]. Interior [DKN^{+}94, RLF09, YIC^{+}09]. Interiors [BPM06]. Interleaved [DDC09]. Intermediate [MBBT00, FLBS07]. Intermixing [CS99a]. Internal [POB^{+}07]. International [Ano95b, Ano06g, Bar06, Kun04, Suz89, van89b, ACM80, Enc81, Van80, WG82, tH83a, Sei88]. Internet [Wat96]. Interpolant [dD85]. Interpolating [Nas03, BBA08, KS92, KFK94]. Interpolation [AW00, BAU05, CLT^{+}08, CK84, DKW94b, DF90, KB89, LAA08, Nar95, RF96, RZLG08, Rok97, RSC01, SW08a, SLAM08, SW04b, YR97, EMK09, FS08, STM93]. Interpolatory [GBP05, Kob96, LG00, LMB05, LM07, SLLW08, WM09]. Interpretation [BP83a]. Interpreted [van89a]. Interreflection [NN89, KJ92]. Interreflections [PLPB07, RK09b]. Interrogation [Elb95, HHB93]. Intersecting [Ska87, Sug94]. Intersection [ML91, NHH97, PT03, SN84, GA93]. Intersections [FB94]. Interval [Büh01, KHK^{+}09]. Intrinsic [DMA02, OSG08, DZM08]. Introducing [LM93]. Introduction [van82]. Intuitive [HK09a, MG09, PSK09, WVVV08]. Invariant [Her89, MS93]. Inventor [WHR97]. Inverse [BLD^{+}09, BS98, CC08, GMW04, LP95, MMP09, McC96, MBBT00, PP05, PHL91, RSC01, BT92]. Inverse-Kinematics [RSC01]. Inversion [CRW09]. Investigation [TSYK01]. Investigative [WMS^{+}08]. Invited [Ano08c, Ano09b, Baj03a, Col05, Coo05, Des04, Dut04, Ede06, Fuc04, Han05, Jen07, Kle06, Kob03b, Pur03a, Šár07, Saw07, Sta06]. iPCA [JZF^{+}09]. Ireland [Che06, HP95]. Iris [LB06]. Irradiance [BGB08b, JZJ08a, SiK– DM05, Ure00]. Irregular [AKSA09, BV09, CK84, GP87, NOS09, WW99]. Irregularly [DLC05]. Islamic [AS92]. Iso [GE98, Ano84b, Bak91a, Bon85, Gal84, tH83b]. Iso-Surface [GE98]. ISO/TC97/SC5/WG2 [Gal84, tH83b]. Isometric [AE97, Fiu95, HAWG08]. Isosurface [CL03a, CL03b, HB94, LCD09a, MMFE08, PWH98, RW08]. Isosurfaces [CWA^{+}08, HSS^{+}05, MJC01, OB01, The02b]. Isothetic [JA95]. Isotopic [DLRW09]. Isotropic [YLL^{+}09]. Issues [Kin95, Pat95, Pur03a, Pur03b, Sco02, vJB85]. Italy [ACM80, SvZ95, SP06, MRS06]. Iteration [SK99, SKCA01]. Iterations [Szy91a]. Iterative [SG08]. IV [BK05b, JKS05, WC05, ZRKS05].

J [Cal07]. Japan [FMK04, Mik82]. JAPE [SNLW01]. Jitter [Kla00]. John [Duc06b, Duc07, JW01]. Join [Ano95t, Ano95u, Ano95v, Ano96m, Ano96n, Ano96o, Ano97-33, Ano97-34, Ano97-35, Ano97-36, Ano98z, Ano98-27, Ano03f, Ano04d]. Joint [yKL08]. Joy [JCJ09, SDC09, WBCG09a]. July [Des06, FMK04, Kon06, LLD05, LLRD07, SCA04]. Junction [RT08a]. June [Ano07j, HP95, HK94, Kei04, KSH04, Kon06, LMD04, MJ98, Oll04, PB95, PS96b, SZAB04, SP06]. Jürgen [Duc06a].

kaleidoscopes [PR93]. Karner [Ano06c]. KD [PGSS07, SSK07]. KD-Tree [PGSS07, SSK07]. Kelemen [Ano04c]. Kernel [BS08, Enc98, ÖGG09]. Keyframe [MAA^{+}09]. Keyword [Owe86, Owe87, Owe88, Owe89b, Owe94, Col93, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. Keyword-Indexed [Owe86, Owe87, Owe88, Owe89b, Owe94, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. kinematic [BT92]. Kinematics [Col05, KOB^{+}08, MBBT00, RSC01]. Kinetic [PB07]. Klingler [SSA^{+}08]. Knit [DS02a]. Knit-Wear [DS02a]. Knits [DPT^{+}08, IIS08, PZY08, WL08]. Knitted [ME98]. Knitting [IIS08]. Knots [ST08]. Knowledge [RA94]. Konrad [Ano06c]. Kubelka [ARC05]. Kubelka-Munk [ARC05]. Kuwahara [KKD09].

L [CBSF07]. L-systems [CBSF07]. Label [BV09]. Labeling [BRL09, YLD07, SNKS09]. Labelling [GM96]. lace [NW91]. Lacerations [GCMS00]. Lady [Duc00]. Lagrangian [GWT^{+}08, KySK08, PB07]. Lambertian [Jen97]. Landscape [WS01]. Landscapes [RPP93, TON^{+}02, BCF^{+}05]. Landscaping [Joo86]. Language [BŠ90, CCP09, GC96, LO95, Mil90, PCS94, PKL88, RB03a, RB03b, WGK88, BH93, Sam93a, Spa85, Cal07]. Languages [Her82, Mil87]. Lansdown [Duc06b, Duc07, JW01]. Laplace [CLB^{+}09]. Laplacian [FAT07, SG08, VMM99]. Large [AG06, BPM06, Bov90, CCI08, CDG^{+}07, Coc83, HK00, IP99, KSN08, LeYTM08, LYP^{+}08, MG87, MKO^{+}08, PD04, PSK09, SHLS02, ILRS03a, ILRS03b, PFC^{+}05]. Large-Scale [HK00, MG87, PD04]. Laser [Zot08]. last [Bak91a]. Latency [WO94]. Latent [BiA06]. Lattice [EMK09]. Lattices [DDKL09, FAVM09]. laws [VD90, Par89]. Layer [SC84]. Layered [NAS07, TG98, DSSD99]. Layers [LKC08]. Layout [Dwy09, GC96, SE02a, SE02b]. Layouts [YM06]. Lazy [DLTD08]. LazyBrush [SDC09]. LCTS [HB00]. Leakproof [RHv95]. Learning [BB07, Dav07, KK07, LCM^{+}09, VBP^{+}09, WHL^{+}04]. Least [SB99a]. Least-Squares [SB99a]. Leeds [Bro90]. length [FND92]. Lenses [KTMN07, LSS98]. Less [OHBKH09]. Letterform [Sch00]. Leuven [LMD04]. Level [BK03a, BK03b, CJC^{+}09, CKSW08, DN08, MB97, MMS07, MBW^{+}05, SW08a, SG03, Ste84, SW04b, Wei04, YSL08, BS02, HK92, MKO^{+}08, Dam91]. Level-of-Detail [SW08a]. Level-Set [BK03a, BK03b, Wei04]. Levels [ASVNB00, OCV^{+}02]. Library [Vel99, WT93]. lichen [DGA04]. LIDAR [BCWR08]. Life [JNX^{+}08, KOB^{+}08, MMHL08, WVKR08]. Lifted [PP09b]. Light [AGG^{+}08, AHL^{+}06, BYP95, BP01, CCC08, CZGF05, DKYN96, EBA^{+}09, GHH01, HVAPB08, KSKAC02, KKK09, KHM09, KB04, LB06, LCM^{+}06, MMP08, NMNP98, NMOT01, PPD98, PP05, RSTK08, SSSK04a, SP01, SEA08, SN08, SSSK04b, TT95a, TT97, TRSKK08, UGLY08, Váz07, WMTG05, Wu90, ZBP99, ZBA^{+}07, FLBS07, KJ92, RCM^{+}01]. Lightcuts [AWB08, HMS09]. Lighting [AMS09, BBP08, BAJ08, CAM08b, DMCP94, DKN^{+}94, DKN^{+}95, FBP08, Kau04, OKP^{+}08, SL01, YK08, ZCG08]. Lightness [KMS05]. Lights [GJW08, MT01, PPD98]. Like [SDG99, KMA05, OM01, TO97]. Limbs [Neb00]. Limited [MTCT84]. Limiting [TPS09]. Line [Che97, CW99, Elb99, FA87, Gos89, GM96, HL02, KPiAS01, Kuz95, LJN02, LSŽ08, Nie95, PGG^{+}09, SVG^{+}08, SC08b, Ska87, VW90, vKB94, Day92, FND92, Kra92, Ska96, SDD^{+}92, WG93]. Line-Picture [Gos89]. Linear [BIMO04, BF09, Ber09, Büh01, HP04, Nar95, NCKG00, ÖGG09, Rok97, SSSB07, TT95a, YR97]. Linearised [Ben94]. Linearization [HD02]. Lines [BBW^{+}09, BB99, CPK09, DS05a, Liu94, MSSK08, Par86, RWP88, SWPL08, SP03a, SP03b, Tam82, WW87b, ZCQ^{+}09, DMP93, Liu93b]. Link [Bak91b]. Linking [IF09]. Linkless [CJC^{+}09]. Linköping [Fah85]. Links [SSSS98]. Lip [KK07]. Lip-Synch [KK07]. Liquid [HK03a, HK03b]. Liquids [MMS07]. Lisbon [Req86]. List [MGAF95]. Listener [BMD^{+}08]. Literature [Owe86, Owe87, Owe88, Owe89b, Owe94, Spe91, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. Live [DJZ^{+}09]. Load [MB99]. Local [AGM^{+}06, BBA08, CCI08, CAM08a, DGV08, EZK08, ITYI09, KRG03, KC08, LAA08, LZX^{+}08, MTR08, Mai00, MKB^{+}05, MSK06, OTH02, OSG08, PZY08, PCR89, RHC08, WHCO08]. Local-to-global [ITYI09]. Local/Global [LZX^{+}08]. locality [KFK94]. Localized [DSC09a, GKS00]. Locally [KLD^{+}09, NP00]. located [IF09]. Locomotion [HKG06, KRFC09]. LOD [GW07, YWY08]. Logarithmic [DMAC03a, DMAC03b]. Logical [vt87]. Longest [HB00]. Lookup [RW08, SZG93]. Loop [LWY08, SLLW08]. Loops [Pic86a]. loose [BR96]. Loosely [BCN03b, BCN03a]. Lossless [TW96, VCP09]. Lossy [DDPL00, DDV^{+}02]. Loughborough [Mum86]. Low [BDA^{+}09, GLHB09, PWP08, RCM^{+}01]. Low-Anisotropy [PWP08]. Lower [GKS00, RGW05]. Lukas [Ano06c]. Lumigraphs [SHS99, SMS01]. Luminance [DKYN96].

Maastricht [TT95b, Ano95j, Ano95k]. Machine [AHM09, AO86, AO87, KK07, NYTN87, Gue82, NW91]. Machines [Bij87, SSB05]. Macintosh [Moh87]. macro [Guo93]. Macroscopic [BG89]. Made [BWPP04, HD95, VR95]. MagicSphere [CMS94]. Magnification [BD07]. Magnitude [RAP08]. Maintainability [BGAM04]. Makes [CAH00]. Making [Bij87, Coo05, Lew94]. Man [van82]. Man-Computer [van82]. Management [BW87, BBS^{+}09, Duc90a, End84c, Fis98, Mac85, PF90, CFT86, DSS90, Sas92]. Managing [HHS01, Her84]. Manchester [WG82, HHA82]. Mancunian [Ano96p]. Manifold [BSJ08, LCM^{+}09, Sad09, SHPS08, VL08, FR92, SAHt91]. Manifold-valued [SHPS08]. Manifolds [DGGP05, GE04, LDR09, MDP08, EZK08]. Manipulating [CCI^{+}07]. Manipulation [AAB09, Ary84, CC00, DQ00, KAAT03a, KAAT03b, PIWB98, SSB08, SD09, SSCO09, SDB97, Sni95, STD09, ZCG98, van90, vL90, FFD93]. Manufacturing [Gol85, RGM85]. Many [HVAPB08, PPD98, Mac94]. Many-Light [HVAPB08]. Map [GKD07, HP02, Mar95, NGB^{+}09, RBB01, TON^{+}02, Déc05]. Mapped [BJCW09, SKMS06, TCRS00]. Mapping [Ara94, AE97, BLD^{+}09, BL08, CC08, DCPS08, DKS01, DMAC03a, DMAC03b, ESG01, FD09, GSGC08, GBP07, HHS05, JZJ08b, JBS^{+}06, KH02, LP95, MK99, PHL91, SSO08, SS96b, SKMS06, SSSK04b, SKU08, WGS04, YFGL09, YD88, YMMS06, Hal99, MS08, MMTH09, NG92]. Maps [AHL^{+}06, BTB02, COS95, DF90, LG95, MSW04, NAS07, PWC^{+}09, Sch01, SEA08, WDR09, WN09, YK08, ZHM08, PFC^{+}05]. March [Ano97z, Ano97-28]. Marching [AG01, DZTS08, HWC^{+}05, RW08, RHv95, SW05, The02b]. Marine [DTA94]. Mario [Ano06c]. Marker [MMS07, YLD07]. Markerless [PG08, SKSK07]. Markers [FBW01]. Market [PP89, tHS90]. Marmitt [Ano07b]. Mask [SL08]. Masking [WPG02]. Masks [DS05a, NMMK05]. Mass [HH98, PP89]. Mass-Spring [HH98]. Massachusetts [LLRD07]. Massive [BGAM04, WLML99]. Massless [SL07]. Matching [AAB09, BD04, RF96, RvBWR04, SPD07, WH04, CCFM08, DOS93]. Material [AGDJ08, ABB^{+}07, ON05, RLW^{+}09, WWG07, XGL^{+}07, XWT^{+}08, Ano99m, BYB09]. Materials [Cal96, LD09, RK09b, SMTG07, SB00]. Math [JCJ09, Pic91a, SDC09, WBCG09a]. Mathematical [Pic86a, TC93]. Mathematics [Kra89]. Matrices [VB00]. Matrix [NB94]. Matting [GVWD06, GCL^{+}06, Wil06b]. Max [DKC00]. Maximal [ABC^{+}04]. Maximum [BG09, CS98a, HS08, MHG00, Váz07]. May [Jan91, Kun04, NSGP06, SvZ95]. Mazes [WT09]. Measure [GMSK09]. Measurement [BPV^{+}09, GCP^{+}09, KMG96, RK09a, SLS04, SW04a]. Measurement-Based [SLS04]. Measurements [RMS^{+}08]. Measures [CCSLT09]. Measuring [CRS98]. Mechanical [NC99, LDB07]. Media [BN08b, Enc98, JZJ08a, PWP08, PP09a, RZLG08]. Medial [BTG95]. Median [FP94]. medical [BG93]. Medicine [Baj03a, Baj03b, VBB^{+}06]. Meeting [Ano97q, Ano97z, Ano04g, Ano05e, Ano06e, Ano07l, Arn84, Bon85, Duc82b, Nas03, tH83b, Gal84]. Meets [SL08]. Megalithic [PL96]. Members [Ano04g, Ano05e, Ano06e, Ano07l]. Membrane [EBV05]. Memoriam [Duc06a, Wil06a]. Memory [Bel87, CPP08, ESC00, KH95, LeYTM08, MW06, MO08, PO02, Jac85]. Memory-Conserving [MW06]. Mental [BAHS06]. Mer [Ano97s]. Mesh [Ale02, AS96a, BS08, CH09, DTTS08, EGKT08, FAT07, GOT^{+}07, GGRZ06, JLCW06, JKS05, KT09, KR05, KCL06, LAD02, LZ07, LZX^{+}08, MK06, MH00, OB01, OZ09, PJ94, RW08, RNV07, SHB07, SSFS06, Sha08, Sor06, VC04, VCP09, VLV^{+}04, WSLG07, YBS07, vKP06, BYB09, CCFM08, HB94]. Meshes [AD01, ATBG08, BV99, BG02, BGI08, BK01, BK03c, BBA08, CCFM08, CCI08, CC06, CL03a, CL03b, CGG^{+}03b, DMA02, GMC^{+}06, GE98, GGG08, HLS96, KBS00, LS09, Lee99, LL02, LCL^{+}06, LMB05, LBH^{+}01, Lov06, LSW09, MGS07, MNP08, NSY09, PSF04, RW08, RSS96, SWPL08, SHB07, SWG08, VG00, VMM99, WDGT01, GGRZ06, Jon96, CGG^{+}03a]. Meshing [BYB09, DSC09b, DSC09a, SSFS06, SNKS09, YLL^{+}09, dCTAD09]. Meshing-Simplex [BYB09, dCTAD09]. Metaballs [Hei95, KSN08, NN94, NIDN97]. Metabolic [WVKR08]. Metafile [BHM87, OF84, Sch84, SM84]. Metafiles [End82, Osl82, SC84]. Metallic [NNSK99b, RMS^{+}08]. Metamorphosis [BMWM01, GA96, LLSL98, RK94, STK02]. Metaphors [SABG^{+}05, ZK09]. Metering [NMNP98]. Method [BF84, BB99, CC08, CK84, Coc83, CDPS09, CL99, DHvOS00, DG95, DKN^{+}95, DKYN96, DMYN08, GBU00, Hei95, HL03c, IDN02, IDN03a, IDN03b, JC08, Kje83, LA06, LW95, LM07, LD98, MPT98, NN94, NIDN97, PH87, Pat89, PB94, SMAB02, SK86, SP01, Sla88, SKFNC97, Tam82, TSYK01, TSP05, TSH01, WBG07, WO94, YKH^{+}09, ZBP99, ZY02, BM93, DH93a, Gui92, LBT92, Sbe93, STM93, VSD09, Vol93]. Methodology [NM91]. Methods [BV96, CRW09, Duc91, HE01, HR87, HHD03a, HHD03b, JL06, LD08a, Ric87b, Sab86, Sch00, SW04a, SRH^{+}09, SW99, Szy91b, Szy91c, TP89, WG09, AS92, KH92, SW92b]. Metis [TLG99]. Metric [BCGB08, EP09, FP04, GMY97, LA06, LZSCO09, LSW09, SL01]. Metrics [vKP06]. Metro [CRS98]. Metropolis [KSKAC02, SIP07]. Michelangelo [Lev99]. Micro [Gol85]. Micro-Based [Gol85]. Microcomputers [NB88]. microfacet [DWL^{+}09]. MICROGRAPHICS [Sch85]. Microprocessor [Pil85]. Microscopic [MMHL08]. Microscopy [JNX^{+}08]. Mie [JW97]. Mie-Scattering [JW97]. Migration [PNR89]. Million [Pic91b, Sil97]. Million-Point [Pic91b]. Miltenberg [BP82]. Miltenberg/Darmstadt [BP82]. Min [DKC00]. Min-Max [DKC00]. mind [PR93]. Minds [Bij87]. Minimal [FBW01, JA95, VW95]. Minimising [ADS06]. minimization [FS08]. minimum [MS93]. minimum-cost [MS93]. Minkowski [GA96]. Minutes [Ano95r, Ano97-30]. Miranda [Par89]. Mireille [Ano05c]. Mirrors [HNR^{+}04]. Mixed [JL06, SLLW08]. Mixing [LD09]. MLS [WSBZ08]. MO [DM92, Mil90]. Mobile [RTN03a, RTN03b]. Modal [HWAG09, KRFC09]. Mode [BG01]. Model [AGM^{+}06, BR97, BR02, BK05a, BPMG04, BBN02, Bur95, CTSO03a, CTSO03b, DAP08, DWL^{+}09, FML06, GCMS00, HSS^{+}09, Hd89, IIS08, IK01a, JW95, KB04, LB06, LSJK09, Mac85, Mil88b, OW91, ÖGG09, PA06, PH87, PNR89, RMN05, RK09b, Sch94b, SDK09, SPH^{+}09, St88, SSCO09, SAHt91, ST08, SH02, Str84, TMO04, TK98, VPLL08, VGSS04, WAH^{+}09, XTLP02, XTJ^{+}07, AHR93, BM93, DMP93, FZP92, HS92, LG92, ML03, TC93, WY92, DP93]. Model-Based [SSCO09]. Model-driven [TMO04]. Model-Predictive [DAP08]. Modeling [AR94, Baj03a, Baj03b, Bak91b, BDS^{+}03, BF09, BGK^{+}96, BW07, Che06, CDPS09, DWL^{+}09, DDKL09, DMCP94, DJZ^{+}09, FCOL00, FG04, GBU00, GMW04, GLW96, HK09a, HMT01, HE94, IOI06, ITYI09, JW97, KB92, KFG09, KT97, LC09, LCWK07, MS01, MG05, MS08, Mil90, MF00, NIDN97, PWC^{+}09, PGGM09a, PGGM09b, RP01, RR00, RLF09, SDG99, SS08, SLSK07, SMTG07, SG04, TNK^{+}93, TZF04, TON^{+}02, VSD09, VW95, VPP^{+}04, WGG99, YCLE09, YSY94, YYPZ07, YIC^{+}09, ZT96, ZR96a, ZPS03, DTKT93, DGA04, DH93a, ESP92, Gro92, ZR96b]. Modeller [NAB86]. Modelling [Ano98d, BPM06, BBT^{+}06, BCF^{+}05, BB91, BAHS06, Cal96, CL92, CTSO03a, CPE92, CBSF07, DJW^{+}06, EKM01, GD09, HL03b, HH90, IOI06, LM96b, Las84, LC99, LD98, LC09, ME98, Nic85, OOI05, PL96, Sei88, SFS05, TD00, Wai88, WSC06, WBEF97, YYPZ07, Zac89, DTG96, FFD93, FG04, HR88, LM96a, SC04, ZY04]. Models [AA09, AHR84, ADJ^{+}01, BB00b, BAHS06, BPF^{+}03a, BPF^{+}03b, BCF94, BMWM01, BL86, BNC96, CATM09, CCI08, CZ09, CK84, CDG^{+}07, DCPS08, DKW94b, DGC^{+}98, EWHS08, FJW^{+}05, GD96, GVV05, GM96, KŽ05, Kle06, KGL^{+}98, LeYTM08, LJN02, LCZ99, LC09, MK05, MG87, MBW^{+}05, NMK^{+}06, Neb00, NK99, NN89, NNSK99b, NNSK99a, OTH02, OZS08, Pai02, Pat89, PZB^{+}09, RGM85, Ros97, Sch94c, SB99b, SB00, SP03a, SFWS03a, SP03b, SFWS03b, WF97, WC05, Wie96, WLML99, ZQK04, AVF04, DPT^{+}08, GK03a, GK03b, Jac85, LM93]. modifiable [ZK08]. Modification [Ska87]. Modified [RHv95]. Modulation [MRT08]. Molecular [Ai98, Dia84, RWP88]. Molecules [PMW86]. Monaco [Gob96]. Monitoring [HKD^{+}08]. Monkey [Pic91a]. Montage [DJZ^{+}09]. Monte [Gob95, Gob96, BBL^{+}09, PWP08, Pic86b, Sbe97, SKFNC97]. morass [Cla92]. Morphing [Ale02, BP98, GLA00, TE99]. Morphological [BE95, Mil90]. Morphology [KMA05, Vel93]. Morphology-independent [KMA05]. Morse [SWPL08]. Mosaic [BDFG07]. Mosaicking [CL99]. Mosaics [PCK09]. Motion [AW00, CYJ02, Coh95, DAP08, EBSC99, FL06, GPD09, GFW^{+}06, HSS^{+}09, HPT89, HK09b, JL08, KAAT03a, hKLS00, KRFC09, LS09, LL05, LLY09, MMS07, MAA^{+}09, MBBT00, NG03a, OM01, Osh08, PG08, PH03b, PH03a, RAP08, RR00, SKSK07, SSK^{+}05, TySK00, TWC^{+}09, UGB^{+}04, YLD07, BT92, SF92]. Motion-based [EBSC99]. motion-motion [SF92]. Motions [KAAT03a, KAAT03b, yKL08, LG96, NJ04, KMA05]. Motorcycle [EGKT08]. mountain [BM93]. Mounted [WO94]. Mouse [STD09]. Movement [GJL^{+}09]. Movements [BB09, JL98, WVV09]. Movie [Coo05]. Moving [CLHL08, Fah85, Los97, LD97, SLAM08, WSBZ08, ZHM08]. MRF [TSP05]. Mud [SOH99]. Multi [AGDJ08, BV09, CS99a, CJC^{+}09, DSSD99, FH09, JBTS08, LLHY09, LSŽ08, MAA^{+}09, MO08, PKG03a, PKG03b, Pil85, PSK09, SGS05, SNKS09, SOG09, TCRS00, WS02, BYB09]. Multi-Agent [SGS05]. Multi-Featured [TCRS00]. Multi-fields [JBTS08]. Multi-GPU [MO08]. Multi-Label [BV09]. Multi-layered [DSSD99]. Multi-Level [CJC^{+}09]. Multi-Material [AGDJ08, BYB09]. Multi-Microprocessor [Pil85]. Multi-objective [SNKS09]. Multi-Resolution [WS02]. Multi-Scale [SOG09, PKG03a, PKG03b]. Multi-Spectral [PSK09]. Multi-step [LSŽ08]. Multi-Variate [FH09]. Multi-Volume [LLHY09, CS99a]. MultiClusterTree [LL09]. Multicolor [MKO^{+}00]. Multidimensional [FR00, GMDW09, GJL^{+}09, KK02, LL09, SNLH09, FT93]. Multifield [JBL^{+}06]. Multilayered [CJW^{+}06]. Multilevel [RA94]. Multimedia [But94, DH98, Kje91a, PH96, AHR93, Cla92, FT93, Kin92, KSH92, ST93, VR95, dBD^{+}92, HK94, Kje92]. Multimedia/hypermedia [HK94]. Multimodal [RRRP08]. MultiOOP [Cla92]. Multipass [SW99]. Multipath [CSN04, MSSK08]. Multiplayer [GC09]. Multiple [BCH^{+}95, Che97, DMKP07, EBA^{+}09, GD96, GA98, Her84, KHIK01, LC99, NK99, SJH08, WHCO08, YLD07]. Multiply [RR94]. Multiply-Connected [RR94]. Multiprocessor [CWBV86, NS93]. Multiresolution [AAB^{+}96, AG06, BK03c, BK03d, CLF^{+}03a, CLF^{+}03b, DKC00, EBV01, GCMS00, hKLS00, KCL06, KBS00, Lee99, MK05, Mey94, PWH98, Sal96, SB99a, SMAB02, SHB07, YP95]. Multiscale [WE97]. Multivariate [FS91, JBTS08, LL09, MKO^{+}08, PC94, PV08, SJH08]. Munk [ARC05]. Mural [LLP00]. Muscle [KSK97a, KSK97b, hZCK98, KMTT92]. Muscle-based [KSK97a, KSK97b]. Muscles [YCLE09]. Music [LL05, SNI06]. Mutation [KSKAC02]. Mutual [CDPS09].

NAG [BFTL82]. Name [Bak88, CJ90]. Nano [Baj03a, Baj03b]. Nano-Medicine [Baj03a, Baj03b]. Narrative [RB03a, RB03b]. Natural [BPMG04, CG07a, DGA04, GMW04, GP06, KB04, KRFC09, PKL88, PA01, Wai88, Wu90, Zar06, KH92]. Nature [BCF^{+}05, Bur95, HK03b, Jen07, OOI05, QNTN03b, SFS05, WY92]. Navigating [MCO97]. Navigation [AVF04, BT98, HDM98, LD04, MSWK02, MDWK08, PPD07, SSCO09, SKKS08, SGC04, ZK08, Lam09a]. NBS [CP88]. NCGA [Kuh82, Mum88]. NCSDCT [Mud83]. Near [ADS06, MMP08, NMMK05, BGB^{+}08a]. Near-Field [MMP08]. Near-Regular [NMMK05]. Nebulas [NGN^{+}01]. Needs [GL94b, GL94a]. Neighborhood [SPD07]. Neighbourhoods [CD08]. Netherlands [TT95b]. Nets [Kob96]. Network [Ano95w, Ano95y, Ano95x, Ano96r, Ano96q, Ano97p, Ano97-37, Ano97-38, Ano97-40, Ano97-39, Ano98-28, Ano98-29, Ano03e, Ano04e, AO89, How90, LLP00, MB99, PCR89, WVKR08, BG93]. Networked [PLT^{+}97]. Networking [ADS90, SGM^{+}93]. Networks [LBD^{+}08]. neural [BG93]. Neurology [CZCE08, RRRP08, SSA^{+}08]. Neuroscience [Neb00]. Neurosurgical [RRRP08]. Neutralization [JD00, JD01]. NeWS [RSD^{+}88]. Nice [DJ88, SZAB04, Van85]. Nicograph [HP84]. Nicosia [Ano07j]. NIL [Mil87]. Ninth [Duc98]. Node [Hea90, Hea89]. Noise [NOS09]. Noisy [VMM99]. Non [BSJ08, CYJ02, CMH^{+}01, DSC09b, DSC09a, DWE03b, DBHM03a, GSA03a, GSA03b, HS99, Hei01, HP04, HHD03b, HH98, HAWG08, Jen97, KE97, KQWM08, LSP08, LBD^{+}08, MSWK02, MDWK08, ÖGG09, SSSB07, SNKS09, SFWS03a, SP03b, SK99, TSM94, VMG09, XLTP03b, XJJ^{+}08, YLK08, BPZ96, Ska96]. non- [BPZ96]. Non-Active [HH98]. non-convex [Ska96]. Non-Diffuse [GSA03a, GSA03b, Hei01, SK99]. Non-Immersive [MSWK02]. Non-Lambertian [Jen97]. Non-Linear [HP04, ÖGG09, SSSB07]. Non-Manifold [BSJ08]. Non-parallel [LBD^{+}08]. Non-Photorealistic [CMH^{+}01, HS99, DWE03b, DBHM03a, HHD03b, KQWM08, MDWK08, SFWS03a, SP03b, XLTP03b, XJJ^{+}08, YLK08]. Non-Regular [TSM94]. Non-Rigid [HAWG08, LSP08, CYJ02]. Non-Simplex [DSC09b, DSC09a, SNKS09]. Non-Uniform [KE97, VMG09]. Nonlinear [Kje83, WSE04, WDGT01]. Nonphotorealistic [Hal99]. nonprogrammers [dBv93]. Nonregularly [Frü94]. Nonsplitting [Guo93]. Nonuniform [Vas97]. Normal [Dan96, PA06, SW92b, TW97a, TW97b, Thü01]. normals [SAE93]. Norrköping [Oll04]. Norwich [Ano97z, Ano97-28]. Note [LR88, Lut02, Hub93]. Notes [Kra89]. Novel [CAH00, HLM97, NMOT01, WO94]. November [CA05, MRS06]. NPR [CJW^{+}09, LLY09, PGG^{+}09, WDR09]. NPR-Style [PGG^{+}09]. Number [KSN08, Nas03]. Numeric [Elb95]. Numerical [BDŠ84, HE01, RSS96, Szy91b, GA93]. NURB [WO92]. NURBS [AES94, AP92, GK03a, GK03b, LLG97, PH94, QSW92].

O.D.E. [Ric87a]. Obituaries [Duc06a, Wil06a]. Object [AA09, BT95, BB91, BWRT96, CS98a, CCI^{+}07, CC00, CMH^{+}01, DR87, DH98, GGW98, HMB08, Hd89, KAAT03a, KAAT03b, KH96, Mar95, MSF00, Ott90, PG08, PIWB98, RPZ02, Sch88, SEASM09, SF83, TC94, TLG99, Vel91, WCB^{+}95, van90, AHR93, DSS90, Sch94a, ZK92]. Object-Oriented [BB91, DR87, DH98, Hd89, KH96, Ott90, Sch88, TLG99, Vel91, ZK92]. Object-Space [HMB08]. Object-Surface [Mar95]. objective [SNKS09]. Objects [AAK^{+}09, AS95, AM95, AB97, BDŠ84, BB08, Büh01, CS00, CBS96, DMYN08, GCMS00, GTS86, HS04, HE01, HHD03a, HHD03b, HE94, HL03a, HL03b, HLL07, HNR^{+}04, IK01b, LGB^{+}03, Los97, LD97, MCL96, NK99, OTSG09, PL94, RGM85, RHv95, SE04, SS96b, SWB01, TKH^{+}05, BLS93, DH93b, Ger92, KWF^{+}01, LDB07, MFT02, OCL96, TC93]. Observational [SB00]. Occluded [COFHZ98]. Occluder [WS99]. Occluders [BKES00, BNRSV01, NRJS03a, NRJS03b]. Occlusion [AGG^{+}08, ASVNB00, BKES00, BWPP04, BWRT96, COFHZ98, ED08, MBW08, MPS05, RMSD^{+}08, SPH^{+}09, SBW06, ZCG08]. Occlusion-Driven [SBW06]. Occlusions [JZJ08a]. Ocean [GSMA08]. Oct [Arn84]. October [ACM80]. Octree [CJC^{+}09, Hea90, CLL^{+}08, GA93, Hea89, PG94]. Octrees [AS95, DKC00, ES94, PG93]. Octtree [NAB86]. Off [HHS01, PSK09]. Offers [Ano95n]. Offset [PK08, MNR94, SW92b]. Omni [ABB^{+}07]. Omni-directional [ABB^{+}07]. on-Surface [TPBC09]. On-the-Fly [SMS01, SLSK07]. One [LŽY04]. One-Pass [LŽY04]. Onlay [SG08]. Online [KQWM08, YLD07]. only [DSC09b]. Ontology [GSGC08]. Opacity [RSK06, YK08]. Open [Kob96, Pur03a, Pur03b, HS92, HK94, KSH92, WHR97]. OpenGL [Cal07]. Operations [AO86, AO89, BE95, HL03c, KP87, Li07, Man83, Tho86, WW88, WGG99, van87, FND92, LM93, PG93]. operative [BSG^{+}95]. Operator [CLB^{+}09, LLSL98, MS08]. Operators [BDS^{+}03, GGRZ06]. Opportunities [Bry96]. Opportunity [tHS90]. Optical [IDN02, JNX^{+}08, PSL98, MG96]. Optics [GSMA08, KMN^{+}05]. Optimal [BGAM04, FAVM09, Sbe97]. Optimality [Got03]. optimisation [Hub93]. Optimised [ZC95]. Optimization [BIMO04, Coh95, Den03a, Den03b, LA06, LSP08, LYP^{+}08, LSS98, MB97, MGAF95, OB01, SWB98, VLV^{+}04, CLL^{+}08]. Optimized [BTB02, JFS09, WK04]. Optimizing [FCH^{+}06, xHMC09, LWZ^{+}09]. Order [CAH00, DHS04, GCP^{+}09, GSA03a, GSA03b, LS08a, ME04, Pic87, SWS09, WTHS04, ZY04, BPZ96]. Order-Independent [ME04]. Ordering [BJCO03a, BJCO03b, HMB08, SS00]. Organ [DKY98]. Organized [HV08]. Organizing [AMT02]. orientation [SW92b]. Oriented [BDŠ84, BB91, DR87, DH98, Hd89, JFSO06, KH96, Ott90, Sch88, TLG99, Vel91, KCB97, ZK92]. Orthogonal [DGF98, KE97]. Orthographic [GTS86]. Other [Kin95, KH92]. Our [Enc98]. Out-of-Core [FCGW02, FK09, BBS^{+}09, ILRS03a, ILRS03b]. outdoor [TNK^{+}93]. Outline [Sch00, SR96]. Outline-Based [Sch00]. Output [BG01, HR87, ND94, Rus01, Ste84, SG96b]. Output-Sensitive [SG96b, BG01]. Outstanding [Dre07]. Overlap [Tho86]. Overlapping [DMYN08, SAA09]. Overlays [BY08]. Overview [BDFG07, End83a, End84a, HGG^{+}84, Kil82].

Pablo [Ano05c]. Pace [Str83]. Packages [Klo87]. Packet [YFGL09]. Packet-based [YFGL09]. page [Sam93a]. Paint [EKM01, RMS^{+}08, CPE92]. Paint-Composition [EKM01]. Painterly [LLY09]. Painting [DWE03b, HHD03b, IK01b, SDC09, TO97, WW87a, WS01, XLTP03b, XTJ^{+}07]. Painting-like [TO97]. Palazzo [ACM80]. palette [Vol93]. Palettes [WVVV08]. Pandora [Kin92]. Panel [HEtH^{+}83, SG97]. Panorama [KH02, KLD^{+}09]. Panoramic [HVM^{+}08, KPiAS01]. Paper [BW07, RSW^{+}97]. Papers [AR06, Ano96t, Ano07i, Ano08f, Arn08, DCPS08, Gob95, Gob96, Zot08]. Para [PSF04]. Para-Graph [PSF04]. Paradigm [AR94, KCB97, YSY94, LG92, WBCG09b, YN93]. Paradigms [Ano95q, Ano96c, Ano98e, AS98, DH98]. Parallel [Ano98b, Ano98c, AH89, CMPS93, COF95, CDD09, HKS09, JC08, KH95, KHH^{+}09, MSSK08, McC96, MM08, MB99, MNP08, Nar95, NG97, PRS89, Raf05, Rei03, RA94, San06, SSK07, SF83, TP89, Wei08, ZYQ^{+}08, ZCQ^{+}09, CD94, HB92, LBD^{+}08, MMAG93, VL93]. Parameter [SL09, STD09, WTHS06]. Parameterization [BCGB08, BF84, EGKT08, KNP07, LS08a, LZX^{+}08, MTAD08, PSF04, SHPS08]. Parameterizations [DMA02]. Parameterized [GGK06, SH02]. Parameters [BCS96, KR05, WVVV08]. Parametric [Bak90, Bou88, Büh01, FB94, KM83, MJC01, VCP09, WR05, dFSV03, Gui92, NN93, STM93, VG96]. Parametrically [PA06]. PARCUM [Jac85]. Parser [Her82]. Part [LZSCO09, RT08a]. Part-aware [LZSCO09]. Part-type [RT08a]. Partial [BV09, TVD09]. Participating [BN08b, JZJ08a, PWP08, PP09a]. Participation [Ano95l, Ano95m, Ano95k, Ano96g, Ano96h, Ano98h, Ano98i, Ano97e, Ano97g]. Particle [CSI09, CKSW08, HE94, JFSO06, KRG03, KBKŠ09, PTB^{+}03a, PTB^{+}03b, VCC98, YKH^{+}09]. Particle-Based [PTB^{+}03a, PTB^{+}03b]. Particles [BCN03b, BCN03a, CSI09, CATM09, DMLG02]. Partition [NOS09]. Partitioning [CN05, COFHZ98, EGKT08, Hea90, JD00, LLHY09, WLML99, JD01]. Parts [SN86]. Pascal [Sla84]. Pass [LŽY04, MPT98, AWB08]. Past [RSVP02]. Pasteur [Arn08]. Pasting [DLC05]. Patch [DRS09, RT08b, TWJ06]. Patch-based [TWJ06]. Patch-type [RT08b]. Patched [HM83]. Patches [DLC05, DRS08, HJ99, IK01a, RDG01, Guo93, NN93, SAE93]. Path [BRS01, BBS^{+}09, Lam09a]. Pathlines [BGB^{+}08a]. Paths [LG96]. Pathways [SKKS08]. Patination [CS00]. Pattern [BBT^{+}06, GPD09, TVD09]. patterning [Bak90]. Patterns [LA06, ME98, Par86, SSK^{+}05, AS92, CC93]. PBL [MGJ06]. PC [IK01b]. PC-based [IK01b]. PCA [JZF^{+}09, VS09]. PCA-based [JZF^{+}09]. PDE [DQ00, ZY04]. PDF [CRW09]. PDLs [Dau90]. Pearlescent [EKM01]. Pedestrian [PPD07]. Peek [SC08a]. Peek-in-the-Pic [SC08a]. Peeling [RSK06]. Pen [FJW^{+}05, FMS01, STD09]. Pen-and-Ink [FJW^{+}05, FMS01]. Pencil [SB99b, SB00]. Pendulum [PGR96]. Penrose [RMA88]. Pent [MNP08]. Penumbra [LA05]. People [SPR^{+}94]. Per-Face [BL08]. Per-Fragment [TRSKK08]. Per-Pixel [CZGF05, LKC08]. Perception [AMS09, BBT99, CWM09, HMDO05, HKMS08, KMS05, MMTH09, McN01, NMP98, RAP08, RIF^{+}09, SABG^{+}05, SLTM08]. Perception-based [HKMS08]. Perceptual [Čad08, GMY97, GMSK09, HMDO05, IP00, May99, May00, RAP08, SL01, WPG02]. Perceptually [FP04, GH97, HMYS01, HBK02, Red96, SLTM08]. Perceptually-Driven [GH97, Red96]. Perfect [CJC^{+}09, WBEF97]. Performance [Bel87, CGG^{+}03a, CGG^{+}03b, Den86, HE01, MSH^{+}92, MRS08, NAS07, PGSS07]. Persistence [CCSG^{+}09, Ede06, WG09]. Personal [Gre85, Mor86, SG97]. Personalised [KQWM08]. Perspective [AS95, KL03]. Pertinent [Cal96]. Perturbation [Day90]. Peucker [VW90]. PEX [WT93]. phase [MMS09]. Phenomena [BR07, BPMG04, CG07a, DGA04, GMW04, KB04, KH92, GP06]. PHIFI [AM92]. PHIGS [AE86, AHR85, AM92, Bak90, Bak91b, BBMR88, Fre90, HM91, Hew84b, How89, How90, Kra89, Pat89, PNR89, WT93, WH89]. Philipp [Ano07b]. Phoneme [DN08]. Phoneme-Level [DN08]. Phong [KB89, PA06]. Photo [BPV^{+}09, CS93, JBS^{+}06, LMS04, Pur03a, Pur03b, QNTN03a, QNTN03b, RMS^{+}08]. Photo-Realistic [QNTN03a, QNTN03b, CS93, Pur03a, Pur03b, RMS^{+}08]. Photograph [Ano97-41]. Photographic [NK99]. Photographs [ELS08, LGMT00, MZT09, WMTG05, DPT^{+}08]. Photography [AAB09, MKV09, SD09, SSCO09, PCF05]. Photometric [GHH01]. Photon [BGB08b, FD09, HHS05, HHK^{+}07, HP02, JZJ08b, SSO08, SJ09a, WGS04]. Photon-driven [BGB08b]. Photorealism [CLF^{+}03a, GSA03b, MTF03b]. Photorealistic [CMH^{+}01, FCH^{+}06, HS99, LB06, MSK06, DWE03b, DBHM03a, HHD03b, KQWM08, MDWK08, SFWS03a, SP03b, XLTP03b, XJJ^{+}08, YLK08, BJ94]. Photosimulation [Bou90, NN89]. Physical [Fuc04, IIS09, OTH02, SW04a, ESP92]. Physically [BPMG04, CS00, DO00, FL06, HKW09, HE94, LW94, LG96, LLA06, Lew94, LTH08, ME98, NMK^{+}06, OW91, PBP96, Sch94b, UWP06]. Physically-Based [BPMG04, CS00, LG96, LLA06, PBP96, Sch94b, LTH08]. Physics [Bou90, LNS05, MAA^{+}09, Ter02]. Physics-based [LNS05]. Physics-driven [MAA^{+}09]. Pic [SC08a]. Pictorial [DOS93, vvT84]. Picture [Ary84, Chr86, Fah85, Gos89, KPiAS01, SK86, SS91, WH91, WT09, YLK08]. Pictures [Bij87, BBDM85, OAIS09, WM85]. Piecewise [BV96, Fiu95, Li07, NW01]. Pigment [XTJ^{+}07]. Pigmented [Cal96]. Piles [PGGM09b]. Pinchmaps [TC05]. pinscreen [LG92]. Pipeline [AEW90, BR02, BBT99, Hea90, Kra89, FS92, Hea89]. pipeline-enabling [FS92]. Pisa [MRS06]. Pixar [Coo05]. Pixel [BJCO03a, BJCO03b, Bel87, CZGF05, HK09c, LKC08, MCO97, Mil84, PO85, SW04b, PWC^{+}09]. Pixel-Accurate [HK09c]. Pixel-Level [SW04b]. Pixels [Sil97]. Placement [BWRT96, CJ90, FCOL00, Pic86b, SE04, Váz07]. PLAN [CY89]. PLAN-I [CY89]. Planar [BV09, Day88, LVT08, ND06, WBCG09a]. Plane [HHD03a, HHD03b, LCM^{+}06, Szy91a, Szy91c]. Planned [LLSS03a, LLSS03b]. planner [Lam09a]. Planning [CWKS00, GSHN94, Joo86, KAAT03a, KAAT03b, ZV09, vvT84]. Plans [GC96]. Plant [BR97, SFS05]. Plants [LD98]. Plasma [BR07]. Plate [GLW96, SHPS08, WW98]. plates [Lei94]. platform [EFGS96]. Plausible [ED08, Lew94]. Playing [WLSG03]. PLUS [Bak90, Bak91b]. Point [AKP^{+}05, AA06, BSJ08, DBG99, GBKG04, GBP04, GBP05, GGG08, HMB08, JWB^{+}06, JFS09, Kei04, KZ04, KB00, LKC08, LSW09, MMG06, ÖGG09, PKG03a, PKG03b, Pic91b, RPZ02, RL09, SWK07, SWG08, VHB08, VMG09, WSG05, WK04, WDR09, ZQK04, AS00, CATM09, SKC01]. Point-Based [BSJ08, GBP05, HMB08, Kei04, GBKG04, GBP04, WK04]. Point-Cloud [SWK07]. Point-in-Polygon [JFS09]. Point-Sampled [AKP^{+}05, PKG03a, PKG03b, WSG05]. Points [AKP^{+}05, AR95, BGI08, CCFM08, CJW^{+}09, DHvOS00, EBV05, FS91, Gro01, GBP05, GGG08, HH98, LLY09, Nas03, SPOK95, TIS^{+}95, WSG05, WDR09, ZLK05]. Pointwise [FS08]. Poisson [LD08a]. Poitiers [Ano97-29, BH96]. Polygon [CY89, GP87, HL02, JFS09, ML91, Red96, SS00, TP88, TP89, WC05, Rap92]. Polygonal [BB00b, BGI08, BK01, Bou88, CFFP84, CCI08, HS04, HBK02, KVLS99, MS01, Mat85, SB99b, VC04, WF97, WB01]. Polygonization [AG01, HJ99]. polygonizations [AGH^{+}93]. Polygonized [OB01]. Polygons [And89, ANF97a, ANF97b, KP87, ML91, Ric87b, Rok97, RR94, Sil97, YR97]. Polyhedra [ANF97a, ANF97b, BG01, EL01, JA95, KP87, NAB86, RR96, RK94, Sug94, SN84, WBG07, BR96, MLP92, Ska96]. Polyhedral [CKM^{+}99, KT97, MKB^{+}08, OPC96, STK02]. Polyhedron [SN86, HMRSK92]. Polylines [Mat85]. Polynomial [Fiu95, HM83, Li07]. Polynomials [FB94]. Polytopes [HS94, HS08]. Portal [JEO00, HDS03b, HDS03a]. Porto [PS96b]. Portugal [PS96b, Req86]. Portuguese [Dun91]. Pose [FL06, HSS^{+}09, RLN06]. Pose-Space [RLN06]. Positional [EP09]. Positions [WMTG05]. PostScript [DR87]. Posture [ZLK05]. Potential [BDS^{+}03, BRS01, PM93]. Powerful [CY89]. Practical [Ano99m, CAM08b, FCP^{+}90, HDSD99, MKV09, MCHAM08, SCD05, GS85]. Practice [Bar05, BPF^{+}03b, Day88, ESK03b, Sei88, SPS94, TCH^{+}03a, VGSS04]. Prague [Gob96]. Pre [JFS09]. Pre-processing [JFS09]. Precise [SFWS03a, SFWS03b]. Precision [SWS09]. Precomputed [BN08b, Gar09, PLPB07]. Preconceptions [ZK09]. Predicting [AMS09]. Prediction [MG95, SSSB07, WO94]. Predictive [DAP08, LB06, LNS05, PPD07, WW09a]. Preeminent [Hec01]. Preface [FG98]. Prefiltering [Lin85]. PREMO [DDtR94, WHR97]. Preparation [Kle06]. Preparatory [RAP08]. Prepress [DDPL00]. Preprocessing [WBP98]. Prescription [BCGB08]. Presence [JZJ08a]. Presentation [Baj03a, Kob03b, Pur03a, SC84]. Presentations [Ano98v, Col05, Coo05, Des04, Dut04, Ede06, Fuc04, Han05, Kle06, KH96, Sta06]. Preservation [PK08, RHC08, SHB07, TRS03a, TRS03b, VPP^{+}04]. Preserving [BK05a, CLDD09, CLME09, CL03a, CL03b, GVWD06, HWC^{+}05, xHMC09, MK05, ÖGG09, SVG^{+}08, TE99, Thü01, Wil87a, WKM^{+}09, XM09, ZCHM09, ZK08, BYB09]. Pressing [CWA^{+}08]. Pressure [Pud94]. Primal [SW05]. Primer [Ede06]. Primitive [BP83a, HR87, IK01a, ND94, SDK09, SM86]. Primitives [Gro01, BR96]. Principal [AM00, TWC^{+}09]. principle [MNR94]. Principles [Ara94, May99, May00, VBB^{+}06]. Printing [DDV^{+}02, MKO^{+}00, MOT99]. Priority [AE86, JD98, MF00, MGAF95]. Priority-Driven [MF00]. PRIP [VL93]. PRIP-a [VL93]. Prism [CC08]. Prism-Free [CC08]. Prize [Ano97-41]. Proactive [But94]. Probabilistic [VLV^{+}04]. Probe [CCC08]. Problem [GW07, Zac89, HR88]. Problematics [MTT89]. Problems [Den03a, Den03b, Mac94, OZ09, SPS94, VW91]. ProcDef [ITYI09]. Procedural [BD04, DG97, GS09, IMIM08, PGGM09b, SS08, SPS95, YKH^{+}09]. Procedurally [JBL^{+}06]. Procedure [MK06, YF85]. proceeding [HHS89]. proceedings [ACM80, BH96, BJ94, DJ88, Enc81, HP95, HK94, PB95, PS96b, Req86, SvZ95, TT95b, TB84, Van80, Van85, WG82, tH83a, Ano96s]. process [Hea89, RPP93]. Processing [AR06, ABC^{+}04, BB08, CKSW08, Des06, DGEG09, Edm83, ESP08, GD85, GE04, GGRZ06, GBP05, HP04, KT09, KO88, Kob03a, Kob03b, KSH04, KM83, KS07, LTH08, LSJK09, MK06, RGW05, SWPL08, SMS01, SSK^{+}05, SSFS06, SZAB04, SP06, Sor06, SSB05, SF83, TSP05, VC04, VL08, WKG85, WMTG05, YGL^{+}09, DPT^{+}08, Jac85, JFS09, SAE93, YN93]. Processor [CY89, Sch88, TP88, VL93]. Processors [BB88, FG88, HB92]. Product [CAM08b, LS08b, PF90, Bar93]. Production [BL08, Coc83, Gol85, Zot08, Jon96]. Professionals [Ano03c]. Professor [Wil06a]. Profiles [DG97]. Profiling [LC09]. Programmable [BIMO04, BS03a, BS03b, Ert02, MRT08, SCD05, SG03]. Programme [BH06, DS09, Ano05g]. Programmer [Hew84b]. Programming [Ano95q, Ano96c, Ano98e, AS98, AHR85, DH98, FK09, Gna82, HM86, MFDA86, Vel99, AHR93, CL92, DM92, VCDF95, WT93]. Programs [Duf88, KM83]. Progressive [BG02, CC06, CLT^{+}08, CL03a, CL03b, DBG99, FP04, GE98, LSS98, NG97, PJ94, PT03, VCP09, WDGT01, BP93]. Project [vL90, Lev99, Mil83]. Projection [BES00, CS98a, DMKP07, Kun04, MMFE08, MHG00, TON^{+}02, DRBR09, WG93]. Projection-Based [DMKP07]. Projections [CZGF05, GTS86, KLD^{+}09, SGCH94]. Projective [Her89, KH02, LJN02, Wil06b]. Projector [BIWG08]. Projector-Camera [BIWG08]. Projectors [KK08, MS08, MRT08]. Prolog [CJ90, HM86, Mil88b, Mil88a]. Propagation [VPLL08, XWT^{+}09, HB94]. Properties [BMD^{+}08, CDPS09, DGV08, EZK08, MG05, OSG08, RWP88]. Proposal [MA85, Mil88b, Ste84]. Proposals [AHR84]. Protein [BGB^{+}08a, PMW86]. Prototype [TZF04, PW93]. Prototyping [MTT89, SNLW01]. Provably [MB08, SOG09]. Proximity [EL01, MRS08]. Proxy [LCD09b]. Pseudo [MTKO02, SSK93]. Pseudo-satellite [SSK93]. Psychophysical [BLD^{+}09]. Psychovisual [DDV^{+}02]. Ptex [BL08]. Purpose [Coc83, OLG^{+}07, Pil85]. Putting [SSA^{+}08]. Puzzles [YLK08]. Pyramid [Dan96, SL08]. Pyramid-Based [SL08]. Pyramidal [GGW98, KS07].

Quad [EGKT08, HW91, LAD02, LZX^{+}08, MK06, MTAD08, MNP08, NSY09, PD04, SL03, LDB07]. Quad-Dominant [MK06]. Quad-encoding [HW91]. Quad/Triangle [SL03, LDB07]. QuadCover [KNP07]. Quadgraph [Par86]. Quadrant [Arn08]. Quadratic [KPS95]. Quadric [FG04, RvBWR04]. Quadrics [VHB08]. Quadrilateral [DSC09b, DSC09a, Kob96, LMB05]. Quadrilateral-only [DSC09b]. Quadrupeds [SRH^{+}09]. Quadtree [CFFP84, PG94]. Quadtrees [MA91]. Quality [BBDM85, CSI09, FAVM09, GBP07, GMSK09, HK09a, HMS09, JC08, KZ08, Kle06, LWZ^{+}09, Los97, LD97, MHG00, NS01, RW08, RPZ02, STM93, SL01, SiK– DM05, ZISS04, DDM03]. Quantization [BFH^{+}98, DMAC03b, RTN03b, CFM93]. Quantized [hKAC07]. Quasi [EMK09, JKS05, NR95, SKFNC97, TPBC09]. Quasi-Affine [NR95]. Quasi-Developable [JKS05, TPBC09]. Quasi-interpolation [EMK09]. Quasi-Monte [SKFNC97]. Queries [BWPP04, EL01, MRS08]. query [Déc05]. Querying [Sal96]. Quick [DKN^{+}95]. Quickens [Str83].

R [Pri85]. R&D [tHS90]. Racing [GC09]. Radial [AR95, BK05b, LWP^{+}04, RSC01, FS08]. Radiance [GKB09, HMS09, HP02, JZJ08b, MSW04, PLPB07]. Radiance-Cache [HMS09]. Radiative [LF00]. Radiosity [Ano99m, AM95, CSN04, CAH00, DHS04, FP94, GH97, GSA03a, GSA03b, HDSD99, Hec92, HS98, KSS97, MPT98, MSSK08, MB99, MCL96, MSF00, NNB97, NSW09, ORDP96, PB94, PJ94, Sbe97, SSS02, SGCH94, SIP07, Sha97, SSS97, SSSS98, SKFNC97, YP95, BP93, GH96, LBT92, NS93, NN93, OCL96]. Radiotherapy [CWKS00]. Radix [HKS09]. Rain [DTA94]. Randi [Cal07].

Random [Coq85, CH09, KCL06, LR88, Sbe97, SKCA01]. Randomized [TWJ06]. Range [AGDJ09, BDA^{+}09, CZ09, HL03a, HL03b, KK08, KMS05, LPK09, MKV09, WO02, YMMS06, PFC^{+}05, RR96]. Range-Space [AGDJ09]. Rank [DDKL09]. Rank-1 [DDKL09]. Ranked [MGS07]. RANSAC [SWK07]. Rapid [GD01, KGL^{+}98, NS01]. Raster [AO86, AO87, AEL^{+}82, HHA82, Lin85, Pil85, Ric87b, RM89, SK86, SF83, Tho86, WW88, vJB85, VL93]. RasterCalc [van87]. Rasterization [LCD09a, MBGS01]. rasterizing [AP92]. Rasters [van87]. Rates [IHS02, PMDS06]. Rational [DF90, JW95, KMTT92]. raw [MK08]. Ray [Bou88, BBDM85, CDP95, CLF^{+}03a, CLF^{+}03b, CDG^{+}07, CWBV86, DHK08, FQK08, FM04, Gar09, GA93, GFW^{+}06, HSS^{+}05, HB00, HHK^{+}07, HMRSK92, KZ08, KSN08, KH95, KHK^{+}09, Kuz94, LD08c, LD08b, LeYTM08, MFS08, MTF03a, MTF03b, NNB97, NG97, PJ94, PGSS07, RS08, RA94, RDG01, She99, SSK07, SD94b, Spe91, SC95, SKALP05, TNF89, UH92, WSBW01, WMG^{+}09, WSE04, YWY08, ZSP98, ZBP99, ES94, Ger92, KJ92, MDC93, MSH^{+}92]. Ray-Casted [CDG^{+}07]. Ray-Casting [FQK08, HSS^{+}05, KZ08, KSN08, RS08]. ray-generators [ES94]. Ray-Tracing [BBDM85, CDP95, SC95, SKALP05, NG97, PJ94]. Raycasting [Frü94, LCD09b, RSTK08]. Rays [DHK08, GGW98, Sun92]. Raytracing [AKP^{+}05, MW06]. Re [HS99, RGW05, SSFS06, VW90]. Re-coloring [RGW05]. Re-evaluation [VW90]. Re-Using [HS99]. Reaching [KAAT03a, KAAT03b]. Reactive [NG03a, NG03b, PPD07]. Real [AHT04, BK05b, BBP08, BNC96, BN08a, CP88, CLH^{+}08, CWK07, CMT02, CMT05, DRS08, FD09, FR00, GS09, GBP05, GJW08, HSS^{+}05, HL01, HS04, HK00, IK01b, Kle06, KC08, KMN^{+}08, LD04, LKC08, LO95, Mai00, MSW04, NK99, NB94, NSW09, NG03a, NG03b, NS09, PD04, PP89, Pud94, RZLG08, RLN06, RIF^{+}09, RH06, RSKN08, RD05, SSSK04a, SW08a, SL07, SDB97, SG03, TCRS00, TLC02, UT02, WS03a, WS03b, WRS01, XGL^{+}07, YWB03, YSL08, YMMS06, ZCP07, hZCK98, BCF^{+}05, DRBR09, HGW92, Kni93, Lam09a, MC02, SKSK07, YNBH09]. Real-Time [BK05b, BN08a, CLH^{+}08, CWK07, CMT05, DRS08, FR00, GS09, GBP05, GJW08, HSS^{+}05, HL01, HK00, LKC08, NG03a, NG03b, PD04, RLN06, RIF^{+}09, RH06, RD05, SDB97, TLC02, UT02, WS03a, WS03b, WRS01, YWB03, AHT04, BNC96, CMT02, HS04, IK01b, Kle06, KC08, LO95, MSW04, RZLG08, RSKN08, SSSK04a, SL07, SG03, XGL^{+}07, hZCK98, BCF^{+}05, DRBR09, Kni93, MC02, SKSK07, YNBH09]. Real-World [YMMS06]. Realism [Bou90, TSP05]. Realistic [AM02a, BCF^{+}05, BPMG04, Cal96, Han05, HGW92, JBB^{+}08, McN01, ME98, Neb00, PMDS06, QNTN03a, QNTN03b, RBG08, RSTK08, TFK^{+}03a, TFK^{+}03b, XLTP03a, XLTP03b, CS93, Pur03a, Pur03b, RMS^{+}08, Vol93]. Realities [Ano98w]. Reality [AKB^{+}95, BES00, BWRT96, Bry96, Fuc97, Haa96, JL06, SS96a, SG96b, SG97, TLG99, WCB^{+}95, Zar06, MG96, MC02]. Realizing [Bro95, VCDF95]. Really [CAH00]. Realtime [DSW09]. Reanimating [BBPV03a, BBPV03b]. Reasoning [Duk95, MCH94].

Recognition [EBSC99]. Reconstruct [KSS97]. Reconstructed [LCM^{+}09, SPOK95]. Reconstructing [DPT^{+}08, DZM08, LFGG08, MPS08, MB08, SWG08, WL08]. Reconstruction [AKSA09, AAK^{+}09, AS96b, AGDJ08, BV09, BB00a, BCS96, BTG95, BG08, BHGS06, DLRW09, EMK09, FCGA97, FAVM09, GHH01, GKS00, GTS86, GMW97, JWB^{+}06, KPS95, LPK09, LBD^{+}08, MPS08, MB08, Men95, NOS09, NP00, NMOT01, OPC96, ÖGG09, PK08, PZY08, PG08, RL09, RK09b, Sad09, SDK09, SLS^{+}06, VHB08, VMA^{+}04, VMG09, WO02, YWB03, ZLK05, AS00, BBA08, BG93, DF93, GJ02, Koc93, WG93]. Recorder [WLSG03]. Recording [WLSG03]. Recovering [AAK^{+}09, PG94]. Recovery [Kob05, Váz07]. Rectangles [SK86]. Recursive [NW01, PN97]. Reduce [KLD^{+}09]. Reduced [BG08, CZ09]. ReduceM [LeYTM08]. Reducing [WO94]. Reduction [Red96]. Reeb [TVD09]. Refereed [Ano96t]. Reference [CPP08, DP93, DGC^{+}98, Pin92, PNR89]. Refinement [BS08, DRS08, GBP05, KT09, VCP09, GH96]. Reflectance [BR97, FLBS07, GKD07, NN89, NNSK99b, NNSK99a, Sch94c, Sch94a]. Reflected [LF97, PMDS06]. Reflected-Scene [PMDS06]. Reflection [KMG96, MTKO02]. Reflections [PMDS06, RH06, YWY08]. Reflective [IDN03a, IDN03b, LL01]. Reflector [MMP09]. Refractive [HLL07, IDN03a, IDN03b]. Region [AMSF08, EWHS08, Ric87b, BM93]. Region-Fill [Ric87b]. Regions [BK01, CFFP84, MZT09, NSY09, PKPH09, Wil87a, WO92]. Registering [CZ08, HAWG08, LSP08, SG08, SWG08, ZSCO^{+}08]. Registration [CZ08, CZ09, CDPS09, HAWG08, LD06, LSP08, PFC^{+}05]. Regression [KHIK01, NCKG00, ÖGG09]. Regular [AAK^{+}09, AM95, DLRW09, HS94, MMG06, NMMK05, Sad09, TSM94, DSC09b]. Regularised [WSC06]. Regularity [VT94]. Regulatory [WVKR08]. related [CDPS09, Mud83]. Relation [Dau90]. Relational [ADF85]. Relationships [PB90]. Relative [LG96]. relativity [GMDW09]. Relaxation [SJ09a]. Relevant [PMW86]. Relief [ABB^{+}07, BJCW09]. Relighting [CCC08, Dut04, HVM^{+}08, MTR08, WDC^{+}08]. Relocations [dCTAD09]. Relying [Gna82]. Remark [Zac89]. Remarkable [Duc00]. remarks [HR88]. Remembering [Enc98]. Remeshing [DSC09b, KVLS99, MK06, VRKS01, YLL^{+}09]. Remote [GSHN94, MCO97, TL01, Wat96, YN00]. Removal [AKSA09, JD98, SL08, Hea89]. Removing [Pas02]. Render [SPR^{+}94]. Render2MPEG [HKMS08]. RenderBots [SGS05]. Rendered [WW09b]. Rendering [ARC05, AGG^{+}08, ABB^{+}07, Ano07j, AWB08, BW09, BSJ08, BB09, BR07, BMD^{+}08, BJCW09, BWPP04, BG02, BPV^{+}09, BPMG04, BPMG08, BBP08, BG07, BG09, BN08a, BL08, BBP09, CCS95, CS99a, CJW^{+}06, Cal96, CSI09, CLH^{+}08, CKB04, CC08, CYJ02, COF95, CKM^{+}99, CDPS09, CMH^{+}01, DWL^{+}09, DDKL09, Déc05, DSW09, DWE03b, DKN^{+}95, DKC00, DJZ^{+}09, DDC09, DBHM03a, DBHM03b, EWHS08, Elb99, EKM01, EBA^{+}09, ESK03a, ESK03b, FCH^{+}06, FQK08, FP04, FW99, Gar09, GKB09, GCP^{+}09, GMC^{+}06, GKD07, GGK06, GGG08, GFW^{+}06, GSMA08, HK09a, HSS^{+}05, HB96, Hal99, HS99, HL01, HVAPB08, HHS05, HKMS08, HMK^{+}95, HH02, HHD03b, HMB08, HK03b, HK09c, HLL07, IP00, IDN02, IDN03a, IDN03b, JW97, JBB^{+}08, Jan91, JZJ08b, Jen97, KZ08, KSN08, KSL^{+}08, Kle06, KHM09, Kon06, KS07, KW05, KQWM08]. Rendering [LW94, LMD04, LA05, LB06, LD07, LKC08, LLY09, LLA06, LGB^{+}03, LC99, LLG97, LMS04, LLHY09, LG95, LGK97, LCM^{+}06, LCD09a, LCD09b, MMFE08, MT01, MBW08, MS98, MSW04, MPS05, MKB^{+}05, ME04, MGAF95, MDWK08, MMS^{+}05, MSK06, NIDN97, OKP^{+}08, Oll04, PLPB07, PWC^{+}09, PPD98, PSL98, PCF05, PMDS06, PA01, Pur03a, Pur03b, QNTN03a, QNTN03b, RBG08, RS08, RAP08, RPZ02, RZLG08, RSK06, RSTK08, RH06, RMSD^{+}08, RK09a, RMS^{+}08, Rus01, RD05, SDG99, SHLS02, SW08a, Sch94b, Sch88, SPH^{+}09, SC08b, SLTM08, SiK– DM05, SCM^{+}09, SMTG07, SB99b, SFWS03a, SP03b, SE02a, SE02b, SBLD03, SvLD03, SKS09, TSM94, TCRS00, TMO04, TP88, TRSKK08, UWP06, VSD09, WSBW01, WS02, WW99, WWG07, WW09a, WSE04, WE97, Wie96, WO92, WS09b, XLTP03b, XGL^{+}07, XJJ^{+}08, YN00, YM06, YLK08, YW97]. Rendering [YIC^{+}09, ZISS04, ZCG08, BCF^{+}05, BLS93, BJ94, DTKT93, DSSD99, KWF^{+}01, TNK^{+}93, Ano95d, Ano95c, Ano97g, Ano98h, Ano98g, Ano98i, BJ94, DS98, HP95, PS96b]. RenderMan [SPS94, SPS95]. Rendition [HS99]. Repair [BK05a]. Replica [KKK09]. Report [Ano95e, Ano96b, Ano97h, Ano05e, Ano06e, Arb90, Arn84, Arn89, ADS90, Aus91, Bla88, Bon85, Bou90, BP82, BP83b, BG85, Bro90, CP88, Cla89, Cre88, Dau90, Duc90a, Duc91, End83b, End84b, End84c, Gal84, Gre84, Heg90, Hég91, Hew90, HP84, Jan89, Jan91, KB90, Kil85, Kje91a, Kui91, Kwi89, Las84, Le 90, Lis90, Mac85, Mae90, Mar82, MTPS08, Mum86, Mum90, Ott90, PH91, Ros82, Sch85, Sei88, SJ09b, Str84, Suz89, SW83, Vel91, Wat88, Wei08, tH83b, van89b, Kid84, Ano04g, Ano07l]. Reports [Ano98y, Ano98x, Ano99j, Ano99k, Ano99l, Ano00h, Ano00i, Ano02f, Ano02g, Ano02h, Ano02i, Ano05g, Ano06g, Ano07h, Ano07i, Ano07g, Ano07j, Ano07k, Ano07l, Ano08a, Ano08e, Ano08f, Ano08d, Ano09c, Ano09d, Bar06, Bot07, BPB^{+}04, BH06, CA05, Che06, CG07a, CDD09, Des06, DHKS05, DL04, DS09, Dre07, DT04, Duc06b, Duc07, FMK04, GP06, HEtH^{+}83, HJL07, Kau07, Kei04, KSH04, Kon06, Kun04, LMD04, LLR^{+}04, LLD05, LLRD07, MRS06, NSGP06, Neu06, Oll04, Pur07, Raf05, SCA04, San06, SZAB04, SP06]. Represent [PC94]. Representation [BCG^{+}96, BK03c, BK03d, BP01, CZCE08, CL99, Cot85, DDPL00, FLW00, Gos86, GBKG04, Her89, IK01a, JA95, MTKO02, MG09, NAB86, OKP^{+}08, Pat89, PG94, SB99a, SPOK95, XS06, KMA05]. Representations [AAB^{+}96, ADF85, CCI^{+}07, GA98, HMDO05, Har97, Kob03a, Kob03b, Par89, SR96, Sor06]. Represented [AS95, VS09]. Representing [AM00, QSW92]. Reproducing [AHKS94, YMMS06]. Reproduction [KMS05, SSS00]. Republic [Gob96]. Resampling [BK01]. Research [Arn08, Bar05].

Researcher [Bot07, Kau07]. Reservoir [Wat87]. Reshape [CLDD09]. Reshaping [GBU00]. Resizing [xHMC09, ZHM08, ZCHM09]. Resolution [BPMG08, LWZ^{+}09, PK08, PGG^{+}09, VG00, WS02, WH04, WHL^{+}04]. Resource [St88]. Resources [BBS^{+}09]. Responses [LC09]. Restoration [DCPS08, hKAC07, KMS07, LLP00]. Restricted [YLL^{+}09]. Restricting [CLB^{+}09]. Restructuring [xHMC09, LWX^{+}09, ZCHM09]. Results [HH90, Joo86, MSM^{+}08, TSP05]. Retargeting [BSVS04, MBBT00, NJ04, WHL^{+}04]. Retargetting [NJ04]. Retrieval [BSG^{+}95, CTSO03a, CTSO03b, DCG87, TVD09]. Reusable [MCL96, OCL96]. Reverse [HHS05, Jen07, SMAB02]. Reversing [SB99a]. Review [Ano87b, Ano08b, Cal07, DDM03, GL94b, Kau04, Lan07, MFS08, OLG^{+}07, PF90, UWP06, VBB^{+}06, ten82b, GL94a]. Reviewers [Ano07h]. Reviews [Ano97i, Ano97j, Ano97k, Ano98n, Ano98o, Ano98p, Ano99c, Ano99d, Ano00b, Ano00c, Ano02a, Bar05, BIWG08, BDG^{+}04, BBC^{+}05, CON08, How91a, How91b, How91c, How97, LHD^{+}04, MMS^{+}05, NMK^{+}06, Owe89a, Owe90a, RD05, Sor06, TKH^{+}05]. revision [DD92]. Revisions [VD90]. Revisited [MBW08, FZP92]. Ricci [YGL^{+}09]. Rid [SSSS98]. Rig [OZS08]. Rigid [BT95, BPWG07, DMYN08, HS04, HAWG08, LSP08, RKC02, ZV09, CYJ02]. Ring [RM89]. rivers [YNBH09]. Rixensart [BP83b]. Robust [AR94, ATBG08, BG01, BB00b, BCS96, JFS09, KSKAC02, hKLS00, LPK09, LD08b, MK06, NOS09, NHH97, ST08, Sug94, WW09b, MS93]. Rock [PGGM09b, WS01]. Role [Str84]. Roman [Ano06c]. Rome [GBU00]. Room [AHM09, BMD^{+}08]. Root [Szy91c]. Root-Finding [Szy91c]. Rost [Cal07]. Rotating [AKSA09, DJZ^{+}09, SF83]. Rotation [AO87]. Rotations [Mai00]. Rough [Har97]. Roughness [GCP^{+}09]. Round [CD08]. Round-Shaped [CD08]. Rounding [Rok97, YR97]. Rounding-Up [Rok97, YR97]. Route [ZK08]. Routines [RSS96]. Rule [CJ90, RT08a]. Rule-Based [CJ90]. Ruled [AM95]. Rules [SB99a]. Run [DS05a, FND92]. run-length [FND92]. Run-Masks [DS05a].

s [HHS89]. S-splines [KFK94]. Saddles [Pic91a]. SADIST [Mun83]. Sail [Haw85]. Salience [GVWD06]. saliency [CCFM08]. Salient [HBK02, WB01, WG09]. Sample [SEA08]. Sampled [AKP^{+}05, AAK^{+}09, CCI^{+}07, PKG03a, PKG03b, WSG05]. samples [PCF05]. Sampling [AKSA09, AAK^{+}09, AHL^{+}06, BV09, CLH^{+}08, CAM08b, CAE08, DLRW09, DDC09, DZM08, ED07b, FP04, FP94, FCGW02, FBP08, GGG08, HAML05, HK09c, JCJ09, KK02, KC08, LLSS03a, LLSS03b, MPS08, MMP08, MB08, NOS09, NNSK99a, RZLG08, Sad09, SSSK04b, SKS09, TSYK01, Ure00, WGS04, WÅ09, WK04, DTG96, DF93, KJ92, NN93, Sun92]. San [SCA04]. Sand [RSKN08, SOH99]. Sand-Water [RSKN08]. Satellite [BPBD08, SSK93]. Satisfaction [HHS01, Pin92]. Saturated [MZT09]. sbm [Che06]. SC21 [Bon85]. SC5 [Gal84, tH83b]. SCA [SCA04]. Scalable [AWB08, BWS03a, BWS03b, Dwy09, SGC04, YNBH09]. Scalar [NGB^{+}09, AM92, ILRS03a, ILRS03b]. Scale [BPM06, CYJ02, HL03c, HK00, Kje83, MG87, MS93, PD04, SOG09, MMS09, PKG03a, PKG03b]. Scale-invariant [MS93]. Scaling [BCGB08, ZK08]. Scan [CZ09, Che97, CW99, DKS01, Lin85, Rok97, TT94, WH04, WW87b, YR97, SDD^{+}92]. Scan-Conversion [CW99, Rok97, YR97, Che97]. scan-line [SDD^{+}92]. Scanline [BG01, PH87]. scanner [RCM^{+}01]. Scanning [BG08, LVT08, LBD^{+}08, PK08, Sco02, PFC^{+}05]. Scans [AKSA09, LSP08]. Scattered [PS96a, SPOK95, SHLS02]. Scattering [BN08b, JW97, MKB^{+}05, PP09a, RIF^{+}09, SJ09a, WN09, RZLG08]. Scatterplots [BW09, JZF^{+}09, SW09, ZCQ^{+}09]. Scene [FdABS99, cLCtLL98, PMDS06, RA94, SBW06, WO02, YWB03, DH93a, ZHC^{+}00]. Scene-Graph-As-Bus [ZHC^{+}00]. Scenery [SDB97]. Scenes [BG08, CLDD09, CDP95, CLF^{+}03a, CLF^{+}03b, COFHZ98, DMAC03a, DMAC03b, FML06, Gar09, GJW08, GFW^{+}06, LC99, Los97, LD97, MB99, NKF09, ORDP96, PLPB07, PPD98, SDG99, SSS02, SHS99, SC08a, SSK07, SG96b, TL01, VB99, WMG^{+}09, WS02, YIC^{+}09, BP93, TNK^{+}93]. Schedule [HTSFP09]. Schematic [Ben94]. Scheme [ADF85, FCGW02, IP99, MS01, SGM^{+}93]. Schemes [HAML05, LM07, SLLW08, WW98, Ger92, MK08]. Schönhut [Duc06a]. Science [Arn08, Dav07, Fuc97, WG82]. Sciences [JNX^{+}08, KOB^{+}08, MMHL08, WVKR08]. Scientific [Ano97s, Ano97f, Ano98j, Duc98, FH09, Haa96, Le 90, ME04, MKO^{+}08, Mum86, PH91, Wat96, WTHS04, Wei04, FS92, Gob96, MM93, SvZ95]. Scratches [BPMG04]. Screen [JL98, JC08, RvBWR04]. Screen-Space [JL98]. Screw [PA06]. ScribbleBoost [LAA08]. Scripting [LG96]. SCROOGE [Red96]. Sculpting [BF09, DQ00, KFG09, MOT99, RLF09, Vas97]. Sculptured [KGMM97]. Sculptures [Pic91b]. Sculpturing [AM02b]. SDViz [WKM^{+}09]. Seamless [IOI06, SDB99]. Search [ML91]. searches [CLL^{+}08]. Second [BJ94, Cla88, End84a, GCP^{+}09, Vel91, Hég91, Jan91, PH91, SCA04, SZAB04]. SecondSkin [VSD09]. Section [Ano96d, Ano00d, Ano02b, Hew84a]. Sections [BV09, MB08, POB^{+}07]. Seeheim [End84c, Mac85]. Segment [Che97, Ste84]. Segmentation [AGDJ09, BK03a, CLT^{+}08, DGV08, JKS05, KT09, LZ07, LCWK07, PKG03a, PH03a, RT08b, RT08a, Sha08, SLSK07, SLLW08, WE97, YYPZ07, ZQQS08, SNKS09]. Seismic [McC83]. Selected [AR06, Arn08, DCPS08, Zot08, Gob95, Gob96]. Selecting [SNLH09]. Selection [AMSF08, KR05, LSJK09, Sbe97]. Selective [MCO97, vP94]. Self [AMT02, BB07, SN08, VT94]. Self-Collision [VT94]. Self-Learning [BB07]. Self-Organizing [AMT02]. Self-Shadowing [SN08]. Semantic [BCWR08, CN05, JBS^{+}06, FS92]. Semanticons [SABG^{+}05]. Semantics [PL94, RBG08]. Semantics-Based [PL94]. Semi [DSC09b, DRS09, KySK08, LCM^{+}09, Nie95, WS09b]. Semi-Automatic [WS09b]. Semi-Homogeneous [Nie95]. Semi-Lagrangian [KySK08]. Semi-regular [DSC09b]. Semi-Supervised [LCM^{+}09]. Semi-Uniform [DRS09]. Sensing [GSHN94]. Sensitive [SG96b, VRKS01, BG01]. Sensor [GJL^{+}09]. Separatrix [WG09]. September [Ano97w, Ano04b, Arn91, BH96, CG07a, Cre88, DJ88, DT04, Enc81, HHS89, HB91, Kil85, Pur07, TT95b, Van80, Van85, WG82, tH83a, TB84]. sequence [PFC^{+}05]. Sequences [BCS96, BG02, HB00, SEASM09, Koc93, van89a]. Sequencing [WS09b]. Sequential [PWP08, SW08a]. Series [WS09b]. Serious [Saw07]. Service [Ano95w, Ano95x, Ano96q, Ano97p, Ano97-37, Ano97-38, Ano97-39, Ano98-28, Ano98-29, Ano03e, Ano04e]. Services [Ano95y, Ano96r, Ano97-40]. Servoing [MC02]. Sessions [HEtH^{+}83]. Set [AA06, BK03a, BK03b, CKSW08, GGG08, Man83, MMS07, MBW^{+}05, NM91, ÖGG09, Wei04]. Sets [ASVNB00, Bak88, Bak91b, CWM09, CPK09, McC83, NNB97, PSK09, SHLS02, SAA09, VMG09, WK04, AS00, FR92, PFC^{+}05]. Several [Szy91c]. SGS [Pas02]. Shaders [Lew94, SPS95, SW08b]. Shading [BYP95, BL86, Cal07, HSS^{+}05, Kau04, KC08, KB89, MTCT84, Nar95, PA06, Sch94c, SPH^{+}09, TT97, WB02, WN09]. Shadow [AHL^{+}06, BS03a, BS03b, CJW^{+}06, FBGP09, GBP07, LLA06, PWC^{+}09, SL08, SEA08, YFGL09]. Shadowing [NKF09, SN08, TT95a, TT97]. Shadows [AHT04, CJW^{+}06, CAM08b, ED07b, ED08, FBP08, FBGP09, GBP07, GJW08, HMS09, Los97, LD97, MSW04, NKF09, SSSK04a, SS07, SEA08, SN08, UGLY08, WÅ09, WS99, YFGL09, YK08]. Shallow [DHK08]. Shape [AHKS94, Ber09, BK05b, CCSG^{+}09, CLE07, CLME09, Coh95, DGV08, ESP08, EZK08, FG04, GBAL09, HSS^{+}09, HWAG09, KL08, KJC^{+}09, KBS00, Kob03a, Kob03b, LZSCO09, LCWK07, MB08, OSG08, SWK07, SC04, SSB05, SOG09, TVD09, VPP^{+}04, VT94, WL08, WSLG07, WBCG09a, ZY04, ZSCO^{+}08, ZCHM09, DOS93, GGRZ06, KS92, SNKS09]. Shape-aware [CLE07]. Shape-Preserving [ZCHM09]. Shape-simplifying [KL08]. Shaped [CD08, DLC05]. Shapes [AMSF08, CZ08, CCSG^{+}09, FCGA97, GMW97, Li07, OMT02, OPC96, OSG08, RF96, RT08b, RT08a, SDK09, SLSK07]. Shared [KH95]. Shared-Memory [KH95]. Sharp [LWY08]. Shattering [SWB01]. Shaving [NLED08]. Sheared [CS98a]. Shearing [HL02]. shell [Fre90]. Shells [CWK07, WSG05]. ShellTrees [KGL^{+}98]. Shepard [BAU05]. ShieldTester [NRJS03b, NRJS03a]. Shooting [HB00, Sbe97]. Short [Ano98v, DAP08, Kje91a, Sab82]. Short-Horizon [DAP08]. Shortest [SS91]. Shrink [KVLS99]. Shrinkability [ZHM08]. Shrinkage [VPLL08]. Shrinking [EBV05]. Shutter [TSY^{+}07]. SIAS [van89a]. SIBGRAPI [AR06]. Sided [NSY09]. SIGGRAPH [Des06, Rob87, Ano95z, Ano97-32, End83b, Hop84, Jan89, LLR^{+}04, Ros82, Ano98-30, LLD05, LLRD07, Sch98]. Siggraph/EG [Ano98-30]. Siggraph/Eurographics [Sch98]. Sight [NYTN87]. Sigma [DM92]. Sign [LO95, RB03a, RB03b, HR92]. Signal [BB88]. Signature [SOG09, XYL09]. Signatures [AA09, CCSG^{+}09]. Signed [CCI08, HR92]. Silhouette [CC08, DRS08, HK09c, IHS02, KL03, OZ06, OZ09]. Silhouettes [CLM09, IHS02, MT01, TZF04, MFT02, WBCG09b]. SIMD [AO86, AO87, BIMO04, DHK08, TP88]. Similarity [CTSO03a, CTSO03b, FAT07]. Simple [Dwy09, FA87, KSKAC02, MKV09, RR94, SLTM08, SN84, BSV92, Rap92]. Simplex [BYB09, DSC09b, DSC09a, SNKS09, YLL^{+}09, dCTAD09]. Simplicial [JFSO06]. Simplification [AS96a, ABC^{+}04, AWB08, CLH^{+}08, CL03a, CL03b, DGGP05, ED07a, ESV99, ESC00, FCGW02, HLS96, LS09, RMS^{+}08, SC08b, VLV^{+}04, VW90, WBP98, vKP06, BC01]. Simplified [CRS98, DSC09b, OPC96, RT08b]. simplifying [KL08]. Simulated [KMN^{+}08, Sch94a]. Simulating [Buc98, DGA04, GLHB09, Sta97, Wu90, Ye08]. Simulation [Ai98, Ano95-27, BNC96, CWKS00, CWK07, CMT05, DAP08, DYN04, DMYN08, Enc82, FBT99, HS04, Heg90, HE94, HH98, HK00, KTMN07, KMTT92, KySK08, KW05, KMN^{+}08, MBCN09, MMS09, OIST91, Pai02, PPD07, PTB^{+}03a, PTB^{+}03b, RP01, RIF^{+}09, RGTC98, SLS04, SL07, SWB01, ST08, TTW90, UT02, Wat87, WMWG09, Wil87b, WDGT01, YKH^{+}09, ZY02, ZCP07, BH96, HPT89, KWF^{+}01, LG92, RPP93, SW92a, SGC04, TT95b, Ano98f, Hég91, TvdP98]. Simulations [ATBG08, BGAM04, OHBKH09, SCD05]. Simulator [PCR89]. Simultaneous [ABC^{+}91]. since [Joo86]. Single [AWB08, AW07, BPV^{+}09, GE04, HP83, PP09a, RZLG08, SC08a, SL89, WL08]. Single-pass [AWB08]. Single-scattering [RZLG08]. Single-Strip [GE04]. Single-Valued [SL89]. Singular [Got94]. Singularities [RRS97b, RRS97a]. Sites [GBU00, PL96]. Sixth [Ano95-27, AS98, Kui91]. Size [BGAM04]. Skeletal [FCGA97, WSLG07, YCLE09]. Skeletally [KŽ05]. Skeleton [DTTS08, HBK02, ITYI09, MBBT00, PWH98, YBS07, SLSK07]. Skeleton-based [DTTS08, HBK02, YBS07]. Skeleton-free [ITYI09]. Skeletonization [BB08]. Skeletons [RT08b]. Sketch [AHKS94, BN08a, Che06, GM96, LFGG08, SSB08, SS08, DH93a, WBCG09b]. Sketch-Based [Che06, LFGG08, BN08a, SSB08, SS08, WBCG09b]. Sketched [KO88, Pud94, TZF04]. Sketches [OOI05]. Sketching [BG95, BCF94, HLH96, IOI06, KQWM08, SSB08, SC04, WSBZ08, KHR02]. Sketchpad [EBSC99]. Skin [KB04]. Skinned [RHC08]. Skinning [GOT^{+}07, GW05, SHB07, WSLG07, YBS07, SW92b]. Skull [DPT^{+}08, IIS08, PZY08, WL08]. Skulls [PZY08]. Sky [DKYN96]. Skylight [DKN^{+}94, TNK^{+}93]. Slab [LAM09b]. Slabs [DWL^{+}09]. Slaved [SS00]. slice [FND92]. Slider [STD09]. Slit [LVT08]. Slope [Thü01]. Slow [Pas02]. Slusallek [Ano07b]. SMAC [SCD05]. Small [VW08, Day92]. SMALLTALK [Moh87]. Smart [Osh08, EZK08]. SMARTPAPER [SC04]. Smoke [DMYN08, RR00]. Smooth [CWA^{+}08, GW07, HTH96, Li07, LWY08, LS08b, MNP08, PG94, SWS09, VMA^{+}04]. Smoothed [KBKŠ09]. Smoothing [NOS09, Thü01, VMM99, EZK08, MLP92]. Smoothly [VT94]. Snakes [LL02]. Snow [NIDN97, SOH99]. Soap [Dur01]. Soccer [BB00a]. Society [Fra83]. Soft [AHT04, AHL^{+}06, BMG99, CJW^{+}06, ED08, FBGP09, GOT^{+}07, GCMS00, GBP07, GJW08, HE94, HL03a, HL03b, LLA06, Los97, LD97, RGTC98, SS07, SEA08, SN08, UGLY08, YFGL09]. Software [BP83b, BG85, BD08, CPP08, Duc82a, Gna83, KH96, Mac84, Mar82, Mum86, TA08]. Solid [BL86, DGF98, DLTD08, GM96, PH87, XTLP02, YSY94, FFD93, Gro92, Jac85, WG93]. Solids [Bel87, KGMM97, NHH97, Nic85, Pai02, PN97, SPOK95, SLS04]. Solution [CDP95, FCP^{+}90, PP09a, Ska87, TCH^{+}03a, TCH^{+}03b]. Solutions [Hei01, Ric87a]. Solvent [BGB^{+}08a]. Solver [HE01, KySK08, Sta06]. Solving [Den03a, Den03b, GW07]. Some [HR88, HH90, Ric87c, vJB85, DDR93]. Sorman [Ano06c]. Sorting [CKM^{+}99, HKS09, SBW06]. Sound [BMD^{+}08, DYN04, HS04, SLS04]. Source [AHL^{+}06, Sbe97, SSSK04b, TA08, Váz07]. Sources [BYP95, GHH01, KHM09, MMP08, TT95a, TT97, ZBP99, KJ92]. Space [AGDJ09, AHT04, AB97, BTB02, BT95, BT98, BiA06, BPWG07, CS98a, CLH^{+}08, CN05, DCOM00, DHS04, GGW98, GW07, HMB08, JD00, JD01, JL98, JC08, LLHY09, LCM^{+}09, Mül86, NSW09, OZ06, PB07, PZB^{+}09, RLW^{+}09, RPZ02, RLN06, RK94, RNtH03, SLAM08, TW97b, ZC95, AS00, MSH^{+}92, TW97a]. Space-Optimized [BTB02]. Spaced [JL00, SLCZ09]. Spaces [KE97, SL09, SC95]. Spain [DT04, Gob95, NSGP06]. Span [BB99]. Spanning [THN93]. Sparse [CCFM08, RK09a]. Spatial [BT95, DG97, GKB09, IK01a, Kuz94, PG93]. Spatially [LLSS03a, LLSS03b, RK09b]. Special [Buc96]. Specific [BMD^{+}08, SJH08]. Specification [Ano98d, Dam91, Duc82a, DF85, Duc85, Gre84, LS89, Mac84, MJ98, ND94, PCS94, PP05, TC94, WGK88, DD92, DP93, FZP92, PB95]. Specifying [van90]. Spectacle [KTMN07]. Spectral [ARC05, DMCP94, DG97, IP00, KB04, LZ07, MTAD08, PSK09, TYK^{+}09, VL08, XS06, Sch94a]. Spectral-Based [TYK^{+}09]. Spectrum [MRT08]. Specular [BP01, GCP^{+}09, RH06, SKUP^{+}09, WB02, YW97]. Speech [BBPV03a, DN08, KMT03a, KMT03b, LO95, RB03a]. Speed [AGM^{+}06, TSY^{+}07, DZTS08]. Speed-up [AGM^{+}06]. Speeds [GLHB09]. Sphere [PG95, HMRSK92]. Sphere-Trees [PG95]. Spheres [LCD09b, Pat93]. Spherical [GP09, GCP^{+}09, GJW08, JCJ09, MSW04, PGR96, PS95, TRSKK08, XS06]. Spines [JNX^{+}08]. Spirals [Pic91a]. Splash [TFK^{+}03a, TFK^{+}03b]. Splatting [CS98a, GBP04, HMYS01, HHK^{+}07, HMS09, LKC08, RPZ02, WW99, WK04, ZCQ^{+}09]. Splicing [HKG06, LWX^{+}09]. Spline [DF90, FS91, GM06, Gre94, KGL^{+}98, NW01, ND06, YP95, Guo93]. Splines [GBS99, KZ08, LS08b, McC96, PS95, PS96a, Sei94, SHPS08, Vas97, WW98, KS92, KFK94, VMG09]. splines/S [KFK94]. Splitting [KySK08]. Spoken [MSWK02]. Spring [HH98, TK98]. Squares [HWC^{+}05, RNtH03, SB99a]. Stability [CCSLT09]. Stabilization [CLHL08]. Stable [CCSG^{+}09]. Stackless [PGSS07]. Stand [ZHC^{+}00]. Stand-alone [ZHC^{+}00]. Standard [DH98, ELM^{+}83, Klo87, WHR97, FZP92]. Standardization [End82, Gna83, ten82a]. Standardized [ADJ^{+}01]. Standards [Duc82a, Hew90, OF84, RSK90, Sab82, Str83, UH92]. Stars [NGN^{+}01].

State [Bar05, BIWG08, BDG^{+}04, BBC^{+}05, CON08, DDM03, LHD^{+}04, MFS08, MMS^{+}05, NMK^{+}06, OLG^{+}07, RD05, Sor06, SKU08, SKUP^{+}09, TKH^{+}05, UWP06, VBB^{+}06, WMG^{+}09, Cai95]. States [Kau04]. Statistical [CLL^{+}08, HSS^{+}09, CCFM08, SDD^{+}92]. Status [Ano84a, KHIK01]. Statutory [Ano04g, Ano05e, Ano06e, Ano07l]. Steering [GMY97]. Steganography [BGI08]. Stellar [VLV^{+}04]. Step [MK06, MG95, Sab82, LSŽ08]. Stereo [CCC08, TGS96]. Stereochemical [DCG87]. Stereoscopic [ABC^{+}91, Koc93]. Still [Mac94, ZR96a, ZR96b]. Stipple [DHvOS00]. Stippling [HHD03a, HHD03b, KSL^{+}08]. Stochastic [RP01, SF92, Sta97, SK99, TSYK01, EZK08]. Stockholm [Kje91a]. Storage [FT93]. Storytelling [Gla99]. Straight [Liu94, Liu93b]. Strain [TPS09]. Strands [Hol94, Pai02]. Strategies [Ano99m, HDSD99]. Strategy [KSKAC02, TMO04]. Stratified [GM06]. Stream [HKD^{+}08, SWS09]. Streaming [MPS08, TL01]. Streamlines [JL00, SLCZ09]. STRETCH [GS85]. Stretcher [HK00]. Stringed [TCH^{+}03a, TCH^{+}03b]. Strip [GE04, Ise01, dFSV03]. Stripification [DGGP05]. Stroke [BBT^{+}06, GCL^{+}06, Sch00]. Stroke-Based [Sch00]. Strokes [IHS02, WHCO08, van89a]. Strong [COFHZ98]. Structural [PZY08, TA08]. Structure [AAK^{+}09, BW00, BK05a, CLDD09, CJC^{+}09, CYJ02, CCP09, Cot85, How90, xHMC09, KHH^{+}09, Kob05, MK05, MH00, PVtHR09, Str84, TPBC09, WBCG09b, DKW94a]. Structure-Preserving [CLDD09]. Structured [DG95, Frü94, LD04, Mun83, DF93, PC92, RCM^{+}01]. Structures [GWT^{+}08, JBTS08, Kuz94, MMHL08, Sta97, Wai88, WTHS06]. Structuring [VMA^{+}04, KH92]. Students [Kol08]. Studies [BRL09, IEGC08, JR08, MSM^{+}08, STD09, ZK09, AS92]. Study [MSM^{+}08]. Style [BG07, FMS01, PGG^{+}09, UGB^{+}04]. Style-Based [UGB^{+}04]. Styles [HS99]. Stylized [CKB04, EWHS08, TMO04, ZISS04]. Stylizing [IHS02]. Sub [BK03a, BK03b, yKL08, PWC^{+}09, WK04, MMS09]. Sub-Joint [yKL08]. Sub-pixel [PWC^{+}09]. Sub-Sampling [WK04]. sub-scale [MMS09]. Sub-Voxel [BK03a, BK03b]. Subdivision [ADS06, BT95, CLT^{+}08, Kob96, KSS97, LG00, LMB05, LM07, LWY08, MS01, MH00, MTF03a, MTF03b, NW01, Nas03, NSY09, Pas02, PN97, RT08b, RT08a, RNV07, SB99a, SMAB02, SHLS02, SLLW08, SD94b, SL03, WM09, WW98, WP04, ZC95, ZQQS08, LDB07, MSH^{+}92]. Subdivisions [LM07, ES94]. Submersible [RL84]. Subscenes [CSN04]. Subsequent [IOI06]. Subspaces [SJH08]. Subsurface [AWB08, MKB^{+}05]. Suggestive [SP03a, SP03b]. Summed [HSC^{+}05]. Summed-Area [HSC^{+}05]. Sums [GA96]. Super [Ber09, LWZ^{+}09]. Super-Helices [Ber09]. Supercomputing [BBC^{+}05]. Supercover [ANF97a, ANF97b]. Superresolution [FLBS07]. Supersampling [GGW98]. Supervised [LCM^{+}09]. Supplement [BFTL82]. Supplementary [Ano99m]. Support [EWHS08, SSB05, WSC06, FT93, ST93]. Supported [NP00]. Supporting [GCMS00, NM91, MK08]. Surface [AG01, AS96b, BV99, BK01, BK03c, BK03d, BHGS06, CCS95, CD08, CK84, Cot85, Dan96, DMA02, DZM08, EL01, EP09, GWT^{+}08, GJ02, GKS00, GLW96, GE98, GMW97, HP04, HWC^{+}05, HJ99, HLH96, JD98, JW95, KNP07, KL03, Kob05, KPS95, KT97, LPK09, LS89, LF00, LSJK09, LBD^{+}08, LZSCO09, MPS08, Mar95, Men95, Mil84, MH00, NOS09, NRJS03a, NRJS03b, OS08, PP05, RPZ02, RT08b, SP97, SWPL08, SB99a, SPOK95, SS08, SSFS06, SLS^{+}06, SG08, SE02a, SE02b, TPBC09, VHB08, VMA^{+}04, VT94, VMM99, WSC06, WK04, WGG99, ZRKS05, ZY04, AS00, BC01, BBA08, BG93, DMP93, HHB93, HB92, KH92, LDB07, SDD^{+}92, THN93]. Surfaces [AES94, AKP^{+}05, AA06, AE97, AM95, BSJ08, BCG^{+}96, BES00, BK03a, BK03b, BTG95, Blo97, BW07, BPMG08, Bou88, CZ08, CLT^{+}08, CRS98, CJW^{+}09, DLRW09, DQ00, DZM08, Elb99, EBV01, FS91, FM04, FG04, FB94, Gre94, GGK06, GGG08, Har97, HM83, HTH96, HBK02, HAWG08, IDN03a, IDN03b, Jen97, JC94, KNP07, KHK^{+}09, KVLS99, LLG97, LSP08, LWY08, LZ07, LS08b, MPS08, MK99, MS98, MB08, MTF03a, MTF03b, MNP08, NOS09, Nas03, NSY09, NGB^{+}09, ÖGG09, PA06, PKG03a, PKG03b, PK08, PN97, POB^{+}07, RS08, RHv95, RRS97b, RL09, SMAB02, SHLS02, SWS09, She99, STK02, SLCZ09, SSS97, SG08, SWG08, TZF04, TSYK01, TCRS00, TW97b, Thü01, WB01, WG09, WP04, YYPZ07, ZSCO^{+}08, Bak90, BP93, CL92, DTG96, Ger92, Guo93, KHR02, RRS97a, STM93, SW92b, TW97a]. surfaces [VG96, Vel92, dS96, Ano95b]. Surgery [BNC96]. Surperspective [TON^{+}02]. Survey [AM02a, Han97, Hei01, Kau04, LD07, OLG^{+}07, PP05, PT88, RP98, RD05, Sab86, Sch94c, Szy91c, Vel92, WH89, Pri85, Sha08, WG93]. SVG [HD02]. SVG.Open [Neu06]. SW [PRS89]. Sweden [Oll04]. Sweep [CBS96, LCM^{+}06, Mül86, YK06, SW92b]. Sweep-based [YK06]. Swept [JW95, WC02, YYPZ07]. Swirling [PKPH09]. Switzerland [Kei04, Kun04, Van80]. SX [KSH92]. SX/Tools [KSH92]. Symbiosis [Haa96]. Symbolic [Elb95]. Symmetric [TSH01, AS92]. Symmetrical [CC93]. Symmetries [OSG08, YM09]. Symmetry [BBW^{+}09, VBP^{+}09, YM09]. Symposium [AR06, Ano07j, BPB^{+}04, CDD09, Des06, GP06, HJL07, Kei04, LMD04, Oll04, Raf05, SCA04, San06, SZAB04, Wei08, HK94, KSH04, SP06]. Synch [KK07]. Synchronisation [MB99]. Synchronization [LG96, LL05]. Synchronized [GA98]. Syntax [Kin95]. Synthesis [BBT^{+}06, CD08, DN08, DGF98, DLC05, DLTD08, ELS08, Fah85, GSHN94, HK09b, JBS^{+}06, KRFC09, LR88, McN01, MAA^{+}09, MMS^{+}05, NMMK05, Osh08, PRS89, PB90, SPD07, SL09, TWJ06, UGB^{+}04, WPG02, WS01, YD88, YKH^{+}09, ZPS03, BG89, Hor90, Pin92, RPP93]. Synthesizing [DSC95, DYN04, FLBS07, HLM97]. Synthetic [NT95, PO02, ST94, SHS99, TSY^{+}07, TTW90]. System [AHKS94, BRS01, Ben94, CKB04, CJ90, FW99, Haa96, HD95, HGG^{+}84, IMIM08, JZF^{+}09, KMG96, KH96, KT97, LM96b, LLP00, LO95, MSWK02, MOT99, ON05, Pil85, Rey86, RSK90, RL84, ST94, Sch88, SC04, SNLW01, Ste85, WH04, WW87a, WKM^{+}09, WLSG03, WGG99, vvT84, AM92, CFT86, DTKT93, DP93, GJ02, HR92, Jac85, JW89, LM96a, OYSO92, VR95, DSS90, Enc82, Hew84b, WO94]. Systematic [AKSA09]. Systems [Ano91b, Ano98d, AHR84, BES00, BIWG08, Dau90, DMKP07, Duc90a, End84c, Fuc97, HM91, HE94, Mac85, MJ98, MG87, Mil90, MSK06, PCS94, PF90, RP01, SGS05, TO97, CBSF07, DKY98, FZP92, HS92, Kje92, PB95, Sam93b, Sas92, Str82, Vol93]. Systolic [KP87, Mil84].

T.Node [MMAG93]. Table [CAE08, EBSC99, HSC^{+}05, RW08, WS09a, SZG93]. Table-based [WS09a]. Table-driven [CAE08]. Tabular [CRW09]. Tactics [ZV09]. Tactics-Based [ZV09]. tailored [EFGS96]. Talk [Bij87]. Talks [Ano08c, Ano09b, Jen07, Šár07, Saw07]. Tangential [OZ09]. Tangible [Zot08]. Tapered [RSVP02]. Targets [YLD07]. Task [KCB97]. Task-oriented [KCB97]. Taxonomy [FGM99]. TC97 [Gal84, tH83b, Bon85]. TC97/SC21/WG2 [Bon85]. Teaching [BB07, Bro90, Duc86, GC09, MGJ06]. Tearing [MBCN09]. Technical [DWE02, Dre07, HGG^{+}84, WKM^{+}09]. Technique [AGM^{+}06, GC96, KQWM08, OM01, van90, SAE93]. Techniques [AM02a, BLD^{+}09, BN08a, GL94b, GBKG04, GBP04, Han97, HTSFP09, HCG08, HW91, LHD^{+}04, LD07, MFS08, MCT01, NB88, NLED08, PIWB98, Ros82, RD05, SP97, SSB08, SS08, Sha08, WK04, XWT^{+}08, ZT96, CFM93, DOS93, FFD93, GL94a, HP95, JR08, KWF^{+}01, MSH^{+}92, PS96b, SDD^{+}92]. Technische [Enc81]. Technologies [BBC^{+}05]. Technology [Coo05, Enc98, Kun04, Sco02, WG82]. Telepresence [Fuc97]. Television [Ste85]. Template [YK92]. Template-based [YK92]. Temporal [MKO^{+}08, RIF^{+}09, WS09b]. Temporally [SiK– DM05]. Tension [dD85]. Tensor [CZCE08, HVAPB08, LS08b, RK09a, TSH01, VMA^{+}04]. Term [Sab82]. Terminal [MTCT84]. Terminals [Ric87b]. Terrain [Ano07b, CLDD09, CGG^{+}03a, CGG^{+}03b, COS95, DSW09, GMC^{+}06, MDWK08, PGGM09a, PD04, WMWG09, DKW94a, KH92]. Terrains [AG06, BN08a, PGGM09a, CS93]. Terrorism [WMS^{+}08]. Tesselation [MH00]. Tessellation [AES94, DRS09, SS09]. Test [JFS09, NRJS03a, NRJS03b]. Testing [Ano88b, CB09]. Tetra [JFSO06]. Tetra-Trees [JFSO06]. Tetrahedra [AMSF08, HJ99, MDP08]. Tetrahedral [CL03a, CL03b, RNV07]. Tetrahedralizations [LD08c]. Texas [Rob87]. Text [AHM09, CFT86, CCP09, GD85, IF09, SE02a, SE02b, HR92, BP83a]. Texture [AE97, BTB02, BD04, BL08, CLH^{+}08, CC06, CD08, DGF98, DMLG02, DLC05, DLTD08, ESG01, ELS08, GD96, HSC^{+}05, IK00, IK01b, KH02, LHD^{+}04, LG95, LWX^{+}09, LDR09, MTCT84, Mar95, MK99, MO08, MMS^{+}05, MCHAM08, NMMK05, SPD07, SWB98, SS96b, SBLD03, SvLD03, TE99, TC05, TWJ06, VB00, WPG02, Wei04, WOBT09, XWT^{+}09, YD88, BG89, BG93, CPE92, GK03a, GK03b, NG92]. Texture-Based [LHD^{+}04, Wei04, VB00]. Textured [CLDD09, FBGP09, MMS07]. Textures [BL08, CYJ02, DDKL09, DWL08, DG95, DG97, DYN04, DLTD08, ED08, EDM^{+}08, ELS08, GK03b, IMIM08, KLD^{+}09, LJN02, LLSS03b, LDR09, OAIS09, PGG^{+}09, SvLD03, WS01, WS09a, WW09b, WOBT09, DSC95, TC05]. Texturing [NK99, POB^{+}07, RLW^{+}09, ZQK04]. Them [Ros83]. theoretic [AS92]. Theoretical [Bou88, LW94]. Theory [ARC05, Bar06, Büh01, Day88, FdABS99, PA06, SWPL08, Sei88, WO94, Bar92]. There [Mac94]. Theses [Kje89, Kje90, Kje91b, Kje91c, Kje91d, Kje95]. Thin [CWK07, DWL^{+}09, GLW96, Pai02, SHPS08, SW08b, WW98, WSG05]. Thin-Plate [SHPS08]. Third [Bla88, Cre88, Des06, Kwi89, Suz89]. Thoughtful [Bur95]. Three [AHKS94, DBG99, HS94, HMK^{+}95, LD06, Sal96, SS96a, TMT86, ZR96a, ZR96b]. Three-Dimensional [LD06, SS96a, TMT86, ZR96a, AHKS94, ZR96b]. Thresholding [LR88]. Throwing [CJW^{+}09]. Tickmarks [Kje83]. Tiled [MG05, PD04]. Tiles [ABC^{+}04]. Tiling [EGKT08, LZX^{+}08, Mey94, MTAD08, PCK09, PGGM09b]. Tilings [RMA88]. Timberline [Bon85]. Time [Ber09, BK05b, BBP08, BN08a, CLH^{+}08, CWK07, CMT05, CPK09, DRS08, FD09, FR00, GS09, Got94, GBP05, GJW08, HSS^{+}05, HL01, HK00, Kel86, LD04, LKC08, Mai00, NB94, NSW09, NG03a, NG03b, NS09, PD04, PP89, Pud94, RLN06, RIF^{+}09, RH06, RD05, SW08a, SDB97, SLAM08, SWG08, TCRS00, TLC02, TSH01, UT02, WS03a, WS03b, WRS01, WAH^{+}09, WS09b, YWB03, YSL08, ZCP07, AHT04, AB97, BCF^{+}05, BNC96, CMT02, DHS04, DRBR09, HGW92, HS04, IK01b, JR08, KB92, Kle06, Kni93, KC08, Lam09a, LO95, MC02, MSW04, PZB^{+}09, RZLG08, RSKN08, SSSK04a, SKSK07, SL07, SG03, XGL^{+}07, YNBH09, hZCK98]. Time-Adaptive [CPK09]. Time-Dependent [TSH01]. Time-Series [WS09b]. Time-Varying [SWG08, WAH^{+}09, JR08]. TimeArcTrees [GBD09]. TimeRadarTrees [BD08]. Tissue [BR97, BMG99, RGTC98]. Today [Mum89]. Together [Edm83]. Token [Rey86]. Tokyo [FMK04]. Tomography [FAVM09]. Tomorrow [Mum89]. Tone [BLD^{+}09, JBS^{+}06, KMS05, MS08, SSS00, SKMS06, YMMS06, MMTH09]. Tone-mapping [MS08]. toning [Buc96]. too [Mac94]. Tool [CMS94, Gol85, HE94, NM91, SDC09, MM93]. Toolkit [DR87, SD00, TLG99, WHR97]. Tools [BY08, Gre84, HV08, RNtH03, WVVV08, CTW92, WT93, KSH92]. Top [Sla88]. TopDraw [How90]. Topological [BGK^{+}96, Ede06, RA94, TIS^{+}95, WTHS04, WTHS06, WG09, Lam09a]. Topologically [BR96]. Topologies [CL03a, CL03b, SMAB02, MLP92]. Topology [BK03a, BK03b, FML06, GE04, HK09b, Kob96, SVG^{+}08, The02a, TRS03a, TRS03b, TSH01]. Topology-Based [FML06]. Topology-Preserving [SVG^{+}08]. TopoPlan [Lam09a]. Touch [ten82b]. tough [Gue82]. Toulouse [PB95]. Tour [KPiAS01]. tours [AVF04]. Town [Joo86]. Trace [FM04]. Traces [CPP08]. Tracing [BMD^{+}08, Bou88, BBDM85, BBS^{+}09, CDP95, CLF^{+}03a, CLF^{+}03b, CDG^{+}07, CWBV86, DHK08, Gar09, GFW^{+}06, KH95, KRG03, KHK^{+}09, LD08c, LD08b, LeYTM08, LSS98, MFS08, MTF03a, MTF03b, PGSS07, RA94, RDG01, She99, SSK07, SD94b, Spe91, SC95, SKALP05, TNF89, WSBW01, WMG^{+}09, WSE04, YWY08, ZSP98, ZBP99, ES94, GA93, Ger92, KJ92, MDC93, MSH^{+}92, NG97, PJ94, Sun92, UH92]. Track [Arn08, Zot08, DCPS08]. Tracked [TGS96]. Tracking [MG95, RSVP02, TSH01, vKB94, MG96, SKSK07]. Tracts [CZCE08]. Trade [HHS01]. Trade-Off [HHS01]. Trajectories [SSSB07]. Transfer [BG07, LF00, PLPB07, RvBWR04, SC84, SL09, SCM^{+}09, WHL^{+}04, WHCO08, WS09b, XM09, VSD09]. Transferring [OZS08]. Transfers [YW97]. Transflective [BES00]. Transform [RM89, XS06, ZBA^{+}07]. Transformation [AEW90, BD04, HL02, LS89, NB94, PA06, TSP05]. Transformations [BCGB08, FAT07, LS08a, NR95, SHPS08, van90]. Transformed [SK86]. Transforms [MA91]. Transillumination [SKFNC97]. Transition [AW00]. Transitions [GW07]. Translating [Fuc97]. Translation [AHM09, HR92]. Translational [ND06]. translator [HR92]. Translucent [CLH^{+}08, EBSC99, LGB^{+}03, XGL^{+}07]. Transmission [MCO97, SG96a]. Transmittance [BR97]. Transparency [DWE02, MTKO02]. Transparent [BG02, DWL^{+}09, HLL07, LG95, Ger92]. Transport [DBG99, KSKAC02, KKK09, LB06, MSW04, PP05]. Transputer [AO89, RM89]. Traversal [HB00, Kuz94, LŽY04, PGSS07]. Traversing [LSŽ08, ŽCO97]. Treatment [CWKS00, RRRP08]. Tree [AE86, BPF^{+}03a, BPF^{+}03b, HDM98, JD98, PD04, PGSS07, WGG99, ZC95, Hol94, Rap92, SSK07]. Trees [AMT02, BHGS06, BBP08, DRBR09, GMW04, HKW09, JFSO06, OOI05, PG95, QNTN03a, QNTN03b, SS08, SG04, WBCG09b, dFSV03, CS92]. Trends [de 97]. Tri [MNP08]. Tri/Quad/Pent [MNP08]. Triangle [Ano07b, Bak91b, Gar09, Ise01, LAD02, MGS07, PSF04, SL03, LDB07]. Triangle-Quad [LAD02]. Triangles [AG01, Nar95, Ure00, dD85]. Triangular [LL02, PS95, PS96a, RDG01, Jon96]. Triangulated [DGGP05, MDP08, RR96]. Triangulating [RR94]. Triangulation [BBA08, CY89, GE04, GKS00, KB00, PB07, RWP88, SW04b, dCTAD09, CMPS93]. Triangulations [FS91]. Trillion [Ano07b]. Trimmed [AES94, GK03a, GK03b]. Trinity [Che06]. Trivariate [KZ08]. Tumor [RRRP08]. Tuning [ADS06, EBMT00]. Tunnel [ANF97a, ANF97b]. Tunnel-Free [ANF97a, ANF97b]. Turbulence [Sta97, SW92a]. Turbulent [DYN04]. Tutorials [BB07, tH84]. Two [BDS^{+}03, DGGP05, FR92, ML91, MK06, MPT98, MG95, Mil87, NHH97, SG97, SN84, SN86, TP89, TGS96, WTHS06, Day92, MMS09, QSW92, SDD^{+}92]. Two-dimensional [BDS^{+}03]. Two-Handed [SG97, TGS96]. Two-manifold [FR92]. Two-Manifolds [DGGP05]. Two-Parameter-Dependent [WTHS06]. Two-Pass [MPT98]. two-phase [MMS09]. Two-Step [MK06, MG95]. type [RT08b, RT08a]. Types [BMWM01, MG87]. Typographic [SR96].

U.K. [WG82]. U.M.I.S.T. [WG82]. UIMS [End84b, Str84]. UK [Ano85, Ano86, Ano87a, Ano88a, Ano89, Ano91a, Ano95p, Ano96k, Ano96p, Ano97z, Ano97-28, Duc89, MJ98, Mat86, Ano95o, Ano96j, Ano97q, Ano97r, Aus91, Kid84, Mum90, Pin85]. Ulicny [Ano05c]. Ultra [PSK09, WW87b]. Ultra-Large [PSK09]. UltraPaint [WW87a]. Unbiased [BAJ08]. Uncertainty [JNX^{+}08]. Underground [CS00]. Undersampled [LCM^{+}06]. Understand [AHM09]. Understanding [DF85, ME04, PKL88, ZK09]. Underwater [GSMA08, IDN02]. unescapable [Gue82]. Unfolding [TVD09]. Unification [PB90]. Unified [KH96, MS01, Sta06, ESP92]. Uniform [DRS09, KE97, VC04, VMG09, ES94]. Unions [RF96]. Unity [NOS09]. Universal [BG02]. University [Ano97z, Ano97-28, Bro90, Van80, WG82, tH83a, AEL^{+}82, Fah85, HHA82, Kil85]. Unlimited [Nas03]. Unorganized [MB08, VHB08]. Unpopping [GW07]. Unregistered [HLM97, NK99]. Unsolved [Mac94]. Unsteady [CKSW08, CPK09, JBTS08, JL00, Wei04]. Unstructured [AS96b, BTG95, CKM^{+}99, KRG03, SCM^{+}09, WÅ09]. Updates [GD01]. Updating [AGH^{+}93]. upon [RA94]. Upper [HKG06, Neb00]. Upper-Body [HKG06]. Urban [BCWR08, CDG^{+}07, CBSF07, FBT99, SDB97, WS99]. USA [Pri85]. Use [AEL^{+}82, Arn08, BBMR88, BB88, Gar09, GLW96, RB03a, RB03b, Ros83]. Use-Inspired [Arn08]. Useful [BWPP04]. User [BRL09, DR87, Duc82b, Duc85, Duc90a, End84c, GL94b, HM91, Hd89, IEGC08, JR08, KO88, Kin95, LFGG08, LD98, Mac85, MSM^{+}08, NM91, PF90, Sas92, SC04, STD09, WKG85, ZK08, ZK09, vD98, vJB85, BH93, EFGS96, GL94a, HP83, KSH92, ZK92, dBv93]. User-Computer [vD98]. User-Interface [DR87]. User-modifiable [ZK08]. user-tailored [EFGS96]. Users [DMKP07]. Using [AG01, AR95, BBW^{+}09, BAU05, Buc98, CS98a, CLH^{+}08, CZ09, CD08, CJC^{+}09, CZGF05, Dan96, DS05a, DGF98, DKN^{+}95, DMYN08, DKC00, DF90, EL01, FP04, FCP^{+}90, GBAL09, GGW98, GD96, GP87, GPD09, GJW08, GMW97, HB96, HS99, HCG08, IDN02, JW97, KOB^{+}08, KSS97, KPS95, KBKŠ09, yKL08, LKC08, LWP^{+}04, LSS98, MTCT84, MPS08, MG87, MKB^{+}08, MBBT00, Nie95, NIDN97, OIST91, Par86, PD04, PZB^{+}09, RW08, RSC01, RL09, RM89, RLF09, ST93, SHLS02, SS08, SSK^{+}05, SK86, SP01, SPS95, SF83, SKC01, SOM04, SE02a, SE02b, VLV^{+}04, VMA^{+}04, VPP^{+}04, WPG02, WDC^{+}08, WG09, WDGT01, WN09, XS06, YWB03, Ye08, ZY04, dFSV03, BCF^{+}05, BGK^{+}96, BES00, BK05b, BNC96, BG09, CJW^{+}06, CSN04, CCSG^{+}09, Cla92, CCP09, DAP08, Day90, Den03a, Den03b, DYN04, DKS01, DF93]. using [DZTS08, DRS08, EBV01, GHH01, GFW^{+}06, HE01, HB00, Hei95, HHK^{+}07, HH98, HWAG09, JFSO06, Jon96, KNP07, KPiAS01, KE97, KHH^{+}09, KHIK01, KB98, KGL^{+}98, LD08c, LLHY09, LCD09b, Mai00, MFT02, MZT09, Moh87, MRS08, MH00, NG92, NCKG00, NN93, NN94, NP00, NW91, OOI05, OAIS09, OPC96, Pai02, PG95, Par89, PCF05, PGGM09b, PH94, QSW92, RF96, RS08, RT08b, RT08a, Rob93, RTN03a, RTN03b, SPD07, SSK93, SCD05, SL01, STK02, SEA08, SNLH09, Tam82, TWJ06, TPBC09, TWC^{+}09, VT94, VMG09, WGS04, WVVV08, WO94, YKH^{+}09, YWY08, ZT96, hZCK98].

Vague [BAHS06]. Valence [AD01, Got03]. Valence-based [Got03]. Valence-Driven [AD01]. Valences [KPRW05]. Validation [BG85, SW04a]. Valuable [Bak91b]. Value [Got94]. Valued [SL89, SHPS08]. Values [CLM09]. Vandoni [Duc00]. Vanishing [KPiAS01]. Variable [EBV05, VG00]. Variate [FCH^{+}06, FH09]. Variation [ADS06, HTH96, NC99, SFS05]. Variational [Ara94, Gre94, HHB93, Kob05, LSJK09, LCWK07, YBS07, KHR02]. Varying [LLSS03a, LLSS03b, SWG08, WAH^{+}09, YMMS06, JR08, RK09b]. VAST [MRS06]. VAST2003 [CA05]. Vector [BS02, BN08a, EWHS08, FLW00, HE94, xHMC09, IEGC08, PA06, SS96b, SSB05, The02a, TRS03a, TRS03b, WTHS04, WTHS06, YM09, ZSP98, vP94, AM92]. Vector-based [BN08a, FLW00]. Vectorization [vKB94]. Vectors [Dan96, PA06, SP01, Thü01]. Vehicle [Arn08]. Verification [Ano98d, MJ98, UWP06, PB95]. Verifying [AA09]. Versatile [Dwy09, EBMT00, MH00]. Version [ten82b]. Vertex [BG02, KBS00]. Vertices [ADS06, LS08b]. Verve [Kni93]. Verve-voxel [Kni93]. Very [IP99, WH04]. vessel [WVV09]. VI [Ano06c]. Via [VCP09, AW00, GSGC08, HHS05, Kob05, KB89, LZ07, LSW09, McC96, RGM85, RK09a, Sad09, WS09b]. VIBE [BSG^{+}95]. Vibrations [KRFC09]. Video [BB00a, BCS96, BBPV03a, BBPV03b, CLHL08, CCC08, DMHS08, DJZ^{+}09, HVM^{+}08, HKMS08, KMN^{+}08, KKD09, LL00, LAA08, LDR09, OAIS09, PG08, RGW05, Sal96, SSK^{+}05, SEASM09, SLTM08, TSP05, WDC^{+}08, WWG07, WS09a, WMTG05, WLSG03, ZHM08]. Video-Based [BB00a, LL00]. Vienna [CG07a, Des06, HB91, Pur07]. View [Ano96p, CKB04, ED07a, ESV99, ESC00, ESK03a, ESK03b, Gre85, How87, LYP^{+}08, MTR08, MN87, Mor86, PNR89, SLAM08, tHS90, ZK92]. View-Dependent [CKB04, ED07a, ESV99, ESC00, MTR08]. Viewfinder [AGP08]. Viewing [AS95, DMKP07, Kra89, SS00, SJH08, SSG^{+}00, TGS96, YK92]. Viewpoint [FM04, HK09c, WLSG03]. Viewpoint-Dependent [FM04]. Views [ABC^{+}91, DGF98, HLM97, NK99, SNLH09, VBP^{+}09]. Viewspace [COFHZ98]. Virtual [Ai98, AVF04, BB09, Bar05, BSG^{+}95, BGAM04, Bro95, Bry96, CWKS00, CN05, CBSF07, CMT02, DJW^{+}06, EBSC99, FBT99, Fuc04, GHH01, Haa96, HMDO05, HK00, HJL07, JL98, KH95, KW05, KRFC09, Kun04, LD04, Lam09a, LL01, LD06, LGMT00, LLP00, LW99, MCO97, MC02, MT01, MG95, MSWK02, MF00, MKO^{+}00, MOT99, PLT^{+}97, PO02, PIWB98, RB03a, RB03b, SS00, SS96a, SG96a, SSA^{+}08, SDM99, SD00, SDB99, SNLW01, SG96b, SGC04, TYK^{+}09, TMO04, TLG99, VGSS04, WBP98, XTLP02, XLTP03a, XLTP03b, Zar06, MG96, TD00, VCDF95, Gob95, Gob96]. Virtually [KCB97]. Viscoelastic [SLS04]. Visibility [AHL^{+}06, AMS09, BNRSV01, CSI09, CATM09, CAM08a, COS95, COFHZ98, ED07b, GKD07, GBP07, HDS03a, HMB08, JEO00, MBGS01, NRJS03a, NRJS03b, ORDP96, SS00, SS07, SEA08, SG96b, WBP98, WWS01]. Visible [ASVNB00, BGK^{+}96, Mil84]. Vision [Cre88, Dut04, GVV05, MTKO02, NT95, PO02, Šár07, TO97, NW91]. Vissym04 [DHKS05]. Visual [BGB^{+}08a, BIWG08, BG08, BBC^{+}05, BCWR08, CTSO03a, CTSO03b, CG07b, Fra83, HKD^{+}08, HV08, IF09, JNM^{+}09, JZF^{+}09, LWZ^{+}09, LL09, MC02, MK08, McN01, MKO^{+}08, PCS94, PV08, SABG^{+}05, TTW90, Váz07, WMS^{+}08, Wil87b, ZYQ^{+}08, ZK09, dBD^{+}92, SW92a, SKSK07, VCDF95]. Visual-Quality [LWZ^{+}09]. Visualisation [DTA94, GJL^{+}09, HDM98, RSVP02, SAA09, Rei03]. Visualiser [How90]. Visualising [KH96]. Visualization [AM02a, AHM09, AGDJ09, Ano97s, Ano97d, Ano97f, Ano97-29, Ano98j, Baj03a, Baj03b, BY08, BMD^{+}08, BGB^{+}08a, BS02, BPBD08, BW01, BDG^{+}04, BAU05, BBN02, BG07, BG09, BD08, BBP09, CLE07, CZCE08, CPP08, CMS94, CGG^{+}03a, CGG^{+}03b, CDG^{+}07, CCP09, CG07b, CH09, CMH^{+}01, CKSW08, CPK09, DMKP07, Duc98, DGC^{+}98, Dwy09, Ert02, FQK08, FK09, FH09, GMW04, GWT^{+}08, GL94b, GSGC08, GBD09, Haa96, HS94, HV08, HV09, HTSFP09, IP99, IEGC08, IF09, JBB^{+}08, JBL^{+}06, JBTS08, JNX^{+}08, JNM^{+}09, JL00, KOB^{+}08, LHD^{+}04, Le 90, LTH08, LSS98, MSM^{+}08, MJC01, ME04, MMHL08, MDWK08, MKO^{+}08, NGB^{+}09, NSGP06, NS01, OHBKH09, PVtHR09, PKPH09, PD04, PH91, PV08, RSK06, RRRP08, RRS97b, RSS96, SP97, SVG^{+}08, SWS09, SSA^{+}08, SDB97, SAA09, SJH08, SKKS08, TSYK01, TA08, TK98]. Visualization [TSH01, VW08, VW90, VMG09, Wat96, WTHS04, WTHS06, Wei04, WVKR08, WVVV08, WVV09, WAH^{+}09, WKM^{+}09, XYL09, YCLE09, ZK08, vP94, AM92, BG93, Col93, FS92, GY93, GL94a, Gob96, GMDW09, HPT89, JR08, Kni93, NKP93, PW93, Rob93, RRS97a, Vol93, Ano98b, Ano98c, CDD09, Raf05, San06, SvZ95, Wei08]. Visualizations [KJC^{+}09]. Visualizing [BD08, CCP09, GBD09, GSMA08, HNR^{+}04, NGN^{+}01, TLC02, TA08, WVKR08]. Visyllable [KMT03a, KMT03b]. VLSI [ELM^{+}83, Mil84, NB94]. Volume [ARC05, AGG^{+}08, AGDJ08, Ano04h, Ano05f, Ano06f, Ano06g, BHGS06, BW01, BG07, BG09, BBP09, CCS95, CT00, CLE07, Coc83, COF95, CMH^{+}01, DHK08, DKC00, FBGP09, FK09, Frü94, GSMA08, HB96, HS04, HMK^{+}95, HWC^{+}05, HMB08, IP99, JBB^{+}08, JZJ08b, KZ08, LAM09b, LLA06, LLHY09, LG95, LGK97, LCD09b, LCWK07, MW06, MFS08, MMFE08, MBW08, ME04, NCKG00, NS01, OHBKH09, PBP96, RBG08, RSK06, RSTK08, RRRP08, RHC08, RMSD^{+}08, SHB07, SPH^{+}09, SCM^{+}09, TSM94, VPLL08, WGS04, WW99, WE97, WRS01, WS09b, YCLE09, YM06, hZCK98, ZCG08, BLS93, CS99a, Jon96, YK92, GBKG04, FMK04]. Volume-Surface [BHGS06]. Volumes [BK03c, BK03d, BS03a, BS03b, BBA08, BBP09, CCI08, PN97, RW08, WC02].

Volumetric [Baj03a, Baj03b, BNC96, DS02a, DN09, FAVM09, GBKG04, HDS03b, HDS03a, JZJ08b, KW05, LC99, ME04, PK08, RSTK08]. Voronoi [DZM08, OPC96, RL09, VC04, YLL^{+}09]. Vortex [PKPH09, RSVP02, SVG^{+}08, YKH^{+}09]. VOTS [GBKG04]. Voxel [BK03a, BK03b, BGAM04, BL86, Buc98, LŽY04, LSŽ08, RT08b, RT08a, SC95, WGS04, ŽCO97, Kni93]. Voxel-Shapes [RT08a]. Voxel-Traversing [ŽCO97]. voxelisation [Jon96]. VR [BSG^{+}95, BG07, CLE07, CG07b, DMKP07, EFGS96, LO95, SDB99]. VR-VIBE [BSG^{+}95]. VRML [cLCtLL98].

Walk [FMS01, Sbe97, SKCA01]. Walk-Through [FMS01]. Walking [VSC01]. walkthrough [AVF04]. Walkthroughs [GS09, GSA03a, GSA03b, SSS00, TL01, WBP98, WS99]. Warping [AGP08, AR95, BPBD08, FLW00, SL08, WGG99]. Washington [CP88]. Water [IDN03a, IDN03b, RSKN08]. Watermarking [hKLS00, OMT02]. Watermarks [BB00b]. Waters [PA01]. Watershed [FBW01]. Watershed-from-Markers [FBW01]. Wave [ATBG08, KMN^{+}05, ZCG08]. Wave-Based [ZCG08]. Waveform [Pic86b]. Wavefront [LSS98]. Wavelength [MRT08]. Wavelet [GMY97, IP99, LF00, LG95, PB94, SGCH94, TSP05, ZQQS08]. Wavelet-Based [IP99]. Wavelets [CAH00, DHS04, EBV01, KE97, MPS08, TWJ06, WM09]. Waving [WMTG05]. Wax [RMN05]. Way [AVF04, HKS09, SS91]. Way-Finder [AVF04]. Wayfinding [MDWK08]. Wear [DS02a]. weather [SSK93]. Weathering [XWT^{+}08]. Web [Cai95, CG02, DHH02, GSGC08, Wat96, YN00]. Web-Based [YN00]. Weighted [LA06, RLN06, SW99]. Well [NNB97, RHv95]. Well-Defined [RHv95]. WG2 [Ano84b, Bon85, Gal84, tH83b]. Wheelchair [OIST91]. Which [Sil97]. While [Thü01]. Wide [Wat96, Cai95]. Widgets [SSB08]. Will [ADJ^{+}01, Hec01]. win [AHT04]. Winchester [HGG^{+}84]. Wind [DRBR09, RR00, SF92]. Window [Ano91b, DSS90]. Windowing [RSK90]. windows [Day92]. Wipe [PSK09]. Wipe-Off [PSK09]. wireframe [THN93]. within [BGAM04, LW99, VCDF95]. Without [GVWD06, Wei04, FND92, JFS09, KySK08]. Wood [Buc98]. Woodblock [MKO^{+}00]. Woodblocks [MKO^{+}00]. Woodcut [MOT99]. Work [Ros97]. Workbench [TCH^{+}03a, TCH^{+}03b, TCH^{+}03b]. Working [Ano84b, End84b, Mac85, Str84]. workplace [PW93]. Workshop [Ano87b, Ano88b, Ano91b, Ano95b, Ano95a, Ano95d, Ano95c, Ano95q, Ano95-27, Ano96a, Ano96c, Ano97s, Ano97e, Ano97d, Ano97g, Ano97f, Ano97-29, Ano98f, Ano98d, Ano98b, Ano98e, Ano98c, Ano98h, Ano98g, Ano98i, Ano98j, Ano98-30, Ano06g, Arb90, AS98, ADS90, Bla88, Bou90, BH96, BP82, BP83b, BG85, Bro90, BJ94, CP88, Che06, Cla88, Cla89, Dau90, DS98, Duc82a, Duc86, Duc90a, Duc91, Duc98, End84c, HP95, Heg90, Hég91, Hew90, Jan91, Kon06, Kui91, Kun04, Kwi89, LMD04, Le 90, Lis90, Mac85, MJ98, Mar82, Mum86, NSGP06, Ott90, PB95, PH91, PS96b, Rei03, Ros82, SvZ95, Sch85, Sch98, TT95b, TvdP98, Vel91, tHS90, CG07a, Kje91a]. Workshops [Gob95, Gob96]. Workspaces [CG07b]. Workstation [Bak88, Wil84]. workstations [HP83]. World [IK01b, MG87, NK99, SDM99, VW08, YMMS06, Cai95, CP88, Wat96]. World-Wide [Cai95]. Worlds [Bro95, cLCtLL98]. WPPG97 [Ano98e, AS98]. Wrapping [KVLS99]. Wrinkling [Lov06, PZB^{+}09]. Writings [XJJ^{+}08]. www.eg.org [Che06]. WYSIWYG [SD94a].

X [Bow92, DSS90]. X3H3 [Arn84]. XGKS [RSK90].

Year [Dav07, Owe88, Owe89b, Owe94, Owe90b, Owe92a, Owe92b, Owe93, Owe95]. Years [Ano84a]. Yes [Mac94]. Young [Bot07, Kau07]. Yugoslavia [tH83a, Suz89, SW83]. Yugraph [Duc90b].

Z [Bow92, Sun92]. Z-buffer [Sun92]. Zach [Ano06c]. Zagreb [SW83, tH83a]. Zebedin [Ano06c]. Zippy [FQK08]. Zones [COS95]. Zoom [DF90]. Zooming [CG02]. Zurich [Kei04, Kun04].

References

Adamson:2006:APS

[AA06] A. Adamson and M. Alexa. Anisotropic point set surfaces. Computer Graphics Forum, 25(4):717–724, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aliaga:2009:CMG

[AA09] Daniel G. Aliaga and Mikhail J. Atallah. Captured models: Genuinity signatures: Designing signatures for verifying 3D object genuinity. Computer Graphics Forum, 28(2):437–446, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:1996:AGM

[AAB^{+}96] C. Andujar, D. Ayala, P. Brunet, R. Joan-Arinyo, and J. Sole. Automatic generation of multiresolution boundary representations. Computer Graphics Forum, 15(3):C87–C96, C463–C464, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ancuti:2009:CPI

[AAB09] Cosmin Ancuti, Codruta Orniana Ancuti, and Philippe Bekaert. Computational photography and image manipulation: Deblurring by matching. Computer Graphics Forum, 28(2):619–628, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aichholzer:2009:RRS

[AAK^{+}09] O. Aichholzer, F. Aurenhammer, B. Kornberger, S. Plantinga, G. Rote, A. Sturm, and G. Vegter. Reconstruction — regular sampling: Recovering structure from r-sampled objects. Computer Graphics Forum, 28(5):1349–1360, July 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aubert:1997:ADS

[AB97] Fabrice Aubert and Dominique Bechmann. Animation by deformation of space–time objects. Computer Graphics Forum, 16 (3):C57–C66, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=142&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Andujar:2007:MIB

[ABB^{+}07] C. Andújar, J. Boo, P. Brunet, M. Fairén, I. Navazo, P. Vázquez, and À. Vinacua. Material and image based rendering: Omnidirectional relief impostors. Computer Graphics Forum, 26(3): 553–560, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adelson:1991:SGS

[ABC^{+}91] Stephen J. Adelson, Jeffrey B. Bentley, In Seok Chong, Larry F. Hodges, and Joseph Winograd. Simultaneous generation of stereoscopic views. Computer Graphics Forum, 10(1):3–10, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2004:GPC

[ABC^{+}04] C. Andujar, P. Brunet, A. Chica, I. Navazo, J. Rossignac, and A. Vinacua. Geometry processing: Computing maximal tiles and application to impostor-based simplification. Computer Graphics Forum, 23(3):401–410, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

ACM:1980:EPI

[ACM80] ACM, editor. EUROGRAPHICS 79: proceedings, international conference and exhibition, Palazzo dei congressi, Bologna, Italy, 25–27 October 1979. ACM Press, New York, NY 10036, USA, 1980. LCCN T385 .E95 1979a.

Alliez:2001:VDC

[AD01] Pierre Alliez and Mathieu Desbrun. Valence-driven connectivity encoding for 3D meshes. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ansaldi:1985:EFR

[ADF85] S. Ansaldi, K. De Floriani, and B. Falcidieno. An edge-face relational scheme for boundary representations. Computer Graphics Forum, 4(4):319–332, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:2001:HDC

[ADJ^{+}01] D. B. Arnold, A. M. Day, V. Jennings, A. Courtenay, and D. A. Duce. Haptic device control — will it fit standardized input models? Computer Graphics Forum, 20(4):??, December 2001. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1990:REA

[ADS90] D. Arnold, D. Duce, and D. Sutcliffe. Report on the EU-ROGRAPHICS/ARGOSI workshop on graphics and networking. Computer Graphics Forum, 9(4):385–387, December 1990. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Augsdorfer:2006:GTS

[ADS06] U. H. Augsdörfer, N. A. Dodgson, and M. A. Sabin. Geometry: Tuning subdivision by minimising Gaussian curvature variation near extraordinary vertices. Computer Graphics Forum, 25(3): 263–272, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abi-Ezzi:1986:PTH

[AE86] S. S. Abi-Ezzi. The priority tree, a HL/HSR approach for PHIGS. Computer Graphics Forum, 5(4):283–289, December 1986. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arad:1997:ITM

[AE97] Nur Arad and Gershon Elber. Isometric texture mapping for free-form surfaces. Computer Graphics Forum, 16(5): 247–256, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=192&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Arnold:1982:EUC

[AEL^{+}82] D. B. Arnold, D. J. Earl, D. Lord, G. P. McKeown, W. J. Milne, and A. Norris-Sherborn. Experiences in the use of colour raster graphics at the University of East Anglia. Computer Graphics Forum, 1(4):185–190, December 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abi-Ezzi:1994:FDT

[AES94] S. S. Abi-Ezzi and S. Subramaniam. Fast dynamic tessellation of trimmed NURBS surfaces. Computer Graphics Forum, 13(3): C/107–C/126, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abi-Ezzi:1990:FHT

[AEW90] Salim S. Abi-Ezzi and Michael J. Wozny. Factoring a homogeneous transformation for a more efficient graphics pipeline. Computer Graphics Forum, 9(3):245–255, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akkouche:2001:AIS

[AG01] Samir Akkouche and Eric Galin. Adaptive implicit surface polygonization using marching triangles. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Atlan:2006:IME

[AG06] Samuel Atlan and Michael Garland. Interactive multiresolution editing and display of large terrains. Computer Graphics Forum, 25(2):211–223, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anderson:2008:FFE

[AGDJ08] J. C. Anderson, C. Garth, M. A. Duchaineau, and K. I. Joy. Flow feature extraction: Discrete multi-material interface reconstruction for volume fraction data. Computer Graphics Forum, 27 (3):1015–1022, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anderson:2009:FIV

[AGDJ09] John C. Anderson, Luke J. Gosink, Mark A. Duchaineau, and Kenneth I. Joy. Functions: Interactive visualization of function fields by range-space segmentation. Computer Graphics Forum, 28(3):727–734, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agus:2008:OVR

[AGG^{+}08] Marco Agus, Enrico Gobbetti, José Antonio Iglesias Guitiàn, Fabio Marton, and Giovanni Pintore. Occlusion and volume rendering: GPU accelerated direct volume rendering on an interactive light field display. Computer Graphics Forum, 27(2):231–240, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abellanas:1993:UP

[AGH^{+}93] M. Abellanas, J. Garcia, G. Hernandez, F. Hurtado, O. Serra, and J. Urrutia. Updating polygonizations. Computer Graphics Forum, 12(3):C143–C152, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Artusi:2006:STL

[AGM^{+}06] A. Artusi, C. Gatta, D. Marini, W. Purgathofer, and A. Rizzi. Speed-up technique for a local automatic colour equalization model. Computer Graphics Forum, 25(1):5–14, March 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adams:2008:IWV

[AGP08] Andrew Adams, Natasha Gelfand, and Kari Pulli. Imaging and warping: Viewfinder alignment. Computer Graphics Forum, 27 (2):597–606, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1989:IPG

[AH89] D. B. Arnold and M. R. Hinds. On implementing parallel GKS. Computer Graphics Forum, 8(1):13–19, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akeo:1994:CGS

[AHKS94] M. Akeo, H. Hashimoto, T. Kobayashi, and T. Shibusawa. Computer graphics system for reproducing three-dimensional shape from idea sketch. Computer Graphics Forum, 13(3):C/477–C/488, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Atty:2006:SSM

[AHL^{+}06] Lionel Atty, Nicolas Holzschuch, Marc Lapierre, Jean-Marc Hasenfratz, Charles Hansen, and François X. Sillion. Soft shadow maps: Efficient sampling of light source visibility. Computer Graphics Forum, 25(4):725–741, December 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Albrecht:2009:TVC

[AHM09] Joshua Albrecht, Rebecca Hwa, and G. Elisabeta Marai. Text visualization: The Chinese room: Visualization and interaction to understand and correct ambiguous machine translation. Computer Graphics Forum, 28(3):1047–1054, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1984:PCM

[AHR84] D. Arnold, G. Hall, and G. Reynolds. Proposals for configurable models of graphics systems. Computer Graphics Forum, 3(3): 201–208, September 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1985:GPP

[AHR85] David B. Arnold, Gillian Hall, and Graham J. Reynolds. GKS programming in a PHIGS environment. Computer Graphics Forum, 4(4):349–358, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arbab:1993:OMM

[AHR93] F. Arbab, I. Herman, and G. J. Reynolds. An object model for multimedia programming. Computer Graphics Forum, 12(3): C101–C113, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arvo:2004:RTI

[AHT04] Jukka Arvo, Mika Hirvikorpi, and Joonas Tyystjärvi. Real-time illumination and shadows: Approximate soft shadows win an image-space flood-fill algorithm. Computer Graphics Forum, 23 (3):271–279, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ai:1998:MDS

[Ai98] Z. Ai. Molecular dynamics simulation in virtual environments. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=273&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Aish:1985:IAC

[Ais85] Robert Aish. 3D icons and architectural CAD. Computer Graphics Forum, 4(3):177–186, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ahlers:1995:DAR

[AKB^{+}95] K. H. Ahlers, A. Kramer, D. E. Breen, P.-Y. Chevalier, C. Crampton, E. Rose, M. Tuceryan, R. T. Whitaker, and D. Greer. Distributed augmented reality for collaborative design applications. Computer Graphics Forum, 14(3):C/3–C/14, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adams:2005:PER

[AKP^{+}05] Bart Adams, Richard Keiser, Mark Pauly, Leonidas J. Guibas, Markus Gross, and Philip Dutré. Points: Efficient raytracing of deforming point-sampled surfaces. Computer Graphics Forum, 24 (3):677–684, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abbasinejad:2009:RIS

[AKSA09] Fatemeh Abbasinejad, Yong Joo Kil, Andrei Sharf, and Nina Amenta. Reconstruction — irregular sampling: Rotating scans for systematic error removal. Computer Graphics Forum, 28(5):1319– 1326, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alexa:2002:RAM

[Ale02] Marc Alexa. Recent advances in mesh morphing. Computer Graphics Forum, 21(2):173–196, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ascani:1992:PPI

[AM92] F. Ascani and L. Moltedo. PHIFI: a PHIGS interactive system for the visualization of scalar and vector fields. Computer Graphics Forum, 11(2):139–149, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arques:1995:NRA

[AM95] D. Arques and S. Michelin. A new radiosity approach for regular objects: Application to ruled surfaces. Computer Graphics Forum, 14(3):C/299–C/310, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alexa:2000:RAP

[AM00] Marc Alexa and Wolfgang Müller. Representing animations by principal components. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=433&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Adabala:2002:TRV

[AM02a] Neeharika Adabala and Swami Manohar. Techniques for realistic visualization of fluids: a survey. Computer Graphics Forum, 21 (1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ayasse:2002:SDD

[AM02b] Jörg Ayasse and Heinrich Müller. Sculpturing on discrete displacement fields. Computer Graphics Forum, 21(3):431–440, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aydin:2009:HPG

[AMS09] Tunç Ozan Aydin, Karol Myszkowski, and Hans-Peter Seidel. Human perception and Green graphics: Predicting display visibility under dynamically changing lighting conditions. Computer Graphics Forum, 28(2):173–182, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Attene:2008:THC

[AMSF08] Marco Attene, Michela Mortara, Michela Spagnuolo, and Bianca Falcidieno. Tetrahedra: Hierarchical convex approximation of 3D shapes for fast region selection. Computer Graphics Forum, 27 (5):1323–1332, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ar:2002:DSO

[AMT02] Sigal Ar, Gil Montag, and Ayellet Tal. Deferred, self-organizing BSP trees. Computer Graphics Forum, 21(3):269–278, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andreev:1989:ACA

[And89] Rumen D. Andreev. Algorithm for clipping arbitrary polygons. Computer Graphics Forum, 8(3):183–191, September 1989. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andres:1997:TFSa

[ANF97a] Eric Andrès, Philippe Nehlig, and Jean Françon. Tunnel-free supercover 3D polygons and polyhedra. Computer Graphics Forum, 16(3):??, September 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=137&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Andres:1997:TFSb

[ANF97b] Eric Andres, Philippe Nehlig, and Jean Francon. Tunnel-free supercover 3D polygons and polyhedra. Computer Graphics Forum, 16(3):C3–C13, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1982:CGE

[Ano82] Anonymous. Computer graphics in Europe. Computer Graphics Forum, 1(4):173–174, December 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1984:EFF

[Ano84a] Anonymous. EUROGRAPHICS — the first five years status and activities of the European Association for Computer Graphics. Computer Graphics Forum, 3(3):239–242, September 1984. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1984:WIW

[Ano84b] Anonymous. WG2 — The ISO working group on computer graphics. Computer Graphics Forum, 3(3):227–231, September 1984.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1985:EUC

[Ano85] Anonymous. EUROGRAPHICS (UK) Conference. Computer Graphics Forum, 4(3):??, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1986:EUC

[Ano86] Anonymous. EUROGRAPHICS (UK) Conference. Computer Graphics Forum, 5(2):151–157 (or 131–157??), June 1986. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1987:EUA

[Ano87a] Anonymous. Eurographics (UK) Annual Conference. Computer Graphics Forum, 6(2):111–155, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1987:GRW

[Ano87b] Anonymous. GKS review workshop. Computer Graphics Forum, 6(4):367–369, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1988:AEU

[Ano88a] Anonymous. 6th Annual EUROGRAPHICS (UK) Conference. Computer Graphics Forum, 7(2):91–134, June 1988. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1988:EWC

[Ano88b] Anonymous. Eurographics Workshop on Conformance Testing of the Computer Graphics Interface (CGI). Computer Graphics Forum, 7(3):231–232, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1989:EUC

[Ano89] Anonymous, editor. 7th EUROGRAPHICS UK Conference, volume 8(2) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1991:AEU

[Ano91a] Anonymous, editor. 9th Annual EUROGRAPHICS (UK) Conference, volume 10(2) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1991:AWD

[Ano91b] Anonymous. ARGOST workshop on distributed window systems. Computer Graphics Forum, 10(3):274, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1991:E

[Ano91c] Anonymous. Eurographics ’92. Computer Graphics Forum, 10 (3):273, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1992:EAC

[Ano92] Anonymous, editor. European Association for Computer Graphics 13th Annual Conference. EUROGRAPHICS 92, volume 11(3) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1993:EAC

[Ano93] Anonymous, editor. European Association for Computer Graphics 14th Annual Conference and Exhibition. EUROGRAPHICS ’93, volume 12(3) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1994:ACE

[Ano94] Anonymous, editor. 15th Annual Conference and Exhibition. EU-ROGRAPHICS ’94, volume 13(3) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EWC

[Ano95a] Anonymous. 10th EUROGRAPHICS Workshop on Computer Graphics Hardware. Computer Graphics Forum, 14(5):333–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EIW

[Ano95b] Anonymous. 1st EUROGRAPHICS International Workshop on Implicit Surfaces. Computer Graphics Forum, 14(5):329–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EWRb

[Ano95c] Anonymous. 6th EUROGRAPHICS Workshop on Rendering. Computer Graphics Forum, 14(5):331–??, December 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EWRa

[Ano95d] Anonymous. 7th EUROGRAPHICS Workshop on Rendering. Computer Graphics Forum, 14(5):317–??, December 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:AR

[Ano95e] Anonymous. Auditor’s report. Computer Graphics Forum, 14 (5):344–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:CCa

[Ano95f] Anonymous. Cover competition. Computer Graphics Forum, 14(2):171–??, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:CCb

[Ano95g] Anonymous. Cover competition. Computer Graphics Forum, 14 (4):251–??, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:CCc

[Ano95h] Anonymous. Cover competition. Computer Graphics Forum, 14 (5):316–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:E

[Ano95i] Anonymous, editor. EUROGRAPHICS ’95, volume 14(3) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EM

[Ano95j] Anonymous. EUROGRAPHICS ’95 Maastricht. Computer Graphics Forum, 14(5):323–??, December 1995. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EMC

[Ano95k] Anonymous. EUROGRAPHICS ’95 Maastricht, call for participation. Computer Graphics Forum, 14(2):164–??, June 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:ECPa

[Ano95l] Anonymous. EUROGRAPHICS ’96, call for participation. Computer Graphics Forum, 14(4):247–??, October 1995. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:ECPb

[Ano95m] Anonymous. EUROGRAPHICS ’96, call for participation. Computer Graphics Forum, 14(5):313–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:ECO

[Ano95n] Anonymous. EUROGRAPHICS ’98 — call for offers to host EG ’98. Computer Graphics Forum, 14(5):318–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EU

[Ano95o] Anonymous. EUROGRAPHICS UK ’95. Computer Graphics Forum, 14(5):327–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:EUC

[Ano95p] Anonymous. EUROGRAPHICS UK Chapter 14th AGM. Computer Graphics Forum, 14(5):345–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:FEW

[Ano95q] Anonymous. Fifth EUROGRAPHICS workshop on programming paradigms in graphics. Computer Graphics Forum, 14(2):167–??, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:GAM

[Ano95r] Anonymous. General assembly minutes. Computer Graphics Forum, 14(5):339–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:GEA

[Ano95s] Anonymous. Gunter Enderle award. Computer Graphics Forum, 14(5):325–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:JNa

[Ano95t] Anonymous. Join now! Computer Graphics Forum, 14(2):176–??, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:JNb

[Ano95u] Anonymous. Join now! Computer Graphics Forum, 14(4):257– ??, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:JNc

[Ano95v] Anonymous. Join now! Computer Graphics Forum, 14(5):347– ??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:NSa

[Ano95w] Anonymous. Network service. Computer Graphics Forum, 14(2): 177–??, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:NSc

[Ano95x] Anonymous. Network service. Computer Graphics Forum, 14(5): 348–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:NSb

[Ano95y] Anonymous. Network services. Computer Graphics Forum, 14 (4):258–??, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:SA

[Ano95z] Anonymous. SIGGRAPH ’95, Los Angeles. Computer Graphics Forum, 14(5):335–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1995:SEW

[Ano95-27] Anonymous. Sixth EUROGRAPHICS workshop on animation and simulation. Computer Graphics Forum, 14(2):169–??, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:EWG

[Ano96a] Anonymous. 11th EG Workshop on Graphics Hardware. Computer Graphics Forum, 15(1):79–??, March 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:AR

[Ano96b] Anonymous. 15th AGM report. Computer Graphics Forum, 15 (4):281–??, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:EWP

[Ano96c] Anonymous. 5th EG Workshop on Programming Paradigms in Graphics. Computer Graphics Forum, 15(1):89–??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:CS

[Ano96d] Anonymous. Colour section. Computer Graphics Forum, 15(3): 453–??, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:CCa

[Ano96e] Anonymous. Cover competition. Computer Graphics Forum, 15(1):87–??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:CCb

[Ano96f] Anonymous. Cover competition. Computer Graphics Forum, 15(2):183–??, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:ECPa

[Ano96g] Anonymous. Eurographics ’96, call for participation. Computer Graphics Forum, 15(1):75–??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:ECPb

[Ano96h] Anonymous. Eurographics ’96, call for participation. Computer Graphics Forum, 15(2):171–??, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:E

[Ano96i] Anonymous. Eurographics ’97. Computer Graphics Forum, 15 (4):277–??, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:EUCa

[Ano96j] Anonymous. Eurographics UK ’97 conference. Computer Graphics Forum, 15(2):175–??, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:EUCb

[Ano96k] Anonymous. Eurographics UK ’97 Conference. Computer Graphics Forum, 15(4):278–??, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:EAC

[Ano96l] Anonymous. European Association for Computer Graphics 17th Annual Conference and Exhibition. EUROGRAPHICS ’96. Computer Graphics Forum, 15(3):??, September 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:JNa

[Ano96m] Anonymous. Join now! Computer Graphics Forum, 15(1):93– ??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:JNb

[Ano96n] Anonymous. Join now! Computer Graphics Forum, 15(2):185–??, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:JNc

[Ano96o] Anonymous. Join now! Computer Graphics Forum, 15(4):283– ??, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:MVE

[Ano96p] Anonymous. A Mancunian view of the 14th EG UK Conference. Computer Graphics Forum, 15(4):279–??, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:NSb

[Ano96q] Anonymous. Network service. Computer Graphics Forum, 15(2): 187, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:NSa

[Ano96r] Anonymous. Network services. Computer Graphics Forum, 15(1): 94–??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:PAC

[Ano96s] Anonymous, editor. Proceedings of the 1996 17th Annual Conference and Exhibition of the European Association for Computer Graphics, EUROGRAPHICS’96, volume 15(3) of Computer Graphics Forum. North-Holland, Amsterdam, The Netherlands, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1996:RP

[Ano96t] Anonymous. Refereed papers. Computer Graphics Forum, 15(1): ??, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:CCa

[Ano97a] Anonymous. 1998 cover competition. Computer Graphics Forum, 16(1):86, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:CCb

[Ano97b] Anonymous. 1998 cover competition. Computer Graphics Forum, 16(2):154–??, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:CCc

[Ano97c] Anonymous. 1998 cover competition. Computer Graphics Forum, 16(2):??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=133&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:EWG

[Ano97d] Anonymous. 4th EG workshop on graphics and visualization education. Computer Graphics Forum, 16(4):237–??, ???? 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:EWD

[Ano97e] Anonymous. 5th EG Workshop on DSV-IS: call for participation. Computer Graphics Forum, 16(5):??, December 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=202&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Anonymous:1997:EWV

[Ano97f] Anonymous. 8th EG workshop on visualization in scientific computing. Computer Graphics Forum, 16(4):235–??, ???? 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:EWR

[Ano97g] Anonymous. 9th EG Workshop on Rendering: call for participation. Computer Graphics Forum, 16(5):??, December 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=201&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Anonymous:1997:AR

[Ano97h] Anonymous. Auditor’s report. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=208&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:BRa

[Ano97i] Anonymous. Book reviews. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=118&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Anonymous:1997:BRb

[Ano97j] Anonymous. Book reviews. Computer Graphics Forum, 16(2):??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=131&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:BRc

[Ano97k] Anonymous. Book reviews. Computer Graphics Forum, 16 (5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=197&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:CEb

[Ano97l] Anonymous. Calendar of events. Computer Graphics Forum, 16(4):??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=183&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Anonymous:1997:CEc

[Ano97m] Anonymous. Calendar of events. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=198&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:CEa

[Ano97n] Anonymous. Calender of events. Computer Graphics Forum, 16(2):??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=132&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:EEA

[Ano97o] Anonymous. EG Eurographics ’97: Annual Conference of the European Association for Computer Graphics. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=120&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Anonymous:1997:ENS

[Ano97p] Anonymous. EG network service. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=123&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Anonymous:1997:EUA

[Ano97q] Anonymous. EG UK 16th annual general meeting. Computer Graphics Forum, 16(4):241–??, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:EUCa

[Ano97r] Anonymous. EG UK ’97 conference. Computer Graphics Forum, 16(4):239–??, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:EEW

[Ano97s] Anonymous. Eighth Eurographics Workshop on Visualization in Scientific Computing: Boulogne-sur-Mer, France, April 28–30, 1997. Computer Graphics Forum, 16(4):??, October 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=185&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=16.

Anonymous:1997:ECCb

[Ano97t] Anonymous. Eurographics 1998 cover competition. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=205&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Anonymous:1997:Ea

[Ano97u] Anonymous. Eurographics ’97. Computer Graphics Forum, 16 (1):85, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:EAa

[Ano97v] Anonymous. Eurographics ’97 Awards. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=204&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:EBH

[Ano97w] Anonymous. Eurographics ’97: Budapest, Hungary, September 4–8, 1997. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=203&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:Ec

[Ano97x] Anonymous. Eurographics ’98. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=199&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:ECCa

[Ano97y] Anonymous. Eurographics ’98: call for contributions. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=200&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Anonymous:1997:EUCb

[Ano97z] Anonymous. Eurographics UK Chapter 16th Annual General Meeting: University of East Anglia, Norwich, March 25, 1997. Computer Graphics Forum, 16(4):??, October 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=188&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=16.

Anonymous:1997:Eb

[Ano97-27] Anonymous. Eurographics’98. Computer Graphics Forum, 16 (4):233–??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=184&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Anonymous:1997:FAC

[Ano97-28] Anonymous. Fifteenth Annual Conference Eurographics UK ’97: University of East Anglia, Norwich, March 24–26, 1997. Computer Graphics Forum, 16(4):??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=187&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=16.

Anonymous:1997:FEW

[Ano97-29] Anonymous. Fourth Eurographics Workshop on Graphics and Visualization Education (GVE): Poitiers, France, August 24–25, 1996. Computer Graphics Forum, 16(4):??, October 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=186&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=16.

Anonymous:1997:GAM

[Ano97-30] Anonymous. General assembly minutes. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=207&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:ICG

[Ano97-31] Anonymous. Image credit to G. Elber. Computer Graphics Forum, 16(2):141–??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=130&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:ISA

[Ano97-32] Anonymous. Impressions of SIGGRAPH ’97: Los Angeles, August 3rd–8th 1997. Computer Graphics Forum, 16 (5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=206&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:JNa

[Ano97-33] Anonymous. Join now! Computer Graphics Forum, 16

(1):87, March 1997.  CODEN CGFODY.  ISSN 0167-

7055 (print), 1467-8659 (electronic). URL aid=122&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Anonymous:1997:JNb

[Ano97-34] Anonymous. Join now! Computer Graphics Forum, 16 (2):155–??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=134&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:JNc

[Ano97-35] Anonymous. Join now! Computer Graphics Forum, 16(4):243–??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-

8659 (electronic). URL http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&iid=4&src=ard&vid=16.

Anonymous:1997:JNd

[Ano97-36] Anonymous. Join now! Computer Graphics Forum, 16 (5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=209&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:NSa

[Ano97-37] Anonymous. Network service. Computer Graphics Forum, 16 (1):88, ???? 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1997:NSb

[Ano97-38] Anonymous. Network service. Computer Graphics Forum, 16(2):156–??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=135&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Anonymous:1997:NSd

[Ano97-39] Anonymous. Network service. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=210&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Anonymous:1997:NSc

[Ano97-40] Anonymous. Network services. Computer Graphics Forum, 16 (4):244–??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=190&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Anonymous:1997:PCC

[Ano97-41] Anonymous. Prize competition: 1998 cover photograph. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=121&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=16.

Anonymous:1998:AIb

[Ano98a] Anonymous. 1998 annual index. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=295&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:EWPa

[Ano98b] Anonymous. 2nd EG Workshop on Parallel Graphics and Visualization — call for contributions. Computer Graphics Forum, 17(1):97–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:EWPc

[Ano98c] Anonymous. 2nd Eurographics Workshop on Parallel Graphics & Visualization: Call for contributions. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=223&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:EWD

[Ano98d] Anonymous. 5th Eurographics Workshop on Design, Specification and Verification of Interactive Systems, DSV-IS ’98: Modelling and Design. Computer Graphics Forum, 17(1):99– ??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=224&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:EWPb

[Ano98e] Anonymous. 6th EG Workshop on Programming Paradigms in Graphics (WPPG97). Computer Graphics Forum, 17(1):107–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:EWC

[Ano98f] Anonymous. 8th EG Workshop on Computer Animation and Simulation. Computer Graphics Forum, 17(1):103–??, ???? 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:EWRb

[Ano98g] Anonymous. 8th EG Workshop on Rendering. Computer Graphics Forum, 17(1):101–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:EWRa

[Ano98h] Anonymous. 9th EG Workshop on Rendering — call for participation. Computer Graphics Forum, 17(1):93–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:EWRc

[Ano98i]  Anonymous.  9th Eurographics Workshop on Rendering:

Call for participation. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=221&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:EWV

[Ano98j] Anonymous. 9th Eurographics Workshop on Visualization in Scientific Computing. Computer Graphics Forum, 17 (1):95–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=222&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:Aa

[Ano98k] Anonymous. Announcements. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=292&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:Ab

[Ano98l] Anonymous. Announcements. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=296&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:AIa

[Ano98m] Anonymous. Author index. Computer Graphics Forum, 17 (3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=244&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Anonymous:1998:BRa

[Ano98n] Anonymous. Book reviews. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=217&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:BRb

[Ano98o] Anonymous. Book reviews. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=237&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Anonymous:1998:BRc

[Ano98p] Anonymous. Book reviews. Computer Graphics Forum, 17 (4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=290&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:CEa

[Ano98q] Anonymous. Calendar of events. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=218&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:CEb

[Ano98r] Anonymous. Calendar of events. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=291&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:CEc

[Ano98s] Anonymous. Calendar of events. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=304&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:CI

[Ano98t] Anonymous. Colour illustrations. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=283&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Anonymous:1998:EA

[Ano98u] Anonymous. Eurographics ’98 announcement. Computer Graphics Forum, 17(1):89–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1998:ECS

[Ano98v] Anonymous. Eurographics ’98: Call for short presentations. Computer Graphics Forum, 17(1):91–??, March 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=220&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=17.

Anonymous:1998:EDN

[Ano98w] Anonymous. Eurographics ’98: Discovering new realities. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=219&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=17.

Anonymous:1998:ERb

[Ano98x] Anonymous. Eurographics reports. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=294&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:ERa

[Ano98y] Anonymous. Event reports. Computer Graphics Forum, 17 (4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=293&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Anonymous:1998:JNa

[Ano98z]  Anonymous. Join now!  Computer Graphics Forum, 17

(1):109–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=229&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:JNb

[Ano98-27] Anonymous. Join now! Computer Graphics Forum, 17(2): ??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=242&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=2&amp; src=ard&amp; vid=17.

Anonymous:1998:NSa

[Ano98-28] Anonymous. Network service. Computer Graphics Forum, 17(1):110–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=230&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Anonymous:1998:NSb

[Ano98-29] Anonymous. Network service. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=241&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Anonymous:1998:SEW

[Ano98-30] Anonymous. Siggraph/EG Workshop on Graphics Hardware. Computer Graphics Forum, 17(1):105–??, ???? 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:1999:Aa

[Ano99a] Anonymous. Announcements. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=305&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Anonymous:1999:Ab

[Ano99b] Anonymous. Announcements. Computer Graphics Forum, 18(4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=380&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Anonymous:1999:BRa

[Ano99c] Anonymous. Book reviews. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=369&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Anonymous:1999:BRb

[Ano99d] Anonymous. Book reviews. Computer Graphics Forum, 18 (4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=379&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Anonymous:1999:CE

[Ano99e] Anonymous. Calendar of events. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=370&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Anonymous:1999:Eb

[Ano99f] Anonymous. Erratum. Computer Graphics Forum, 18(2): ??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=368&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=2&amp; src=ard&amp; vid=18.

Anonymous:1999:Ea

[Ano99g] Anonymous. Eurographics. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=307&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Anonymous:1999:Ec

[Ano99h] Anonymous. Eurographics. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=372&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Anonymous:1999:Ed

[Ano99i] Anonymous. Eurographics. Computer Graphics Forum, 18 (4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=382&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Anonymous:1999:ERa

[Ano99j] Anonymous. Event reports. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=306&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Anonymous:1999:ERb

[Ano99k] Anonymous. Event reports. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=371&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Anonymous:1999:ERc

[Ano99l] Anonymous. Event reports. Computer Graphics Forum, 18 (4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=381&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Anonymous:1999:PAC

[Ano99m] Anonymous. A practical analysis of clustering strategies for hierarchical radiosity (supplementary material). Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=343&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Anonymous:2000:A

[Ano00a] Anonymous. Announcements. Computer Graphics Forum, 19(4): ??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2000:BRa

[Ano00b] Anonymous. Book reviews. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=389&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Anonymous:2000:BRb

[Ano00c] Anonymous. Book reviews. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2000:CS

[Ano00d] Anonymous. Colour section. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=444&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Anonymous:2000:Ea

[Ano00e] Anonymous. Eurographics. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=391&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Anonymous:2000:Eb

[Ano00f] Anonymous. Eurographics. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2000:Ec

[Ano00g] Anonymous. Eurographics. Computer Graphics Forum, 19(4): ??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2000:ERa

[Ano00h] Anonymous. Event reports. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=390&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Anonymous:2000:ERb

[Ano00i] Anonymous. Event reports. Computer Graphics Forum, 19(4): ??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2001:Ea

[Ano01a] Anonymous. Editorial. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2001:Ec

[Ano01b] Anonymous. Editorial. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2001:Eb

[Ano01c] Anonymous. Eurographics. Computer Graphics Forum, 20(1): ??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2001:Ed

[Ano01d] Anonymous. Eurographics. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2001:Ee

[Ano01e] Anonymous. Eurographics. Computer Graphics Forum, 20(4): ??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:BR

[Ano02a] Anonymous. Book reviews. Computer Graphics Forum, 21(2): 197–201, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:CS

[Ano02b] Anonymous. Color section. Computer Graphics Forum, 21(3): 615–656, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:Ea

[Ano02c] Anonymous. Eurographics. Computer Graphics Forum, 21(1): ??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:Eb

[Ano02d] Anonymous. Eurographics. Computer Graphics Forum, 21(2): 207–208, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:Ec

[Ano02e] Anonymous. Eurographics. Computer Graphics Forum, 21(4): 816–832, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:ERa

[Ano02f] Anonymous. Event reports. Computer Graphics Forum, 21(1): ??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:ERb

[Ano02g] Anonymous. Event reports. Computer Graphics Forum, 21(1): ??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:ERc

[Ano02h] Anonymous. Event reports. Computer Graphics Forum, 21(2): 203–206, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2002:ERd

[Ano02i] Anonymous. Event reports. Computer Graphics Forum, 21(4): 803–815, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:A

[Ano03a] Anonymous. Announcement. Computer Graphics Forum, 22(2): 217, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:CGF

[Ano03b] Anonymous. Computer Graphics Forum. Computer Graphics Forum, 22(2):222, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:CGP

[Ano03c] Anonymous. Computer graphics professionals. Computer Graphics Forum, 22(2):221, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:E

[Ano03d] Anonymous. Editorial. Computer Graphics Forum, 22(2):106, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:ENS

[Ano03e] Anonymous. EG network service. Computer Graphics Forum, 22 (1):104, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2003:JN

[Ano03f] Anonymous. Join now! Computer Graphics Forum, 22(1):103, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:E

[Ano04a] Anonymous. Editorial. Computer Graphics Forum, 23(1):i, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:EEG

[Ano04b] Anonymous. Eurographics: 25th EUROGRAPHICS General Assembly: Grenoble, France, 4 September 2004. Computer Graphics Forum, 23(4):848–852, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:ECC

[Ano04c] Anonymous. Eurographics: CGForum 2004 Cover Image: “Blue Cones” by Csaba Kelemen. Computer Graphics Forum, 23(1): 124–125, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:EJN

[Ano04d] Anonymous. Eurographics: Join now! Computer Graphics Forum, 23(1):126, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:ENS

[Ano04e] Anonymous. Eurographics: Network service. Computer Graphics Forum, 23(1):127, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:ENE

[Ano04f] Anonymous. Eurographics: New Eurographics Fellows. Computer Graphics Forum, 23(4):846–847, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:ERS

[Ano04g] Anonymous. Eurographics: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 23(4):853, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2004:IAI

[Ano04h] Anonymous. Index: Author index volume 23 (2004). Computer Graphics Forum, 23(4):854–858, December 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:EEE

[Ano05a] Anonymous. Editorial: EG editorial. Computer Graphics Forum, 24(1):i, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:EEG

[Ano05b] Anonymous. Eurographics: 26th EUROGRAPHICS General Assembly. Computer Graphics Forum, 24(4):851–857, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:ECC

[Ano05c] Anonymous. Eurographics: CGForum 2005 Cover Image: “CrowdBrush” by Branislav Ulicny, Pablo De Heras Ciechomski and Mireille Clavien. Computer Graphics Forum, 24(1):134–135, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:ENE

[Ano05d] Anonymous. Eurographics: New EUROGRAPHICS Fellows. Computer Graphics Forum, 24(4):858–859, December 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:ERS

[Ano05e] Anonymous. Eurographics: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 24(4):860, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:IAI

[Ano05f] Anonymous. Index: Author index volume 24 (2005). Computer Graphics Forum, 24(4):ii–vi, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2005:REA

[Ano05g] Anonymous. Reports: Eurographics Awards Programme. Computer Graphics Forum, 24(1):127–129, March 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:E

[Ano06a] Anonymous. Errata. Computer Graphics Forum, 25(2):261, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:EEG

[Ano06b] Anonymous. Eurographics: 27th EUROGRAPHICS General Assembly. Computer Graphics Forum, 25(4):844–849, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:ECC

[Ano06c] Anonymous. Eurographics: CGForum 2006 Cover Image “Charles VI, Holy Roman Emperor” by Mario Sorman, Christopher Zach, Lukas Zebedin and Konrad Karner. Computer Graphics Forum, 25(1):147–148, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:ENE

[Ano06d] Anonymous. Eurographics: New EUROGRAPHICS Fellows. Computer Graphics Forum, 25(4):850–851, December 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:ERS

[Ano06e] Anonymous. Eurographics: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 25(4):852, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:IAI

[Ano06f] Anonymous. Index: Author index volume 25 (2006). Computer Graphics Forum, 25(4):853–857, December 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2006:RIW

[Ano06g] Anonymous. Reports: 4th International Workshop on Volume Graphics. Computer Graphics Forum, 25(1):143, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:E

[Ano07a] Anonymous. Erratum. Computer Graphics Forum, 26(4):877, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:ECC

[Ano07b] Anonymous. Eurographics: CGForum 2007 Cover Image “Trillion Triangle Terrain” by Andreas Dietrich, Gerd Marmitt and Philipp Slusallek. Computer Graphics Forum, 26(1):129–130, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:EEF

[Ano07c] Anonymous. Eurographics: Eurographics Fellow. Computer Graphics Forum, 26(1):127–128, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:EEHa

[Ano07d] Anonymous. Eurographics: Eurographics Honorary Fellowship. Computer Graphics Forum, 26(1):123–124, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:EEHb

[Ano07e] Anonymous. Eurographics: Eurographics Honorary Fellowship. Computer Graphics Forum, 26(1):125–126, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:II

[Ano07f] Anonymous. Index: Index. Computer Graphics Forum, 26(4): 871–876, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:REG

[Ano07g] Anonymous. Reports: 28th EUROGRAPHICS General Assembly. Computer Graphics Forum, 26(4):860–866, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:RCR

[Ano07h] Anonymous. Reports: CGF reviewers. Computer Graphics Forum, 26(4):870, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:REC

[Ano07i] Anonymous. Reports: EG2008 — call for papers. Computer Graphics Forum, 26(2):240, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:RES

[Ano07j] Anonymous. Reports: Eurographics Symposium on Rendering Nicosia, Cyprus, June 26–28, 2006. Computer Graphics Forum, 26 (2):241, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:RNE

[Ano07k] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 26(4):867–868, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2007:RRS

[Ano07l] Anonymous. Reports: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 26(4):869, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:ARN

[Ano08a] Anonymous. Auditor’s reports: New EUROGRAPHICS fellows. Computer Graphics Forum, 27(6):xi–xiv, September 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:BR

[Ano08b] Anonymous. Book review. Computer Graphics Forum, 27(8): 2247, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:ITF

[Ano08c] Anonymous. Invited talks: Frontmatter. Computer Graphics Forum, 27(2):i–xxii, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:REG

[Ano08d] Anonymous. Reports: 29th EUROGRAPHICS General Assembly. Computer Graphics Forum, 27(6):iii–x, September 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:RCAa

[Ano08e] Anonymous. Reports: Computational Aesthetics 2008. Computer Graphics Forum, 27(6):i, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2008:RCAb

[Ano08f] Anonymous. Reports: Computational Aesthetics 2009 call for papers. Computer Graphics Forum, 27(6):ii, September 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2009:E

[Ano09a] Anonymous. Erratum. Computer Graphics Forum, 28(8):2396, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2009:ITF

[Ano09b] Anonymous. Invited talks: Frontmatter. Computer Graphics Forum, 28(2):i–xix, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2009:REG

[Ano09c] Anonymous. Reports: 30th EUROGRAPHICS General Assembly. Computer Graphics Forum, 28(6):1726–1731, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2009:RNE

[Ano09d] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 28(6):1733–1734, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arabnia:1986:FOR

[AO86] H. R. Arabnia and M. A. Oliver. Fast operations on raster images with SIMD machine architectures. Computer Graphics Forum, 5 (3):179–188, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arabnia:1987:ARR

[AO87] H. R. Arabnia and M. A. Oliver. Arbitrary rotation of raster images with SIMD machine architectures. Computer Graphics Forum, 6(1):3–11, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arabnia:1989:TNF

[AO89] H. R. Arabnia and M. A. Oliver. A Transputer network for fast operations on digitised images. Computer Graphics Forum, 8(1): 3–11, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anantakrishnan:1992:ICA

[AP92] N. Anantakrishnan and L. A. Piegl. Integer de Casteljau algorithm for rasterizing NURBS curves. Computer Graphics Forum, 11(2):151–162, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agrawal:1994:PRD

[AR94] A. Agrawal and A. A. G. Requicha. A paradigm for the robust design of algorithms for geometric modeling. Computer Graphics Forum, 13(3):C/33–C/44, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arad:1995:IWU

[AR95] Nur Arad and Daniel Reisfeld. Image warping using few anchor points and radial functions. Computer Graphics Forum, 14(1): 35–46, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andreia:2006:GES

[AR06] Maria Andréia and Formico Rodrigues. Guest editorial: Selected papers from the 18th Brazilian Symposium on Computer Graphics and Image Processing (SIBGRAPI’ 2005). Computer Graphics Forum, 25(4):673–674, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arad:1994:DIG

[Ara94] N. Arad. Designing and implementing a grid-distortion mapping based on variational principles. Computer Graphics Forum, 13(3): C/259–C/270, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arbab:1990:RFE

[Arb90] Farhad Arbab. Report on the Fourth Eurographics Workshop on Intelligent CAD. Computer Graphics Forum, 9(3):283–284, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abdul-Rahman:2005:RIS

[ARC05] Alfie Abdul-Rahman and Min Chen. Rendering II: Spectral volume rendering based on the kubelka-munk theory. Computer Graphics Forum, 24(3):413–422, September 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1984:RAX

[Arn84] D. B. Arnold. Report on the ANSC X3H3 meeting at Carmel, Oct 1st–5th ’84. Computer Graphics Forum, 3(4):321–322, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1989:REC

[Arn89] David Arnold. Report on the EUROGRAPHICS ’89 conference. Computer Graphics Forum, 8(4):361–366, December 1989. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:1991:ECE

[Arn91] David Arnold. EUROGRAPHICS ’92: Cambridge, England — 7–11 September 1992. Computer Graphics Forum, 10(4):367–372, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnold:2008:SPE

[Arn08] David Arnold. Selected papers from the EG 2007 cultural heritage track: Cultural heritage as a vehicle for basic research in computing science: Pasteur’s quadrant and a use-inspired basic research agenda. Computer Graphics Forum, 27(8):2188–2196, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arya:1984:FAP

[Ary84] K. Arya. A functional approach to picture manipulation. Computer Graphics Forum, 3(1):35–45, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arya:1986:FAA

[Ary86] Kavi Arya. A functional approach to animation. Computer Graphics Forum, 5(4):297–311, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abas:1992:GGT

[AS92] S. J. Abas and A. Salman. Geometric and group-theoretic methods for computer graphic studies of Islamic symmetric patterns. Computer Graphics Forum, 11(1):43–53, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aref:1995:APV

[AS95] Walid G. Aref and Hanan Samet. An algorithm for perspective viewing of objects represented by octrees. Computer Graphics Forum, 14(1):59–66, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Algorri:1996:MS

[AS96a] Maria-Elena Algorri and Francis Schmitt. Mesh simplification. Computer Graphics Forum, 15(3):C77–C86, September 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Algorri:1996:SRU

[AS96b] Maria-Elena Algorri and Francis Schmitt. Surface reconstruction from unstructured 3D data. Computer Graphics Forum, 15(1): 47–60, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arbab:1998:SEW

[AS98] Farhad Arbab and Philipp Slusallek. Sixth Eurographics Workshop on Programming Paradigms in Graphics (WPPG97). Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=228&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=17.

Attene:2000:ASR

[AS00] Marco Attene and Michela Spagnuolo. Automatic surface reconstruction from point sets in space. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=438&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Andujar:2000:IOC

[ASVNB00] Carlos Andújar, Carlos Saona-Vázquez, Isabel Navazo, and Pere Brunet. Integrating occlusion culling and levels of detail through hardly-visible sets. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=442&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Angst:2008:ARE

[ATBG08] Roland Angst, Nils Thuerey, Mario Botsch, and Markus Gross. Animation: Robust and efficient wave simulations on deforming meshes. Computer Graphics Forum, 27(7):1895–1900, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Austen:1991:REU

[Aus91] Malcolm Austen. Report of the 9th Eurographics UK conference. Computer Graphics Forum, 10(2):179–180, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2004:NCB

[AVF04] C. Andújar, P. Vázquez, and M. Fairén. Navigation and crowd behaviour in virtual environments: Way-finder: guided tours through complex walkthrough models. Computer Graphics Forum, 23(3):499–508, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ashraf:2000:GCM

[AW00] G. Ashraf and K. C. Wong. Generating consistent motion transition via decoupled framespace interpolation. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=437&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Assa:2007:ICD

[AW07] J. Assa and L. Wolf. Image and color: Diorama construction from a single image. Computer Graphics Forum, 26(3):599–608, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arbree:2008:RSS

[AWB08] Adam Arbree, Bruce Walter, and Kavita Bala. Rendering and simplification: Single-pass scalable subsurface rendering with lightcuts. Computer Graphics Forum, 27(2):507–516, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Badler:1993:GBA

[Bad93] N. I. Badler. Graphical behaviors and animated agents. Computer Graphics Forum, 12(3):C513–??, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blanz:2006:MIC

[BAHS06] V. Blanz, I. Albrecht, J. Haber, and H.-P. Seidel. Modelling II: Creating face models from vague mental images. Computer Graphics Forum, 25(3):645–654, September 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bajaj:2003:IPV

[Baj03a] Chandrajit Bajaj. Invited presentation: Volumetric filtering, modeling and visualization for nano-medicine. Computer Graphics Forum, 22(3):xvii, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bajaj:2003:VFM

[Baj03b] Chandrajit Bajaj. Volumetric filtering, modeling and visualization for nano-medicine. Computer Graphics Forum, 22(3):??, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Budge:2008:ICF

[BAJ08] B. C. Budge, J. C. Anderson, and K. I. Joy. Illumination: Caustic forecasting: Unbiased estimation of caustic lighting for global illumination. Computer Graphics Forum, 27(7):1963–1970, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bakker:1988:ANS

[Bak88] Miente Bakker. Adding name sets and workstation filters to GKS: a future extension. Computer Graphics Forum, 7(1):3–7, March 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bakker:1990:PSP

[Bak90] Miente Bakker. Parametric surfaces in PHIGS PLUS. A new chance for patterning and hatching? Computer Graphics Forum, 9(1):3–8, March 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bakker:1991:LIC

[Bak91a] Miente Bakker. At last an ISO C binding of GKS. Computer Graphics Forum, 10(3):249–254, September 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bakker:1991:TSP

[Bak91b] Miente Bakker. Triangle sets in PHIGS PLUS: a valuable link with finite element modeling. Computer Graphics Forum, 10(1): 61–65, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barsky:1992:GCB

[Bar92] B. A. Barsky. Geometric continuity: basic theory and applications. Computer Graphics Forum, 11(3):C459–??, ???? 1992. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barr:1993:GPF

[Bar93] R. Barr. GIS-product or function? Computer Graphics Forum, 12(3):C510–??, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bartz:2005:SAR

[Bar05] Dirk Bartz. State of the art reviews: Virtual endoscopy in research and clinical practice. Computer Graphics Forum, 24(1):111–126, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barhak:2006:RGI

[Bar06] Jacob Barhak. Reports: GRAPP 2006 International Conference on Computer Graphics Theory and Applications. Computer Graphics Forum, 25(4):837, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:2005:CVU

[BAU05] K. W. Brodlie, M. R. Asim, and K. Unsworth. Constrained visualization using the Shepard interpolation family. Computer Graphics Forum, 24(4):809–820, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bland:1988:UDS

[BB88] D. Bland and T. Bonnet. The use of digital signal processors in computer graphics. Computer Graphics Forum, 7(2):91–96, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhalla:1991:OOD

[BB91] Neelam Bhalla and S. Balasundaram. Object-oriented data modelling for graphics databases: a declarative approach. Computer Graphics Forum, 10(4):297–308, December 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boyer:1999:FLS

[BB99] V. Boyer and J. J. Bourdin. Fast lines: a span by span method. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=358&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Bebie:2000:VBR

[BB00a] T. Bebie and H. Bieri. A video-based 3D-reconstruction of soccer games. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=431&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Benedens:2000:TBD

[BB00b] Oliver Benedens and Christoph Busch. Towards blind detection of robust watermarks in polygonal models. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=412&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Beckhaus:2007:TEL

[BB07] S. Beckhaus and K. J. Blom. Teaching, exploring, learning— developing tutorials for in-class teaching and self-learning. Computer Graphics Forum, 26(4):725–736, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brunner:2008:GPF

[BB08] D. Brunner and G. Brunnett. Geometry processing: Fast force field approximation and its application to skeletonization of discrete 3D objects. Computer Graphics Forum, 27(2):261–270, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banf:2009:VHE

[BB09] Michael Banf and Volker Blanz. Virtual humans: Example-based rendering of eye movements. Computer Graphics Forum, 28(2): 659–666, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchart:2008:MVG

[BBA08] C. Buchart, D. Borro, and A. Amundarain. Meshes and volumes: GPU local triangulation: an interpolating surface reconstruction algorithm. Computer Graphics Forum, 27(3):807–814, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:2005:SAR

[BBC^{+}05] Ken Brodlie, John Brooke, Min Chen, David Chisnall, Ade Fewings, Chris Hughes, Nigel W. John, Mark W. Jones, Mark Riding, and Nicolas Roard. State of the art reviews: Visual supercomputing: Technologies, applications and challenges. Computer Graphics Forum, 24(2):217–245, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouville:1985:GHQ

[BBDM85] C. Bouville, R. Brusq, J. L. Dubois, and I. Marchal. Generating high quality pictures by ray-tracing. Computer Graphics Forum, 4(2):87–99, June 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brouillat:2009:BMC

[BBL^{+}09] Jonathan Brouillat, Christian Bouville, Brad Loos, Charles Hansen, and Kadi Bouatouch. A Bayesian Monte Carlo approach to global illumination. Computer Graphics Forum, 28(8): 2315–2329, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bettels:1988:GDW

[BBMR88] Jürgen Bettels, Peter R. Bono, Eileen McGinnis, and Joachim Rix. Guidelines for determining when to use GKS and when to use PHIGS. Computer Graphics Forum, 7(4):347–354, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bronsvoort:2002:FMV

[BBN02] Willem F. Bronsvoort, Rafael Bidarra, and Alex Noort. Feature model visualization. Computer Graphics Forum, 21(4):661–673, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulanger:2008:RTR

[BBP08] K. Boulanger, K. Bouatouch, and S. Pattanaik. Real time: Rendering trees with indirect lighting in real time. Computer Graphics Forum, 27(4):1189–1198, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Busking:2009:VRD

[BBP09] Stef Busking, Charl P. Botha, and Frits H. Post. Volume rendering: Direct visualization of deformation in volumes. Computer Graphics Forum, 28(3):799–806, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blanz:2003:FSR

[BBPV03a] V. Blanz, C. Basso, T. Poggio, and T. Vetter. Faces and speech: Reanimating faces in images and video. Computer Graphics Forum, 22(3):641–650, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blanz:2003:RFI

[BBPV03b] V. Blanz, C. Basso, T. Poggio, and T. Vetter. Reanimating faces in images and video. Computer Graphics Forum, 22(3):641–650, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Budge:2009:GCD

[BBS^{+}09] Brian Budge, Tony Bernardin, Jeff A. Stuart, Shubhabrata Sengupta, Kenneth I. Joy, and John D. Owens. GPU: Out-of-core data management for path tracing on hybrid resources. Computer Graphics Forum, 28(2):385–396, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bordeux:1999:EFP

[BBT99] Christophe Bordeux, Ronan Boulic, and Daniel Thalmann. An efficient and flexible perception pipeline for autonomous agents. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=324&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Barla:2006:MIS

[BBT^{+}06] Pascal Barla, Simon Breslav, Joëlle Thollot, François Sillion, and Lee Markosian. Modelling II: Stroke pattern analysis and synthesis. Computer Graphics Forum, 25(3):663–671, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bokeloh:2009:SDI

[BBW^{+}09] M. Bokeloh, A. Berner, M. Wand, H.-P. Seidel, and A. Schilling. Symmetry detection and intelligent galleries: Symmetry detection using feature lines. Computer Graphics Forum, 28(2):697–706, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boissonnat:2001:CFS

[BC01] Jean-Daniel Boissonnat and Frédéric Cazals. Coarse-to-fine surface simplification with geometric guarantees. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bourguignon:2001:DIA

[BCD01] David Bourguignon, Marie-Paule Cani, and George Drettakis. Drawing for illustration and annotation in 3D. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Branco:1994:SMI

[BCF94] V. Branco, A. Costa, and F. Nunes Ferreira. Sketching 3D models with 2D interaction devices. Computer Graphics Forum, 13(3): C/489–C/502, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Behrendt:2005:MNR

[BCF^{+}05] S. Behrendt, C. Colditz, O. Franzke, J. Kopf, and O. Deussen. Modelling nature: Realistic real-time rendering of landscapes using billboard clouds. Computer Graphics Forum, 24(3):507–516, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barequet:1996:BHR

[BCG^{+}96] Gill Barequet, Bernard Chazelle, Leonidas J. Guibas, Joseph S. B. Mitchell, and Ayellet Tal. BOXTREE: a hierarchical representation for surfaces in 3D. Computer Graphics Forum, 15(3):C387– C396, C484, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ben-Chen:2008:PTC

[BCGB08] Mirela Ben-Chen, Craig Gotsman, and Guy Bunin. Parameterization and transformations: Conformal flattening by curvature prescription and metric scaling. Computer Graphics Forum, 27 (2):449–458, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulic:1995:HEI

[BCH^{+}95] R. Boulic, T. Capin, Z. Huang, P. Kalra, B. Lintermann, N. Magnenat-Thalmann, L. Moccozet, T. Molet, I. Pandzic, K. Saar, A. Schmitt, J. Shen, and D. Thalmann. The HU-MANOID environment for interactive animation of multiple deformable human characters. Computer Graphics Forum, 14(3): C/337–C/348, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bando:2003:AHL

[BCN03a] Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Animating hair with loosely connected particles. Computer Graphics Forum, 22(3):411–418, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bando:2003:AFH

[BCN03b] Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Animation of fluids and hair: Animating hair with loosely connected particles. Computer Graphics Forum, 22(3):411–418, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berger:1996:CCV

[BCS96] M. O. Berger, C. Chevrier, and G. Simon. Compositing computer and video image sequences: Robust algorithms for the reconstruction of the camera parameters. Computer Graphics Forum, 15 (3):C23–C32, C457–C458, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Butkiewicz:2008:VAV

[BCWR08] Thomas Butkiewicz, Remco Chang, Zachary Wartell, and William Ribarsky. Visual analytics: Visual analysis and semantic exploration of urban LIDAR change detection. Computer Graphics Forum, 27(3):903–910, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bourque:2004:ICM

[BD04] Eric Bourque and Gregory Dudek. Image calibration and matching: Procedural texture matching and transformation. Computer Graphics Forum, 23(3):461–468, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bae:2007:IDM

[BD07] Soonmin Bae and Frédo Durand. Imaging (1): Defocus magnification. Computer Graphics Forum, 26(3):571–579, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burch:2008:SVT

[BD08] M. Burch and S. Diehl. Software visualization: TimeRadarTrees: Visualizing dynamic compound digraphs. Computer Graphics Forum, 27(3):823–830, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banterle:2009:HDR

[BDA^{+}09] Francesco Banterle, Kurt Debattista, Alessandro Artusi, Sumanta Pattanaik, Karol Myszkowski, Patrick Ledda, and Alan Chalmers. High dynamic range imaging and low dynamic range expansion for generating HDR content. Computer Graphics Forum, 28(8): 2343–2367, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Battiato:2007:DMF

[BDFG07] S. Battiato, G. Di Blasi, G. M. Farinella, and G. Gallo. Digital Mosaic frameworks — an overview. Computer Graphics Forum, 26(4):794–812, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:2004:SAR

[BDG^{+}04] K. W. Brodlie, D. A. Duce, J. R. Gallop, J. P. R. B. Walton, and J. D. Wood. State of the art reviews: Distributed and collaborative visualization. Computer Graphics Forum, 23(2):223–251, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barbic:1984:OGO

[BDŠ84] J. Barbič, F. Dacar, and M. Špegel. Oriented geometric objects in computer graphics and numerical control. Computer Graphics Forum, 3(1):103–110, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barthe:2003:TDP

[BDS^{+}03] L. Barthe, N. A. Dodgson, M. A. Sabin, B. Wyvill, and V. Gaildrat. Two-dimensional potential fields for advanced implicit modeling operators. Computer Graphics Forum, 22(1):23–33, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Busch:1995:MOC

[BE95] Chr Busch and M. Eberle. Morphological operations for colorcoded images. Computer Graphics Forum, 14(3):C/193–C/204, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Belch:1987:EHP

[Bel87] F. R. Belch. An economic high performance pixel memory with 3-D solids capability. Computer Graphics Forum, 6(1):13–25, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benest:1994:SED

[Ben94] I. D. Benest. A schematic entry drawing capability in a linearised hypermedia system. Computer Graphics Forum, 13(5):293–303, December 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bertails:2009:SAH

[Ber09] Florence Bertails. Shape analysis and helices: Linear time superhelices. Computer Graphics Forum, 28(2):417–426, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bimber:2000:ARB

[BES00] Oliver Bimber, L. Miguel Encarnação, and Dieter Schmalstieg. Augmented reality with back-projection systems using transflective surfaces. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=408&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Bez:1984:NMC

[BF84] H. E. Bez and A. Filipe. A new method of curve parameterization with applications in computer aided design. Computer Graphics Forum, 3(1):29–34, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernstein:2009:MSF

[BF09] Gilbert Bernstein and Don Fussell. Modeling and sculpting: Fast, exact, linear Booleans. Computer Graphics Forum, 28(5):1269– 1278, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buhmann:1998:DCQ

[BFH^{+}98] J. M. Buhmann, D. W. Fellner, M. Held, J. Ketterer, and J. Puzicha. Dithered color quantization. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=269&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Brodlie:1982:DNG

[BFTL82] K. W. Brodlie, D. L. Fisher, G. G. Tolton, and T. W. Lambert. The development of the NAG Graphical Supplement. Computer Graphics Forum, 1(3):133–142, September 1982. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1985:RWV

[BG85] K. W. Brodlie and M. Gobel. Report on workshop on validation of graphics software. Computer Graphics Forum, 4(2):149–150, June 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bennis:1989:DMT

[BG89] Chakib Bennis and Andre Gagalowicz. 2-D macroscopic texture synthesis. Computer Graphics Forum, 8(4):291–300, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Busch:1993:INN

[BG93] C. Busch and M. H. Gross. Interactive neural network texture analysis and visualization for surface reconstruction in medical imaging. Computer Graphics Forum, 12(3):C49–C60, ???? 1993.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Balaguer:1995:SA

[BG95] J.-F. Balaguer and E. Gobbetti. Sketching 3D animations. Computer Graphics Forum, 14(3):C/241–C/258, September 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barkan:2001:CRO

[BG01] Ella Barkan and Dan Gordon. CP3: Robust, output-sensitive display of convex polyhedra in scanline mode. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bogomjakov:2002:URS

[BG02] A. Bogomjakov and C. Gotsman. Universal rendering sequences for transparent vertex caching of progressive meshes. Computer Graphics Forum, 21(2):137–148, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckner:2007:IVV

[BG07] S. Bruckner and M. E. Gröller. Interaction, VR, and visualization: Style transfer functions for illustrative volume rendering. Computer Graphics Forum, 26(3):715–724, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bogomjakov:2008:SRR

[BG08] Alexander Bogomjakov and Craig Gotsman. Scanning and reconstruction: Reduced depth and visual hulls of complex 3D scenes. Computer Graphics Forum, 27(2):175–182, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckner:2009:VRI

[BG09] Stefan Bruckner and M. Eduard Gröller. Volume rendering: Instant volume visualization using maximum intensity difference accumulation. Computer Graphics Forum, 28(3):775–782, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Borro:2004:AOV

[BGAM04] D. Borro, A. García-Alonso, and L. Matey. Approximation of optimal voxel size for collision detection in maintainability simulations within massive virtual environments. Computer Graphics Forum, 23(1):13–23, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bidmon:2008:VAV

[BGB^{+}08a] Katrin Bidmon, Sebastian Grottel, Fabian Bös, Jürgen Pleiss, and Thomas Ertl. Visualization applications: Visual abstractions of solvent pathlines near protein cavities. Computer Graphics Forum, 27(3):935–942, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brouillat:2008:IPD

[BGB08b] J. Brouillat, P. Gautron, and K. Bouatouch. Illumination: Photon-driven irradiance cache. Computer Graphics Forum, 27 (7):1971–1978, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bogomjakov:2008:PMD

[BGI08] Alexander Bogomjakov, Craig Gotsman, and Martin Isenburg. Points and meshes: Distortion-free steganography for polygonal meshes. Computer Graphics Forum, 27(2):637–642, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beylot:1996:ITM

[BGK^{+}96] P. Beylot, F. Gingins, P. Kalra, N. Magnenat Thalmann, W. Maurel, D. Thalmann, and J. Fasel. 3D interactive topological modeling using visible human dataset. Computer Graphics Forum, 15 (3):C33–C44, C459–C460, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bordegoni:1993:DGL

[BH93] M. Bordegoni and M. Hemmje. A dynamic gesture language and graphical feedback for interaction in a 3D user interface. Computer Graphics Forum, 12(3):C1–C11, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulic:1996:CAS

[BH96] R. Boulic and Gerard Hegron, editors. Computer animation and simulation ’96: proceedings of the Eurographics Workshop in Poitiers, France, August 31–September 1, 1996, Springer Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-211-82885-0. LCCN ????

Brown:2006:REP

[BH06] Judy Brown and Werner Hansmann. Reports: Education Programme at Eurographics 2006. Computer Graphics Forum, 25(4): 840–841, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boubekeur:2006:GRV

[BHGS06] Tamy Boubekeur, Wolfgang Heidrich, Xavier Granier, and Christophe Schlick. Geometry reconstruction: Volume-surface trees. Computer Graphics Forum, 25(3):399–406, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1987:CMG

[BHM87] Ken Brodlie, Lofton Henderson, and Anne Mumford. The CGM — a metafile for GKS? Computer Graphics Forum, 6(2):87–90, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baxter:2006:DLD

[BiA06] William Baxter and Ken ichi Anjyo. Deformations: Latent doodle space. Computer Graphics Forum, 25(3):477–485, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bijl:1987:MDT

[Bij87] Aart Bijl. Making drawings talk: Pictures in minds and machines. Computer Graphics Forum, 6(4):289–298, December 1987. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bajaj:2004:SOL

[BIMO04] Chandrajit Bajaj, Insung Ihm, Jungki Min, and Jinsang Oh. SIMD optimization of linear expressions for programmable graphics hardware. Computer Graphics Forum, 23(4):697–714, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bimber:2008:SAR

[BIWG08] Oliver Bimber, Daisuke Iwai, Gordon Wetzstein, and Anselm Grundhöfer. State of the art reviews: The visual computing of projector-camera systems. Computer Graphics Forum, 27(8):

2219–2245, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brunet:1994:PRC

[BJ94] P. Brunet and F. W. Jansen, editors. Photorealistic rendering in computer graphics: proceedings of the Second Eurographics Workshop on Rendering, Focus on computer graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56449-7 (Berlin), 0-387-56449-7 (New York). LCCN T385 .E98 1991. Workshop held in May 1991 at the University of Catalunya in Barcelona, Spain.

Bar-Joseph:2003:CHC

[BJCO03a] Ziv Bar-Joseph and Daniel Cohen-Or. Compression: Hierarchical context-based pixel ordering. Computer Graphics Forum, 22(3): 349–358, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bar-Joseph:2003:HCB

[BJCO03b] Ziv Bar-Joseph and Daniel Cohen-Or. Hierarchical contextbased pixel ordering. Computer Graphics Forum, 22(3):349–358, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhagvat:2009:GRR

[BJCW09] D. Bhagvat, S. Jeschke, D. Cline, and P. Wonka. GPU rendering of relief mapped conical frusta. Computer Graphics Forum, 28(8): 2131–2139, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2001:RFB

[BK01] Mario Botsch and Leif Kobbelt. Resampling feature and blend regions in polygonal meshes for surface anti-aliasing. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bischoff:2003:CSS

[BK03a] Stephan Bischoff and Leif Kobbelt. Capture and segmentation: Sub-voxel topology control for level-set surfaces. Computer Graphics Forum, 22(3):273–280, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bischoff:2003:SVT

[BK03b] Stephan Bischoff and Leif Kobbelt. Sub-voxel topology control for level-set surfaces. Computer Graphics Forum, 22(3):273–280, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2003:MMS

[BK03c] Mario Botsch and Leif Kobbelt. Meshes: Multiresolution surface representation based on displacement volumes. Computer Graphics Forum, 22(3):483–491, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2003:MSR

[BK03d] Mario Botsch and Leif Kobbelt. Multiresolution surface representation based on displacement volumes. Computer Graphics Forum, 22(3):483–491, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bischoff:2005:GIS

[BK05a] Stephan Bischoff and Leif Kobbelt. Geometry III: Structure preserving CAD model repair. Computer Graphics Forum, 24(3): 527–536, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2005:GIR

[BK05b] Mario Botsch and Leif Kobbelt. Geometry IV: Real-time shape editing using radial basis functions. Computer Graphics Forum, 24(3):611–621, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernardini:2000:DDO

[BKES00] Fausto Bernardini, James T. Klosowski, and Jihad El-Sana. Directional discretized occluders for accelerated occlusion culling. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=443&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Bright:1986:SSV

[BL86] Stephen Bright and Susan Laflin. Shading of solid voxel models. Computer Graphics Forum, 5(2):131–137, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burley:2008:TPP

[BL08] Brent Burley and Dylan Lacewell. Textures: Ptex: Per-face texture mapping for production rendering. Computer Graphics Forum, 27(4):1155–1164, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blake:1988:RTE

[Bla88] E. Blake. Report on the third Eurographics workshop on graphics hardware. Computer Graphics Forum, 7(4):361–362, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banterle:2009:PEI

[BLD^{+}09] Francesco Banterle, Patrick Ledda, Kurt Debattista, Marina Bloj, Alessandro Artusi, and Alan Chalmers. A psychophysical evaluation of inverse tone mapping techniques. Computer Graphics Forum, 28(1):13–25, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bloomenthal:1997:BEC

[Blo97] Jules Bloomenthal. Bulge elimination in convolutional surfaces. Computer Graphics Forum, 16(1):31–41, March 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=114&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=16.

Blasi:1993:RAD

[BLS93] P. Blasi, B. Le Saec, and C. Schlick. A rendering algorithm for discrete volume density objects. Computer Graphics Forum, 12 (3):C201–C210, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brivio:1993:FMD

[BM93] P. A. Brivio and D. Marini. A fractal method for digital elevation model construction and its application to a mountain region.

Computer Graphics Forum, 12(5):297–309, December 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bellmann:2008:VAS

[BMD^{+}08] Jens Bellmann, Frank Michel, Eduard Deines, Martin Hering-Bertram, Jan Mohring, and Hans Hagen. Visualization applications: Sound tracing: Rendering listener specific acoustic Room properties. Computer Graphics Forum, 27(3):943–950, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bielser:1999:ICT

[BMG99] Daniel Bielser, Volker A. Maiwald, and Markus H. Gross. Interactive cuts through 3-dimensional soft tissue. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=325&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Breen:2001:MBD

[BMWM01] David E. Breen, Sean Mauch, Ross T. Whitaker, and Jia Mao. 3D metamorphosis between different types of geometric models. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruneton:2008:ISB

[BN08a] Eric Bruneton and Fabrice Neyret. Interactive and sketch-based techniques: Real-time rendering and editing of vector-based terrains. Computer Graphics Forum, 27(2):311–320, April 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruneton:2008:PMP

[BN08b] Eric Bruneton and Fabrice Neyret. Participating media: Precomputed atmospheric scattering. Computer Graphics Forum, 27 (4):1079–1086, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bro-Nielsen:1996:RTV

[BNC96] Morten Bro-Nielsen and Stephane Cotin. Real-time volumetric deformable models for surgery simulation using finite elements and condensation. Computer Graphics Forum, 15(3):C57–C66, C461, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brunet:2001:HCC

[BNRSV01] Pere Brunet, Isabel Navazo, Jarek Rossignac, and Carlos Saona-Vázquez. Hoops: 3D curves as conservative occluders for cellvisibility. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bono:1985:RTM

[Bon85] Peter R. Bono. Report of the timberline meeting of ISO TC97/ SC21/WG2 ‘Computer Graphics’. Computer Graphics Forum, 4 (4):383–386, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2007:RYR

[Bot07] Mario Botsch. Reports: Young Researcher Award 2007. Computer Graphics Forum, 26(4):856, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouatouch:1988:TDP

[Bou88] Kadi Bouatouch. Theoretical developments on polygonal approximation of parametric surfaces for ray tracing. Computer Graphics Forum, 7(4):257–264, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouatouch:1990:REW

[Bou90] K. Bouatouch. Report on the EUROGRAPHICS workshop on photosimulation, realism and physics in computer graphics. Computer Graphics Forum, 9(4):377–378, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bovey:1990:BLD

[Bov90] J. D. Bovey. A browser for large directed graphs. Computer Graphics Forum, 9(3):195–204, September 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bowen:1992:XWZ

[Bow92] Jonathan P. Bowen. X: Why Z? Computer Graphics Forum, 11 (4):221–234, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1982:REW

[BP82] K. W. Brodlie and G. E. Pfaff. Report on the EEC workshop on graphics certification defining a device for certification Miltenberg/Darmstadt, FRG, 25–28 April 1982. Computer Graphics Forum, 1(3):88–90, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1983:AIG

[BP83a] K. W. Brodlie and G. Pfaff. An algorithmic interpretation of the GKS TEXT primitive. Computer Graphics Forum, 2(4):233–241, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1983:REW

[BP83b] K. W. Brodlie and G. E. Pfaff. Report on the EEC workshop on certification of graphics software — Rixensart, Belgium. Computer Graphics Forum, 2(1):23–24, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bao:1993:PRA

[BP93] Hujun Bao and Qunsheng Peng. A progressive radiosity algorithm for scenes containing curved surfaces. Computer Graphics Forum, 12(3):C399–C408, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bao:1998:IM

[BP98] Hujun Bao and Qunsheng Peng. Interactive 3D morphing. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=250&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Briere:2001:ARS

[BP01] Normand Brière and Pierre Poulin. Adaptive representation of specular light. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulic:2004:RAE

[BPB^{+}04] Ronan Boulic, Dinesh K. Pai, Norman Badler, Mathieu Desbrun, and Lionel Reveret. Reports: ACM/EG Symposium on Computer Animation 2004. Computer Graphics Forum, 23(4):841, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bottger:2008:IWD

[BPBD08] Joachim Böttger, Martin Preiser, Michael Balzer, and Oliver Deussen. Imaging and warping: Detail-in-context visualization for satellite imagery. Computer Graphics Forum, 27(2):587–596, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boudon:2003:IDB

[BPF^{+}03a] Frédéric Boudon, Przemysław Prusinkiewicz, Pavol Federl, Christophe Godin, and Radosław Karwowski. Interactive design of bonsai tree models. Computer Graphics Forum, 22(3):591–599, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boudon:2003:PEI

[BPF^{+}03b] Frédéric Boudon, Przemysław Prusinkiewicz, Pavol Federl, Christophe Godin, and Radosław Karwowski. Practice and experience: Interactive design of bonsai tree models. Computer Graphics Forum, 22(3):591–599, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bahmutov:2006:MIE

[BPM06] Gleb Bahmutov, Voicu Popescu, and Mihai Mudure. Modelling II: Efficient large scale acquisition of building interiors. Computer Graphics Forum, 25(3):655–662, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bosch:2004:NPA

[BPMG04] Carles Bosch, Xavier Pueyo, Stéphane Mérillou, and Djamchid Ghazanfarpour. Natural phenomena and aging: a physicallybased model for rendering realistic scratches. Computer Graphics Forum, 23(3):361–370, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bosch:2008:RRI

[BPMG08] C. Bosch, X. Pueyo, S. Mérillou, and D. Ghazanfarpour. Rendering: a resolution independent approach for the accurate rendering of grooved surfaces. Computer Graphics Forum, 27(7):1937–1944, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneel:2009:AMR

[BPV^{+}09] Nicolas Bonneel, Sylvain Paris, Michiel Van De Panne, Frédo Durand, and George Drettakis. Appearance measurement and rendering: Single photo estimation of hair appearance. Computer Graphics Forum, 28(4):1171–1180, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Botsch:2007:DAS

[BPWG07] Mario Botsch, Mark Pauly, Martin Wicke, and Markus Gross. Deformation: Adaptive space deformations based on rigid cells. Computer Graphics Forum, 26(3):339–347, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouatouch:1996:CHO

[BPZ96] Kadi Bouatouch, S. N. Pattanaik, and Eric Zeghers. Computation of higher order illumination with a non- deterministic approach. Computer Graphics Forum, 15(3):C327–C337, C480, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banerjee:1996:TEE

[BR96] Raja P. K. Banerjee and Jarek R. Rossignac. Topologically exact evaluation of polyhedra defined in CSG with loose primitives. Computer Graphics Forum, 15(4):205–217, October 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baranoski:1997:ART

[BR97] Gladimir V. G. Baranoski and Jon G. Rokne. An algorithmic reflectance and transmittance model for plant tissue. Computer Graphics Forum, 16(3):C141–C150, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=150&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Bernardini:2002:MAP

[BR02] Fausto Bernardini and Holly Rushmeier. The 3D model acquisition pipeline. Computer Graphics Forum, 21(2):149–172, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baranoski:2007:RPP

[BR07] G. V. G. Baranoski and J. G. Rokne. Rendering plasma phenomena: Applications and challenges. Computer Graphics Forum, 26 (4):743–768, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bertini:2009:USI

[BRL09] Enrico Bertini, Maurizio Rigamonti, and Denis Lalanne. User studies and interaction: Extended excentric labeling. Computer Graphics Forum, 28(3):927–934, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brodlie:1990:WRF

[Bro90] K. W. Brodlie. Workshop report — future developments in computer graphics teaching. Eurographics Workshop, University of Leeds, Easter 1989. Computer Graphics Forum, 9(3):277–281, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brooks:1995:RVW

[Bro95] Fred P. Brooks, Jr. Realizing virtual worlds. Computer Graphics Forum, 14(3):C/481–??, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beckhaus:2001:GED

[BRS01] Steffi Beckhaus, Felix Ritter, and Thomas Strothotte. Guided exploration with dynamic potential fields: the cubical path system. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bryson:1996:OCV

[Bry96] S. T. Bryson. Opportunities and challenges in virtual reality applications. Computer Graphics Forum, 15(3):1, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barbic:1990:LDG

[BŠ90] Janez Barbič and Marjan Špegel. Language for description of geometric configurations. Computer Graphics Forum, 9(1):9–13, March 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brun:1998:FAI

[BS98] L. Brun and C. Secroun. A fast algorithm for inverse colormap computation. Computer Graphics Forum, 17(4): ??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=289&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Bordoloi:2002:HAI

[BS02] Udeepta Bordoloi and Han-Wei Shen. Hardware accelerated interactive vector field visualization: a level of detail approach. Computer Graphics Forum, 21(3):605–614, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brabec:2003:HBG

[BS03a] Stefan Brabec and Hans-Peter Seidel. Hardware-based graphics: Shadow volumes on programmable graphics hardware. Computer Graphics Forum, 22(3):433–440, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brabec:2003:SVP

[BS03b] Stefan Brabec and Hans-Peter Seidel. Shadow volumes on programmable graphics hardware. Computer Graphics Forum, 22(3): 433–440, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boubekeur:2008:FKA

[BS08] T. Boubekeur and C. Schlick. A flexible kernel for adaptive mesh refinement on GPU. Computer Graphics Forum, 27(1):102–113, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benford:1995:VVV

[BSG^{+}95] S. Benford, D. Snowdon, C. Greenhalgh, R. Ingram, L. Knox, and C. Brown. VR-VIBE: a virtual environment for co-operative information retrieval. Computer Graphics Forum, 14(3):C/349– C/360, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Balsys:2008:PBR

[BSJ08] Ron J. Balsys, K. G. Suffern, and Huw Jones. Point-based rendering of non-manifold surfaces. Computer Graphics Forum, 27(1):

63–72, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouras:1992:SDC

[BSV92] A. Bouras, B. Shariat, and D. Vandorpe. A simple description of complex curves. Computer Graphics Forum, 11(3):C425–C433, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blanz:2004:RAF

[BSVS04] Volker Blanz, Kristina Scherbaum, Thomas Vetter, and Hans-Peter Seidel. Retargeting and animating faces: Exchanging faces in images. Computer Graphics Forum, 23(3):669–676, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulic:1992:CDI

[BT92] Ronan Boulic and Daniel Thalmann. Combined direct and inverse kinematic control for articulated figure motion editing. Computer Graphics Forum, 11(4):189–202, October 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bandi:1995:ASS

[BT95] S. Bandi and D. Thalmann. An adaptive spatial subdivision of the object space for fast collision detection of animated rigid bodies. Computer Graphics Forum, 14(3):C/259–C/270, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bandi:1998:SDE

[BT98] Srikanth Bandi and Daniel Thalmann. Space discretization for efficient human navigation. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=267&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Balmelli:2002:SOT

[BTB02] Laurent Balmelli, Gabriel Taubin, and Fausto Bernardini. Spaceoptimized texture maps. Computer Graphics Forum, 21(3):411– 420, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bittar:1995:ARU

[BTG95] Eric Bittar, Nicolas Tsingos, and Marie-Paule Gascuel. Automatic reconstruction of unstructured 3D data: Combining medial axis and implicit surfaces. Computer Graphics Forum, 14(3):C/457– C/468, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchanan:1996:SEH

[Buc96] John W. Buchanan. Special effects with half-toning. Computer Graphics Forum, 15(3):C97–C108, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchanan:1998:SWU

[Buc98] John W. Buchanan. Simulating wood using a voxel approach. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=258&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Buhler:2001:LIE

[Büh01] Katja Bühler. Linear interval estimations for parametric objects theory and application. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burton:1995:TDC

[Bur95] E. Burton. Thoughtful drawings: a computational model of the cognitive nature of children’s drawing. Computer Graphics Forum, 14(3):C/159–C/170, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buttenfield:1994:MPG

[But94] B. P. Buttenfield. Multimedia and proactive graphics in GIS. Computer Graphics Forum, 13(3):C/513–??, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bercovier:1996:EDM

[BV96] M. Bercovier and O. Volpin. Energy duality methods for piecewise Bézier curve construction. Computer Graphics Forum, 15(2):143– 154, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bercovier:1999:HBS

[BV99] M. Bercovier and O. Volpin. G^{1} hierarchical Bézier surface over arbitrary meshes. Computer Graphics Forum, 18 (4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=376&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Barequet:2009:RIS

[BV09] Gill Barequet and Amir Vaxman. Reconstruction — irregular sampling: Reconstruction of multi-label domains from partial planar cross-sections. Computer Graphics Forum, 28(5):1327–1337, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barn:1987:GIM

[BW87] Balbir S. Barn and Philip Willis. Graphical interaction management. Computer Graphics Forum, 6(2):119–124, May 1987. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baciu:2000:IGN

[BW00] George Baciu and Sai Keung Wong. The impulse graph: a new dynamic structure for global collisions. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=415&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Brodlie:2001:RAV

[BW01] Ken Brodlie and Jason Wood. Recent advances in volume visualization. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bo:2007:SGC

[BW07] Pengbo Bo and Wenping Wang. Surfaces: Geodesic-controlled developable surfaces for modeling paper bending. Computer Graphics Forum, 26(3):365–374, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bachthaler:2009:SEA

[BW09] Sven Bachthaler and Daniel Weiskopf. Scatterplots: Efficient and adaptive rendering of 2-D continuous scatterplots. Computer Graphics Forum, 28(3):743–750, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bittner:2004:GBR

[BWPP04] Jiří Bittner, Michael Wimmer, Harald Piringer, and Werner Purgathofer. GPU-based rendering: Coherent hierarchical culling: Hardware occlusion queries made useful. Computer Graphics Forum, 23(3):615–624, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Breen:1996:IOA

[BWRT96] David E. Breen, Ross T. Whitaker, Eric Rose, and Mihran Tuceryan. Interactive occlusion and automatic object placement for augmented reality. Computer Graphics Forum, 15(3):C11– C22, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benthin:2003:GIS

[BWS03a] Carsten Benthin, Ingo Wald, and Philipp Slusallek. Global illumination: a scalable approach to interactive global illumination. Computer Graphics Forum, 22(3):621–630, September 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benthin:2003:SAI

[BWS03b] Carsten Benthin, Ingo Wald, and Philipp Slusallek. A scalable approach to interactive global illumination. Computer Graphics Forum, 22(3):621–630, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brunet:2005:EEC

[BWS05] Pere Brunet, Phil Willis, and Hans-Peter Seidel. Editorial of the EG Chairs: EG editorial. Computer Graphics Forum, 24(1):1, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bartram:2008:VTA

[BY08] Lyn Bartram and Miao Yao. Visualization tools: Animating causal overlays. Computer Graphics Forum, 27(3):751–758, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boltcheva:2009:MSF

[BYB09] Dobrina Boltcheva, Mariette Yvinec, and Jean-Daniel Boissonnat. Meshing-simplex: Feature preserving Delaunay mesh generation from 3D multi-material images. Computer Graphics Forum, 28 (5):1455–1464, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bao:1995:SCL

[BYP95] Hujun Bao, Jianguo Ying, and Qunsheng Peng. Shading with curve light sources. Computer Graphics Forum, 14(3):C/217– C/227, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chalmers:2005:RVN

[CA05] Alan Chalmers and David Arnold. Reports: VAST2003 5– 7 November 2003. Computer Graphics Forum, 24(1):130–131, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cadik:2008:IEP

[Čad08] M. Čadík. Image editing: Perceptual evaluation of color-to-grayscale image conversions. Computer Graphics Forum, 27(7): 1745–1754, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cline:2008:GIT

[CAE08] David Cline, Daniel Adams, and Parris Egbert. Global illumination: Table-driven adaptive importance sampling. Computer Graphics Forum, 27(4):1115–1123, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cuny:2000:NAM

[CAH00] François Cuny, Laurent Alonso, and Nicolas Holzschuch. A novel approach makes higher order wavelets really efficient for radiosity. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=402&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Cailliau:1995:SWW

[Cai95] Robert Cailliau. The state of the World-Wide Web. Computer Graphics Forum, 14(3):C/395–??, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Callet:1996:PDM

[Cal96] Patrick Callet. Pertinent data for modelling pigmented materials in realistic rendering. Computer Graphics Forum, 15(2):119–127, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Callieri:2007:BRO

[Cal07] M. Callieri. Book review: OpenGL Shading Language by Randi J. Rost. Computer Graphics Forum, 26(4):853–854, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clarberg:2008:LIE

[CAM08a] Petrik Clarberg and Tomas Akenine-Möller. Local illumination: Exploiting visibility correlation in direct illumination. Computer Graphics Forum, 27(4):1125–1136, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clarberg:2008:SLP

[CAM08b] Petrik Clarberg and Tomas Akenine-Möllery. Shadows and lighting: Practical product importance sampling for direct illumination. Computer Graphics Forum, 27(2):681–690, April 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chandak:2009:VPF

[CATM09] Anish Chandak, Lakulish Antani, Micah Taylor, and Dinesh Manocha. Visibility and particles: FastV: From-point visibility culling on complex models. Computer Graphics Forum, 28 (4):1237–1246, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cheng:2009:IGC

[CB09] I. Cheng and A. Basu. Interactive graphics for computer adaptive testing. Computer Graphics Forum, 28(8):2033–2045, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Crespin:1996:ISO

[CBS96] Benoit Crespin, Carole Blanc, and Christophe Schlick. Implicit sweep objects. Computer Graphics Forum, 15(3):C165–C174, C468, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coelho:2007:EMV

[CBSF07] A. Coelho, M. Bessa, A. Augusto Sousa, and F. Nunes Ferreira. Expeditious modelling of virtual urban environments with geospatial L-systems. Computer Graphics Forum, 26(4):769–782, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chung:1993:SPD

[CC93] K. W. Chung and H. S. Y. Chan. Symmetrical patterns from dynamics. Computer Graphics Forum, 12(1):33–40, March 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2000:GAI

[CC00] Min-Hyung Choi and James F. Cremer. Geometrically-aware interactive object manipulation. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=388&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Chen:2006:GCD

[CC06] Chih-Chun Chen and Jung-Hong Chuang. Geometry compression and decompression: Texture adaptation for progressive meshes. Computer Graphics Forum, 25(3):343–350, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2008:RPF

[CC08] Ying-Chieh Chen and Chun-Fa Chang. Rendering: a prism-free method for silhouette rendering in inverse displacement mapping. Computer Graphics Forum, 27(7):1929–1936, October 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Corsini:2008:VRS

[CCC08] Massimiliano Corsini, Marco Callieri, and Paolo Cignoni. Video and relighting: Stereo light probe. Computer Graphics Forum, 27(2):291–300, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Castellani:2008:PMS

[CCFM08] U. Castellani, M. Cristani, S. Fantoni, and V. Murino. Points and meshes: Sparse points matching by combining 3D mesh saliency with statistical descriptors. Computer Graphics Forum, 27(2): 643–652, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2007:MDA

[CCI^{+}07] M. Chen, C. Correa, S. Islam, M. W. Jones, P.-Y. Shen, D. Silver, S. J. Walton, and P. J. Willis. Manipulating, deforming and animating sampled object representations. Computer Graphics Forum, 26(4):824–852, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2008:MVC

[CCI08] B. Chang, D. Cha, and I. Ihm. Meshes and volumes: Computing local signed distance fields for large polygonal models. Computer Graphics Forum, 27(3):799–806, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Collins:2009:TVD

[CCP09] Christopher Collins, Sheelagh Carpendale, and Gerald Penn. Text visualization: DocuBurst: Visualizing document content using language structure. Computer Graphics Forum, 28(3):1039–1046, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:1995:RSV

[CCS95] Wenli Cai, Tianzhou Chen, and Jiaoying Ski. Rendering of surface and volume details in volume data. Computer Graphics Forum, 14(3):C/421–C/430, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chazal:2009:SAG

[CCSG^{+}09] Frédéric Chazal, David Cohen-Steiner, Leonidas J. Guibas, Facundo Mémoli, and Steve Y. Oudot. Shape analysis: Gromovhausdorff stable signatures for shapes using persistence. Computer Graphics Forum, 28(5):1393–1403, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chazal:2009:DSC

[CCSLT09] F. Chazal, D. Cohen-Steiner, A. Lieutier, and B. Thibert. Differentials: Stability of curvature measures. Computer Graphics Forum, 28(5):1485–1496, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clare:1994:EPC

[CD94] A. R. Clare and A. M. Day. Experiments in the parallel computation of 3D convex hulls. Computer Graphics Forum, 13(1): 21–36, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:2000:Ea

[CD00a] Sabine Coquillart and David Duke. Editorial. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:2000:Eb

[CD00b] Sabine Coquillart and David Duke. Editorial. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:2000:Ec

[CD00c] Sabine Coquillart and David Duke. Editorial. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2008:AST

[CD08] D. Chen and F. Dong. Adaptive surface texture synthesis using round-shaped neighbourhoods. Computer Graphics Forum, 27(8): 2128–2142, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Comba:2009:RES

[CDD09] João Comba, Weiskopf Daniel, and Kurt Debattista. Reports: 2009 Eurographics Symposium on Parallel Graphics and Visualization. Computer Graphics Forum, 28(6):1725, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:2007:RTR

[CDG^{+}07] P. Cignoni, M. Di Benedetto, F. Ganovelli, E. Gobbetti, F. Marton, and R. Scopigno. Ray tracing: Ray-casted BlockMaps for large urban models visualization. Computer Graphics Forum, 26 (3):405–413, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cazals:1995:FCH

[CDP95] F. Cazals, G. Drettakis, and C. Puech. Filtering, clustering and hierarchy construction: a new solution for ray-tracing complex scenes. Computer Graphics Forum, 14(3):C/371–C/382, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Corsini:2009:IBM

[CDPS09] Massimiliano Corsini, Matteo Dellepiane, Federico Ponchio, and Roberto Scopigno. Image-based modeling and rendering: Imageto-geometry registration: a mutual information method exploiting illumination-related geometric properties. Computer Graphics Forum, 28(7):1755–1764, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casciani:1984:ACQ

[CFFP84] T. Casciani, B. Falcidieno, G. Fasciolo, and C. Pienovi. An algorithm for constructing a quadtree from polygonal regions. Computer Graphics Forum, 3(4):269–274, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coltelli:1993:AQD

[CFM93] P. Coltelli, G. Faconti, and F. Marfori. On the application of quantization and dithering techniques to history of arts. Computer Graphics Forum, 12(3):C351–C362, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cavaliere:1986:TIG

[CFT86] S. Cavaliere, M. Fantini, and A. Turtur. Text, images and graphics management system for editorial environment. Computer Graphics Forum, 5(1):51–56, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cecconi:2002:AZW

[CG02] Alesandro Cecconi and Martin Galanda. Adaptive zooming in Web cartography. Computer Graphics Forum, 21(4):787–799, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chiba:2007:REW

[CG07a] Norishige Chiba and Eric Galin. Reports: 2nd EG workshop on Natural Phenomena September 5th, 2007 Vienna (Austria). Computer Graphics Forum, 26(1):116, March 2007. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cotting:2007:IVV

[CG07b] Daniel Cotting and Markus Gross. Interaction, VR, and visualization: Interactive visual workspaces with dynamic foveal areas and adaptive composite interfaces. Computer Graphics Forum, 26(3): 685–694, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:2003:BBD

[CGG^{+}03a] P. Cignoni, F. Ganovelli, E. Gobbetti, F. Marton, F. Ponchio, and R. Scopigno. BDAM — Batched Dynamic Adaptive Meshes for high performance terrain visualization. Computer Graphics Forum, 22(3):505–514, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:2003:MBB

[CGG^{+}03b] P. Cignoni, F. Ganovelli, E. Gobbetti, F. Marton, F. Ponchio, and R. Scopigno. Meshes: BDAM — batched dynamic adaptive meshes for high performance terrain visualization. Computer Graphics Forum, 22(3):505–514, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Courbet:2009:HCR

[CH09] Clement Courbet and Celine Hudelot. Hierarchies and compression: Random accessible hierarchical mesh compression for interactive visualization. Computer Graphics Forum, 28(5):1311–1318, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:1997:MSL

[Che97] Jim X. Chen. Multiple segment line scan-conversion. Computer Graphics Forum, 16(5):257–268, December 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=193&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Chen:2006:REW

[Che06] Baoquan Chen. Reports: 2nd Eurographics Workshop on Sketch-Based Interfaces and Modeling Trinity College Dublin, Ireland, August 28–29, 2005 http://www.eg.org/sbm/. Computer Graphics Forum, 25(1):139–140, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chryssafis:1986:AAC

[Chr86] A. Chryssafis. Anti-aliasing of computer-generated images: a picture independent approach. Computer Graphics Forum, 5(2):125– 129, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cook:1990:PRB

[CJ90] Anthony C. Cook and Christopher B. Jones. A Prolog rule-based system for cartographic name placement. Computer Graphics Forum, 9(2):109–126, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2009:DSC

[CJC^{+}09] Myung Geol Choi, Eunjung Ju, Jung-Woo Chang, Jehee Lee, and Young J. Kim. Data structure and collision detection: Linkless octree using multi-level perfect hashing. Computer Graphics Forum, 28(7):1773–1780, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2006:RSS

[CJW^{+}06] Xiao-Hua Cai, Yun-Tao Jia, Xi Wang, Shi-Min Hu, and Ralph R. Martin. Rendering soft shadows using multilayered shadow fins. Computer Graphics Forum, 25(1):15–28, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cline:2009:IPN

[CJW^{+}09] D. Cline, S. Jeschke, K. White, A. Razdan, and P. Wonka. Important points and NPR: Dart throwing on surfaces. Computer Graphics Forum, 28(4):1217–1226, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chiyokura:1984:NSI

[CK84] H. Chiyokura and F. Kimura. A new surface interpolation method for irregular curve models. Computer Graphics Forum, 3(3): 209–218, September 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chaudhuri:2004:SAR

[CKB04] Parag Chaudhuri, Prem Kalra, and Subhashis Banerjee. Stylized animation and rendering: a system for view-dependent animation. Computer Graphics Forum, 23(3):411–420, September 2004. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Comba:1999:FPC

[CKM^{+}99] João Comba, James T. Klosowski, Nelson Max, Joseph S. B. Mitchell, Cláudio T. Silva, and Peter L. Williams. Fast polyhedral cell sorting for interactive rendering of unstructured grids. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=357&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Cuntz:2008:FAG

[CKSW08] Nicolas Cuntz, Andreas Kolb, Robert Strzodka, and Daniel Weiskopf. Flow advection and geometry processing: Particle level set advection for the interactive visualization of unsteady 3D flow. Computer Graphics Forum, 27(3):719–726, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:1992:MPE

[CL92] X. Chen and P. Lienhardt. Modelling and programming evolutions of surfaces. Computer Graphics Forum, 11(5):323–341, December 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cortelazzo:1999:NMI

[CL99] G. M. Cortelazzo and L. Lucchese. A new method of image mosaicking and its application to cultural heritage representation. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=347&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Chiang:2003:MPS

[CL03a] Yi-Jen Chiang and Xiang Lu. Meshes: Progressive simplification of tetrahedral meshes preserving all isosurface topologies. Computer Graphics Forum, 22(3):493–504, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chiang:2003:PST

[CL03b] Yi-Jen Chiang and Xiang Lu. Progressive simplification of tetrahedral meshes preserving all isosurface topologies. Computer Graphics Forum, 22(3):493–504, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Claussen:1988:SEW

[Cla88] U. Claussen. The second Eurographics workshop on graphics hardware. Computer Graphics Forum, 7(1):61, March 1988. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Claussen:1989:RFE

[Cla89] U. Claussen. Report on the Fourth EUROGRAPHICS workshop on graphics hardware. Computer Graphics Forum, 8(4):367–368, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clark:1992:IMU

[Cla92] D. R. Clark. InterHyper MultiOOP using ClothesPEG: an assessment of the multimedia morass. Computer Graphics Forum, 11(3):C461–??, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chuang:2009:DEL

[CLB^{+}09] Ming Chuang, Linjie Luo, Benedict J. Brown, Szymon Rusinkiewicz, and Michael Kazhdan. Differentials: Estimating the Laplace– Beltrami operator by restricting 3D functions. Computer Graphics Forum, 28(5):1475–1484, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:1998:CSE

[cLCtLL98] Tain chi Lu, Chuanwen Chiang, Ming tang Lin, and Chungnan Lee. A collaborative scene editor for VRML worlds. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=253&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Cabral:2009:ICT

[CLDD09] Marcio Cabral, Sylvain Lefebvre, Carsten Dachsbacher, and George Drettakis. Interactive cities and terrain: Structurepreserving reshape for textured architectural scenes. Computer Graphics Forum, 28(2):469–480, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2007:IVV

[CLE07] Wei Chen, Aidong Lu, and David S. Ebert. Interaction, VR, and visualization: Shape-aware volume illustration. Computer Graphics Forum, 26(3):705–714, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Christensen:2003:PRD

[CLF^{+}03a] Per H. Christensen, David M. Laur, Julia Fong, Wayne L. Wooten, and Dana Batali. Photorealism: Ray differentials and multiresolution geometry caching for distribution ray tracing in complex scenes. Computer Graphics Forum, 22(3):543–552, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Christensen:2003:RDM

[CLF^{+}03b] Per H. Christensen, David M. Laur, Julia Fong, Wayne L. Wooten, and Dana Batali. Ray differentials and multiresolution geometry caching for distribution ray tracing in complex scenes. Computer Graphics Forum, 22(3):543–552, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2008:RSR

[CLH^{+}08] Chih-Wen Chang, Wen-Chieh Lin, Tan-Chi Ho, Tsung-Shian Huang, and Jung-Hong Chuang. Rendering and simplification:

Real-time translucent rendering using GPU-based texture space importance sampling. Computer Graphics Forum, 27(2):517–526, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2008:MIC

[CLHL08] Bing-Yu Chen, Ken-Yi Lee, Wei-Ting Huang, and Jong-Shan Lin. Moving images: Capturing intention-based full-frame video stabilization. Computer Graphics Forum, 27(7):1805–1814, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Castro:2008:SOO

[CLL^{+}08] Rener Castro, Thomas Lewiner, Hélio Lopes, Geovan Tavares, and Alex Bordignon. Statistical optimization of octree searches. Computer Graphics Forum, 27(6):1557–1566, September 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chazal:2009:EDC

[CLM09] F. Chazal, A. Lieutier, and N. Montana. Edges: Discrete critical values: a general framework for silhouettes computation. Computer Graphics Forum, 28(5):1509–1518, July 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2009:SCP

[CLME09] Wei Chen, Xiao Liang, Ross Maciejewski, and David S. Ebert. Shape context preserving deformation of 2D anatomical illustrations. Computer Graphics Forum, 28(1):114–126, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2008:SSP

[CLT^{+}08] Zhongxian Chen, Xiaonan Luo, Le Tan, Binghong Ye, and Jiapeng Chen. Subdivision and segmentation: Progressive interpolation based on Catmull–Clark subdivision surfaces. Computer Graphics Forum, 27(7):1823–1827, October 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Csebfalvi:2001:FVO

[CMH^{+}01] Balázs Csébfalvi, Lukas Mroz, Helwig Hauser, Andreas König, and Eduard Gröler. Fast visualization of object contours by nonphotorealistic volume rendering. Computer Graphics Forum, 20 (3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:1993:PDT

[CMPS93] P. Cignoni, C. Montani, R. Perego, and R. Scopigno. Parallel 3D Delaunay triangulation. Computer Graphics Forum, 12(3):C129– C142, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:1994:MIT

[CMS94] P. Cignoni, C. Montani, and R. Scopigno. MagicSphere: an insight tool for 3D data visualization. Computer Graphics Forum, 13(3): C/317–C/328, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cordier:2002:RTA

[CMT02] Frederic Cordier and Nadia Magnenat-Thalmann. Real-time animation of dressed virtual humans. Computer Graphics Forum, 21(3):327–336, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cordier:2005:DDA

[CMT05] Frederic Cordier and Nadia Magnenat-Thalmann. A data-driven approach for real-time clothes simulation. Computer Graphics Forum, 24(2):173–183, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Christie:2005:ISS

[CN05] Marc Christie and Jean-Marie Normand. Interaction: a semantic space partitioning approach to virtual camera composition. Computer Graphics Forum, 24(3):247–256, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cockrell:1983:NGP

[Coc83] P. R. Cockrell. A new general purpose method for large volume production of contour charts. Computer Graphics Forum, 2(1): 35–47, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen-Or:1995:IAA

[COF95] D. Cohen-Or and S. Fleishman. An incremental alignment algorithm for parallel volume rendering. Computer Graphics Forum, 14(3):C/123–C/133, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen-Or:1998:CVS

[COFHZ98] Daniel Cohen-Or, Gadi Fibich, Dan Halperin, and Eyal Zadicario. Conservative visibility and strong occlusion for viewspace partitioning of densely occluded scenes. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=271&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Cohen:1995:ODS

[Coh95] Michael F. Cohen. Optimization in the design of shape and motion. Computer Graphics Forum, 14(3):C/1–??, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Collins:1993:DVK

[Col93] B. M. Collins. Data visualization: a keyword bibliography. Computer Graphics Forum, 12(1):47–73, March 1993. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Collins:2005:IPK

[Col05] Steve Collins. Invited presentations: Kinematics, dynamics, biomechanics: Evolution of autonomy in game animation. Computer Graphics Forum, 24(3):xx, September 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Christie:2008:SAR

[CON08] Marc Christie, Patrick Olivier, and Jean-Marie Normand. State of the art reviews: Camera control in computer graphics. Computer Graphics Forum, 27(8):2197–2218, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cook:2005:IPM

[Coo05] Rob Cook. Invited presentations: Movie making at Pixar: a collaboration of art and technology. Computer Graphics Forum, 24 (3):xix, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1985:DRF

[Coq85] Sabine Coquillart. Displaying random fields. Computer Graphics Forum, 4(1):11–19, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen-Or:1995:VDZ

[COS95] D. Cohen-Or and A. Shaked. Visibility and dead-zones in digital terrain maps. Computer Graphics Forum, 14(3):C/171–C/180, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cottingham:1985:CDS

[Cot85] Marion S. Cottingham. A compressed data structure for surface representation. Computer Graphics Forum, 4(3):217–228, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carr:1988:RCR

[CP88] R. Carr and J. Pink. Report on the “CGM in the Real World” workshop (NBS, Washington). Computer Graphics Forum, 7(1): 59–60, March 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cockshott:1992:MTP

[CPE92] T. Cockshott, J. Patterson, and D. England. Modelling the texture of paint. Computer Graphics Forum, 11(3):C217–C226, C476, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cuntz:2009:TAL

[CPK09] N. Cuntz, A. Pritzkau, and A. Kolb. Time-adaptive lines for the interactive visualization of unsteady flow data sets. Computer Graphics Forum, 28(8):2165–2175, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choudhury:2008:SVI

[CPP08] A. N. M. Imroz Choudhury, Kristin C. Potter, and Steven G. Parker. Software visualization: Interactive visualization for memory reference traces. Computer Graphics Forum, 27(3):815–822, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Crennell:1988:RTA

[Cre88] K. M. Crennell. Report on the Third Alvey Vision Conference, Cambridge, September 1987. Computer Graphics Forum, 7(2): 135–137, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cignoni:1998:MME

[CRS98] P. Cignoni, C. Rocchini, and R. Scopigno. Metro: Measuring error on simplified surfaces. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=236&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Combremont:1983:ICG

[CRW83] J. J. Combremont, M. Roch, and J. Weber. Interactive computer graphics applied to chemistry: experiences and new developments. Computer Graphics Forum, 2(2-3):89–96, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cline:2009:CTP

[CRW09] D. Cline, A. Razdan, and P. Wonka. A comparison of tabular PDF inversion methods. Computer Graphics Forum, 28(1):154– 160, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chrysanthou:1992:CDC

[CS92] Y. Chrysanthou and M. Slater. Computing dynamic changes to BSP trees. Computer Graphics Forum, 11(3):C321–C332, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen:1993:PRI

[CS93] D. Cohen and A. Shaked. Photo-realistic imaging of digital terrains. Computer Graphics Forum, 12(3):C363–C373, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1997:Ea

[CS97a] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 16(2):??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1997:Eb

[CS97b] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 16(4):??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1997:Ec

[CS97c] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:1998:MIP

[CS98a] Wenli Cai and Georgios Sakas. Maximum intensity projection using splatting in sheared object space. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=259&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Coquillart:1998:Ea

[CS98b] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1998:Eb

[CS98c] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1998:Ec

[CS98d] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:1999:DIM

[CS99a] Wenli Cai and Georgios Sakas. Data intermixing and multivolume rendering. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=356&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=18.

Coquillart:1999:Ea

[CS99b] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1999:Eb

[CS99c] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coquillart:1999:Ec

[CS99d] Sabine Coquillart and Hans-Peter Seidel. Editorial. Computer Graphics Forum, 18(4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2000:PBP

[CS00] Yao-Xun Chang and Zen-Chung Shih. Physically-based patination for underground objects. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=403&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Cha:2009:VPG

[CSI09] Deukhyun Cha, Sungjin Son, and Insung Ihm. Visibility and particles: GPU-assisted high quality particle rendering. Computer Graphics Forum, 28(4):1247–1255, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Castro:2004:FMR

[CSN04] F. Castro, M. Sbert, and L. Neumann. Fast multipath radiosity using hierarchical subscenes. Computer Graphics Forum, 23(1): 43–53, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2000:CVG

[CT00] Min Chen and John V. Tucker. Constructive volume geometry. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2003:MVS

[CTSO03a] Ding-Yun Chen, Xiao-Pei Tian, Yu-Te Shen, and Ming Ouhyoung. Modelling: On visual similarity based 3D model retrieval. Computer Graphics Forum, 22(3):223–232, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2003:VSB

[CTSO03b] Ding-Yun Chen, Xiao-Pei Tian, Yu-Te Shen, and Ming Ouhyoung. On visual similarity based 3D model retrieval. Computer Graphics Forum, 22(3):223–232, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:1992:DEC

[CTW92] M. Chen, P. Townsend, and C. Y. Wang. A development environment for constructing graph-based editing tools. Computer Graphics Forum, 11(3):C345–C355, ???? 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:1999:ALS

[CW99] Jim X. Chen and Xusheng Wang. Approximate line scanconversion and antialiasing. Computer Graphics Forum, 18 (1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=303&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Chica:2008:PSI

[CWA^{+}08] A. Chica, J. Williams, C. Andujar, P. Brunet, I. Navazo, J. Rossignac, and A. Vinacua. Pressing: Smooth isosurfaces with flats from binary grids. Computer Graphics Forum, 27(1):36–46, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cleary:1986:MRT

[CWBV86] John G. Cleary, Brian M. Wyvill, Graham M. Birtwistle, and Reddy Vatti. Multiprocessor ray tracing. Computer Graphics Forum, 5(1):3–12, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2007:DRT

[CWK07] Min Gyu Choi, Seung Yong Woo, and Hyeong-Seok Ko. Deformation: Real-time simulation of thin shells. Computer Graphics Forum, 26(3):349–354, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2000:CVS

[CWKS00] Wenli Cai, Stefan Walter, Grigorios Karangelis, and Georgios Sakas. Collaborative virtual simulation environment for radiotherapy treatment planning. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=430&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Chuang:2009:HPG

[CWM09] Johnson Chuang, Daniel Weiskopf, and Torsten Möller. Human perception and Green graphics: Energy aware color sets. Computer Graphics Forum, 28(2):203–211, April 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:1989:PTA

[CY89] Xiangping Chen and Daoning Ying. Polygon triangulation algorithm as a powerful core processor of PLAN-I. Computer Graphics Forum, 8(3):193–198, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cobzas:2002:DTI

[CYJ02] Dana Cobzas, Keith Yerex, and Martin Jägersand. Dynamic textures for image-based rendering of fine-scale 3D structure and animation of non-rigid motion. Computer Graphics Forum, 21(3): 493–502, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2008:RDS

[CZ08] Will Chang and Matthias Zwicker. Registering deforming surfaces II: Automatic registration for articulated shapes. Computer Graphics Forum, 27(5):1459–1468, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2009:CMR

[CZ09] W. Chang and M. Zwicker. Captured models: Range scan registration using reduced deformable models. Computer Graphics Forum, 28(2):447–456, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2008:VNA

[CZCE08] Weri Chen, Song Zhang, Stephfan Correia, and David S. Ebert. Visualization for neurology: Abstractive representation and exploration of hierarchically clustered diffusion tensor fiber tracts. Computer Graphics Forum, 27(3):1071–1078, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cotting:2005:GHA

[CZGF05] Daniel Cotting, Remo Ziegler, Markus Gross, and Henry Fuchs. GPU & hardware: Adaptive instant displays: Continuously calibrated projections using per-pixel light control. Computer Graphics Forum, 24(3):705–714, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Damnjanovic:1991:FSL

[Dam91] Ljiljana B. Damnjanovic. The formal specification of Level 1a of GKS. Computer Graphics Forum, 10(1):11–25, March 1991. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daniel:1996:UCP

[Dan96] M. Daniel. Using a convex pyramid to bound surface normal vectors. Computer Graphics Forum, 15(4):219–227, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DaSilva:2008:CAS

[DAP08] M. Da Silva, Y. Abe, and J. Popović. Character animation: Simulation of human motion data using short-horizon model-predictive control. Computer Graphics Forum, 27(2):371–380, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daun:1990:REW

[Dau90] S. Daun. Report on the EUROGRAPHICS workshop on relation of PDLs and graphics systems. Computer Graphics Forum, 9(4): 375–376, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Davis:2007:GBL

[Dav07] T. A. Davis. Graphics-based learning in first-year computer science. Computer Graphics Forum, 26(4):737–742, December 2007.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Day:1988:PCH

[Day88] A. M. Day. Planar convex hull algorithms in theory and practice. Computer Graphics Forum, 7(3):177–193, September 1988. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Day:1990:ICH

[Day90] A. M. Day. The implementation of a 2D convex hull algorithm using perturbation. Computer Graphics Forum, 9(4):309–316, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Day:1992:NTD

[Day92] J. D. Day. A new two dimensional line clipping algorithm for small windows. Computer Graphics Forum, 11(4):241–245, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deMey:1992:VCM

[dBD^{+}92] V. de Mey, C. Breiteneder, L. Dami, S. Gibbs, and D. Tsichritzis. Visual composition and multimedia. Computer Graphics Forum, 11(3):C9–C22, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dumont:1999:PAT

[DBG99] Reynald Dumont, Kadi Bouatouch, and Philippe Gosselin. A progressive algorithm for three point transport. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=301&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Duke:2003:NPR

[DBHM03a] D. J. Duke, P. J. Barnard, N. Halper, and M. Mellin. Nonphotorealistic rendering: Rendering and affect. Computer Graphics Forum, 22(3):359–368, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2003:RA

[DBHM03b] D. J. Duke, P. J. Barnard, N. Halper, and M. Mellin. Rendering and affect. Computer Graphics Forum, 22(3):359–368, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deBruin:1993:DGU

[dBv93] H. de Bruin, P. Bouwman, and J. van den Bos. DIGIS-a graphical user interface design environment for nonprogrammers. Computer Graphics Forum, 12(3):C13–C24, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dubois:1987:GAR

[DCG87] J. E. Dubois, P. Cayzergues, and G. Georgoulis. Graphic acquisition and retrieval of stereochemical information. Computer Graphics Forum, 6(4):309–316, December 1987. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dafner:2000:CBS

[DCOM00] Revital Dafner, Daniel Cohen-Or, and Yossi Matias. Contextbased space filling curves. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=413&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Dellepiane:2008:SPE

[DCPS08] M. Dellepiane, M. Callieri, F. Ponchio, and R. Scopigno. Selected papers from the EG 2007 Cultural Heritage Track: Mapping highly detailed colour information on extremely dense 3D models: The case of David’s restoration. Computer Graphics Forum, 27(8):2178–2187, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deCastro:2009:MSF

[dCTAD09] Pedro Machado Manhães de Castro, Jane Tournois, Pierre Alliez, and Olivier Devillers. Meshing-simplex: Filtering relocations on a Delaunay triangulation. Computer Graphics Forum, 28(5):1465– 1474, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deFloriani:1985:ITD

[dD85] Leila de Floriani and Giuliana Dettori. An interpolant with tension defined over triangles. Computer Graphics Forum, 4(4):359– 362, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1992:FSR

[DD92] D. A. Duce and L. B. Damnjanovic. Formal specification in the revision of GKS: an illustrative example. Computer Graphics Forum, 11(1):17–30, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Debattista:2009:ICI

[DDB^{+}09] K. Debattista, P. Dubla, F. Banterle, L. P. Santos, and A. Chalmers. Instant caching for interactive global illumination. Computer Graphics Forum, 28(8):2216–2228, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dubla:2009:AIS

[DDC09] P. Dubla, K. Debattista, and A. Chalmers. Adaptive interleaved sampling for interactive high-fidelity rendering. Computer Graphics Forum, 28(8):2117–2130, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dammertz:2009:RAM

[DDKL09] Sabrina Dammertz, Holger Dammertz, Alexander Keller, and Hendrik P. A. Lensch. Rendering and appearance modeling: Textures on rank-1 lattices. Computer Graphics Forum, 28(7):1945– 1954, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Damez:2003:SAR

[DDM03] Cyrille Damez, Kirill Dmitriev, and Karol Myszkowski. State of the art review: State of the art in global illumination for interactive applications and high-quality animations. Computer Graphics Forum, 22(1):55–77, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeNeve:2000:ACR

[DDPL00] P. De Neve, K. Denecker, W. Philips, and I. Lemahieu. An advanced color representation for lossy compression of CMYK prepress images. Computer Graphics Forum, 19 (1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=384&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Damnjanovic:1993:GSI

[DDR93] L. B. Damnjanovic, D. A. Duce, and S. K. Robinson. GKS-9x: some implementation considerations. Computer Graphics Forum, 12(3):C295–C313, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1994:PIA

[DDtR94] D. A. Duce, D. J. Duke, P. J. W. ten Hagen, and G. J. Reynolds. PREMO — an initial approach to a formal definition. Computer Graphics Forum, 13(3):C/393–C/406, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Denecker:2002:PEL

[DDV^{+}02] M. Denecker, P. De Neve, S. Van Assche, R. Van de Walle, I. Lemahieu, and W. Philips. Psychovisual evaluation of lossy CMYK image compression for printing applications. Computer Graphics Forum, 21(1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deBerg:1997:TDC

[de 97] Mark de Berg. Trends and developments in computational geometry. Computer Graphics Forum, 16(1):3–30, March 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=113&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=16.

Decoret:2005:RIB

[Déc05] Xavier Décoret. Rendering II: N-buffers for efficient depth map query. Computer Graphics Forum, 24(3):393–400, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Denham:1986:HPC

[Den86] Den Denham. High performance colour displays for computer graphics. Computer Graphics Forum, 5(1):57–63, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Denny:2003:HBG

[Den03a] Markus Denny. Hardware-based graphics: Solving geometric optimization problems using graphics hardware. Computer Graphics Forum, 22(3):441–451, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Denny:2003:SGO

[Den03b] Markus Denny. Solving geometric optimization problems using graphics hardware. Computer Graphics Forum, 22(3):441–451, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desbrun:2004:IPA

[Des04] Mathieu Desbrun. Invited presentations: Applied geometry: Discrete differential calculus for graphics. Computer Graphics Forum, 23(3):269, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desbrun:2006:RTE

[Des06] Mathieu Desbrun. Reports: Third Eurographics Symposium on Geometry Processing (in cooperation with ACM SIGGRAPH): Vienna, Austria, July 4–6, 2005. Computer Graphics Forum, 25 (2):257, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1985:BUT

[DF85] D. A. Duce and E. V. C. Fielding. Better understanding through formal specification. Computer Graphics Forum, 4(4):333–348, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Durand:1990:RZB

[DF90] Charles X. Durand and Daniel Faguy. Rational zoom of bit maps using B-spline interpolation in computerized 2-D animation. Computer Graphics Forum, 9(1):27–37, March 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Drettakis:1993:ACR

[DF93] G. Drettakis and E. Fiume. Accurate and consistent reconstruction of illumination functions using structured sampling. Computer Graphics Forum, 12(3):C273–C284, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deFigueiredo:2003:APC

[dFSV03] Luiz Henrique de Figueiredo, Jorge Stolfi, and Luiz Velho. Approximating parametric curves with strip trees using affine arithmetic. Computer Graphics Forum, 22(2):171–179, June 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dischler:1995:GBM

[DG95] J.-M. Dischler and D. Ghazanfarpour. A geometrical based method for highly complex structured textures generation. Computer Graphics Forum, 14(4):203–215, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dischler:1997:PDG

[DG97] J.-M. Dischler and D. Ghazanfarpour. A procedural description of geometric textures by spectral and spatial analysis of profiles. Computer Graphics Forum, 16(3):C129–C141, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=149&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Desbenoit:2004:NPA

[DGA04] Brett Desbenoit, Eric Galin, and Samir Akkouche. Natural phenomena and aging: Simulating and modeling lichen growth. Computer Graphics Forum, 23(3):341–350, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1998:RMD

[DGC^{+}98] D. A. Duce, D. Giorgetti, C. S. Cooper, J. R. Gallop, I. J. Johnson, K. Robinson, and C. D. Seelig. Reference models for distributed cooperative visualization. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=286&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Diaz-Gutierrez:2009:DGP

[DGEG09] P. Diaz-Gutierrez, D. Eppstein, and M. Gopi. Digital geometry processing: Curvature aware fundamental cycles. Computer Graphics Forum, 28(7):2015–2024, October 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dischler:1998:AST

[DGF98] J. M. Dischler, D. Ghazanfarpour, and R. Freydier. Anisotropic solid texture synthesis using orthogonal 2D views. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=256&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Diaz-Gutierrez:2005:GIH

[DGGP05] Pablo Diaz-Gutierrez, M. Gopi, and Renato Pajarola. Geometry II: Hierarchyless simplification, stripification and compression of triangulated two-manifolds. Computer Graphics Forum, 24(3): 457–467, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeGoes:2008:SPL

[DGV08] Fernando De Goes, Siome Goldenstein, and Luiz Velho. Shape properties: Local and global: a hierarchical segmentation of articulated bodies. Computer Graphics Forum, 27(5):1349–1356, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Donikian:1993:DDM

[DH93a] S. Donikian and G. Hegron. A declarative design method for 3D scene sketch modeling. Computer Graphics Forum, 12(3):C223– C236, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:1993:AIO

[DH93b] D. J. Duke and M. D. Harrison. Abstract interaction objects. Computer Graphics Forum, 12(3):C25–C36, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:1998:PPO

[DH98] D. J. Duke and I. Herman. Programming paradigms in an object-oriented multimedia standard. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=288&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Duce:2002:WGF

[DHH02] David Duce, Ivan Herman, and Bob Hopgood. Web 2D graphics file formats. Computer Graphics Forum, 21(1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dammertz:2008:RTS

[DHK08] H. Dammertz, J. Hanika, and A. Keller. Ray tracing 1: Shallow bounding volume hierarchies for fast SIMD ray tracing of incoherent rays. Computer Graphics Forum, 27(4):1225–1233, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deussen:2005:RV

[DHKS05] Co-Chairs Oliver Deussen, Charles Hanson, Daniel Keim, and Dietmar Saupe. Reports: Vissym04. Computer Graphics Forum, 24 (1):132, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Damez:2004:STH

[DHS04] Cyrille Damez, Nicolas Holzschuch, and François X. Sillion. Space–time hierarchical radiosity with clustering and higher-order wavelets. Computer Graphics Forum, 23(2):129–141, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deussen:2000:FPM

[DHvOS00] Oliver Deussen, Stefan Hiller, Cornelius van Overveld, and Thomas Strothotte. Floating points: a method for computing stipple drawings. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=396&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Diamond:1984:ACG

[Dia84] R. Diamond. Applications of computer graphics in molecular biology. Computer Graphics Forum, 3(1):3–11, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1988:EPE

[DJ88] David A. Duce and Pierre Jancene, editors. Eurographics ’88: proceedings of the European Computer Graphics Conference and Exhibition, Nice, France, 12–16 September, 1988. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1988. ISBN 0-444-70499-X. LCCN T 385 E97 1988.

Decaudin:2006:MVG

[DJW^{+}06] Philippe Decaudin, Dan Julius, Jamie Wither, Laurence Boissieux, Alla Sheffer, and Marie-Paule Cani. Modelling I: Virtual garments: a fully geometric approach for clothing design. Computer Graphics Forum, 25(3):625–634, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dong:2009:IBM

[DJZ^{+}09] Zilong Dong, Lei Jiang, Guofeng Zhang, Qing Wang, and Hujun Bao. Image-based modeling and rendering: Live video Montage with a rotating camera. Computer Graphics Forum, 28(7):1745– 1753, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dong:2000:FVR

[DKC00] Feng Dong, Meleagros Krokos, and Gordon Clapworthy. Fast volume rendering and data classification using multiresolution in min-max octrees. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=428&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Dobashi:1994:SIL

[DKN^{+}94] Y. Dobashi, K. Kaneda, T. Nakashima, H. Yamashita, T. Nishita, and K. Tadamura. Skylight for interior lighting design. Computer Graphics Forum, 13(3):C/85–C/96, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dobashi:1995:QRM

[DKN^{+}95] Y. Dobashi, K. Kaneda, H. Nakatani, H. Yamashita, and T. Nishita. A quick rendering method using basis functions for interactive lighting design. Computer Graphics Forum, 14(3): C/229–C/240, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Doggett:2001:DMU

[DKS01] Michael Doggett, Anders Kugler, and Wolfgang Straßer. Displacement mapping using scan conversion hardware architectures. Computer Graphics Forum, 20(1):??, March 2001. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dixon:1994:DSA

[DKW94a] A. R. Dixon, G. H. Kirby, and D. P. M. Wills. A data structure for artificial terrain generation. Computer Graphics Forum, 13(1): 37–47, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dixon:1994:TCD

[DKW94b] A. R. Dixon, G. H. Kirby, and D. P. M. Wills. Towards context dependent interpolation of digital elevation models. Computer Graphics Forum, 13(3):C/23–C/32, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Durikovic:1998:ABO

[DKY98] Roman Durikovic, Kazufumi Kaneda, and Hideo Yamashita. Animation of biological organ growth based on L-systems. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=248&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Dobashi:1996:MCS

[DKYN96] Yoshinori Dobashi, Kazufumi Kaneda, Hideo Yamashita, and Tomoyuki Nishita. Method for calculation of sky light luminance aiming at an interactive architectural design. Computer Graphics Forum, 15(3):C109–C118, C465–C466, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Doggett:2004:RGH

[DL04] Michael Doggett and Anselmo Lastra. Reports: Graphics Hardware 2004: Grenoble, France, 29–30 August 2004. Computer Graphics Forum, 23(4):845, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dong:2005:CPI

[DLC05] Feng Dong, Hai Lin, and Gordon Clapworthy. Cutting and pasting irregularly shaped patches for texture synthesis. Computer Graphics Forum, 24(1):17–26, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dey:2009:RRS

[DLRW09] Tamal K. Dey, Kuiyu Li, Edgar A. Ramos, and Rephael Wenger. Reconstruction — regular sampling: Isotopic reconstruction of surfaces with boundaries. Computer Graphics Forum, 28(5):1371– 1382, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dong:2008:TLS

[DLTD08] Yue Dong, Sylvain Lefebvre, Xin Tong, and George Drettakis. Textures: Lazy solid texture synthesis. Computer Graphics Forum, 27(4):1165–1174, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dulli:1992:GPE

[DM92] S. Dulli and V. Milanese. A graphic programming environment based on KA Delta MO Sigma. Computer Graphics Forum, 11(1): 3–16, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desbrun:2002:IPS

[DMA02] Mathieu Desbrun, Mark Meyer, and Pierre Alliez. Intrinsic parameterizations of surface meshes. Computer Graphics Forum, 21 (3):209–218, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Drago:2003:ALM

[DMAC03a] F. Drago, K. Myszkowski, T. Annen, and N. Chiba. Adaptive logarithmic mapping for displaying high contrast scenes. Computer Graphics Forum, 22(3):419–426, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Drago:2003:IQA

[DMAC03b] F. Drago, K. Myszkowski, T. Annen, and N. Chiba. Image quantization: Adaptive logarithmic mapping for displaying high contrast scenes. Computer Graphics Forum, 22(3):419–426, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deville:1994:SDM

[DMCP94] P. M. Deville, S. Merzouk, D. Cazier, and J. C. Paul. Spectral data modeling for a lighting application. Computer Graphics Forum, 13 (3):C/97–C/106, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Didyk:2008:CIE

[DMHS08] P. Didyk, R. Mantiuk, M. Hein, and H. P. Seidel. Computational imaging: Enhancement of bright video features for HDR displays. Computer Graphics Forum, 27(4):1265–1274, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeHaan:2007:IVV

[DMKP07] Gerwin De Haan, René Molenaar, Michal Koutek, and Frits H. Post. Interaction, VR, and visualization: Consistent viewing and interaction for multiple users in projection-based VR systems. Computer Graphics Forum, 26(3):695–704, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dischler:2002:TP

[DMLG02] Jean-Michel Dischler, Karl Maritaud, Bruno Lévy, and Djamchid Ghazanfarpour. Texture particles. Computer Graphics Forum, 21 (3):401–410, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeFloriani:1993:ECL

[DMP93] L. De Floriani, D. Mirra, and E. Puppo. Extracting contour lines from a hierarchical surface model. Computer Graphics Forum, 12 (3):C249–C260, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dobashi:2008:FSF

[DMYN08] Yoshinori Dobashi, Yasuhiro Matsuda, Tsuyoshi Yamamoto, and Tomoyuki Nishita. Fluids and simulation: a fast simulation method using overlapping grids for interactions between smoke and rigid objects. Computer Graphics Forum, 27(2):477–486, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2008:ESA

[DN08] Z. Deng and U. Neumann. Expressive speech animation synthesis with phoneme-level controls. Computer Graphics Forum, 27(8): 2096–2113, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Decaudin:2009:VB

[DN09] Philippe Decaudin and Fabrice Neyret. Volumetric billboards. Computer Graphics Forum, 28(8):2079–2089, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dingliana:2000:GDC

[DO00] John Dingliana and Carol O’Sullivan. Graceful degradation of collision handling in physically based animation. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=416&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

DellaVentura:1993:PES

[DOS93] A. Della Ventura, P. Ongaro, and R. Schettini. Pictorial editing by shape matching techniques. Computer Graphics Forum, 12(2): 111–122, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1993:FSG

[DP93] D. A. Duce and F. Paterno. A formal specification of a graphics system in the framework of the Computer Graphics Reference Model. Computer Graphics Forum, 12(1):3–20, March 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dellepiane:2008:SFK

[DPT^{+}08] M. Dellepiane, N. Pietroni, N. Tsingos, M. Asselot, and R. Scopigno. Skull, face, and knits: Reconstructing head models from photographs for individualized 3D-audio processing. Computer Graphics Forum, 27(7):1719–1727, October 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2000:DMI

[DQ00] Haixia Du and Hong Qin. Direct manipulation and interactive sculpting of PDE surfaces. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=418&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Densmore:1987:UIT

[DR87] Owen M. Densmore and David S. H. Rosenthal. A user-interface toolkit in object-oriented PostScript. Computer Graphics Forum, 6(3):171–179, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diener:2009:AAT

[DRBR09] Julien Diener, Mathieu Rodriguez, Lionel Baboud, and Lionel Reveret. All about trees: Wind projection basis for real-time animation of trees. Computer Graphics Forum, 28(2):533–540, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Drettakis:2007:ROT

[Dre07] George Drettakis. Reports: Outstanding technical contributions award 2007. Computer Graphics Forum, 26(4):855, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dyken:2008:RTG

[DRS08] C. Dyken, M. Reimers, and J. Seland. Real-time GPU silhouette refinement using adaptively blended Bézier patches. Computer Graphics Forum, 27(1):1–12, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dyken:2009:SUA

[DRS09] C. Dyken, M. Reimers, and J. Seland. Semi-uniform adaptive patch tessellation. Computer Graphics Forum, 28(8):2255–2263, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deFigueiredo:1996:AEI

[dS96] Luiz Henrique de Figueiredo and Jorge Stolfi. Adaptive enumeration of implicit surfaces with affine arithmetic. Computer Graphics Forum, 15(5):287–296, December 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dorsey:1998:EWR

[DS98] Julie Dorsey and Philipp Slusallek. 8th Eurographics Workshop on Rendering. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=225&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=1&amp; src=ard&amp; vid=17.

Duke:2001:E

[DS01] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daubert:2002:HBV

[DS02a] Katja Daubert and Hans-Peter Seidel. Hardware-based volumetric knit-wear. Computer Graphics Forum, 21(3):575–584, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2002:Ea

[DS02b] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 21(1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2002:Eb

[DS02c] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 21(2):98, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2002:Ec

[DS02d] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 21(4):659, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2003:E

[DS03] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 22(1):3–4, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2004:E

[DS04a] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 23(4):i, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2004:ECG

[DS04b] David Duke and Roberto Scopigno. Editorial: Computer Graphics Forum. Computer Graphics Forum, 23(2):128, June 2004. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diakopoulos:2005:AAL

[DS05a] Nicholas A. Diakopoulos and Peter D. Stephenson. Anti-aliased lines using run-masks. Computer Graphics Forum, 24(2):165–172, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2005:Ea

[DS05b] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 24(2):i, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2005:Eb

[DS05c] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 24(4):i, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2006:Ea

[DS06a] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 25(1):i, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2006:Eb

[DS06b] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 25(2):i, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2006:Ec

[DS06c] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 25(4):i, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2007:Ec

[DS07a] David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 26(4):i, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2007:Ea

[DS07b] Editors David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 26(1):i, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:2007:Eb

[DS07c] Editors David Duke and Roberto Scopigno. Editorial. Computer Graphics Forum, 26(2):i, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Domik:2009:REP

[DS09] G. Domik and R. Scateni. Reports: Education Programme at Eurographics 2009. Computer Graphics Forum, 28(6):1723–1724, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dai:1995:SFT

[DSC95] Wen-Kai Dai, Zen-Chung Shih, and Ruei-Chuan Chang. Synthesizing feather textures in Galliformes. Computer Graphics Forum, 14(3):C/407–C/420, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daniels:2009:MNSb

[DSC09a] Joel Daniels II, Claudio T. Silva, and Elaine Cohen. Meshing — non-simplex: Localized quadrilateral coarsening. Computer Graphics Forum, 28(5):1437–1444, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daniels:2009:MNSa

[DSC09b] Joel Daniels II, Claudio T. Silva, and Elaine Cohen. Meshing — non-simplex: Semi-regular quadrilateral-only remeshing from simplified base domains. Computer Graphics Forum, 28(5):1427– 1435, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Davison:1990:GOM

[DSS90] Allan Davison, Mel Slater, and Mark Smith. Graphics object management in the X Window System. Computer Graphics Forum, 9(2):93–99, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Decoret:1999:MLI

[DSSD99] Xavier Decoret, François Sillion, Gernot Schaufler, and Julie Dorsey. Multi-layered impostors for accelerated rendering. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=328&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Dick:2009:EGC

[DSW09] Christian Dick, Jens Schneider, and Rüdiger Westermann. Efficient geometry compression for GPU-based decoding in realtime terrain rendering. Computer Graphics Forum, 28(1):67–83, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:2004:REG

[DT04] David Duce and Juan Carlos Torres. Reports: Eurographics 2003: Granada, Spain, 1–5 September 2003. Computer Graphics Forum, 23(1):117–118, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Day:1994:IVM

[DTA94] A. M. Day, S. M. Turner, and D. B. Arnold. Improved visualisation of marine data: Algae and acid rain. Computer Graphics Forum, 13(3):C/361–C/370, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desbrun:1996:ASI

[DTG96] Mathieu Desbrun, Nicolas Tsingos, and Marie-Paule Gascuel. Adaptive sampling of implicit surfaces for interactive modelling and animation. Computer Graphics Forum, 15(5):319–325, December 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daldegan:1993:ISM

[DTKT93] A. Daldegan, N. M. Thalmann, T. Kurihara, and D. Thalmann. An integrated system for modeling, animating and rendering hair. Computer Graphics Forum, 12(3):C211–C221, ???? 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeAguiar:2008:CAA

[DTTS08] Edilson De Aguiar, Christian Theobalt, Sebastian Thrun, and Hans-Peter Seidel. Character animation: Automatic conversion of mesh animations into skeleton-based animations. Computer Graphics Forum, 27(2):389–397, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1982:EWF

[Duc82a] D. A. Duce. The EEC workshop on formal specification of graphics software standards. Computer Graphics Forum, 1(3):92–95, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1982:GUG

[Duc82b] D. A. Duce. Gino users’ group meeting. Computer Graphics Forum, 1(2):42–46, June 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1985:CSU

[Duc85] D. A. Duce. Concerning the specification of user interfaces. Computer Graphics Forum, 4(3):251–258, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1986:TWD

[Duc86] D. A. Duce. Teaching workshop — December 1985. Computer Graphics Forum, 5(4):349–352, December 1986. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1989:EUC

[Duc89] D. A. Duce. EUROGRAPHICS UK Chapter Conference — CGM Demonstration. Computer Graphics Forum, 8(3):266, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1990:REW

[Duc90a] D. A. Duce. Report on the EUROGRAPHICS workshop on user interface management systems. Computer Graphics Forum, 9(4): 381–384, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1990:Y

[Duc90b] D. A. Duce. Yugraph ’90. Computer Graphics Forum, 9(3): 275–276, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1991:REW

[Duc91] D. A. Duce. Report on the EUROGRAPHICS workshop on formal methods in computer graphics. Computer Graphics Forum, 10(4): 359–361, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1998:NEW

[Duc98] D. A. Duce. Ninth Eurographics Workshop on Visualization in Scientific Computing, Blaubeuren, Germany, April 20–22 1998. Computer Graphics Forum, 17(2):??, June 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=239&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=2&amp; src=
ard&amp; vid=17.

Duce:2000:CVR

[Duc00] David Duce. Carla Vandoni — a remarkable lady. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:2006:OMJ

[Duc06a] David Duce. Obituaries: In memoriam: Jürgen Schönhut. Computer Graphics Forum, 25(1):3–4, March 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:2006:RJL

[Duc06b] David Duce. Reports: The John Lansdown Award 2006. Computer Graphics Forum, 25(4):842–843, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:2007:RJL

[Duc07] David Duce. Reports: The John Lansdown Award 2007. Computer Graphics Forum, 26(4):858–859, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dufourd:1988:CIP

[Duf88] Jean-François F. Dufourd. Construction of interactive programs in computer graphics. Computer Graphics Forum, 7(3):161–176, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duke:1995:RAG

[Duk95] D. J. Duke. Reasoning about gestural interaction. Computer Graphics Forum, 14(3):C/55–C/66, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dunha:1991:PCG

[Dun91] João Duarte Dunha. 3rd Portuguese Computer Graphics Conference. Computer Graphics Forum, 10(1):75, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Durikovic:2001:ASB

[Dur01] Roman Durikovic. Animation of soap bubble dynamics, cluster formation and collision. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dutre:2004:IPI

[Dut04] Philip Dutré. Invited presentations: Image-based relighting, and computer vision. Computer Graphics Forum, 23(3):270, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duce:1990:AHI

[Dvt90] D. A. Duce, R. van Liere, and P. J. W. ten Hagen. An approach to hierarchical input devices. Computer Graphics Forum, 9(1): 15–26, March 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diepstraten:2002:TIT

[DWE02] Joachim Diepstraten, Daniel Weiskopf, and Thomas Ertl. Transparency in interactive technical illustrations. Computer Graphics Forum, 21(3):317–326, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diepstraten:2003:ICI

[DWE03a] J. Diepstraten, D. Weiskopf, and T. Ertl. Interactive cutaway illustrations. Computer Graphics Forum, 22(3):523–532, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diepstraten:2003:NPR

[DWE03b] J. Diepstraten, D. Weiskopf, and T. Ertl. Non-photorealistic rendering and painting: Interactive cutaway illustrations. Computer Graphics Forum, 22(3):523–532, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeToledo:2008:GTA

[DWL08] Rodrigo De Toledo, Bin Wang, and Bruno Lévy. Geometry textures and applications. Computer Graphics Forum, 27(8):2053– 2065, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dai:2009:RAM

[DWL^{+}09] Qiang Dai, Jiaping Wang, Yiming Liu, John Snyder, Enhua Wu, and Baining Guo. Rendering and appearance modeling: The dualmicrofacet model for capturing thin transparent slabs. Computer Graphics Forum, 28(7):1917–1925, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dwyer:2009:GVS

[Dwy09] Tim Dwyer. Graph visualization: Scalable, versatile and simple constrained graph layout. Computer Graphics Forum, 28(3):991– 998, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dobashi:2004:SHS

[DYN04] Yoshinori Dobashi, Tsuyoshi Yamamoto, and Tomoyuki Nishita. Simulation with haptics and sound: Synthesizing sound from turbulent field using sound textures for interactive fluid simulation. Computer Graphics Forum, 23(3):539–545, September 2004. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dyer:2008:SRS

[DZM08] Ramsay Dyer, Hao Zhang, and Torsten Möller. Sampling and reconstructing surfaces: Surface sampling and the intrinsic Voronoi diagram. Computer Graphics Forum, 27(5):1393–1402, July 2008.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dyken:2008:HSM

[DZTS08] Christopher Dyken, Gernot Ziegler, Christian Theobalt, and Hans-Peter Seidel. High-speed marching cubes using HistoPyramids. Computer Graphics Forum, 27(8):2028–2039, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Escriva:2009:ARM

[EBA^{+}09] M. Escrivá, J. Blasco, F. Abad, E. Camahort, and R. Vivó. Autostereoscopic rendering of multiple light fields. Computer Graphics Forum, 28(8):2057–2067, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Emering:2000:VTH

[EBMT00] Luc Emering, Ronan Boulic, Tom Molet, and Daniel Thalmann. Versatile tuning of humanoid agent activity. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Encarnacao:1999:TSV

[EBSC99] L. M. Encarnação, O. Bimber, D. Schmalstieg, and S. D. Chandler. A translucent sketchpad for the virtual table exploring motion-based gesture recognition. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=348&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Esteve:2001:MAC

[EBV01] Jordi Esteve, Pere Brunet, and Alvar Vinacua. Multiresolution for algebraic curves and surfaces using wavelets. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Esteve:2005:AVD

[EBV05] Jordi Esteve, Pere Brunet, and Àlvar Vinacua. Approximation of a variable density cloud of points by shrinking a discrete membrane. Computer Graphics Forum, 24(4):791–807, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2007:EEB

[ED07a] E. Eisemann and X. Décoret. On exact error bounds for viewdependent simplification. Computer Graphics Forum, 26(2):202– 213, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2007:SVV

[ED07b] Elmar Eisemann and Xavier Décoret. Shadows and visibility: Visibility sampling on GPU and applications. Computer Graphics Forum, 26(3):535–544, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2008:OTP

[ED08] E. Eisemann and X. Décoret. Occlusion textures for plausible soft shadows. Computer Graphics Forum, 27(1):13–23, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Edelsbrunner:2006:IPP

[Ede06] Herbert Edelsbrunner. Invited presentations: a primer on topological persistence. Computer Graphics Forum, 25(3):xvii, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Edmonds:1983:IHB

[Edm83] E. A. Edmonds. Image handling — the bringing together of computer graphics and image processing. Computer Graphics Forum, 2(1):30–34, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2008:TAF

[EDM^{+}08] M. Eisemann, B. De Decker, M. Magnor, P. Bekaert, E. De Aguiar, N. Ahmed, C. Theobalt, and A. Sellent. Textures and arrangements: Floating textures. Computer Graphics Forum, 27 (2):409–418, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Encarnacao:1996:PUT

[EFGS96] L. Miguel Encarnacao, Jürgen Fechter, Thomas Grunert, and Wolfgang Strasser. A platform for user-tailored interaction development in 2D, 3D, and VR. Computer Graphics Forum, 15(3):

C433–C442, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eppstein:2008:PQT

[EGKT08] David Eppstein, Michael T. Goodrich, Ethan Kim, and Rasmus Tamstorf. Parameterization and quad tiling: Motorcycle graphs: Canonical quad mesh partitioning. Computer Graphics Forum, 27(5):1477–1486, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ershov:2001:RPA

[EKM01] Sergey Ershov, Konstantin Kolchin, and Karol Myszkowski. Rendering pearlescent appearance based on paint-composition modelling. Computer Graphics Forum, 20(3):??, September 2001. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ehmann:2001:AFP

[EL01] Stephen A. Ehmann and Ming C. Lin. Accurate and fast proximity queries between polyhedra using convex surface decomposition. Computer Graphics Forum, 20(3):??, September 2001. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elber:1995:SNC

[Elb95] Gershon Elber. Symbolic and numeric computation in curve interrogation. Computer Graphics Forum, 14(1):25–34, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elber:1999:ILA

[Elb99] Gershon Elber. Interactive line art rendering of freeform surfaces. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=322&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Encarnacao:1983:VIG

[ELM^{+}83] J. L. Encarnacao, R. Lindner, M. Mehl, G. Pfaff, and W. Strasser. A VLSI implementation of the graphics standard GKS. Computer Graphics Forum, 2(2-3):115–121, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisenacher:2008:TAT

[ELS08] C. Eisenacher, S. Lefebvre, and M. Stamminger. Textures and arrangements: Texture synthesis from photographs. Computer Graphics Forum, 27(2):419–428, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Entezari:2009:GRQ

[EMK09] A. Entezari, M. Mirzargar, and L. Kalantari. Geometry & reconstruction: Quasi-interpolation on the body centered cubic lattice. Computer Graphics Forum, 28(3):1015–1022, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Encarnacao:1981:EPI

[Enc81] Jose Luis Encarnacao, editor. EUROGRAPHICS ’81: proceedings of the international conference and exhibition, Technische Hochschule Darmstadt, F.R. of Germany, 9–11 September 1981. Elsevier North-Holland, Inc., New York, NY, USA, 1981. ISBN 0-444-86284-6. LCCN T 385 E97 1981. Proceedings of the Second Eurographics Conference.

Encarnacao:1982:ACS

[Enc82] J. Encarnacao. An approach for CAD-System simulation. Computer Graphics Forum, 1(3):116–123, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Encarnacao:1998:ARO

[Enc98] José Encarnação. Anno 2010 — remembering our future: Challenges and frontiers of human-media technology as the kernel for human-centered computing. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=247&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Enderle:1982:SGM

[End82] G. Enderle. Standardization of graphics metafiles. Computer Graphics Forum, 1(2):50–57, June 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Enderle:1983:GIO

[End83a] G. Enderle. GKS-implementations. overview and inquiry. Computer Graphics Forum, 2(1):15–22, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Enderle:1983:SR

[End83b] G. Enderle. SIGGRAPH ’83 report. Computer Graphics Forum, 2(4):193–195, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Enderle:1984:GIO

[End84a] G. Enderle. GKS implementations overview — second edition. Computer Graphics Forum, 3(2):181–189, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Enderle:1984:IUA

[End84b] G. Enderle. The interface of the UIMS to the application — working group report. Computer Graphics Forum, 3(2):175–179, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Enderle:1984:SWU

[End84c] G. Enderle. Seeheim workshop on user interface management systems — first report. Computer Graphics Forum, 3(2):169–170, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eigensatz:2009:DPM

[EP09] Michael Eigensatz and Mark Pauly. Deformations: Positional, metric, and curvature control for constraint-based surface deformation. Computer Graphics Forum, 28(2):551–558, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ertl:2002:IVP

[Ert02] Thomas Ertl. Interactive visualization with programmable graphics hardware. Computer Graphics Forum, 21(3):??, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endl:1994:CRG

[ES94] R. Endl and M. Sommer. Classification of ray-generators in uniform subdivisions and octrees for ray tracing. Computer Graphics Forum, 13(1):3–19, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

El-Sana:2000:EMV

[ESC00] Jihad El-Sana and Yi-Jen Chiang. External memory viewdependent simplification. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=406&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Eckstein:2001:TMH

[ESG01] Ilya Eckstein, Vitaly Surazhsky, and Craig Gotsman. Texture mapping with hard constraints. Computer Graphics Forum, 20 (3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Evers-Senne:2003:IBI

[ESK03a] J.-F. Evers-Senne and R. Koch. Image based interactive rendering with view dependent geometry. Computer Graphics Forum, 22(3): 573–582, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Evers-Senne:2003:PEI

[ESK03b] J.-F. Evers-Senne and R. Koch. Practice and experience: Image based interactive rendering with view dependent geometry. Computer Graphics Forum, 22(3):573–582, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Essa:1992:UAP

[ESP92] I. A. Essa, S. Sclaroff, and A. Pentland. A unified approach for physical and geometric modeling for graphics and animation. Computer Graphics Forum, 11(3):C129–C138, C470–C471, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eigensatz:2008:GPC

[ESP08] Michael Eigensatz, Robert W. Sumner, and Mark Pauly. Geometry processing: Curvature-domain shape processing. Computer Graphics Forum, 27(2):241–250, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

El-Sana:1999:GVD

[ESV99] Jihad El-Sana and Amitabh Varshney. Generalized viewdependent simplification. Computer Graphics Forum, 18 (3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=330&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Eisemann:2008:ARS

[EWHS08] Elmar Eisemann, Holger Winnemöller, John C. Hart, and David Salesin. Artistic rendering: Stylized vector art from 3D models with region support. Computer Graphics Forum, 27(4):1199–1207, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

ElOuafdi:2008:SPL

[EZK08] A. F. El Ouafdi, D. Ziou, and H. Krim. Shape properties: Local and global: a smart stochastic approach for manifolds smoothing. Computer Graphics Forum, 27(5):1357–1364, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Franklin:1987:SEH

[FA87] W. R. Franklin and V. Akman. A simple and efficient haloed line algorithm for hidden line elimination. Computer Graphics Forum, 6(2):103–109, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fahlander:1985:MPS

[Fah85] Olov Fahlander. Moving picture synthesis at Linköping University. Computer Graphics Forum, 4(4):371–374, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2007:EDS

[FAT07] Hongbo Fu, Oscar Kin-Chung Au, and Chiew-Lan Tai. Effective derivation of similarity transformations for implicit Laplacian mesh editing. Computer Graphics Forum, 26(1):34–45, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Faure:1996:EBA

[Fau96] François Faure. An energy-based approach for contact force computation. Computer Graphics Forum, 15(3):C357–C366, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Finkbeiner:2009:GRH

[FAVM09] Bernhard Finkbeiner, Usman R. Alim, Dimitri Van De Ville, and Torsten Möller. Geometry & reconstruction: High-quality volumetric reconstruction on optimal lattices for computed tomography. Computer Graphics Forum, 28(3):1023–1030, June 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fournier:1994:CPB

[FB94] A. Fournier and J. Buchanan. Chebyshev polynomials for boxing and intersections of parametric curves and surfaces. Computer Graphics Forum, 13(3):C/127–C/142, ???? 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Forest:2009:SST

[FBGP09] Vincent Forest, Loïc Barthe, Gaël Guennebaud, and Mathias Paulin. Shadows: Soft textured shadow volume. Computer Graphics Forum, 28(4):1111–1120, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Forest:2008:SLA

[FBP08] Vincent Forest, Loïc Barthe, and Mathias Paulin. Shadows and lighting: Accurate shadows by depth complexity sampling. Computer Graphics Forum, 27(2):663–674, April 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Farenc:1999:IED

[FBT99] Nathalie Farenc, Ronan Boulic, and Daniel Thalmann. An informed environment dedicated to the simulation of virtual humans in urban context. Computer Graphics Forum, 18 (3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=351&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Felkel:2001:ICW

[FBW01] Petr Felkel, Mario Bruckschwaiger, and Rainer Wegenkittl. Implementation and complexity of the watershed-from-markers algorithm computed as a minimal cost forest. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferley:1997:SRB

[FCGA97] Eric Ferley, Marie-Paule Cani-Gascuel, and Dominique Attali. Skeletal reconstruction of branching shapes. Computer Graphics Forum, 16(5):283–293, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=195&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=5&amp; src=ard&amp; vid=16.

Fei:2002:ASS

[FCGW02] Guangzheng Fei, Kangying Cai, Baining Guo, and Enhua Wu. An adaptive sampling scheme for out-of-core simplification. Computer Graphics Forum, 21(2):111–119, June 2002. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2006:PIB

[FCH^{+}06] Shaohua Fan, Stephen Chenney, Bo Hu, Kam-Wah Tsui, and Yu chi Lai. Photorealistic and image based rendering: Optimizing control variate estimators for rendering. Computer Graphics Forum, 25(3):351–357, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fleishman:2000:ACP

[FCOL00] Shachar Fleishman, Daniel Cohen-Or, and Dani Lischinski. Automatic camera placement for image-based modeling. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Flores:1990:UGC

[FCP^{+}90] J. Zamorano Flores, F. Pérez Costoya, J. Carretero Pérez, L. M. Gomez Henriquez, and A. Pérez Ambite. Using GKS concurrently: a practical solution. Computer Graphics Forum, 9(3): 239–244, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fabianowski:2009:RTG

[FD09] B. Fabianowski and J. Dingliana. Real time global illumination: Interactive global photon mapping. Computer Graphics Forum, 28(4):1151–1159, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Feixas:1999:ITF

[FdABS99] Miquel Feixas, Esteve del Acebo, Philippe Bekaert, and Mateu Sbert. An information theory framework for the analysis of scene complexity. Computer Graphics Forum, 18 (3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=331&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Fa:1993:DMT

[FFD93] Mingxian Fa, T. Fernando, and P. M. Dew. Direct 3D manipulation techniques for interactive constraint-based solid modelling. Computer Graphics Forum, 12(3):C237–C248, ???? 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fontenier:1988:AGP

[FG88] Guy Fontenier and Pascal Gros. Architectures of graphic processors for interactive 2D graphics. Computer Graphics Forum, 7(2): 79–89, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferreira:1998:P

[FG98] F. Nunes Ferreira and Martin Göbel. Preface. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Foufou:2004:SMD

[FG04] Sebti Foufou and Lionel Garnier. Shape modelling: Dupin cyclide blends between quadric surfaces for shape modeling. Computer Graphics Forum, 23(3):321–330, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fontana:1999:FFF

[FGM99] M. Fontana, F. Giannini, and M. Meirana. A free form feature taxonomy. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=332&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=18.

Fuchs:2009:VMV

[FH09] R. Fuchs and H. Hauser. Visualization of multi-variate scientific data. Computer Graphics Forum, 28(6):1670–1690, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fuchs:2008:CIC

[FHL^{+}08] Christian Fuchs, Michael Heinz, Marc Levoy, Hans-Peter Seidel, and Hendrik P. A. Lensch. Computational imaging: Combining confocal imaging and descattering. Computer Graphics Forum, 27 (4):1245–1253, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fischer:1998:ADC

[Fis98] Martin Fischer. An algorithm for dynamic color management. Computer Graphics Forum, 17(1):17–??, March 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=213&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=17.

Fiume:1995:IPP

[Fiu95] Eugene Fiume. Isometric piecewise polynomial curves. Computer Graphics Forum, 14(1):47–58, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fiume:2001:AFA

[Fiu01] E. Fiume. Alain Fournier, 1943–2000: an appreciation. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Foster:2005:IPI

[FJW^{+}05] K. Foster, P. Jepp, B. Wyvill, M. C. Sousa, C. Galbraith, and J. A. Jorge. Interaction: Pen-and-ink for BlobTree implicit models. Computer Graphics Forum, 24(3):267–276, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frank:2009:CDP

[FK09] S. Frank and A. Kaufman. Out-of-core and dynamic programming for data distribution on a volume visualization cluster. Computer Graphics Forum, 28(1):141–153, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fattal:2006:PCP

[FL06] Raanan Fattal and Dani Lischinski. Pose controlled physically based motion. Computer Graphics Forum, 25(4):777–787, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fuchs:2007:GIS

[FLBS07] Martin Fuchs, Hendrik P. A. Lensch, Volker Blanz, and Hans-Peter Seidel. Global illumination (1): Superresolution reflectance fields: Synthesizing images for intermediate light directions. Computer Graphics Forum, 26(3):447–456, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Froumentin:2000:VBR

[FLW00] Max Froumentin, Frédéric Labrosse, and Philip Willis. A vectorbased representation for image warping. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=434&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Formella:2004:VDA

[FM04] A. Formella and K. Müller. A viewpoint-dependent approach to ray trace free-form surfaces. Computer Graphics Forum, 23(2): 143–155, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Faconti:2000:HCI

[FMB^{+}00] G. Faconti, M. Massink, M. Bordegoni, F. De Angelis, and S. Booth. Haptic cues for image disambiguation. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=409&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Fujishiro:2004:RVG

[FMK04] Issei Fujishiro, Klaus Mueller, and Arie Kaufman. Reports: Volume Graphics 2003: Tokyo, Japan, 7–8 July 2003. Computer Graphics Forum, 23(1):121–122, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fradin:2006:HTB

[FML06] D. Fradin, D. Meneveaux, and P. Lienhardt. A hierarchical topology-based model for handling complex indoor scenes. Computer Graphics Forum, 25(2):149–162, June 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Freudenberg:2001:WTI

[FMS01] Bert Freudenberg, Maic Masuch, and Thomas Strothotte. Walkthrough illustrations: Frame-coherent pen-and-ink style in a game engine. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fung:1992:RLS

[FND92] Khun Yee Fung, T. M. Nicholl, and A. K. Dewdney. A run-length slice line drawing algorithm without division operations. Computer Graphics Forum, 11(3):C267–C277 (or C267–277??), ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Foley:1998:CGI

[Fol98] James D. Foley. The convergence of graphics and imaging. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=246&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Feda:1994:MCA

[FP94] M. Feda and W. Purgathofer. A median cut algorithm for efficient sampling of radiosity functions. Computer Graphics Forum, 13(3): C/433–C/442, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Farrugia:2004:SGI

[FP04] Jean-Philippe Farrugia and Bernard Péroche. Sampling for global illumination: a progressive rendering algorithm using an adaptive perceptually based image metric. Computer Graphics Forum, 23 (3):605–614, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2008:GRR

[FQK08] Zhe Fan, Feng Qiu, and Arie E. Kaufman. GPU rendering and ray-casting: Zippy: a framework for computation and visualization on a GPU cluster. Computer Graphics Forum, 27(2):341–350, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Falcidieno:1992:TMC

[FR92] B. Falcidieno and O. Ratto. Two-manifold cell-decomposition of r-sets. Computer Graphics Forum, 11(3):C391–C404, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frohlich:2000:IMI

[FR00] Torsten Fröhlich and Marcus Roth. Integration of multidimensional interaction devices in real-time computer graphics applications. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=423&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Franke:1983:NVA

[Fra83] H. W. Franke. The new visual age: The influence of computer graphics on art and society. Computer Graphics Forum, 2(2-3):145–152, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

French:1990:GGS

[Fre90] Michael French. GosiP: a GKS-3D shell for PHIGS. Computer Graphics Forum, 9(4):317–326, December 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fruhauf:1994:RNS

[Frü94] T. Frühauf. Raycasting of nonregularly structured volume data. Computer Graphics Forum, 13(3):C/293–C/303, ???? 1994. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Firby:1985:ICG

[FS85] P. A. Firby and D. J. Stone. Interference in computer graphics. Computer Graphics Forum, 4(3):209–216, September 1985. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fong:1991:CPM

[FS91] Philip Fong and Hans-Peter P. Seidel. Control points for multivariate B-spline surfaces over arbitrary triangulations. Computer Graphics Forum, 10(4):309–317, December 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Felger:1992:VIP

[FS92] W. Felger and F. Schroder. The visualization input pipelineenabling semantic interaction in scientific visualization. Computer Graphics Forum, 11(3):C139–C151, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Floater:2008:BCP

[FS08] M. S. Floater and C. Schulz. Bases and coordinates: Pointwise radial minimization: Hermite interpolation on arbitrary domains. Computer Graphics Forum, 27(5):1505–1512, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Furtado:1993:SSM

[FT93] P. Furtado and J. C. Teixeira. Storage support for multidimensional discrete data in multimedia databases. Computer Graphics Forum, 12(3):C89–C100, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fuchs:1997:BTS

[Fuc97] Henry Fuchs. Building telepresence systems: Translating science fiction ideas into reality. Computer Graphics Forum, 16 (3):C189–??, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=155&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Fuchs:2004:IPI

[Fuc04] Henry Fuchs. Invited presentations: Immersive integration of physical and virtual environments. Computer Graphics Forum, 23(3):268, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Froumentin:1999:ERC

[FW99] M. Froumentin and P. Willis. An efficient 2\frac{1}{2}D rendering and compositing system. Computer Graphics Forum, 18 (3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=359&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Faconti:1992:IMS

[FZP92] G. P. Faconti, N. Zani, and F. Paterno’. The input model of standard graphics systems revisited by formal specification. Computer Graphics Forum, 11(3):C237–C251, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gargantini:1993:RTO

[GA93] I. Gargantini and H. H. Atkinson. Ray tracing an octree: numerical evaluation of the first intersection. Computer Graphics Forum, 12(4):199–210, October 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:1996:BMB

[GA96] E. Galin and S. Akkouche. Blob metamorphosis based on Minkowski sums. Computer Graphics Forum, 15(3):C143–C153, C467, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grimm:1998:FSE

[GA98] Cindy Grimm and Matthew Ayers. A framework for synchronized editing of multiple curve representations. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=251&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Gallop:1984:RMI

[Gal84] J. R. Gallop. Report on the 7th meeting of ISO/TC97/SC5/WG2 — computer graphics. Computer Graphics Forum, 3(3):233–238, September 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garanzha:2009:RIU

[Gar09] Kirill Garanzha. Rendering II: The use of precomputed triangle clusters for accelerated ray tracing in dynamic scenes. Computer Graphics Forum, 28(4):1199–1206, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gebal:2009:SAS

[GBAL09] K. Gebal, J. A. Bærentzen, H. Aanæs, and R. Larsen. Shape analysis: Shape analysis using the auto diffusion function. Computer Graphics Forum, 28(5):1405–1413, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Greilich:2009:GVV

[GBD09] Martin Greilich, Michael Burch, and Stephan Diehl. Graph visualization: Visualizing the evolution of compound digraphs with TimeArcTrees. Computer Graphics Forum, 28(3):975–982, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grimm:2004:PBT

[GBKG04] Sören Grimm, Stefan Bruckner, Armin Kanitsar, and Eduard Gröller. Point-based techniques: VOTS: VOlume doTS as a pointbased representation of volumetric data. Computer Graphics Forum, 23(3):661–668, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guennebaud:2004:PBT

[GBP04] Gaël Guennebaud, Loïc Barthe, and Mathias Paulin. Point-based techniques: Deferred splatting. Computer Graphics Forum, 23(3): 653–660, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guennebaud:2005:PIR

[GBP05] G. Guennebaud, L. Barthe, and M. Paulin. Points: Interpolatory refinement for real-time processing of point-based geometry. Computer Graphics Forum, 24(3):657–666, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guennebaud:2007:SVH

[GBP07] Gaël Guennebaud, Loïc Barthe, and Mathias Paulin. Shadows and visibility: High-quality adaptive soft shadow mapping. Computer Graphics Forum, 26(3):525–533, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grisoni:1999:HBS

[GBS99] Laurent Grisoni, Carole Blanc, and Christophe Schlick. Hermitian B-splines. Computer Graphics Forum, 18(4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=377&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=4&amp; src=ard&amp; vid=18.

Gaiani:2000:RCR

[GBU00] Marco Gaiani, Marcello Balzani, and Federico Uccelli. Reshaping the Coliseum in Rome: an integrated data capture and modeling method at heritage sites. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=429&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Grimsdale:1996:LDL

[GC96] R. L. Grimsdale and C. W. Chang. The layout design language: a technique for generating layout plans. Computer Graphics Forum, 15(2):97–106, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganovelli:2009:EMC

[GC09] F. Ganovelli and M. Corsini. eNVyMyCar: a multiplayer car racing game for teaching computer graphics. Computer Graphics Forum, 28(8):2025–2032, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guan:2006:MCE

[GCL^{+}06] Yu Guan, Wei Chen, Xiao Liang, Zi’ang Ding, and Qunsheng Peng. Matting and compositing: Easy matting — a stroke based approach for continuous image matting. Computer Graphics Forum, 25(3):567–576, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganovelli:2000:MMS

[GCMS00] Fabio Ganovelli, Paolo Cignoni, Claudio Montani, and Roberto Scopigno. A multiresolution model for soft objects supporting interactive cuts and lacerations. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=419&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Ghosh:2009:AMR

[GCP^{+}09] Abhijeet Ghosh, Tongbo Chen, Pieter Peers, Cyrus A. Wilson, and Paul Debevec. Appearance measurement and rendering: Estimating specular roughness and anisotropy from second order spherical gradient illumination. Computer Graphics Forum, 28 (4):1161–1170, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gibson:1985:GGT

[GD85] D. R. Gibson and D. A. Duce. GKS graphics and text processing. Computer Graphics Forum, 4(3):259–270, September 1985. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghazanfarpour:1996:GTU

[GD96] Djamchid Ghazanfarpour and Jean-Michel M. Dischler. Generation of 3D texture using multiple 2D models analysis. Computer Graphics Forum, 15(3):C311–C323, C478–C479, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Granier:2001:IUR

[GD01] Xavier Granier and George Drettakis. Incremental updates for rapid glossy global illumination. Computer Graphics Forum, 20 (3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gilet:2009:FIH

[GD09] G. Gilet and J. M. Dischler. A framework for interactive hypertexture modelling. Computer Graphics Forum, 28(8):2229–2243, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grosso:1998:PIS

[GE98] Roberto Grosso and Thomas Ertl. Progressive iso-surface extraction from hierarchical 3D meshes. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=260&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Gopi:2004:GPS

[GE04] M. Gopi and David Eppstein. Geometry processing: Single-strip triangulation of manifolds with arbitrary topology. Computer Graphics Forum, 23(3):371–379, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gervautz:1992:CSA

[Ger92] Michael Gervautz. Consistent schemes for addressing surfaces when ray tracing transparent CSG objects. Computer Graphics Forum, 11(4):203–211, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2006:ARR

[GFW^{+}06] Johannes Günther, Heiko Friedrich, Ingo Wald, Hans-Peter Seidel, and Philipp Slusallek. Animation rendering: Ray tracing animated scenes using motion decomposition. Computer Graphics Forum, 25(3):517–525, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guennebaud:2008:PMD

[GGG08] Gaël Guennebaud, Marcel Germann, and Markus Gross. Points and meshes: Dynamic sampling and rendering of algebraic point set surfaces. Computer Graphics Forum, 27(2):653–662, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gress:2006:ARG

[GGK06] Alexander Greß, Michael Guthe, and Reinhard Klein. Animation rendering: GPU-based collision detection for deformable parameterized surfaces. Computer Graphics Forum, 25(3):497–506, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grinspun:2006:MPC

[GGRZ06] Eitan Grinspun, Yotam Gingold, Jason Reisman, and Denis Zorin. Mesh processing: Computing discrete shape operators on general meshes. Computer Graphics Forum, 25(3):547–556, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Genetti:1998:ASO

[GGW98] Jon Genetti, Dan Gordon, and Glen Williams. Adaptive supersampling in object space using pyramidal rays. Computer Graphics Forum, 17(1):29–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=214&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Gibson:1996:EHR

[GH96] S. Gibson and R. J. Hubbold. Efficient hierarchical refinement and clustering for radiosity in complex environments. Computer Graphics Forum, 15(5):297–310, December 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gibson:1997:PDR

[GH97] S. Gibson and R. J. Hubbold. Perceptually-driven radiosity. Computer Graphics Forum, 16(2):129–141, June 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0167-7055&iid=2&src=ard&vid=16.

Gibson:2001:FIB

[GHH01] Simon Gibson, Toby Howard, and Roger Hubbold. Flexible imagebased photometric reconstruction using virtual light sources. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giesen:2002:SRB

[GJ02] Joachim Giesen and Matthias John. Surface reconstruction based on a dynamical system. Computer Graphics Forum, 21(3):363– 372, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grundy:2009:MDV

[GJL^{+}09] Edward Grundy, Mark W. Jones, Robert S. Laramee, Rory P. Wilson, and Emily L. C. Shepard. Multidimensional data: Visualisation of sensor data from animal movement. Computer Graphics Forum, 28(3):815–822, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerrero:2008:RTI

[GJW08] P. Guerrero, S. Jeschke, and M. Wimmer. Real-time indirect illumination and soft shadows in dynamic scenes using spherical lights. Computer Graphics Forum, 27(8):2154–2168, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gluth:2003:ATA

[GK03a] M. Gluth and R. Klein. Automatic texture atlas generation from trimmed NURBS models. Computer Graphics Forum, 22 (3):453–461, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guthe:2003:TBA

[GK03b] M. Guthe and R. Klein. Textures and BRDFs: Automatic texture atlas generation from trimmed NURBS models. Computer Graphics Forum, 22(3):453–461, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gassenbauer:2009:RIS

[GKB09] Václav Gassenbauer, Jaroslav Křivánek, and Kadi Bouatouch. Rendering II: Spatial directional radiance caching. Computer Graphics Forum, 28(4):1189–1198, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Green:2007:GIE

[GKD07] Paul Green, Jan Kautz, and Frédo Durand. Global illumination (2): Efficient reflectance and visibility approximations for environment map rendering. Computer Graphics Forum, 26(3):495–502, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gopiy:2000:SRB

[GKS00] M. Gopiy, S. Krishnan, and C. T. Silva. Surface reconstruction based on lower dimensional localized Delaunay triangulation. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=439&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Giertse:1994:VGA

[GL94a] C. Giertse and A. Lucas. 3D visualization for 2D GIS: an analysis of the users’ needs and a review of techniques. Computer Graphics Forum, 13(3):C/1–C/12, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giertsen:1994:VGA

[GL94b] C. Giertsen and A. Lucas. 3D visualization for 2D GIS: an analysis of the users’ needs and a review of techniques. Computer Graphics Forum, 13(3):C/01–??, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Glassner:1999:AS

[Gla99] Andrew Glassner. Active storytelling. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=321&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Galin:2000:MB

[GLA00] Eric Galin, Antoine Leclercq, and Samir Akkouche. Morphing the BlobTree. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2009:ASG

[GLHB09] Yue Gao, Chen-Feng Li, Shi-Min Hu, and Brian A. Barsky. Animation: Simulating gaseous fluids with low and high speeds. Computer Graphics Forum, 28(7):1845–1852, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Greiner:1996:DDT

[GLW96] Guenther Greiner, Joachim Loos, and Wieger Wesselink. Data dependent thin plate energy and its use in interactive surface modeling. Computer Graphics Forum, 15(3):C175–C185, C469, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grimstead:1996:ILL

[GM96] I. J. Grimstead and R. R. Martin. Incremental line labelling for sketch input of solid models. Computer Graphics Forum, 15(2): 155–166, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gamito:2006:AAS

[GM06] Manuel N. Gamito and Steve C. Maddock. Anti-aliasing with stratified B-spline filters of arbitrary degree. Computer Graphics Forum, 25(2):163–172, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobbetti:2006:GCD

[GMC^{+}06] E. Gobbetti, F. Marton, P. Cignoni, M. Di Benedetto, and F. Ganovelli. Geometry compression and decompression: C-BDAM — compressed batched dynamic adaptive meshes for terrain rendering. Computer Graphics Forum, 25(3):333–342, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grave:2009:MDG

[GMDW09] Frank Grave, Thomas Müller, Carsten Dachsbacher, and Günter Wunner. Multidimensional data: The Gödel engine — an interactive approach to visualization in general relativity. Computer Graphics Forum, 28(3):807–814, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guthe:2009:BCP

[GMSK09] Michael Guthe, Gero Müller, Martin Schneider, and Reinhard Klein. BTF-CIELab: a perceptual difference measure for quality assessment and compression of BTFs. Computer Graphics Forum, 28(1):101–113, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:1997:SRU

[GMW97] Baining Guo, Jai Menon, and Brian Willette. Surface reconstruction using alpha shapes. Computer Graphics Forum, 16 (4):177–190, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=178&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Galbraith:2004:NPA

[GMW04] Callum Galbraith, Lars Mündermann, and Brian Wyvill. Natural phenomena and aging: Implicit visualization and inverse modeling of growing trees. Computer Graphics Forum, 23(3):351–360, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gaddipatti:1997:SIG

[GMY97] Ajeetkumar Gaddipatti, Raghu Machiraju, and Roni Yagel. Steering image generation with wavelet based perceptual metric. Computer Graphics Forum, 16(3):C241–C251, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=161&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=16.

Gnatz:1982:RAF

[Gna82] R. Gnatz. Relying on abstraction to facilitate graphics programming. Computer Graphics Forum, 1(3):124–128, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gnatz:1983:AAS

[Gna83] R. Gnatz. An algebraic approach to the standardization and the certification of graphics software. Computer Graphics Forum, 2 (2-3):153–166, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobel:1995:VES

[Gob95] M. Gobel, editor. Virtual environments ’95: selected papers of the Eurographics Workshops in Barcelona, Spain, 1993, and Monte Carlo, 1995. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82737-4 (paperback). ISSN 0946-2767. LCCN HD30.2122 .E89 1995.

Gobel:1996:VES

[Gob96] M. Gobel, editor. Virtual environments and scientific visualization ’96: selected papers of the Eurographics Workshops in Monte Carlo, Monaco, February 19–20, 1996, and in Prague, Czech Republic, April 23–25, 1996, Springer Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-211-82886-9.

Goldman:1985:MBC

[Gol85] Gary S. Goldman. Micro-based CAD: a production tool for manufacturing engineering. Computer Graphics Forum, 4(4):287–300, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gossling:1986:BHR

[Gos86] T. H. Gossling. Beyond Hershey: a representation of fonts for computer graphics. Computer Graphics Forum, 5(4):291–295, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gossling:1989:LPA

[Gos89] T. H. Gossling. The line-picture attribute. Computer Graphics Forum, 8(2):129–132, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gotsman:1994:CTF

[Got94] C. Gotsman. Constant-time filtering by singular value decomposition. Computer Graphics Forum, 13(2):153–163, June 1994. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gotsman:2003:OVB

[Got03] Craig Gotsman. On the optimality of valence-based connectivity coding. Computer Graphics Forum, 22(1):99–102, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galoppo:2007:SMD

[GOT^{+}07] Nico Galoppo, Miguel A. Otaduy, Serhat Tekin, Markus Gross, and Ming C. Lin. Skinning and mesh deformation: Soft articulated characters with fast contact handling. Computer Graphics Forum, 26(3):243–253, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gourret:1987:IPF

[GP87] J. P. Gourret and J. Paille. Irregular polygon fill using contour encoding. Computer Graphics Forum, 6(4):317–325, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:2006:RES

[GP06] Eric Galin and Pierre Poulin. Reports: Eurographics Symposium on Natural Phenomena. Computer Graphics Forum, 25(1):144, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gauthier:2009:GIA

[GP09] M. Gauthier and P. Poulin. Geometry images of arbitrary genus in the spherical domain. Computer Graphics Forum, 28(8):2201– 2215, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gu:2009:CHM

[GPD09] Qin Gu, Jingliang Peng, and Zhigang Deng. Compression of human motion capture data using motion pattern indexing. Computer Graphics Forum, 28(1):1–12, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Green:1984:RDS

[Gre84] Mark Green. Report on dialogue specification tools. Computer Graphics Forum, 3(4):305–313, December 1984. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Greenaway:1985:EPV

[Gre85] D. S. Greenaway. EUROGRAPHICS ’85 — a personal view. Computer Graphics Forum, 4(4):399–406, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Greiner:1994:VDF

[Gre94] G. Greiner. Variational design and fairing of spline surfaces. Computer Graphics Forum, 13(3):C/143–C/154, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Greiner:1997:CE

[Gre97] Günther Greiner. Calendar of events. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=119&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Greiner:1998:CE

[Gre98] Günther Greiner. Calendar of events. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=238&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Groller:1992:FSM

[Gro92] E. Groller. Fractals and solid modeling. Computer Graphics Forum, 11(3):C415–C424, C484–C485, ???? 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gross:2001:PBG

[Gro01] Markus Gross. Are points the better graphics primitives? Computer Graphics Forum, 20(3):??, September 2001. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gossling:1985:ASP

[GS85] T. H. Gossling and S. T. S. Stokes. APEX and STRETCH: a practical form for the conic arc. Computer Graphics Forum, 4(3): 229–237, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gain:2006:ACG

[GS06] J. E. Gain and W. Straßer. AFRIGRAPH: Computer graphics in Africa. Computer Graphics Forum, 25(4):697–698, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Germer:2009:PAF

[GS09] T. Germer and M. Schwarz. Procedural arrangement of furniture for real-time walkthroughs. Computer Graphics Forum, 28(8): 2068–2078, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobbetti:2003:HHO

[GSA03a] Enrico Gobbetti, Leonardo Spanò, and Marco Agus. Hierarchical higher order face cluster radiosity for global illumination walkthroughs of complex non-diffuse environments. Computer Graphics Forum, 22(3):563–572, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobbetti:2003:PHH

[GSA03b] Enrico Gobbetti, Leonardo Spanò, and Marco Agus. Photorealism: Hierarchical higher order face cluster radiosity for global illumination walkthroughs of complex non-diffuse environments. Computer Graphics Forum, 22(3):563–572, September 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gilson:2008:GVW

[GSGC08] O. Gilson, N. Silva, P. W. Grant, and M. Chen. Graph visualization: From Web data to visualization via ontology mapping. Computer Graphics Forum, 27(3):959–966, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Graf:1994:CGR

[GSHN94] K. C. Graf, M. Suter, J. Hagger, and D. Nuesch. Computer graphics and remote sensing — a synthesis for environmental planning and civil engineering. Computer Graphics Forum, 13(3):C/13– C/22, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gutierrez:2008:RVE

[GSMA08] Diego Gutierrez, Francisco J. Seron, Adolfo Munoz, and Oscar Anson. Rendering volume effects: Visualizing underwater ocean optics. Computer Graphics Forum, 27(2):547–556, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gu:1986:ROO

[GTS86] Kaining Gu, Zesheng Tang, and Jiaguang Sun. Reconstruction of 3D objects from orthographic projections. Computer Graphics Forum, 5(4):317–323, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guedj:1982:BHM

[Gue82] R. A. Guedj. Better human-machine interaction. A tough and unescapable challenge. Computer Graphics Forum, 1(1):20–25, March 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). Check author: conflict between [Gue82, Str82].

GuimaraesBaranoski:1992:PDM

[Gui92] G. V. Guimaraes Baranoski. The parametric differential method: an alternative to the calculation of form factors. Computer Graphics Forum, 11(3):C193–C204, C474, ???? 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:1993:NMP

[Guo93] B. Guo. Nonsplitting macro patches for implicit cubic spline surfaces. Computer Graphics Forum, 12(3):C433–C445, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goldenstein:2005:ADM

[GVV05] Siome Goldenstein, Christian Vogler, and Luiz Velho. Adaptive deformable models for graphics and vision. Computer Graphics Forum, 24(4):729–741, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grundland:2006:MCC

[GVWD06] Mark Grundland, Rahul Vohra, Gareth P. Williams, and Neil A. Dodgson. Matting and compositing: Cross dissolve without cross fade: Preserving contrast, color and salience in image compositing. Computer Graphics Forum, 25(3):577–586, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2005:ASD

[GW05] Zheng Guo and Kok Cheong Wong. Animation: Skinning with deformable chunks. Computer Graphics Forum, 24(3):373–381, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giegl:2007:USI

[GW07] Markus Giegl and Michael Wimmer. Unpopping: Solving the image-space blend problem for smooth discrete LOD transitions. Computer Graphics Forum, 26(1):46–49, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garth:2008:FFE

[GWT^{+}08] Christoph Garth, Alexander Wiebel, Xavier Tricoche, Ken Joy, and Gerik Scheuermann. Flow feature extraction: Lagrangian visualization of flow-embedded surface structures. Computer Graphics Forum, 27(3):1007–1014, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gerald-Yamasaki:1993:CVC

[GY93] M. J. Gerald-Yamasaki. Cooperative visualization of computational fluid dynamics. Computer Graphics Forum, 12(3):C497– C508, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haase:1996:SVR

[Haa96] Helmut Haase. Symbiosis of virtual reality and scientific visualization system. Computer Graphics Forum, 15(3):C443–C451, C486, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hall:1999:NRM

[Hal99] P. Hall. Nonphotorealistic rendering by Q-mapping. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=300&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Hasselgren:2005:FIS

[HAML05] Jon Hasselgren, Tomas Akenine-Möller, and Samuli Laine. A family of inexpensive sampling schemes. Computer Graphics Forum, 24(4):843–848, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hand:1997:SIT

[Han97] Chris Hand. A survey of 3D interaction techniques. Computer Graphics Forum, 16(5):269–281, December 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=194&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Hanrahan:2005:IPR

[Han05] Pat Hanrahan. Invited presentations: Realistic or abstract imagery: The future of computer graphics? Computer Graphics Forum, 24(3):xxi, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hart:1997:IRR

[Har97] John C. Hart. Implicit representations of rough surfaces. Computer Graphics Forum, 16(2):91–99, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=125&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=2&amp; src=
ard&amp; vid=16.

Haw:1985:AGC

[Haw85] R. J. Haw. An application of geodesic curves to sail design. Computer Graphics Forum, 4(2):137–139, June 1985. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2008:RDS

[HAWG08] Qi-Xing Huang, Bart Adams, Martin Wicke, and Leonidas J. Guibas. Registering deforming surfaces II: Non-rigid registration under isometric deformations. Computer Graphics Forum, 27(5):1449–1457, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1991:EVA

[HB91] Toby Howard and Judith R. Brown. Eurographics ’91: Vienna, Austria — 2–6 September 1991. Computer Graphics Forum, 10 (4):345–352, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Highfield:1992:HSE

[HB92] J. C. Highfield and H. E. Bez. Hidden surface elimination on parallel processors. Computer Graphics Forum, 11(5):293–307, December 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howie:1994:MPA

[HB94] C. T. Howie and E. H. Blake. The Mesh Propagation algorithm for isosurface construction. Computer Graphics Forum, 13(3):C/65– C/74, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haley:1996:IVR

[HB96] Michael B. Haley and Edwin H. Blake. Incremental volume rendering using hierarchical compression. Computer Graphics Forum, 15(3):C45–C55, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havran:2000:LRS

[HB00] V. Havran and J. Bittner. LCTS: Ray shooting using longest common traversal sequences. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=398&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Hisada:2002:SBA

[HBK02] Masayuki Hisada, Alexander G. Belyaev, and Tosiyasu L. Kunii. A skeleton-based approach for detection of perceptually salient features on polygonal surfaces. Computer Graphics Forum, 21(4): 689–700, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubball:2008:IBT

[HCG08] Daniel Hubball, Min Chen, and Phil W. Grant. Image-based techniques: Image-based aging using evolutionary computing. Computer Graphics Forum, 27(2):607–616, April 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsu:1993:NAG

[HCL93] S. Y. Hsu, L. R. Chow, and H. C. Liu. A new approach for the generation of circles. Computer Graphics Forum, 12(2):105–109, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hopgood:1986:A

[HD86] F. R. A. Hopgood and D. A. Duce. Ausgraph ’86. Computer Graphics Forum, 5(4):355–356, December 1986. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubner:1989:TOO

[Hd89] Wolfgang Hübner and Manuel de Lancastre. Towards an objectoriented interaction model for graphics user interfaces. Computer Graphics Forum, 8(3):207–217, September 1989. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haindl:1995:MHS

[HD95] M. Haindl and B. De Ruiter. The MADE help system. Computer Graphics Forum, 14(3):C/149–C/157, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:2002:SLA

[HD02] Ivan Herman and Daniel Dardailler. SVG linearization and accessibility. Computer Graphics Forum, 21(4):777–786, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1998:TVN

[HDM98] Ivan Herman, Maylis Delest, and Guy Melancon. Tree visualisation and navigation clues for information visualisation. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=235&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Haumont:2003:VVC

[HDS03a] D. Haumont, O. Debeir, and F. Sillion. Visibility: Volumetric celland-portal generation. Computer Graphics Forum, 22(3):303–312, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haumont:2003:VCP

[HDS03b] D. Haumont, O. Debeir, and F. Sillion. Volumetric cell-and-portal generation. Computer Graphics Forum, 22(3):303–312, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hasenfratz:1999:PAC

[HDSD99] Jean-Marc Hasenfratz, Cyrille Damez, François Sillion, and George Drettakis. A practical analysis of clustering strategies for hierarchical radiosity. Computer Graphics Forum, 18 (3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=342&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Hilton:1994:VFI

[HE94] T. L. Hilton and P. K. Egbert. Vector fields: an interactive tool for animation, modeling and simulation with physically based 3D particle systems and soft objects. Computer Graphics Forum, 13 (3):C/329–C/338, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hauth:2001:HPS

[HE01] Michael Hauth and Olaf Etzmuss. A high performance solver for the animation of deformable objects using advanced numerical methods. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heal:1989:HON

[Hea89] B. W. Heal. Hidden octree node removal as a pipeline process. Computer Graphics Forum, 8(3):199–206, September 1989. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heal:1990:NPO

[Hea90] B. W. Heal. Node partitioning in an octree display pipeline. Computer Graphics Forum, 9(3):205–211, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heckbert:1992:RF

[Hec92] P. S. Heckbert. Radiosity in flatland. Computer Graphics Forum, 11(3):C181–C192, C474, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hecker:2001:WGW

[Hec01] Chris Hecker. Why games will be the preeminent art form of the 21^{st} century. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hegron:1990:RFE

[Heg90] G. Hegron. Report on the first EUROGRAPHICS workshop on animation and simulation. Computer Graphics Forum, 9(4):373– 374, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hegron:1991:RSE

[Hég91] Gérad Hégron. Report on the Second Eurographics Workshop on Animation and Simulation. Computer Graphics Forum, 10(4): 365–366, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heinze:1995:MDM

[Hei95] R. Heinze. A method for displaying metaballs by using Bézier clipping. Computer Graphics Forum, 14(5):337–??, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heidrich:2001:IDG

[Hei01] Wolfgang Heidrich. Interactive display of global illumination solutions for non-diffuse environments — a survey. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1982:PGI

[Her82] Ivan Herman. A parser generator for interactive command languages. Computer Graphics Forum, 1(3):106–115, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1984:MMC

[Her84] I. Herman. Managing multiple context-frames through GKS. Computer Graphics Forum, 3(1):79–82, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1989:PIR

[Her89] Ivan Herman. On the projective invariant representation of conics in computer graphics. Computer Graphics Forum, 8(4):301–314, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hopgood:1983:REP

[HEtH^{+}83] F. R. A. Hopgood, J. Encarnacao, P. J. W. ten Hagen, R. D. Bergeron, W. T. Hewitt, C. D. Osland, and D. A. Duce. Reports of EUROGRAPHICS ’83 panel sessions. Computer Graphics Forum, 2(4):203–206, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hewitt:1984:CGF

[Hew84a] W. T. Hewitt. Computer graphics forum — algorithm section. Computer Graphics Forum, 3(1):111–114, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hewitt:1984:PPH

[Hew84b] W. T. Hewitt. PHIGS — Programmer’s Hierarchical Interactive Graphics System. Computer Graphics Forum, 3(4):299–300, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hewitt:1990:CCG

[Hew90] W. T. Hewitt. Curves for computer graphics standards — workshop report. Computer Graphics Forum, 9(2):165–169, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heywood:1984:WGS

[HGG^{+}84] T. R. Heywood, B. N. Galton, J. Gillett, A. J. Morffew, P. Quarendon, S. J. P. Todd, and W. V. Wright. The Winchester graphics system: a technical overview. Computer Graphics Forum, 3(1): 61–69, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haritsis:1992:RGR

[HGW92] A. Haritsis, D. Gillies, and C. William. Realistic generation and real time animation of images of the human colon. Computer Graphics Forum, 11(3):C367–C379, C482, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubl:1990:MCS

[HH90] Josef Hübl and Ivan Herman. Modelling clip: Some more results. Computer Graphics Forum, 9(2):101–107, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howlett:1998:MSS

[HH98] P. Howlett and W. T. Hewitt. Mass-spring simulation using adaptive non-active points. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=281&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Hidalgo:2002:HGI

[HH02] Eduardo Hidalgo and Roger J. Hubbold. Hybrid geometric — image based rendering. Computer Graphics Forum, 21(3):471–482, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hewitt:1982:RGU

[HHA82] W. T. Hewitt, R. J. Hubbold, and A. C. Arnold. Raster graphics at the University of Manchester. Computer Graphics Forum, 1(2): 58–70, June 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hagen:1993:VSD

[HHB93] H. Hagen, St. Hahmann, and G.-P. Bonneau. Variational surface design and surface interrogation. Computer Graphics Forum, 12 (3):C447–C459, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hiller:2003:BSM

[HHD03a] Stefan Hiller, Heino Hellwig, and Oliver Deussen. Beyond stippling — methods for distributing objects on the plane. Computer Graphics Forum, 22(3):515–522, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hiller:2003:NPR

[HHD03b] Stefan Hiller, Heino Hellwig, and Oliver Deussen. Nonphotorealistic rendering and painting: Beyond stippling — methods for distributing objects on the plane. Computer Graphics Forum, 22(3):515–522, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herzog:2007:GIG

[HHK^{+}07] Robert Herzog, Vlastimil Havran, Shinichi Kinuwaki, Karol Myszkowski, and Hans-Peter Seidel. Global illumination (2): Global illumination using photon ray splatting. Computer Graphics Forum, 26(3):503–513, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hansmann:1989:EPE

[HHS89] W. Hansmann, F. R. A. Hopgood, and Wolfgang Strasser, editors. Eurographics ’89: proceeding[s] of the European Computer Graphics Conference and Exhibition, Hamburg, F.R.G., 4–8 September, 1989. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1989. ISBN 0-444-88013-5. LCCN T 385 E97 1989.

Halper:2001:CEC

[HHS01] Nicolas Halper, Ralf Helbing, and Thomas Strothotte. A camera engine for computer games: Managing the trade-off between constraint satisfaction and frame coherence. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havran:2005:RFF

[HHS05] Vlastimil Havran, Robert Herzog, and Hans-Peter Seidel. Rendering I: Fast final gathering via reverse photon mapping. Computer Graphics Forum, 24(3):323–332, September 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hui:1999:TBA

[HJ99] K. C. Hui and Z. H. Jiang. Tetrahedra based adaptive polygonization of implicit surface patches. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=302&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Hubbold:2007:RES

[HJL07] Roger Hubbold, Joaquim Jorge, and Ming Lin. Reports: 12^{th} Eurographics Symposium on Virtual Environments. Computer Graphics Forum, 26(1):114–115, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsu:1992:HHL

[HK92] Y. H. Hsu and Y. S. Kuo. HAGI, a high-level application/graphics interface. Computer Graphics Forum, 11(3):C71–C80, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herzner:1994:MHO

[HK94] Wolfgang Herzner and Frank Kappe, editors. Multimedia/ hypermedia in open distributed environments: proceedings of the Eurographics symposium in Graz, Austria, June 6–9, 1994. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-82587-8. LCCN QA76.575 .M833 1994.

Hubbold:2000:RTS

[HK00] Roger Hubbold and Martin Keates. Real-time simulation of a stretcher evacuation in a large-scale virtual environment. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hong:2003:ABL

[HK03a] Jeong-Mo Hong and Chang-Hun Kim. Animation of bubbles in liquid. Computer Graphics Forum, 22(3):253–262, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hong:2003:RAN

[HK03b] Jeong-Mo Hong and Chang-Hun Kim. Rendering and animating nature: Animation of bubbles in liquid. Computer Graphics Forum, 22(3):253–262, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Habbecke:2009:IBM

[HK09a] Martin Habbecke and Leif Kobbelt. Image-based modeling and rendering: an intuitive interface for interactive high quality imagebased modeling. Computer Graphics Forum, 28(7):1765–1772, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ho:2009:MSC

[HK09b] Edmond S. L. Ho and Taku Komura. Motion synthesis: Character motion synthesis by topology coordinates. Computer Graphics Forum, 28(2):299–308, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hornung:2009:IPA

[HK09c] A. Hornung and L. Kobbelt. Interactive pixel-accurate free viewpoint rendering from images with silhouette aware sampling. Computer Graphics Forum, 28(8):2090–2103, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2007:ICI

[hKAC07] Tae hoon Kim, Jongwoo Ahn, and Min Gyu Choi. Image and color: Image dequantization: Restoration of quantized colors.

Computer Graphics Forum, 26(3):619–626, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hao:2008:VAD

[HKD^{+}08] Ming Hao, Daniel A. Keim, Umeshwar Dayal, Daniela Oelke, and Chantal Tremblay. Visual analytics: Density displays for data stream monitoring. Computer Graphics Forum, 27(3):895–902, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heck:2006:ASU

[HKG06] Rachel Heck, Lucas Kovar, and Michael Gleicher. Animation: Splicing upper-body actions with locomotion. Computer Graphics Forum, 25(3):459–466, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2000:RMW

[hKLS00] Tae hoon Kim, Jehee Lee, and Sung Yong Shin. Robust motion watermarking based on multiresolution analysis. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=411&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Herzog:2008:PRA

[HKMS08] Robert Herzog, Shinichi Kinuwaki, Karol Myszkowski, and Hans-Peter Seidel. Perception for rendering and animation: Render2MPEG: a perception-based framework towards integrating rendering and video compression. Computer Graphics Forum, 27 (2):183–192, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ha:2009:FFW

[HKS09] Linh Ha, Jens Krüger, and Cláudio T. Silva. Fast four-way parallel radix sorting on GPUs. Computer Graphics Forum, 28(8): 2368–2378, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Habel:2009:AAT

[HKW09] Ralf Habel, Alexander Kusternig, and Michael Wimmer. All about trees: Physically guided animation of trees. Computer Graphics Forum, 28(2):523–532, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harris:2001:RTC

[HL01] Mark J. Harris and Anselmo Lastra. Real-time cloud rendering. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2002:ALC

[HL02] Y. Q. Huang and Y. K. Liu. An algorithm for line clipping against a polygon based on shearing transformation. Computer Graphics Forum, 21(4):683–688, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsu:2003:FFB

[HL03a] P. C. Hsu and C. Lee. Field functions for blending range controls on soft objects. Computer Graphics Forum, 22(3):233–242, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsu:2003:MFF

[HL03b] P. C. Hsu and C. Lee. Modelling: Field functions for blending range controls on soft objects. Computer Graphics Forum, 22(3): 233–242, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsu:2003:SMB

[HL03c] P.-C. Hsu and C. Lee. The scale method for blending operations in functionally-based constructive geometry. Computer Graphics Forum, 22(2):143–158, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hutchinson:1996:SGS

[HLH96] D. Hutchinson, F. Lin, and W. T. Hewitt. Surface graph sketching. Computer Graphics Forum, 15(3):C301–C310, C477, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hui:2007:ARR

[HLL07] K. C. Hui, A. H. C. Lee, and Y. H. Lai. Accelerating refractive rendering of transparent objects. Computer Graphics Forum, 26(1):24–33, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havaldar:1997:SNV

[HLM97] Parag Havaldar, Mi-Suen Lee, and Gérard Medioni. Synthesizing novel views from unregistered 2-D images. Computer Graphics Forum, 16(1):65–73, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=117&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=16.

Hughes:1996:SGI

[HLS96] Merlin Hughes, Anselmo A. Lastra, and Edward Saxe. Simplification of global-illumination meshes. Computer Graphics Forum, 15(3):C339–C345, C481–C482, September 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haw:1983:GCP

[HM83] R. J. Haw and R. C. Munchmeyer. Geodesic curves on patched polynomial surfaces. Computer Graphics Forum, 2(4):225–232, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubner:1986:GBG

[HM86] W. Hübner and Z. I. Markov. GKS based graphics programming in Prolog. Computer Graphics Forum, 5(1):41–50, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hardenbergh:1991:IPU

[HM91] Jan Hardenbergh and James Michener. Integrating PHIGS and user interface systems. Computer Graphics Forum, 10(1):27–36, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofsetz:2008:OSV

[HMB08] C. Hofsetz, N. Max, and R. Bastos. Object-space visibility ordering for point-based and volume rendering. Computer Graphics Forum, 27(1):91–101, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hamill:2005:PPE

[HMDO05] J. Hamill, R. McDonnell, S. Dobbyn, and C. O’Sullivan. Perception: Perceptual evaluation of impostor representations for virtual humans and buildings. Computer Graphics Forum, 24(3):

623–633, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hesser:1995:TAV

[HMK^{+}95] J. Hesser, R. Männer, G. Knittel, W. Strasser, H. Pfister, and A. Kaufman. Three architectures for volume rendering. Computer Graphics Forum, 14(3):C/111–C/122, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Horvath:1992:RCB

[HMRSK92] T. Horvath, G. Marton, P. Risztics, and L. Szirmay-Kalos. Ray coherence between a sphere and a convex polyhedron. Computer Graphics Forum, 11(2):163–172, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herzog:2009:SGI

[HMS09] Robert Herzog, Karol Myszkowski, and Hans-Peter Seidel. Shadows and global illumination: Anisotropic radiance-cache splatting for efficiently computing high-quality global illumination with lightcuts. Computer Graphics Forum, 28(2):259–268, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hadap:2001:MDH

[HMT01] Sunil Hadap and Nadia Magnenat-Thalmann. Modeling dynamic hair as a continuum. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haber:2001:PGC

[HMYS01] Jörg Haber, Karol Myszkowski, Hitoshi Yamauchi, and Hans-Peter Seidel. Perceptually guided corrective splatting. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hurtado:2004:VOM

[HNR^{+}04] F. Hurtado, M. Noy, J.-M. Robert, V. Sacristán, and S. Skiena. Visualizing objects with mirrors. Computer Graphics Forum, 23 (2):157–166, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holton:1994:SGB

[Hol94] M. Holton. Strands, gravity and botanical tree imagery. Computer Graphics Forum, 13(1):57–67, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hopgood:1984:S

[Hop84] F. R. A. Hopgood. SIGGRAPH ’84. Computer Graphics Forum, 3(4):319–320, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Horn:1990:III

[Hor90] Alastair N. Horn. IFSs and interactive image synthesis. Computer Graphics Forum, 9(2):127–137, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1987:VE

[How87] T. Howard. A view of Eurographics ’87. Computer Graphics Forum, 6(4):365–366, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1989:APB

[How89] T. L. J. Howard. An annotated PHIGS bibliography. Computer Graphics Forum, 8(3):262–265, September 1989. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1990:TSN

[How90] T. L. J. Howard. TopDraw: a structure network visualiser for PHIGS. Computer Graphics Forum, 9(2):139–147, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1991:BRa

[How91a] Toby Howard. Book reviews. Computer Graphics Forum, 10(2): 175–178, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1991:BRb

[How91b] Toby Howard. Book reviews. Computer Graphics Forum, 10(3): 265–269, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1991:BRc

[How91c] Toby Howard. Book reviews. Computer Graphics Forum, 10(4): 337–342, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Howard:1997:BR

[How97] Toby Howard. Book reviews. Computer Graphics Forum, 16 (4):225–??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=182&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Hopgood:1983:SUW

[HP83] F. R. A. Hopgood and C. Prosser. Single user workstations. Computer Graphics Forum, 2(4):219–223, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubbold:1984:NCR

[HP84] R. J. Hubbold and I. Page. Nicograph ’83 conference report. Computer Graphics Forum, 3(1):115–116, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanrahan:1995:RTP

[HP95] P. M. Hanrahan and W. Purgathofer, editors. Rendering techniques ’95: proceedings of the Eurographics Workshop in Dublin, Ireland, June 12–14, 1995. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82733-1. LCCN T385 .R396 1995.

Hey:2002:ARE

[HP02] Heinrich Hey and Werner Purgathofer. Advanced radiance estimation for photon map global illumination. Computer Graphics Forum, 21(3):541–546, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hildebrandt:2004:GPA

[HP04] Klaus Hildebrandt and Konrad Polthier. Geometry processing: Anisotropic filtering of non-linear surface features. Computer Graphics Forum, 23(3):391–400, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hegron:1989:MCA

[HPT89] G. Hegron, P. Palamidese, and D. Thalmann. Motion control of animation, simulation and visualization. Computer Graphics Forum, 8(4):347–352, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1985:GDD

[HR85] I. Herman and J. Reviczky. A general device driver for GKS. Computer Graphics Forum, 4(3):245–250, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1987:NMI

[HR87] Ivan Herman and Janos Reviczky. New methods for improving the GKS fill area output primitive. Computer Graphics Forum, 6(3): 195–201, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herman:1988:SRM

[HR88] I. Herman and J. Reviczky. Some remarks on the modelling clip problem. Computer Graphics Forum, 7(4):265–271, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holden:1992:GTE

[HR92] E. J. Holden and G. G. Roy. The graphical translation of English text into Signed English in the hand sign translator system. Computer Graphics Forum, 11(3):C357–C366, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hall:1992:OMH

[HS92] W. Hall and D. Simmons. An open model for hypermedia and its application to geographical information systems. Computer Graphics Forum, 11(3):C1–C7, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hausmann:1994:VRP

[HS94] B. Hausmann and H.-P. Seidel. Visualization of regular polytopes in three and four dimensions. Computer Graphics Forum, 13(3): C/305–C/316, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holzschuch:1998:EEB

[HS98] Nicolas Holzschuch and François Sillion. An exhaustive errorbounding algorithm for hierarchical radiosity. Computer Graphics Forum, 17(4):??, December 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=285&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=17.

Hamel:1999:CRU

[HS99] J. Hamel and T. Strothotte. Capturing and re-using rendition styles for non-photorealistic rendering. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=338&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Hasegawa:2004:SHS

[HS04] S. Hasegawa and M. Sato. Simulation with haptics and sound: Real-time rigid body simulation for haptic interactions based on contact volume of polygonal objects. Computer Graphics Forum, 23(3):529–538, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hormann:2008:BCM

[HS08] K. Hormann and N. Sukumar. Bases and coordinates: Maximum entropy coordinates for arbitrary polytopes. Computer Graphics Forum, 27(5):1513–1520, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hensley:2005:TFS

[HSC^{+}05] Justin Hensley, Thorsten Scheuermann, Greg Coombe, Montek Singh, and Anselmo Lastra. Texture: Fast summed-area table generation and its applications. Computer Graphics Forum, 24 (3):547–555, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hadwiger:2005:RRT

[HSS^{+}05] Markus Hadwiger, Christian Sigg, Henning Scharsach, Khatja Bühler, and Markus Gross. Rendering I: Real-time ray-casting and advanced shading of discrete isosurfaces. Computer Graphics Forum, 24(3):303–312, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hasler:2009:MCS

[HSS^{+}09] N. Hasler, C. Stoll, M. Sunkel, B. Rosenhahn, and H.-P. Seidel. Motion compression: a statistical model of human pose and body shape. Computer Graphics Forum, 28(2):337–346, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Higashi:1996:GSS

[HTH96] Masatake Higashi, Hideo Tsutamori, and Mamoru Hosaka. Generation of smooth surfaces by controlling curvature variation. Computer Graphics Forum, 15(3):C187–C196, September 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2009:AVT

[HTSFP09] Dandan Huang, Melanie Tory, Sheryl Staub-French, and Rachel Pottinger. Applications: Visualization techniques for schedule comparison. Computer Graphics Forum, 28(3):951–958, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hubl:1993:NCO

[Hub93] J. Hubl. A note on 3D-clip optimisation. Computer Graphics Forum, 12(2):159–160, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holten:2008:VTV

[HV08] Danny Holten and Jarke J. Van Wijk. Visualization tools: Visual comparison of hierarchically organized data. Computer Graphics Forum, 27(3):759–766, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holten:2009:GVF

[HV09] Danny Holten and Jarke J. Van Wijk. Graph visualization: Forcedirected edge bundling for graph visualization. Computer Graphics Forum, 28(3):983–990, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hasan:2008:GIT

[HVAPB08] Miloš Hašan, Edgar Velázquez-Armendáriz, Fabio Pellacini, and Kavita Bala. Global illumination: Tensor clustering for rendering many-light animations. Computer Graphics Forum, 27(4):1105– 1114, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hermans:2008:VRA

[HVM^{+}08] C. Hermans, C. Vanaken, T. Mertens, F. Van Reeth, and P. Bekaert. Video and relighting: Augmented panoramic video. Computer Graphics Forum, 27(2):281–290, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hunter:1991:CQE

[HW91] Andrew Hunter and Philip Willis. Classification of quad-encoding techniques. Computer Graphics Forum, 10(2):97–112, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2009:SAH

[HWAG09] Qi-Xing Huang, Martin Wicke, Bart Adams, and Leonidas Guibas. Shape analysis and helices: Shape decomposition using modal analysis. Computer Graphics Forum, 28(2):407–416, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ho:2005:GIC

[HWC^{+}05] Chien.-Chang Ho, Fu-Che Wu, Bing-Yu Chen, Yung-Yu Chuang, and Ming Ouhyoung. Geometry III: Cubical marching squares: Adaptive feature preserving surface extraction from volume data. Computer Graphics Forum, 24(3):537–545, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:1998:RTB

[hZCK98] Qing hong Zhu, Yan Chen, and Arie Kaufman. Real-time biomechanically-based muscle volume deformation using FEM. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=274&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Iwasaki:2002:EMR

[IDN02] Kei Iwasaki, Yoshinori Dobashi, and Tomoyuki Nishita. An efficient method for rendering underwater optical effects using graphics hardware. Computer Graphics Forum, 21(4):701–711, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwasaki:2003:FRM

[IDN03a] Kei Iwasaki, Yoshinori Dobashi, and Tomoyuki Nishita. A fast rendering method for refractive and reflective caustics due to water surfaces. Computer Graphics Forum, 22(3):601–609, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwasaki:2003:GIF

[IDN03b] Kei Iwasaki, Yoshinori Dobashi, and Tomoyuki Nishita. Global illumination: a fast rendering method for refractive and reflective caustics due to water surfaces. Computer Graphics Forum, 22(3): 601–609, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isenberg:2008:USU

[IEGC08] Tobias Isenberg, Maarten H. Everts, Jens Grubert, and Sheelagh Carpendale. User studies and user interaction: Interactive exploratory visualization of 2D vector fields. Computer Graphics Forum, 27(3):983–990, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isenberg:2009:TVC

[IF09] Petra Isenberg and Danyel Fisher. Text visualization: Collaborative brushing and linking for co-located visual analytics of document collections. Computer Graphics Forum, 28(3):1031–1038, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isenberg:2002:SSI

[IHS02] Tobias Isenberg, Nick Halper, and Thomas Strothotte. Stylizing silhouettes at interactive rates: From silhouette edges to silhouette strokes. Computer Graphics Forum, 21(3):249–258, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Igarashi:2008:SFK

[IIS08] Yuki Igarashi, Takeo Igarashi, and Hiromasa Suzuki. Skull, face, and knits: Knitting a 3D model. Computer Graphics Forum, 27 (7):1737–1743, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Igarashi:2009:ADI

[IIS09] Yuki Igarashi, Takeo Igarashi, and Hiromasa Suzuki. Abstraction and design: Interactive cover design considering physical constraints. Computer Graphics Forum, 28(7):1965–1973, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ivanov:2000:CDN

[IK00] Denis V. Ivanov and Yevgeniy P. Kuzmin. Color distribution — a new approach to texture compression. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=420&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Ivanov:2001:SPP

[IK01a] D. Ivanov and Ye. Kuzmin. Spatial patches — a primitive for 3D model representation. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwakiri:2001:PBR

[IK01b] Yuya Iwakiri and Toyohisa Kaneko. PC-based real-time texture painting on real world objects. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ibarria:2003:CCC

[ILRS03a] Lawrence Ibarria, Peter Lindstrom, Jarek Rossignac, and Andrzej Szymczak. Compression: Out-of-core compression and decompression of large n-dimensional scalar fields. Computer Graphics Forum, 22(3):343–348, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ibarria:2003:CCD

[ILRS03b] Lawrence Ibarria, Peter Lindstrom, Jarek Rossignac, and Andrzej Szymczak. Out-of-core compression and decompression of large n-dimensional scalar fields. Computer Graphics Forum, 22(3): 343–348, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ijiri:2008:TAE

[IMIM08] Takashi Ijiri, Radomír Mêch, Takeo Igarashi, and Gavin Miller. Textures and arrangements: an example-based procedural system for element arrangement. Computer Graphics Forum, 27(2):429– 436, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ijiri:2006:MSI

[IOI06] T. Ijiri, S. Owada, and T. Igarashi. Modelling I: Seamless integration of initial sketching and subsequent detail editing in flower modeling. Computer Graphics Forum, 25(3):617–624, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ihm:1999:WBC

[IP99] Insung Ihm and Sanghun Park. Wavelet-based 3D compression scheme for interactive visualization of very large volume data. Computer Graphics Forum, 18(1):??, March 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=298&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=18.

Iehl:2000:ASR

[IP00] Jean Claude Iehl and Bernard Péroche. An adaptive spectral rendering with a perceptual control. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=421&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Isenburg:2001:TSC

[Ise01] Martin Isenburg. Triangle strip compression. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ijiri:2009:MDP

[ITYI09] Takashi Ijiri, Kenshi Takayama, Hideo Yokota, and Takeo Igarashi. Modeling and deformation: ProcDef: Local-to-global deformation for skeleton-free character animation. Computer Graphics Forum, 28(7):1821–1828, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Juan-Arinyo:1995:DEI

[JA95] Robert Juan-Arinyo. Domain extension of isothetic polyhedra with minimal CSG representation. Computer Graphics Forum, 14(5):281–293, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jackel:1985:GPS

[Jac85] D. Jackel. The graphics PARCUM system: a 3D memory based computer architecture for processing and display of solid models. Computer Graphics Forum, 4(1):21–32, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jansen:1989:WNR

[Jan89] F. W. Jansen. What’s new? — report on SIGGRAPH ’89. Computer Graphics Forum, 8(4):369, December 1989. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jansen:1991:RSE

[Jan91] Frederik W. Jansen. Report on the Second Eurographics Workshop on Rendering, Barcelona, 13–15 May 1991. Computer Graphics Forum, 10(4):353–355, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jainek:2008:IRV

[JBB^{+}08] W. M. Jainek, S. Born, D. Bartz, W. Straßer, and J. Fischer. Illustrative and realistic volume rendering: Illustrative hybrid visualization and exploration of anatomical and functional brain data. Computer Graphics Forum, 27(3):855–862, May 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jang:2006:VEI

[JBL^{+}06] Yun Jang, Ralf P. Botchen, Andreas Lauser, David S. Ebert, Kelly P. Gaither, and Thomas Ertl. Visualization: Enhancing the interactive visualization of procedurally encoded multifield data with ellipsoidal basis functions. Computer Graphics Forum, 25(3): 587–596, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Johnson:2006:ITM

[JBS^{+}06] M. Johnson, G. J. Brostow, J. Shotton, O. Arandjelovic, V. Kwatra, and R. Cipolla. Images and tone mapping: Semantic photo synthesis. Computer Graphics Forum, 25(3):407–413, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2008:MVA

[JBTS08] H. Jänicke, M. Böttinger, X. Tricoche, and G. Scheuermann. Multivariate visualization: Automatic detection and visualization of distinctive structures in 3D unsteady multi-fields. Computer Graphics Forum, 27(3):767–774, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jones:1994:NAC

[JC94] M. W. Jones and Min Chen. A new approach to the construction of surfaces from contour data. Computer Graphics Forum, 13(3): C/75–C/84, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Johansson:2008:PCS

[JC08] J. Johansson and M. Cooper. Parallel coordinates: a screen space quality method for data abstraction. Computer Graphics Forum, 27(3):1039–1046, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarosz:2009:JMC

[JCJ09] Wojciech Jarosz, Nathan A. Carr, and Henrik Wann Jensen. The joy of math and cartoon drawing: Importance sampling spherical harmonics. Computer Graphics Forum, 28(2):577–586, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

James:1998:PFD

[JD98] A. James and A. M. Day. The priority face determination tree for hidden surface removal. Computer Graphics Forum, 17(1):55–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=215&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

James:2000:CNB

[JD00] A. James and A. M. Day. Conflict neutralization on binary space partitioning. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See erratum [JD01].

James:2001:ECN

[JD01] A. James and A. M. Day. Erratum: Conflict Neutralization on Binary Space Partitioning. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See [JD00].

Jensen:1997:RCN

[Jen97] Henrik Wann Jensen. Rendering caustics on non-Lambertian surfaces. Computer Graphics Forum, 16(1):57–64, March 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=116&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=16.

Jensen:2007:ITR

[Jen07] Henrik Wann Jensen. Invited talks: Reverse engineering nature. Computer Graphics Forum, 26(3):xvii, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jimenez:2000:EAC

[JEO00] W. F. H. Jiménez, C. Esperança, and A. A. F. Oliveira. Efficient algorithms for computing conservative portal visibility information. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=441&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Jimenez:2009:ROA

[JFS09] J. J. Jiménez, F. R. Feito, and R. J. Segura. Robust and optimized algorithms for the point-in-polygon inclusion test without pre-processing. Computer Graphics Forum, 28(8):2264–2274, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jimenez:2006:POC

[JFSO06] Juan José Jiménez, Francisco R. Feito, Rafael J. Segura, and Carlos J. Ogáyar. Particle oriented collision detection using simplicial coverings and tetra-trees. Computer Graphics Forum, 25(1):53– 68, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Julius:2005:GID

[JKS05] Dan Julius, Vladislav Kraevoy, and Alla Sheffer. Geometry IV: Dcharts: Quasi-developable mesh segmentation. Computer Graphics Forum, 24(3):581–590, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jardillier:1998:SSC

[JL98] Frank Jardillier and Eric Languénou. Screen-space constraints for camera movements: the virtual cameraman. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=265&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Jobard:2000:UFV

[JL00] Bruno Jobard and Wilfrid Lefer. Unsteady flow visualization by animating evenly-spaced streamlines. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=395&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Jacobs:2006:CIM

[JL06] Katrien Jacobs and Céline Loscos. Classification of illumination methods for mixed reality. Computer Graphics Forum, 25(1):29– 51, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ju:2008:CAE

[JL08] Eunjung Ju and Jehee Lee. Character animation: Expressive facial gestures from motion capture data. Computer Graphics Forum, 27(2):381–388, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ji:2006:GEM

[JLCW06] Zhongping Ji, Ligang Liu, Zhonggui Chen, and Guojin Wang. Geometry: Easy mesh cutting. Computer Graphics Forum, 25 (3):283–291, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janoos:2009:BVV

[JNM^{+}09] Firdaus Janoos, Boonthanome Nouanesengsy, Raghu Machiraju, Han Wei Shen, Steffen Sammet, Michael Knopp, and István Á. Mórocz. Biomedical visualization: Visual analysis of brain activity from fMRI data. Computer Graphics Forum, 28(3):903–910, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janoos:2008:VAL

[JNX^{+}08] Firdaus Janoos, Boonthanome Nouansengsy, Xiaoyin Xu, Raghu Machiraju, and Stephen T. C. Wong. Visualization applications in life sciences: Classification and uncertainty visualization of dendritic spines from optical microscopy imaging. Computer Graphics Forum, 27(3):879–886, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jones:1990:DGB

[Jon90] Huw Jones. Dürer, gaskets and Barnsley’s chaos game. Computer Graphics Forum, 9(4):327–332, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jones:1996:PVD

[Jon96] Mark W. Jones. The production of volume data from triangular meshes using voxelisation. Computer Graphics Forum, 15(5):311– 318, December 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Joosten:1986:ITP

[Joo86] Roy Joosten. Integration of town planning, landscaping and 3D architecture: Results and future enhancements since CAMP ’83. Computer Graphics Forum, 5(1):65–69, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Joshi:2008:USU

[JR08] Alark Joshi and Penny Rheingans. User studies and user interaction: Evaluation of illustration-inspired techniques for timevarying data visualization. Computer Graphics Forum, 27(3): 999–1006, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

John:1989:CAS

[JW89] Nigel W. John and Philip J. Willis. The Controller animation system. Computer Graphics Forum, 8(2):133–138, June 1989. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Johnstone:1995:RMS

[JW95] J. K. Johnstone and J. P. Williams. A rational model of the surface swept by a curve. Computer Graphics Forum, 14(3): C/77–C/88, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jackel:1997:MRA

[JW97] D. Jackèl and B. Walter. Modeling and rendering of the atmosphere using Mie-scattering. Computer Graphics Forum, 16 (4):201–210, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=180&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Jones:2001:JLA

[JW01] Huw Jones and Phil Willis. The John Lansdown Award. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jenke:2006:GRB

[JWB^{+}06] P. Jenke, M. Wand, M. Bokeloh, A. Schilling, and W. Straßer. Geometry reconstruction: Bayesian point cloud reconstruction. Computer Graphics Forum, 25(3):379–388, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2009:SII

[JZF^{+}09] Dong Hyun Jeong, Caroline Ziemkiewicz, Brian Fisher, William Ribarsky, and Remco Chang. Scatterplots: iPCA: an interactive system for PCA-based visual analytics. Computer Graphics Forum, 28(3):767–774, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarosz:2008:PMI

[JZJ08a] Wojciech Jarosz, Matthias Zwicker, and Henrik Wann Jensen. Participating media: Irradiance gradients in the presence of participating media and occlusions. Computer Graphics Forum, 27 (4):1087–1096, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarosz:2008:RVE

[JZJ08b] Wojciech Jarosz, Matthias Zwicker, and Henrik Wann Jensen. Rendering volume effects: The beam radiance estimate for volumetric photon mapping. Computer Graphics Forum, 27(2):557– 566, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kallmann:2003:CMP

[KAAT03a] Marcelo Kallmann, Amaury Aubel, Tolga Abaci, and Daniel Thalmann. Collision and motion: Planning collision-free reaching motions for interactive object manipulation and grasping. Computer Graphics Forum, 22(3):313–322, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kallmann:2003:PCF

[KAAT03b] Marcelo Kallmann, Amaury Aubel, Tolga Abaci, and Daniel Thalmann. Planning collision-free reaching motions for interactive object manipulation and grasping. Computer Graphics Forum, 22 (3):313–322, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kautz:2004:SAR

[Kau04] Jan Kautz. States of the art review: Hardware lighting and shading: a survey. Computer Graphics Forum, 23(1):85–112, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kautz:2007:RYR

[Kau07] Jan Kautz. Reports: Young Researcher Award 2007. Computer Graphics Forum, 26(4):857, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuijk:1989:FPS

[KB89] A. A. M. Kuijk and E. H. Blake. Faster Phong shading via angular interpolation. Computer Graphics Forum, 8(4):315–324, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kent:1990:RE

[KB90] P. Kent and R. D. Bergeron. Report on EUROGRAPHICS’90. Computer Graphics Forum, 9(4):365–370, December 1990. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalra:1992:MTE

[KB92] D. Kalra and A. H. Barr. Modeling with time and events in computer animation. Computer Graphics Forum, 11(3):C45–C58, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koch:1998:EEU

[KB98] Rolf M. Koch and Albert A. Bosshard. Emotion editing using finite elements. Computer Graphics Forum, 17 (3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=276&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Kobbelt:2000:IAP

[KB00] Leif P. Kobbelt and Mario Botsch. An interactive approach to point cloud triangulation. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=440&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Krishnaswamy:2004:NPA

[KB04] Aravind Krishnaswamy and Gladimir V. G. Baranoski. Natural phenomena and aging: a biophysically-based spectral model of light interaction with human skin. Computer Graphics Forum, 23 (3):331–340, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kristof:2009:FBH

[KBKŠ09] P. Krištof, B. Beneš, J. Křivánek, and O. Št’ava. Fluids and beyond: Hydraulic erosion using smoothed particle hydrodynamics. Computer Graphics Forum, 28(2):219–228, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:2000:MSD

[KBS00] Leif P. Kobbelt, Thilo Bareuther, and Hans-Peter Seidel. Multiresolution shape deformations for meshes with dynamic vertex connectivity. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=417&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=19.

Krivanek:2008:LIR

[KC08] Jaroslav Křivánek and Mark Colbert. Local illumination: Realtime shading with filtered importance sampling. Computer Graphics Forum, 27(4):1147–1154, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakez:1997:VDE

[KCB97] Samory Kakez, Vania Conan, and Pascal Bisson. Virtually documented environments: a new interface paradigm for taskoriented access to information. Computer Graphics Forum, 16 (3):C319–C327, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=169&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Kim:2006:GCD

[KCL06] Junho Kim, Sungyul Choe, and Seungyong Lee. Geometry compression and decompression: Multiresolution random accessible mesh compression. Computer Graphics Forum, 25(3):323–331, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kazinnik:1997:ODN

[KE97] Roman Kazinnik and Gershon Elber. Orthogonal decomposition of non-uniform bspline spaces using wavelets. Computer Graphics Forum, 16(3):C27–C38, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=139&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Keiser:2004:REI

[Kei04] Richard Keiser. Reports: EG/IEEE Symposium on Point-Based Graphics: Zurich, Switzerland, 2–4 June 2004. Computer Graphics Forum, 23(4):844, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kelly:1986:FTC

[Kel86] Mary B. Kelly. First time classroom computer graphics. Computer Graphics Forum, 5(3):217–220, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karni:2009:MSE

[KFG09] Z. Karni, D. Freedman, and C. Gotsman. Modeling and sculpting: Energy-based image deformation. Computer Graphics Forum, 28 (5):1257–1268, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuroda:1994:CLI

[KFK94] M. Kuroda, S. Furukawa, and F. Kimura. Controllable locality in C^{2} interpolating curves by B2-splines/S-splines. Computer Graphics Forum, 13(1):49–55, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krishnan:1998:RAC

[KGL^{+}98] S. Krishnan, M. Gopi, M. Lin, D. Manocha, and A. Pattekar. Rapid and accurate contact determination between spline models using ShellTrees. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=278&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Krishnan:1997:IBC

[KGMM97] S. Krishnan, M. Gopi, D. Manocha, and M. Mine. Interactive boundary computation of Boolean combinations of sculptured solids. Computer Graphics Forum, 16(3):C67– C78, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=143&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Koh:1992:FGS

[KH92] Eng-Kiat Koh and D. D. Hearn. Fast generation and surface structuring methods for terrain and other natural phenomena. Computer Graphics Forum, 11(3):C169–C180, C472–C473, ????

1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Keates:1995:IRT

[KH95] M. J. Keates and R. J. Hubbold. Interactive ray tracing on a virtual shared-memory parallel computer. Computer Graphics Forum, 14(4):189–202, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kochhar:1996:UOO

[KH96] Sandeep Kochhar and Jim Hall. A unified, object-oriented graphics system and software architecture for visualising CAD/CAM presentations. Computer Graphics Forum, 15(4):229–248, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2002:PTM

[KH02] Dongho Kim and James K. Hahn. Projective texture mapping with full panorama. Computer Graphics Forum, 21(3):421–430, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2009:DSC

[KHH^{+}09] Duksu Kim, Jae-Pil Heo, Jaehyuk Huh, John Kim, and Sung eui Yoon. Data structure and collision detection: HPCCD: Hybrid parallel continuous collision detection using CPUs and GPUs. Computer Graphics Forum, 28(7):1791–1800, October 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kitamura:2001:ICA

[KHIK01] Yoshifumi Kitamura, Tomohiko Higashi, Takayuki Iida, and Fumio Kishino. Interactive computer animation of hand gestures using status estimation with multiple regression analysis. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Knoll:2009:FRT

[KHK^{+}09] A. Knoll, Y. Hijazi, A. Kensler, M. Schott, C. Hansen, and H. Hagen. Fast ray tracing of arbitrary implicit surfaces with interval and affine arithmetic. Computer Graphics Forum, 28(1):26–40, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kniep:2009:REA

[KHM09] S. Kniep, S. Häring, and M. Magnor. Rendering I: Efficient and accurate rendering of complex light sources. Computer Graphics Forum, 28(4):1073–1081, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karpenko:2002:FFS

[KHR02] Olga Karpenko, John F. Hughes, and Ramesh Raskar. Free-form sketching with variational implicit surfaces. Computer Graphics Forum, 21(3):585–594, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kidd:1982:IE

[Kid82] R. M. Kidd. Impressions of EUROGRAPHICS ’82. Computer Graphics Forum, 1(4):177–178, December 1982. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kidd:1984:EUC

[Kid84] R. M. Kidd. EUROGRAPHICS 2nd UK conference (report). Computer Graphics Forum, 3(2):191–192, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kilgour:1982:OG

[Kil82] A. C. Kilgour. An overview of 3D graphics. Computer Graphics Forum, 1(3):143–152, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kilgour:1985:RCA

[Kil85] A. Kilgour. Report on the CREST advanced course on 3D computer graphics University of Glasgow, 27th August — 7th September, 1984. Computer Graphics Forum, 4(1):59–60, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

King:1992:PED

[Kin92] T. King. Pandora: an experiment in distributed multimedia. Computer Graphics Forum, 11(3):C23–C34, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

King:1995:SCO

[Kin95] M. King. Syntax channelling and other issues affecting innovation in the graphical user interface. Computer Graphics Forum, 14(3):

C/43–C/54, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kok:1992:ASA

[KJ92] A. J. F. Kok and F. W. Jansen. Adaptive sampling of area light sources in ray tracing including diffuse interreflection. Computer Graphics Forum, 11(3):C289–C298, C479, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karnick:2009:SGD

[KJC^{+}09] P. Karnick, S. Jeschke, D. Cline, A. Razdan, E. Wentz, and P. Wonka. A shape grammar for developing glyph-based visualizations. Computer Graphics Forum, 28(8):2176–2188, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1983:MCT

[Kje83] L. Kjelldahl. A method for choosing tickmarks on a nonlinear scale. Computer Graphics Forum, 2(4):243–248, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1989:DTC

[Kje89] L. Kjelldahl. Doctoral theses in computer graphics. Computer Graphics Forum, 8(4):353–358, December 1989. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1990:DTC

[Kje90] Lars Kjelldahl. Doctoral theses in computer graphics. Computer Graphics Forum, 9(3):269–271, September 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1991:SRE

[Kje91a] L. Kjelldahl. A short report from the Eurographics Multimedia Workshop, 18–19 April, 1991, held in Stockholm. Computer Graphics Forum, 10(3):260, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1991:DTCa

[Kje91b] Lars Kjelldahl. Doctoral theses in computer graphics. Computer Graphics Forum, 10(1):71–74, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1991:DTCb

[Kje91c] Lars Kjelldahl. Doctoral theses in computer graphics. Computer Graphics Forum, 10(3):270–272, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1991:DTCc

[Kje91d] Lars Kjelldahl. Doctoral theses in computer graphics. Computer Graphics Forum, 10(4):343–344, December 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kjelldahl:1992:MSI

[Kje92] L. Kjelldahl, editor. Multimedia: systems, interaction, and applications, EurographicSeminars. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55201-4. LCCN QA76.76.H92 M85 1992. Record of the first Eurographics Workshop on Multimedia, held at the center for Numerical Analysis and Computing Science … Royal Institute of Technology.

Kjelldahl:1995:DT

[Kje95] L. Kjelldahl. Doctoral theses. Computer Graphics Forum, 14(1): 67–??, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kollig:2002:EMS

[KK02] Thomas Kollig and Alexander Keller. Efficient multidimensional sampling. Computer Graphics Forum, 21(3):557–564, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2007:FLS

[KK07] Ig-Jae Kim and Hyeong-Seok Ko. Faces: 3D lip-synch generation with data-faithful machine learning. Computer Graphics Forum, 26(3):295–301, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2008:IPC

[KK08] Min H. Kim and Jan Kautz. Imaging and projectors: Characterization for high dynamic range imaging. Computer Graphics Forum, 27(2):691–697, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kyprianidis:2009:ADI

[KKD09] Jan Eric Kyprianidis, Henry Kang, and Jürgen Döllner. Abstraction and design: Image and video abstraction by anisotropic Kuwahara filtering. Computer Graphics Forum, 28(7):1955–1963, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kitaoka:2009:REL

[KKK09] Shinya Kitaoka, Yoshifumi Kitamura, and Fumio Kishino. Replica exchange light transport. Computer Graphics Forum, 28(8):2330– 2342, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2003:PSC

[KL03] Ku-Jin Kim and In-Kwon Lee. The perspective silhouette of a canal surface. Computer Graphics Forum, 22(1):15–22, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kang:2008:IES

[KL08] Henry Kang and Seungyong Lee. Image editing: Shapesimplifying image abstraction. Computer Graphics Forum, 27(7): 1773–1780, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klassen:2000:FJ

[Kla00] R. Victor Klassen. Filtered jitter. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kopf:2009:ITL

[KLD^{+}09] Johannes Kopf, Dani Lischinski, Oliver Deussen, Daniel Cohen-Or, and Michael Cohen. Images and textures: Locally adapted projections to reduce panorama distortions. Computer Graphics Forum, 28(4):1083–1089, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klein:2006:IPD

[Kle06] Reinhard Klein. Invited presentations: Data preparation for realtime high quality rendering of complex models. Computer Graphics Forum, 25(3):xviii, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klok:1987:ESG

[Klo87] Fopke Klok. Evaluation of standard graphics packages. Computer Graphics Forum, 6(4):299–307, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koparkar:1983:DPP

[KM83] P. A. Koparkar and S. P. Mudur. The development of programs for the processing of parametric curves. Computer Graphics Forum, 2 (2-3):135–144, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kulpa:2005:AMI

[KMA05] R. Kulpa, F. Multon, and B. Arnaldi. Animation: Morphologyindependent representation of motions for interactive human-like animation. Computer Graphics Forum, 24(3):343–351, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurtenbach:1994:CAG

[KMB94] G. Kurtenbach, T. P. Moran, and W. Buxton. Contextual animation of gestural commands. Computer Graphics Forum, 13(5): 305–314, December 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karner:1996:IBM

[KMG96] Konrad F. Karner, Heinz Mayer, and Michael Gervautz. An image based measurement system for anisotropic reflection. Computer Graphics Forum, 15(3):C119–C128, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakimoto:2005:GGB

[KMN^{+}05] Masanori Kakimoto, Kaoru Matsuoka, Tomoyuki Nishita, Takeshi Naemura, and Hiroshi Harashima. Glare generation based on wave optics. Computer Graphics Forum, 24(2):185–193, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kwatra:2008:FSF

[KMN^{+}08] Vivek Kwatra, Philippos Mordohai, Rahul Narain, Sashi Kumar Penta, Mark Carlsonk, Marc Pollefeys, and Ming C. Lin. Fluids and simulation: Fluid in video: Augmenting real video with simulated fluids. Computer Graphics Forum, 27(2):487–496, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krawczyk:2005:PLP

[KMS05] Grzegorz Krawczyk, Karol Myszkowski, and Hans-Peter Seidel. Perception: Lightness perception in tone reproduction for high dynamic range images. Computer Graphics Forum, 24(3):635–645, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krawczyk:2007:ICR

[KMS07] Grzegorz Krawczyk, Karol Myszkowski, and Hans-Peter Seidel. Imaging (1): Contrast restoration by adaptive countershading. Computer Graphics Forum, 26(3):581–590, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kshirsagar:2003:FSV

[KMT03a] Sumedha Kshirsagar and Nadia Magnenat-Thalmann. Faces and speech: Visyllable based speech animation. Computer Graphics Forum, 22(3):631–639, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kshirsagar:2003:VBS

[KMT03b] Sumedha Kshirsagar and Nadia Magnenat-Thalmann. Visyllable based speech animation. Computer Graphics Forum, 22(3): 631–639, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalra:1992:SFM

[KMTT92] P. Kalra, A. Mangili, N. M. Thalmann, and D. Thalmann. Simulation of facial muscle actions based on rational free form deformations. Computer Graphics Forum, 11(3):C59–C69, C466, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Knittel:1993:VVE

[Kni93] G. Knittel. Verve-voxel engine for real-time visualization and examination. Computer Graphics Forum, 12(3):C37–C48, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalberer:2007:SQS

[KNP07] Felix Kälberer, Matthias Nieser, and Konrad Polthier. Surfaces: QuadCover — surface parameterization using branched coverings. Computer Graphics Forum, 26(3):375–384, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaya:1988:CPI

[KO88] Aydin Kaya and Bulent Ozguc. Continuous processing of images through user sketched functional blocks. Computer Graphics Forum, 7(4):273–280, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:1996:ISO

[Kob96] Leif Kobbelt. Interpolatory subdivision on open quadrilateral nets with arbitrary topology. Computer Graphics Forum, 15(3):C409– C420, C485, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:2003:FSR

[Kob03a] Leif Kobbelt. Freeform shape representations for efficient geometry processing. Computer Graphics Forum, 22(3):??, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:2003:IPF

[Kob03b] Leif Kobbelt. Invited presentation: Freeform shape representations for efficient geometry processing. Computer Graphics Forum, 22(3):xviii, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:2005:GSR

[Kob05] Jianhua Wu Leif Kobbelt. Geometry I: Structure recovery via hybrid variational surface approximation. Computer Graphics Forum, 24(3):277–284, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Keefe:2008:VAL

[KOB^{+}08] D. F. Keefe, T. M. O’Brien, D. B. Baier, S. M. Gatesy, E. L. Brainerd, and D. H. Laidlaw. Visualization applications in life sciences: Exploratory visualization of animal kinematics using instantaneous helical axes. Computer Graphics Forum, 27(3):863– 870, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koch:1993:ARB

[Koc93] R. Koch. Automatic reconstruction of buildings from stereoscopic image sequences. Computer Graphics Forum, 12(3):C339–C350, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kolingerova:2008:CGE

[Kol08] I. Kolingerovà. Computational geometry education for computer graphics students. Computer Graphics Forum, 27(6):1531–1538, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Konstanz:2006:REW

[Kon06] Oliver Deussen Konstanz. Reports: 16th Eurographics Workshop on Rendering Konstanz, Germany, 29th June–1st July 2005. Computer Graphics Forum, 25(1):141, March 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krishnan:1987:SAB

[KP87] D. Krishnan and L. M. Patnaik. Systolic architecture for Boolean operations on polygons and polyhedra. Computer Graphics Forum, 6(3):203–210, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kang:2001:TPU

[KPiAS01] Hyung Woo Kang, Soon Hyoung Pyo, Ken ichi Anjyo, and Sung Yong Shin. Tour into the picture using a vanishing line and its extension to panoramic images. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalberer:2005:GIF

[KPRW05] Felix Kälberer, Konrad Polthier, Ulrich Reitebuch, and Max Wardetzky. Geometry II: FreeLence — coding with free valences. Computer Graphics Forum, 24(3):469–478, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kolb:1995:FSR

[KPS95] A. Kolb, H. Pottmann, and H.-P. Seidel. Fair surface reconstruction using quadratic functionals. Computer Graphics Forum, 14 (3):C/469–C/479, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ku:2008:NPR

[KQWM08] Daychyi Ku, Shengfeng Qin, David K. Wright, and Cuixia Ma. Non-photorealistic rendering: Online personalised nonphotorealistic rendering technique for 3D geometry from incremental sketching. Computer Graphics Forum, 27(7):1861–1868, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2005:GGG

[KR05] ByungMoon Kim and Jarek Rossignac. Geometry I: GeoFilter: Geometric selection of mesh filter parameters. Computer Graphics Forum, 24(3):295–302, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krammer:1989:NMP

[Kra89] Gergely Krammer. Notes on the mathematics of the PHIGS viewing pipeline. Computer Graphics Forum, 8(3):219–226, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krammer:1992:LCA

[Kra92] G. Krammer. A line clipping algorithm and its analysis. Computer Graphics Forum, 11(3):C253–C266, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kry:2009:MSM

[KRFC09] P. G. Kry, L. Reveret, F. Faure, and M.-P. Cani. Motion synthesis: Modal locomotion: Animating virtual characters with natural vibrations. Computer Graphics Forum, 28(2):289–298, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kipfer:2003:LEP

[KRG03] Peter Kipfer, Frank Reck, and Günther Greiner. Local exact particle tracing on unstructured grids. Computer Graphics Forum, 22(2):133–142, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krokos:1992:ISC

[KS92] M. A. Krokos and M. Slater. Interactive shape control of interpolating B-splines. Computer Graphics Forum, 11(3):C435–C447, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kraus:2007:IDF

[KS07] M. Kraus and M. Strengert. Imaging (2): Depth-of-field rendering by pyramidal image processing. Computer Graphics Forum, 26(3): 645–654, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuhme:1992:STO

[KSH92] T. Kuhme and M. Schneider-Hufschmidt. SX/Tools — an open design environment for adaptable multimedia user interfaces. Computer Graphics Forum, 11(3):C93–C105, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:2004:RSG

[KSH04] Leif Kobbelt, Peter Schöder, and Hugues Hoppe. Reports: Symposium on Geometry Processing: Aachen, Germany, 23–25 June 2003. Computer Graphics Forum, 23(1):119–120, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Komura:1997:MBFa

[KSK97a] Taku Komura, Yoshihisa Shinagawa, and Tosiyasu L. Kunii. A muscle-based feed-forward controller of the human body. Computer Graphics Forum, 16(3):C165–??, September 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=153&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Komura:1997:MBFb

[KSK97b] Taku Komura, Yoshihisa Shinagawa, and Tosiyasu L. Kunii. Muscle-based feed-forward controller of the human body. Computer Graphics Forum, 16(3):C165–C176, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kelemen:2002:SRM

[KSKAC02] Csaba Kelemen, László Szirmay-Kalos, György Antal, and Ferenc Csonka. A simple and robust mutation strategy for the Metropolis light transport algorithm. Computer Graphics Forum, 21(3): 531–540, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2008:ARF

[KSL^{+}08] Dongyeon Kim, Minjung Son, Yunjin Lee, Henry Kang, and Seungyong Lee. Artistic rendering: Feature-guided image stippling. Computer Graphics Forum, 27(4):1209–1216, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kanamori:2008:GRR

[KSN08] Yoshihiro Kanamori, Zoltan Szego, and Tomoyuki Nishita. GPU rendering and ray-casting: GPU-based fast ray casting for a large number of metaballs. Computer Graphics Forum, 27(2):351–360, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:1997:USH

[KSS97] Leif Kobbelt, Marc Stamminger, and Hans-Peter Seidel. Using subdivision on hierarchical data to reconstruct radiosity distribution. Computer Graphics Forum, 16(3):C347–C355, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=172&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Kuriyama:1997:PSM

[KT97] Shigeru Kuriyama and Keihachiro Tachibana. Polyhedral surface modeling with a diffusion system. Computer Graphics Forum, 16 (3):C39–C46, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=140&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Kaplansky:2009:DGP

[KT09] Lotan Kaplansky and Ayellet Tal. Digital geometry processing: Mesh segmentation refinement. Computer Graphics Forum, 28(7): 1995–2003, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakimoto:2007:IIS

[KTMN07] Masanori Kakimoto, Tomoaki Tatsukawa, Yukiteru Mukai, and Tomoyuki Nishita. Imaging (2): Interactive simulation of the human eye depth of field and its correction by spectacle lenses. Computer Graphics Forum, 26(3):627–636, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuhlmann:1982:N

[Kuh82] H. W. Kuhlmann. NCGA ’82. Computer Graphics Forum, 1(4): 175, December 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuijk:1991:RSE

[Kui91] A. A. M. Kuijk. Report on the Sixth Eurographics Workshop on Graphics Hardware. Computer Graphics Forum, 10(4):363–364, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kunz:2004:RIW

[Kun04] Andreas Kunz. Reports: 7^{th} International Workshop on Immersive Projection Technology and 9^{th} Eurographics Workshop on Virtual Environments: Zurich, Switzerland, 22–23 May, 2003. Computer Graphics Forum, 23(1):113–114, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuzmin:1990:ECD

[Kuz90] Yevgeni P. Kuzmin. An efficient circle-drawing algorithm. Computer Graphics Forum, 9(4):333–336, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuzmin:1994:RTS

[Kuz94] Y. P. Kuzmin. Ray traversal of spatial structures. Computer Graphics Forum, 13(4):223–227, October 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuzmin:1995:BLG

[Kuz95] Yevgeny P. Kuzmin. Bresenham’s line generation algorithm with built-in clipping. Computer Graphics Forum, 14(5):275–280, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobbelt:1999:SWA

[KVLS99] Leif P. Kobbelt, Jens Vorsatz, Ulf Labsik, and Hans-Peter Seidel. A shrink wrapping approach to remeshing polygonal surfaces. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=333&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Kruger:2005:GHG

[KW05] Jens Krüger and Rüdiger Westermann. GPU & hardware: GPU simulation and rendering of volumetric effects for computer games and virtual environments. Computer Graphics Forum, 24(3): 685–693, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kunimatsu:2001:FSR

[KWF^{+}01] A. Kunimatsu, Y. Watanabe, H. Fujii, T. Saito, K. Hiwada, T. Takahashi, and H. Ueki. Fast simulation and rendering techniques for fluid objects. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kwiers:1989:RTE

[Kwi89] R. S. S. Pieters Kwiers. Report on the Third EUROGRAPHICS Workshop on Intelligent CAD. Computer Graphics Forum, 8(3): 267–268, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2008:FSS

[KySK08] Doyub Kim, Oh young Song, and Hyeong-Seok Ko. Fluids and simulation: a semi-Lagrangian CIP fluid solver without dimensional splitting. Computer Graphics Forum, 27(2):467–475, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klein:2004:DFC

[KZ04] Jan Klein and Gabriel Zachmann. Distance fields and collision detection: Point cloud collision detection. Computer Graphics Forum, 23(3):567–576, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kavan:2005:AFC

[KŽ05] L. Kavan and J. Žára. Animation: Fast collision detection for skeletally deformable models. Computer Graphics Forum, 24(3): 363–372, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalbe:2008:GRR

[KZ08] Thomas Kalbe and Frank Zeilfelder. GPU rendering and raycasting: Hardware-accelerated, high-quality rendering based on trivariate splines approximating volume data. Computer Graphics Forum, 27(2):331–340, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laine:2005:RHP

[LA05] Samuli Laine and Timo Aila. Rendering I: Hierarchical penumbra casting. Computer Graphics Forum, 24(3):313–322, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laine:2006:WEM

[LA06] S. Laine and T. Aila. A weighted error metric and optimization method for antialiasing patterns. Computer Graphics Forum, 25(1):83–94, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2008:CIS

[LAA08] Y. Li, E. Adelson, and A. Agarwala. Computational imaging: ScribbleBoost: Adding classification to edge-aware interpolation of local image and video adjustments. Computer Graphics Forum, 27(4):1255–1264, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2002:AAT

[LAD02] Haeyoung Lee, Pierre Alliez, and Mathieu Desbrun. Angleanalyzer: a triangle-quad mesh codec. Computer Graphics Forum, 21(3):383–392, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lamarche:2009:VHT

[Lam09a] F. Lamarche. Virtual humans: TopoPlan: a topological path planner for real time human navigation under floor and ceiling constraints. Computer Graphics Forum, 28(2):649–658, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Larsson:2009:BVH

[LAM09b] T. Larsson and T. Akenine-Möller. Bounding volume hierarchies of slab cut balls. Computer Graphics Forum, 28(8):2379–2395, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Langdon:2007:BR

[Lan07] W. B. Langdon. Book review. Computer Graphics Forum, 26(4): 853, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lastra:1984:RGM

[Las84] G. Léon Lastra. Report on a geometric modelling course. Computer Graphics Forum, 3(1):117–118, March 1984. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lam:2006:PIB

[LB06] Michael W. Y. Lam and Gladimir V. G. Baranoski. Photorealistic and image based rendering: a predictive light transport model for the human Iris. Computer Graphics Forum, 25(3):359–368, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2008:SRS

[LBD^{+}08] L. Liu, C. Bajaj, J. O. Deasy, D. A. Low, and T. Ju. Scanning and reconstruction: Surface reconstruction from non-parallel curve networks. Computer Graphics Forum, 27(2):155–163, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2001:CFM

[LBH^{+}01] Xinguo Liu, Hujun Bao, PhengAnn Heng, TienTsin Wong, and Qunsheng Peng. Constrained fairing for meshes. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Languenou:1992:ADM

[LBT92] E. Languenou, K. Bouatouch, and P. Tellier. An adaptive discretization method for radiosity. Computer Graphics Forum, 11 (3):C205–C216, C475, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leu:1999:MRG

[LC99] Adrian Leu and Min Chen. Modelling and rendering graphics scenes composed of multiple volumetric datasets. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=366&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Lo:2009:MDA

[LC09] Cheng-Hung Lo and Chih-Hsing Chu. Modeling and deformation: Affective modelling: Profiling geometrical models with human emotional responses. Computer Graphics Forum, 28(7):1811– 1820, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:FIR

[LCD09a] B. Liu, G. J. Clapworthy, and F. Dong. Fast isosurface rendering on a GPU by Cell rasterization. Computer Graphics Forum, 28(8): 2151–2164, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:VRH

[LCD09b] B. Liu, G. J. Clapworthy, and F. Dong. Volume rendering & hardware acceleration: Accelerating volume raycasting using proxy spheres. Computer Graphics Forum, 28(3):839–846, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lewiner:2006:GGD

[LCL^{+}06] Thomas Lewiner, Marcos Craizer, Hélio Lopes, Sinésio Pesco, Luiz Velho, and Esdras Medeiros. GEncode: Geometry-driven compression for general meshes. Computer Graphics Forum, 25 (4):685–695, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lerner:2007:CAC

[LCL07] Alon Lerner, Yiorgos Chrysanthou, and Dani Lischinski. Crowds and animation: Crowds by example. Computer Graphics Forum, 26(3):655–664, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2006:ULF

[LCM^{+}06] Yang Liu, George Chen, Nelson Max, Christian Hofsetz, and Peter McGuinness. Undersampled light field rendering by a plane sweep. Computer Graphics Forum, 25(2):225–236, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:SSL

[LCM^{+}09] Junfa Liu, Yiqiang Chen, Chunyan Miao, Jinjing Xie, Charles X. Ling, Xingyu Gao, and Wen Gao. Semi-supervised learning in reconstructed manifold space for 3D caricature generation. Computer Graphics Forum, 28(8):2104–2116, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2007:MSV

[LCWK07] Lin Lu, Yi-King Choi, Wenping Wang, and Myung-Soo Kim. Modeling and segmentation: Variational 3D shape segmentation for bounding volume computation. Computer Graphics Forum, 26(3):329–338, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liebowitz:1999:CAM

[LCZ99] David Liebowitz, Antonio Criminisi, and Andrew Zisserman. Creating architectural models from images. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=326&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Loscos:1997:IHQb

[LD97] Celine Loscos and George Drettakis. Interactive high-quality soft shadows in scenes with moving objects. Computer Graphics Forum, 16(3):C219–C230, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lintermann:1998:MMU

[LD98] Bernd Lintermann and Oliver Deussen. A modelling method and user interface for creating plants. Computer Graphics Forum, 17(1):73–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=216&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Laycock:2003:RDA

[LD03] S. D. Laycock and A. M. Day. Recent developments and applications of haptic devices. Computer Graphics Forum, 22(2): 117–132, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lamarche:2004:NCB

[LD04] Fabrice Lamarche and Stéphane Donikian. Navigation and crowd behaviour in virtual environments: Crowd of virtual humans: a new approach for real time navigation in complex and structured environments. Computer Graphics Forum, 23(3):509–518, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laycock:2006:IRC

[LD06] R. G. Laycock and A. M. Day. Image registration in a coarse three-dimensional virtual environment. Computer Graphics Forum, 25(1):69–82, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laycock:2007:SHR

[LD07] S. D. Laycock and A. M. Day. A survey of haptic rendering techniques. Computer Graphics Forum, 26(1):50–65, March 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lagae:2008:CMG

[LD08a] Ares Lagae and Philip Dutré. A comparison of methods for generating Poisson disk distributions. Computer Graphics Forum, 27 (1):114–129, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lagae:2008:RTC

[LD08b] Ares Lagae and Philip Dutré. Ray tracing 1: Compact, fast and robust grids for ray tracing. Computer Graphics Forum, 27 (4):1235–1244, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lagae:2008:RTA

[LD08c] Ares Lagae and Philip Dutré. Ray tracing 2: Accelerating ray tracing using constrained tetrahedralizations. Computer Graphics Forum, 27(4):1303–1312, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lenaerts:2009:FBM

[LD09] Toon Lenaerts and Philip Dutré. Fluids and beyond: Mixing fluids and granular materials. Computer Graphics Forum, 28(2): 213–218, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lavoue:2007:FQT

[LDB07] Guillaume Lavoué, Florent Dupont, and Atilla Baskurt. A framework for quad/triangle subdivision surface fitting: Application to mechanical objects. Computer Graphics Forum, 26(1):1–14, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2009:TVD

[LDR09] Jianye Lu, Julie Dorsey, and Holly Rushmeier. Textures and video: Dominant texture and diffusion distance manifolds. Computer Graphics Forum, 28(2):667–676, April 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

LeLous:1990:RFE

[Le 90] Y. Le Lous. Report on the first EUROGRAPHICS workshop on visualization in scientific computing. Computer Graphics Forum, 9(4):371–372, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:1999:IME

[Lee99] Seungyong Lee. Interactive multiresolution editing of arbitrary meshes. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=329&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=18.

Leister:1994:CGC

[Lei94] W. Leister. Computer generated copper plates. Computer Graphics Forum, 13(1):69–77, March 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Levoy:1999:DMP

[Lev99] Marc A. Levoy. The Digital Michelangelo Project. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lewis:1994:MSM

[Lew94] R. R. Lewis. Making shaders more physically plausible. Computer Graphics Forum, 13(2):109–120, June 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lauterbach:2008:RTR

[LeYTM08] Christian Lauterbach, Sung eui Yoon, Ming Tang, and Dinesh Manocha. Ray tracing 2: ReduceM: Interactive and memory efficient ray tracing of large models. Computer Graphics Forum, 27 (4):1313–1321, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lalonde:1997:GRD

[LF97] Paul Lalonde and Alain Fournier. Generating reflected directions from BRDF data. Computer Graphics Forum, 16(3): C293–C300, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=166&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Lewis:2000:WRT

[LF00] Robert R. Lewis and Alain Fournier. Wavelet radiative transfer and surface interaction. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2008:SBU

[LFGG08] Sangwon Lee, David Feng, Cindy Grimm, and Bruce Gooch. A sketch-based user interface for reconstructing architectural drawings. Computer Graphics Forum, 27(1):81–90, March 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lopes:1992:CMP

[LG92] P. F. Lopes and M. R. Gomes. A computer model for pinscreen simulation: a new animation paradigm. Computer Graphics Forum, 11(1):31–42, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lippert:1995:FWB

[LG95] L. Lippert and M. H. Gross. Fast wavelet based volume rendering by accumulation of transparent texture maps. Computer Graphics Forum, 14(3):C/431–C/443, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lamouret:1996:SIP

[LG96] Alexis Lamouret and Marie-Paule Gascuel. Scripting interactive physically-based motions with relative paths and synchronization. Computer Graphics Forum, 15(1):25–34, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Labsik:2000:IS

[LG00] U. Labsik and G. Greiner. Interpolatory √3-subdivision. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=405&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Lensch:2003:IRT

[LGB^{+}03] Hendrik P. A. Lensch, Michael Goesele, Philippe Bekaert, Jan Kautz, Marcus A. Magnor, Jochen Lang, and Hans-Peter Seidel. Interactive rendering of translucent objects. Computer Graphics Forum, 22(2):195–205, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lippert:1997:CDV

[LGK97] L. Lippert, M. H. Gross, and C. Kurmann. Compression domain volume rendering for distributed environments. Computer Graphics Forum, 16(3):C95–C107, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=146&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Lee:2000:GAV

[LGMT00] WonSook Lee, Jin Gu, and Nadia Magnenat-Thalmann. Generating animatable 3D virtual humans from photographs. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lauterbach:2009:GFB

[LGS^{+}09] C. Lauterbach, M. Garland, S. Sengupta, D. Luebke, and D. Manocha. GPU: Fast BVH construction on GPUs. Computer Graphics Forum, 28(2):375–384, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laramee:2004:SAR

[LHD^{+}04] Robert S. Laramee, Helwig Hauser, Helmut Doleisch, Benjamin Vrolijk, Frits H. Post, and Daniel Weiskopf. State of the art reviews: The state of the art in flow visualization: Dense and texture-based techniques. Computer Graphics Forum, 23(2):203– 221, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2007:SPP

[Li07] Q. Li. Smooth piecewise polynomial blending operations for implicit shapes. Computer Graphics Forum, 26(2):157–171, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindner:1985:PRS

[Lin85] R. Lindner. A prefiltering raster scan algorithm. Computer Graphics Forum, 4(2):101–110, June 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lister:1990:RFE

[Lis90] P. F. Lister. Report on the fifth EUROGRAPHICS workshop on graphics hardware. Computer Graphics Forum, 9(4):379–380, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:1993:GCA

[Liu93a] Yong-Kui Liu. The generation of circular arcs on hexagonal grids. Computer Graphics Forum, 12(1):21–26, March 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:1993:GSL

[Liu93b] Yong-Kui Liu. The generation of straight lines on hexagonal grids. Computer Graphics Forum, 12(1):27–31, March 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:1994:AIA

[Liu94] Yong-Kui Liu. An all-integer algorithm for drawing anti-aliased straight lines. Computer Graphics Forum, 13(4):219–221, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2002:AIF

[LJN02] Sung Chun Lee, Soon Ki Jung, and Ram Nevatia. Automatic integration of facade textures into 3D building models with a projective geometry based line clustering. Computer Graphics Forum, 21(3):511–519, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2008:RRT

[LKC08] Sungkil Lee, Gerard Jounghyun Kim, and Seungmoon Choi. Rendering: Real-time depth-of-field rendering using point splatting on per-pixel layers. Computer Graphics Forum, 27(7):1955–1962, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lao:2000:VBA

[LL00] Zhiqiang Lao and Ling Li. Video-based approach to human animation. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=432&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Lane:2001:RIV

[LL01] J. R. T. Lane and V. Lalioti. Reflective interaction in virtual environments. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2002:GST

[LL02] Yunjin Lee and Seungyong Lee. Geometric snakes for triangular meshes. Computer Graphics Forum, 21(3):229–238, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2005:AAS

[LL05] Hyun-Chul Lee and In-Kwon Lee. Animation: Automatic synchronization of background music and motion in computer animation. Computer Graphics Forum, 24(3):353–361, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Long:2009:MDM

[LL09] Tran Van Long and Lars Linsen. Multidimensional data: Multi-ClusterTree: Interactive visual exploration of hierarchical clusters in multidimensional multivariate data. Computer Graphics Forum, 28(3):823–830, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lehtinen:2006:RIP

[LLA06] Jaakko Lehtinen, Samuli Laine, and Timo Aila. Rendering: an improved physically-based soft shadow volume algorithm. Computer Graphics Forum, 25(3):303–312, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laycock:2005:RSA

[LLD05] S. D. Laycock, R. G. Laycock, and A. M. Day. Reports: Siggraph 2005 Los Angeles, California, 31st July–4th August 2005. Computer Graphics Forum, 24(4):849–850, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:1997:IRD

[LLG97] Frederick W. B. Li, Rynson W. H. Lau, and Mark Green. Interactive rendering of deforming NURBS surfaces. Computer Graphics Forum, 16(3):C47–C56, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=141&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Lindholm:2009:VRH

[LLHY09] Stefan Lindholm, Patric Ljung, Markus Hadwiger, and Anders Ynnerman. Volume rendering & hardware acceleration: Fused multi-volume DVR using binary space partitioning. Computer Graphics Forum, 28(3):847–854, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2000:VDM

[LLP00] Xiangyang Li, Dongming Lu, and Yunhe Pan. Virtual Dunhuang mural restoration system in collaborative network environment. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=425&amp;   http://www.blackwellpublishers.co.

uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Laycock:2004:RSA

[LLR^{+}04] S. D. Laycock, R. G. Laycock, G. Ryder, J. Juarez-Comboni, and C. Montes de Oca. Reports: SIGGRAPH 2004: Los Angeles, California, 8–12 August 2004. Computer Graphics Forum, 23(4): 839–840, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laycock:2007:RSB

[LLRD07] R. G. Laycock, S. D. Laycock, G. Ryder, and A. M. Day. Reports: Siggraph 2006 Boston, Massachusetts, 30th July–3rd August 2006. Computer Graphics Forum, 26(1):117–118, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:1998:FFB

[LLSL98] Tong-Yee Lee, Young-Ching Lin, Y. N. Sun, and Leeween Lin. Fast feature-based metamorphosis and operator design. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=249&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Lensch:2003:PSS

[LLSS03a] Hendrik P. A. Lensch, Jochen Lang, Asla M. Sá, and Hans-Peter Seidel. Planned sampling of spatially varying BRDFs. Computer Graphics Forum, 22(3):473–482, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lensch:2003:TBP

[LLSS03b] Hendrik P. A. Lensch, Jochen Lang, Asla M. Sá, and Hans-Peter Seidel. Textures and BRDFs: Planned sampling of spatially varying BRDFs. Computer Graphics Forum, 22(3):473–482, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2009:IPN

[LLY09] H. Lee, C. H. Lee, and K. Yoon. Important points and NPR: Motion based painterly rendering. Computer Graphics Forum, 28 (4):1207–1215, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laakko:1993:IBO

[LM93] T. Laakko and M. Mantyla. Introducing blending operations in feature models. Computer Graphics Forum, 12(3):C165–C176, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laakko:1996:ICM

[LM96a] Timo Laakko and Martti Mantyla. Incremental constraint modelling in a feature modelling system. Computer Graphics Forum, 15(3):C367–C376, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laako:1996:ICM

[LM96b] T. Laako and M. Mantyla. Incremental constraint modelling in a feature modelling system. Computer Graphics Forum, 15(3): 367–376, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2007:MCI

[LM07] G. Li and W. Ma. A method for constructing interpolatory subdivision schemes and blending subdivisions. Computer Graphics Forum, 26(2):185–201, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2005:NIS

[LMB05] G. Li, W. Ma, and H. Bao. A new interpolatory subdivision for quadrilateral meshes. Computer Graphics Forum, 24(1):3–16, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lagae:2004:RES

[LMD04] Ares Lagae, Vincent Masselus, and Philip Dutre. Reports: Eurographics Symposium on Rendering 2003: (14th Eurographics Workshop on Rendering) Leuven, Belgium, 25–27 June 2003. Computer Graphics Forum, 23(1):123, March 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2004:GBR

[LMS04] Ming Li, Marcus Magnor, and Hans-Peter Seidel. GPU-based rendering: Hardware-accelerated rendering of photo hulls. Computer Graphics Forum, 23(3):635–642, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laszlo:2005:IPF

[LNS05] Joe Laszlo, Michael Neff, and Karan Singh. Interaction: Predictive feedback for interactive control of physics-based characters. Computer Graphics Forum, 24(3):257–265, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:1995:RTC

[LO95] Rung-Huei Liang and Ming Ouhyoung. A real-time continuous alphabetic sign language to speech conversion VR system. Computer Graphics Forum, 14(3):C/67–C/76, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loscos:1997:IHQa

[Los97] Céline Loscos. Interactive high-quality soft shadows in scenes with moving objects. Computer Graphics Forum, 16(3): ??, September 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=159&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Loviscach:2006:DWC

[Lov06] J. Loviscach. Deformations: Wrinkling coarse meshes on the GPU. Computer Graphics Forum, 25(3):467–476, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Logie:1995:IDM

[LP95] J. R. Logie and J. W. Patterson. Inverse displacement mapping in the general case. Computer Graphics Forum, 14(5):261–273, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Labatut:2009:RES

[LPK09] P. Labatut, J.-P. Pons, and R. Keriven. Robust and efficient surface reconstruction from range data. Computer Graphics Forum, 28(8):2275–2290, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lastra:1988:NRT

[LR88] G. L. Lastra and E. N. Rüdinger. A note on random thresholding for colour synthesis. Computer Graphics Forum, 7(4):343–346, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lakshminarasimhan:1989:FFS

[LS89] A. L. Lakshminarasimhan and Mandayam Srivas. A framework for functional specification and transformation of hidden surface elimination algorithms. Computer Graphics Forum, 8(2):75–98, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Langer:2008:PTH

[LS08a] Torsten Langer and Hans-Peter Seidel. Parameterization and transformations: Higher order barycentric coordinates. Computer Graphics Forum, 27(2):459–466, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loop:2008:SST

[LS08b] Charles Loop and Scott Schaefer. Smooth surfaces: G^{2} tensor product splines over extraordinary vertices. Computer Graphics Forum, 27(5):1373–1382, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Landreneau:2009:MCS

[LS09] Eric Landreneau and Scott Schaefer. Motion compression: Simplification of articulated meshes. Computer Graphics Forum, 28 (2):347–353, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2009:DGP

[LSJK09] Bao Li, Ruwen Schnabel, Shiyao Jin, and Reinhard Klein. Digital geometry processing: Variational surface approximation and model selection. Computer Graphics Forum, 28(7):1985–1994, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2008:RDS

[LSP08] Hao Li, Robert W. Sumner, and Mark Pauly. Registering deforming surfaces I: Global correspondence optimization for nonrigid registration of depth scans. Computer Graphics Forum, 27 (5):1421–1430, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loos:1998:UWT

[LSS98] J. Loos, Ph. Slusallek, and H.-P. Seidel. Using wavefront tracing for the visualization and optimization of progressive lenses. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=272&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Luo:2009:DAG

[LSW09] Chuanjiang Luo, Issam Safa, and Yusu Wang. Differentials: Approximating gradients for meshes and point clouds via diffusion metric. Computer Graphics Forum, 28(5):1497–1508, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2008:GMS

[LSŽ08] Y. K. Liu, H. Y. Song, and B. Žalik. A general multi-step algorithm for voxel traversing along a line. Computer Graphics Forum, 27(1):73–80, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2008:FAG

[LTH08] Guo-Shi Li, Xavier Tricoche, and Charles Hansen. Flow advection and geometry processing: Physically-based dye advection for flow visualization. Computer Graphics Forum, 27(3):727–734, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lutz:2002:NCI

[Lut02] Bernd Lutz. Note on cover image. Computer Graphics Forum, 21 (1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leotta:2008:SSP

[LVT08] Matthew J. Leotta, Austin Vandergon, and Gabriel Taubin. 3D slit scanning with planar constraints. Computer Graphics Forum, 27(8):2066–2080, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lau:1992:AIG

[LW92] Wing Hung Lau and N. Wiseman. Accurate image generation and interactive image editing with the A-buffer. Computer Graphics Forum, 11(3):C279–C288, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lafortune:1994:TFP

[LW94] E. P. Lafortune and Y. D. Willems. A theoretical framework for physically based rendering. Computer Graphics Forum, 13(2): 97–107, June 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lau:1995:CBF

[LW95] Wing Hung Lau and N. Wiseman. The compositing buffer: a flexible method for image generation and image editing. Computer Graphics Forum, 14(4):229–238, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lutz:1999:VDA

[LW99] B. Lutz and M. Weintke. Virtual Dunhuang art cave: a cave within a CAVE. Computer Graphics Forum, 18(3): ??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=346&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Li:2004:IFU

[LWP^{+}04] Q. Li, D. Wills, R. Phillips, W. J. Viant, J. G. Griffiths, and J. Ward. Implicit fitting using radial basis functions with ellipsoid constraint. Computer Graphics Forum, 23(1):55–69, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:IRT

[LWX^{+}09] Yiming Liu, Jiaping Wang, Su Xue, Xin Tong, Sing Bing Kang, and Baining Guo. Image restructuring: Texture splicing. Computer Graphics Forum, 28(7):1907–1915, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ling:2008:SSF

[LWY08] Ruotian Ling, Wenping Wang, and Dongming Yan. Smooth surfaces: Fitting sharp features with loop subdivision surfaces. Computer Graphics Forum, 27(5):1383–1391, July 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:VQO

[LWZ^{+}09] F. Liu, J. Wang, S. Zhu, M. Gleicher, and Y. Gong. Visualquality optimizing super resolution. Computer Graphics Forum, 28(1):127–140, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2008:IED

[LYP^{+}08] Chunxiao Liu, Yingzhen Yang, Qunsheng Peng, Jin Wang, and Wei Chen. Image editing: Distortion optimization based image completion from a large displacement view. Computer Graphics Forum, 27(7):1755–1764, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2007:SMS

[LZ07] Rong Liu and Hao Zhang. Surfaces: Mesh segmentation via spectral embedding and contour analysis. Computer Graphics Forum, 26(3):385–394, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2009:SAH

[LZSCO09] Rong Liu, Hao Zhang, Ariel Shamir, and Daniel Cohen-Or. Shape analysis and helices: a part-aware surface metric for shape analysis. Computer Graphics Forum, 28(2):397–406, April 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2008:PQT

[LZX^{+}08] Ligang Liu, Lei Zhang, Yin Xu, Craig Gotsman, and Steven J. Gortler. Parameterization and quad tiling: a local/global approach to mesh parameterization. Computer Graphics Forum, 27 (5):1495–1504, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2004:IOP

[LŽY04] Y. K. Liu, B. Žalik, and H. Yang. An integer one-pass algorithm for voxel traversal. Computer Graphics Forum, 23(2):167–172, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:1985:PAB

[MA85] R. R. Martin and C. Anderson. A proposal for an ALGOL 68 binding of GKS. Computer Graphics Forum, 4(1):43–57, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:1991:QTI

[MA91] R. R. Martin and M. M. Anguh. Quadtrees, transforms and image coding. Computer Graphics Forum, 10(2):91–96, June 1991. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mitake:2009:MSP

[MAA^{+}09] Hironori Mitake, Kazuyuki Asano, Takafumi Aoki, Salvati Marc, Makoto Sato, and Shoichi Hasegawa. Motion synthesis: Physicsdriven multi dimensional keyframe animation for artist-directable interactive character. Computer Graphics Forum, 28(2):279–287, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

MacanAirchinnigh:1984:SIG

[Mac84] M. Mac an Airchinnigh. The specification and implementation of GKS application software in ADA. Computer Graphics Forum, 3(2):153–167, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

MacanAirchinnigh:1985:SWU

[Mac85] M. Mac an Airchinnigh. Seeheim workshop on user interface management systems: Report of the working group on the user’s conceptual model. Computer Graphics Forum, 4(1):63–69, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Machover:1994:YTS

[Mac94] C. Machover. Yes… there are still too many unsolved problems in computer graphics! Computer Graphics Forum, 13(3):C/515–??, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maeder:1990:RAA

[Mae90] A. Maeder. Report on AUSGRAPH’90, the Australian Graphics Conference. Computer Graphics Forum, 9(4):389, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maillot:2000:RTL

[Mai00] Jerome Maillot. Real time local approximation of deformations using rotations. Computer Graphics Forum, 19 (3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=435&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Mantyla:1983:SOG

[Man83] M. Mantyla. Set operations of GWB. Computer Graphics Forum, 2(2-3):122–134, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:1982:CGS

[Mar82] J. W. Martin. Computer graphics software workshop report. Computer Graphics Forum, 1(1):10–13, March 1982. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maruya:1995:GTM

[Mar95] M. Maruya. Generating a texture map from object-surface texture data. Computer Graphics Forum, 14(3):C/397–C/405, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matthew:1985:PCP

[Mat85] A. J. Matthew. Polygonal clipping of polylines. Computer Graphics Forum, 4(4):407–414, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matthew:1986:EUG

[Mat86] A. J. Matthew. EUROGRAPHICS UK 1986 Glasgow Conference. Computer Graphics Forum, 5(3):225–232, September 1986. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

May:1999:PPC

[May99] Jon May. Perceptual principles and computer graphics. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=320&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

May:2000:PPC

[May00] Jon May. Perceptual principles and computer graphics. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mason:1997:AHL

[MB97] Ashton E. W. Mason and Edwin H. Blake. Automatic hierarchical level of detail optimization in computer animation. Computer Graphics Forum, 16(3):C191–C199, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=156&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Meneveaux:1999:SLB

[MB99] Daniel Meneveaux and Kadi Bouatouch. Synchronisation and load balancing for parallel hierarchical radiosity of complex scenes on a heterogeneous computer network. Computer Graphics Forum, 18(4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=374&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Memari:2008:SRS

[MB08] Pooran Memari and Jean-Daniel Boissonnat. Sampling and reconstructing surfaces: Provably good 2D shape reconstruction from unorganized cross-sections. Computer Graphics Forum, 27 (5):1403–1410, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Monzani:2000:UIS

[MBBT00] Jean-Sébastien Monzani, Paolo Baerlocher, Ronan Boulic, and Daniel Thalmann. Using an intermediate skeleton and inverse kinematics for motion retargeting. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=393&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Metaaphanon:2009:AST

[MBCN09] Napaporn Metaaphanon, Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Animation: Simulation of tearing cloth with frayed edges. Computer Graphics Forum, 28(7):1837–1844, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meissner:2001:VDR

[MBGS01] M. Meißner, D. Bartz, R. Günther, and W. Straßer. Visibility driven rasterization. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Museth:2005:AIE

[MBW^{+}05] Ken Museth, David E. Breen, Ross T. Whitaker, Sean Mauch, and David Johnson. Algorithms for interactive editing of level set models. Computer Graphics Forum, 24(4):821–841, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2008:OVR

[MBW08] Oliver Mattausch, Jiří Bittner, and Michael Wimmer. Occlusion and volume rendering: CHC++: Coherent hierarchical culling revisited. Computer Graphics Forum, 27(2):221–230, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marchand:2002:VVS

[MC02] Éric Marchand and François Chaumette. Virtual visual servoing: a framework for real-time augmented reality. Computer Graphics Forum, 21(3):289–298, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McCullagh:1983:ACF

[McC83] M. J. McCullagh. Automatic contouring of faulted seismic data sets. Computer Graphics Forum, 2(1):57–66, March 1983. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McCool:1996:AEB

[McC96] Michael D. McCool. Accelerated evaluation of box splines via a parallel inverse FFT. Computer Graphics Forum, 15(1):35–45, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mah:1994:CBR

[MCH94] Sang Mah, T. W. Calvert, and W. Havens. A constraint-based reasoning framework for behavioural animation. Computer Graphics Forum, 13(5):315–324, December 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munkberg:2008:PHT

[MCHAM08] J. Munkberg, P. Clarberg, J. Hasselgren, and T. Akenine-Möller. Practical HDR texture compression. Computer Graphics Forum, 27(6):1664–1676, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ming:1996:RRO

[MCL96] Ouhyoung Ming, Yung-Yu Chuang, and Rung-Huei Liang. Reusable radiosity objects. Computer Graphics Forum, 15(3): C347–C356, C483, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNamara:2001:VPR

[McN01] Ann McNamara. Visual perception in realistic image synthesis. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mann:1997:SPT

[MCO97] Yair Mann and Daniel Cohen-Or. Selective pixel transmission for navigating in remote virtual environments. Computer Graphics Forum, 16(3):C201–C206, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=157&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Monzani:2001:IBA

[MCT01] Jean-Sébastien Monzani, Angela Caicedo, and Daniel Thalmann. Integrating behavioural animation techniques. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maurel:1993:RT

[MDC93] H. Maurel, Y. Duthen, and R. Caubet. A 4D ray tracing. Computer Graphics Forum, 12(3):C285–C294, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mesmoudi:2008:TDD

[MDP08] Mohammed Mostefa Mesmoudi, Leila De Floriani, and Umberto Port. Tetrahedra: Discrete distortion in triangulated 3-manifolds. Computer Graphics Forum, 27(5):1333–1340, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moser:2008:NPR

[MDWK08] Sebastian Möser, Patrick Degener, Roland Wahl, and Reinhard Klein. Non-photorealistic rendering: Context aware terrain visualization for wayfinding and navigation. Computer Graphics Forum, 27(7):1853–1860, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meissner:1998:AKF

[ME98] M. Meißner and B. Eberhardt. The art of knitted fabrics, realistic & physically based modelling of knitted patterns. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=282&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Mora:2004:SVI

[ME04] Benjamin Mora and David S. Evert. Scientific visualization: Instant volumetric understanding with order-independent volume rendering. Computer Graphics Forum, 23(3):489–497, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mencl:1995:GBA

[Men95] R. Mencl. A graph-based approach to surface reconstruction. Computer Graphics Forum, 14(3):C/445–C/456, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meyers:1994:MT

[Mey94] D. Meyers. Multiresolution tiling. Computer Graphics Forum, 13 (5):325–340, December 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Min:2000:PDA

[MF00] Patrick Min and Thomas Funkhouser. Priority-driven acoustic modeling for virtual environments. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=410&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Morrison:1986:IGP

[MFDA86] R. Morrison, A. L. Florianis, A. Dearle, and M. P. Atkinson. An integrated graphics programming environment. Computer Graphics Forum, 5(2):147–157, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marmitt:2008:SAR

[MFS08] Gerd Marmitt, Heiko Friedrich, and Philipp Slusallek. State of the art review: Efficient CPU-based volume ray tracing techniques. Computer Graphics Forum, 27(6):1687–1709, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2002:FOU

[MFT02] D. Martín, J. D. Fekete, and J. C. Torres. Flattening 3D objects using silhouettes. Computer Graphics Forum, 21(3):239–248, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marsh:1987:UAD

[MG87] S. C. Marsh and R. L. Grimsdale. Using abstract data types to define large-scale world models for image generation systems. Computer Graphics Forum, 6(2):77–85, May 1987. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mazuryk:1995:TSP

[MG95] T. Mazuryk and M. Gervautz. Two-step prediction and image deflection for exact head tracking in virtual environments. Computer Graphics Forum, 14(3):C/29–C/41, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Madritsch:1996:CCB

[MG96] Franz Madritsch and Michael Gervautz. CCD-camera based optical beacon tracking for virtual and augmented reality. Computer Graphics Forum, 15(3):C207–C216, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Majumder:2005:MCP

[MG05] Aditi Majumder and M. Gopi. Modeling color properties of tiled displays. Computer Graphics Forum, 24(2):149–163, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Menzel:2009:GBI

[MG09] N. Menzel and M. Guthe. g-BRDFs: an intuitive and editable BTF representation. Computer Graphics Forum, 28(8):2189– 2200, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morer:1995:OPL

[MGAF95] P. Morer, A. M. Garcia-Alonso, and J. Flaquer. Optimization of a priority list algorithm for 3-D rendering of buildings. Computer Graphics Forum, 14(4):217–227, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marti:2006:PET

[MGJ06] E. Martí, D. Gil, and C. Julià. A PBL experience in the teaching of computer graphics. Computer Graphics Forum, 25(1):95–103, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marais:2007:DRC

[MGS07] P. Marais, J. Gain, and D. Shreiner. Distance-ranked connectivity compression of triangle meshes. Computer Graphics Forum, 26(4): 813–823, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2000:SST

[MH00] Kerstin Müller and Sven Havemann. Subdivision surface tesselation on the fly using a versatile mesh data structure.

Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=407&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Masuko:2007:FHE

[MH07] Soh Masuko and Junichi Hoshino. Faces: Head-eye animation corresponding to a conversation for CG characters. Computer Graphics Forum, 26(3):303–312, September 2007. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mroz:2000:IHQ

[MHG00] Lukas Mroz, Helwig Hauser, and Eduard Gröller. Interactive high-quality maximum intensity projection. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=426&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Mikami:1982:CGA

[Mik82] T. Mikami. Computer graphics applications in Japan. Computer Graphics Forum, 1(4):191–193, December 1982. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milne:1983:GAP

[Mil83] W. J. Milne. Geography Algorithms Project. Computer Graphics Forum, 2(1):48–56, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mills:1984:SPV

[Mil84] P. L. Mills. The systolic pixel: a visible surface algorithm for VLSI. Computer Graphics Forum, 3(1):47–59, March 1984. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milanese:1987:ANT

[Mil87] V. Milanese. ADA and NIL: Two concurrent languages for GKS. Computer Graphics Forum, 6(3):219–233, September 1987. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milanese:1988:PEG

[Mil88a] V. Milanese. A Prolog environment for GKS-based graphics. Computer Graphics Forum, 7(1):9–20, March 1988. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milanese:1988:PDM

[Mil88b] V. Milanese. A proposal for a distributed model of GKS based on Prolog. Computer Graphics Forum, 7(3):203–213, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milanese:1990:CLC

[Mil90] Vitaliano Milanese. KAΔMOΣ: a clausal language for CAD modeling systems with morphological constraints. Computer Graphics Forum, 9(1):39–51, March 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Markopoulos:1998:FEW

[MJ98] P. Markopoulos and P. Johnson. Fifth Eurographics Workshop on Design Specification and Verification of Interactive Systems, Abingdon, UK, June 3–5 1998. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=240&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Mora:2001:VIP

[MJC01] B. Mora, J. P. Jessel, and R. Caubet. Visualization of isosurfaces with parametric cubes. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matsushita:1999:EHT

[MK99] Kenji Matsushita and Toyohisa Kaneko. Efficient and handy texture mapping on 3D surfaces. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=355&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Marinov:2005:GIA

[MK05] M. Marinov and L. Kobbelt. Geometry II: Automatic generation of structure preserving multiresolution models. Computer Graphics Forum, 24(3):479–486, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marinov:2006:MPR

[MK06] Martin Marinov and Leif Kobbelt. Mesh processing: a robust twostep procedure for quad-dominant remeshing. Computer Graphics Forum, 25(3):537–546, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

May:2008:VAT

[MK08] T. May and J. Kohlhammer. Visual analytics: Towards closing the analysis gap: Visual generation of decision supporting schemes from raw data. Computer Graphics Forum, 27(3):911–918, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mertens:2005:ERL

[MKB^{+}05] Tom Mertens, Jan Kautz, Philippe Bekaert, Frank Van Reeth, and Hans-Peter Seidel. Efficient rendering of local subsurface scattering. Computer Graphics Forum, 24(1):41–49, March 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2008:BCP

[MKB^{+}08] Sebastian Martin, Peter Kaufmann, Mario Botsch, Martin Wicke, and Markus Gross. Bases and coordinates: Polyhedral finite elements using harmonic basis functions. Computer Graphics Forum, 27(5):1521–1529, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mizuno:2000:AGV

[MKO^{+}00] S. Mizuno, T. Kasaura, T. Okouchi, S. Yamamoto, M. Okada, and J. Toriwaki. Automatic generation of virtual woodblocks and multicolor woodblock printing. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=397&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Muigg:2008:MVF

[MKO^{+}08] Philipp Muigg, Johannes Kehrer, Steffen Oeltze, Harald Piringer, Helmut Doleisch, Bernhard Preim, and Helwig Hauser. Multivariate visualization: a four-level focus+context approach to interactive visual analysis of temporal features in large scientific data. Computer Graphics Forum, 27(3):775–782, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mertens:2009:EFS

[MKV09] T. Mertens, J. Kautz, and F. Van Reeth. Exposure fusion: a simple and practical alternative to high dynamic range photography. Computer Graphics Forum, 28(1):161–171, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mangen:1991:SIP

[ML91] Alain Mangen and Nadine Lasudry. Search for the intersection polygon of any two polygons: Application to the garment industry. Computer Graphics Forum, 10(3):195–208, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Masso:2003:AHH

[ML03] J. P. Molina Massó and P. González López. Automatic hybrid hierarchy creation: a cost-model based approach. Computer Graphics Forum, 22(1):5–13, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:1992:ESP

[MLP92] Lizhuang Ma, Youdong Liang, and Qunsheng Peng. Equidistant smoothing of polyhedra with arbitrary topologies. Computer Graphics Forum, 11(3):C405–C414, ???? 1992. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moltedo:1993:AIT

[MM93] L. Moltedo and S. Morigi. ANIMA: an interactive tool for scientific data animation. Computer Graphics Forum, 12(5):277–288, December 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McDonnell:2008:PCI

[MM08] K. T. McDonnell and K. Mueller. Parallel coordinates: Illustrative parallel coordinates. Computer Graphics Forum, 27(3):1031–1038, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Michelin:1993:FFC

[MMAG93] S. Michelin, G. Maffeis, D. Arques, and J. C. Grossetie. Form factor calculation: a new expression with implementations on a parallel T.node computer. Computer Graphics Forum, 12(3):C421– C432, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marroquim:2008:VIR

[MMFE08] R. Marroquim, A. Maximo, R. Farias, and C. Esperança. Volume and isosurface rendering with GPU-accelerated Cell projection. Computer Graphics Forum, 27(1):24–35, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Merry:2006:CDR

[MMG06] Bruce Merry, Patrick Marais, and James Gain. Compression of dense and regular point clouds. Computer Graphics Forum, 25 (4):709–716, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mosaliganti:2008:VAL

[MMHL08] Kishore Mosaliganti, Raghu Machiraju, Kun Huang, and Gustavo Leone. Visualization applications in life sciences: Geometrydriven visualization of microscopic structures in biology. Computer Graphics Forum, 27(3):871–878, May 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mas:2008:CIS

[MMP08] Albert Mas, Ignacio Martín, and Gustavo Patow. Compression and importance sampling of near-field light sources. Computer Graphics Forum, 27(8):2013–2027, December 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mas:2009:FIR

[MMP09] A. Mas, I. Martín, and G. Patow. Fast inverse reflector design (FIRD). Computer Graphics Forum, 28(8):2046–2056, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2005:SAR

[MMS^{+}05] G. Müller, J. Meseth, M. Sattler, R. Sarlette, and R. Klein. State of the art reviews: Acquisition, synthesis, and rendering of bidirectional texture functions. Computer Graphics Forum, 24(1):83– 109, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mihalef:2007:FMC

[MMS07] V. Mihalef, D. Metaxas, and M. Sussman. Fluid and motion capture: Textured liquids based on the marker level set. Computer Graphics Forum, 26(3):457–466, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mihalef:2009:FBS

[MMS09] Viorel Mihalef, Dimitris Metaxas, and Mark Sussman. Fluids and beyond: Simulation of two-phase flow with sub-scale droplet and bubble effects. Computer Graphics Forum, 28(2):229–238, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mantiuk:2009:HPG

[MMTH09] R. Mantiuk, R. Mantiuk, A. Tomaszewska, and W. Heidrich. Human perception and Green graphics: Color correction for tone mapping. Computer Graphics Forum, 28(2):193–202, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moltedo:1987:IVC

[MN87] L. Moltedo and A. Noferini. An implementer’s view of CGM in GKS environment. Computer Graphics Forum, 6(2):91–101, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Myles:2008:SSF

[MNP08] A. Myles, T. Ni, and J. Peters. Smooth surfaces: Fast parallel construction of smooth surfaces from meshes with tri/quad/ pent facets. Computer Graphics Forum, 27(5):1365–1372, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mikheev:1994:COC

[MNR94] A. Mikheev, M. Nozik, and J. Rubinstein. Computation of offset curves by the Huygens principle. Computer Graphics Forum, 13 (4):249–252, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moerschell:2008:DTM

[MO08] Adam Moerschell and John D. Owens. Distributed texture memory in a multi-GPU environment. Computer Graphics Forum, 27 (1):130–151, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mohamed:1987:EAA

[Moh87] Ramzan Mohamed. An experiment in algorithm animation using SMALLTALK on a Macintosh. Computer Graphics Forum, 6(2): 151–155, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morrow:1986:EPV

[Mor86] T. M. Morrow. EUROGRAPHICS ’86 — a personal view. Computer Graphics Forum, 5(4):353–354, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mizunoy:1999:IDS

[MOT99] S. Mizunoy, M. Okadayy, and J. Toriwakiy. An interactive designing system with virtual sculpting and virtual woodcut printing. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=339&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Mei:2005:ROC

[MPS05] Chunhui Mei, Voicu Popescu, and Elisha Sacks. Rendering I: The occlusion camera. Computer Graphics Forum, 24(3):335–342, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2008:SRS

[MPS08] J. Manson, G. Petrova, and S. Schaefer. Sampling and reconstructing surfaces: Streaming surface reconstruction using wavelets. Computer Graphics Forum, 27(5):1411–1420, July 2008.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:1998:TPH

[MPT98] I. Martín, X. Pueyo, and D. Tost. A two-pass hardwarebased method for hierarchical radiosity. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=263&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Mudge:2006:RVN

[MRS06] Mark Mudge, Nick Ryan, and Roberto Scopigno. Reports: VAST 2005 November 7–11, Pisa, Italy. Computer Graphics Forum, 25 (1):142, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morvan:2008:HPG

[MRS08] T. Morvan, M. Reimers, and E. Samset. High performance GPUbased proximity queries using distance fields. Computer Graphics Forum, 27(8):2040–2052, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mohan:2008:IPA

[MRT08] Ankit Mohan, Ramesh Raskar, and Jack Tumblin. Imaging and projectors: Agile spectrum imaging: Programmable wavelength modulation for cameras and projectors. Computer Graphics Forum, 27(2):709–717, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moreton:1993:SIM

[MS93] H. P. Moreton and C. H. Sequin. Scale-invariant minimum-cost curves: fair and robust design implements. Computer Graphics Forum, 12(3):C473–C484, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muench:1996:ICH

[MS96a] Stefan Muench and Martin Stangenberg. Intelligent control for haptic displays. Computer Graphics Forum, 15(3):217–226, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munch:1996:ICH

[MS96b] S. Munch and M. Stangenberg. Intelligent control for haptic displays. Computer Graphics Forum, 15(3):C217–C226, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McCormack:1998:CRC

[MS98] Jon McCormack and Andrei Sherstyuk. Creating and rendering convolution surfaces. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=232&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Maillot:2001:USS

[MS01] Jérôme Maillot and Jos Stam. A unified subdivision scheme for polygonal modeling. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mantiuk:2008:IPM

[MS08] Rafał Mantiuk and Hans-Peter Seidel. Imaging and projectors: Modeling a generic tone-mapping operator. Computer Graphics Forum, 27(2):699–708, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2000:ACO

[MSF00] Gordon Müller, Stephan Schäfer, and W. Dieter Fellner. Automatic creation of object hierarchies for radiosity clustering. Computer Graphics Forum, 19(4):??, December 2000. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNeill:1992:PSS

[MSH^{+}92] M. D. J. McNeill, B. C. Shah, M.-P. Hebert, P. F. Lister, and R. L. Grimsdale. Performance of space subdivision techniques in ray tracing. Computer Graphics Forum, 11(4):213–220, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2006:PIB

[MSK06] Gero Müller, Ralf Sarlette, and Reinhard Klein. Photorealistic and image based rendering: Data-driven local coordinate systems for image-based rendering. Computer Graphics Forum, 25(3): 369–378, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2008:USU

[MSM^{+}08] Joel P. Martin, J. Edward Swan II, Robert J. Moorhead II, Zhanping Liu, and Shangshu Cai. User studies and user interaction: Results of a user study on 2D hurricane visualization. Computer Graphics Forum, 27(3):991–998, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martinez:2008:IMR

[MSSK08] Roel Martínez, Mateu Sbert, and László Szirmay-Kalos. Improving multipath radiosity with bundles of parallel lines. Computer Graphics Forum, 27(6):1632–1646, September 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mei:2004:RTI

[MSW04] Chunhui Mei, Jiaoying Shi, and Fuli Wu. Real-time illumination and shadows: Rendering with spherical radiance transport maps. Computer Graphics Forum, 23(3):281–290, September 2004. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNeill:2002:SDS

[MSWK02] M. D. J. McNeill, H. Sayers, S. Wilson, and P. Mc Kevitt. A spoken dialogue system for navigation in non-immersive virtual environments. Computer Graphics Forum, 21(4):713–723, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2001:RSV

[MT01] Domingo Martín and Juan Carlos Torres. Rendering silhouettes with virtual lights. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mullen:2008:PQT

[MTAD08] Patrick Mullen, Yiying Tong, Pierre Alliez, and Mathieu Desbrun. Parameterization and quad tiling: Spectral conformal parameterization. Computer Graphics Forum, 27(5):1487–1494, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Magnenat-Thalmann:1984:CGS

[MTCT84] N. Magnenat-Thalmann, N. Chourot, and D. Thalmann. Colour gradation, shading and texture using a limited terminal. Computer Graphics Forum, 3(1):83–90, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2003:ART

[MTF03a] Kerstin Müller, Torsten Techmann, and Dieter Fellner. Adaptive ray tracing of subdivision surfaces. Computer Graphics Forum, 22(3):553–562, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2003:PAR

[MTF03b] Kerstin Müller, Torsten Techmann, and Dieter Fellner. Photorealism: Adaptive ray tracing of subdivision surfaces. Computer Graphics Forum, 22(3):553–562, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matsuoka:2002:RPI

[MTKO02] Hiroto Matsuoka, Tatsuto Takeuchi, Hitoshi Kitazawa, and Akira Onozawa. Representation of pseudo inter-reflection and transparency by considering characteristics of human vision. Computer Graphics Forum, 21(3):503–509, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mezger:2008:RCC

[MTPS08] Johannes Mezger, Bernhard Thomaszewski, Simon Pabst, and Wolfgang Strasser. Report: CGForum 2008 cover image. Computer Graphics Forum, 27(1):152–153, March 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mahajan:2008:LIA

[MTR08] Dhruv Mahajan, Yu-Ting Tseng, and Ravi Ramamoorthi. Local illumination: an analysis of the in-out BRDF factorization for view-dependent relighting. Computer Graphics Forum, 27 (4):1137–1145, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Magnenat-Thalmann:1989:PHP

[MTT89] Nadia Magnenat-Thalmann and Daniel Thalmann. The problematics of human prototyping and animation. Computer Graphics Forum, 8(2):115–123, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mudur:1983:GRA

[Mud83] S. P. Mudur. Graphics related activities at NCSDCT, Bombay. Computer Graphics Forum, 2(1):25–29, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:1986:IGS

[Mül86] H. Müller. Image generation by space sweep. Computer Graphics Forum, 5(3):189–195, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mumford:1986:LWR

[Mum86] A. Mumford. Loughborough workshop report: Scientific software based on GKS. Computer Graphics Forum, 5(1):71–72, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mumford:1988:IN

[Mum88] A. Mumford. Integrating at NCGA. Computer Graphics Forum, 7(3):229–230, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mumford:1989:CTT

[Mum89] A. M. Mumford. The CGM today and tomorrow. Computer Graphics Forum, 8(2):125–128, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mumford:1990:REU

[Mum90] Anne Mumford. Report of the Eurographics UK conference. Computer Graphics Forum, 9(3):273, September 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munro:1983:SIE

[Mun83] A. T. D. Munro. SADIST: an interactive editor for structured analysis. Computer Graphics Forum, 2(2-3):104–114, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Murch:1985:CGB

[Mur85] Gerald M. Murch. Colour graphics — blessing or ballyhoo? Computer Graphics Forum, 4(2):127–135, June 1985. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mahovsky:2006:MCB

[MW06] J. Mahovsky and B. Wyvill. Memory-conserving bounding volume hierarchies with coherent raytracing. Computer Graphics Forum, 25(2):173–182, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Masood:2009:IEA

[MZT09] Syed Z. Masood, Jiejie Zhu, and Marshall F. Tappen. Image editing: Automatic correction of Saturated regions in photographs using cross-channel correlation. Computer Graphics Forum, 28(7): 1861–1869, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Navazo:1986:GMB

[NAB86] I. Navazo, D. Ayala, and P. Brunet. A geometric modeller based on the exact octtree representation of polyhedra. Computer Graphics Forum, 5(2):91–104, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Narayanaswami:1995:EPG

[Nar95] Chandrasekhar Narayanaswami. Efficient parallel Gouraud shading and linear interpolation over triangles. Computer Graphics Forum, 14(1):17–24, March 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nasri:2003:IUN

[Nas03] Ahmad H. Nasri. Interpolating an unlimited number of curves meeting at extraordinary points on subdivision surfaces. Computer Graphics Forum, 22(1):87–97, March 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neff:2007:CAL

[NAS07] Michael Neff, Irene Albrecht, and Hans-Peter Seidel. Crowds and animation: Layered performance animation with correlation maps. Computer Graphics Forum, 26(3):675–684, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neelamkavil:1988:TAM

[NB88] F. Neelamkavil and L. Beare. Techniques for animation on microcomputers. Computer Graphics Forum, 7(1):21–27, March 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ng:1994:NRT

[NB94] C. M. Ng and D. W. Bustard. A new real time geometric transformation matrix and its efficient VLSI implementation. Computer Graphics Forum, 13(5):285–292, December 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nelson:1999:IMD

[NC99] Donald D. Nelson and Elaine Cohen. Interactive mechanical design variation for haptics and CAD. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=349&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Neumann:2000:GEV

[NCKG00] László Neumann, Balázs Csébfalvi, Andreas König, and Eduard Gröller. Gradient estimation in volume data using 4D linear regression. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=427&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Nehlig:1994:GDO

[ND94] P. W. Nehlig and D. A. Duce. GKS-9x: The design output primitive, an approach to a specification. Computer Graphics Forum, 13(3):C/381–C/392, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neacsu:2006:TCC

[ND06] Cristina Neacsu and Karen Daniels. Translational covering of closed planar cubic B-spline curves. Computer Graphics Forum, 25(4):743–757, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nebel:2000:RCA

[Neb00] Jean-Christophe Nebel. Realistic collision avoidance of upper limbs based on neuroscience models. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=414&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Neumann:2006:RSO

[Neu06] Andreas Neumann. Reports: SVG.Open 2005 Conference. Computer Graphics Forum, 25(2):258–259, June 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nehlig:1992:ATM

[NG92] P. Nehlig and D. Ghazanfarpour. Affine texture mapping and antialiasing using integer arithmetic. Computer Graphics Forum, 11(3):C227–C236, C477, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Notkin:1997:PPR

[NG97] Irena Notkin and Craig Gotsman. Parallel progressive ray-tracing. Computer Graphics Forum, 16(1):43–55, March 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=115&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=16.

Niederberger:2003:CMH

[NG03a] C. Niederberger and M. Gross. Collision and motion: Hierarchical and heterogeneous reactive agents for real-time applications. Computer Graphics Forum, 22(3):323–331, September 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Niederberger:2003:HHR

[NG03b] C. Niederberger and M. Gross. Hierarchical and heterogeneous reactive agents for real-time applications. Computer Graphics Forum, 22(3):323–331, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neugebauer:2009:BVM

[NGB^{+}09] Mathias Neugebauer, Rocco Gasteiger, Oliver Beuing, Volker Diehl, Martin Skalej, and Bernhard Preim. Biomedical visualization: Map displays for the analysis of scalar data on cerebral aneurysm surfaces. Computer Graphics Forum, 28(3):895–902, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nadeau:2001:VSE

[NGN^{+}01] David R. Nadeau, Jon D. Genetti, Steve Napear, Bernard Pailthorpe, Carter Emmart, Erik Wesselak, and Dennis Davidson. Visualizing stars and emission nebulas. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nakamura:1997:RCI

[NHH97] Hiroyuki Nakamura, Masatake Higashi, and Mamoru Hosaka. Robust computation of intersection graph between two solids. Computer Graphics Forum, 16(3):C79–C88, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=144&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Nicol:1985:MSF

[Nic85] C. J. Nicol. Modelling solids in four dimensions. Computer Graphics Forum, 4(3):239–244, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishita:1997:MRM

[NIDN97] Tomoyuki Nishita, Hiroshi Iwasaki, Yoshinori Dobashi, and Eihachiro Nakamae. A modeling and rendering method for snow by using metaballs. Computer Graphics Forum, 16(3): C357–C364, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=173&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Nielsen:1995:LCU

[Nie95] Hans Peter Nielsen. Line clipping using semi-homogeneous coordinates. Computer Graphics Forum, 14(1):3–16, March 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Na:2004:RAF

[NJ04] Kyunggun Na and Moonryul Jung. Retargeting and animating faces: Hierarchical retargetting of fine facial motions. Computer Graphics Forum, 23(3):687–695, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neugebauer:1999:TMR

[NK99] Peter J. Neugebauer and Konrad Klein. Texturing 3D models of real world objects from multiple unregistered photographic views. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=345&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Nowrouzezahrai:2009:SGI

[NKF09] Derek Nowrouzezahrai, Evangelos Kalogerakis, and Eugene Fiume. Shadows and global illumination: Shadowing dynamic scenes with arbitrary BRDFs. Computer Graphics Forum, 28(2): 249–258, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nielson:1993:DV

[NKP93] G. M. Nielson, A. Kaufman, and F. Post. Data visualization. Computer Graphics Forum, 12(3):C509–??, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2008:IBT

[NLED08] Minh Hoai Nguyen, Jean-François Lalonde, Alexei A. Efros, and Fernando De la Torre. Image-based techniques: Image-based shaving. Computer Graphics Forum, 27(2):627–635, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neelamkavil:1991:MTS

[NM91] F. Neelamkavil and O. Mullarney. A methodology and tool set for supporting the development of graphical user interfaces. Computer Graphics Forum, 10(1):37–47, March 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nealen:2006:SAR

[NMK^{+}06] Andrew Nealen, Matthias Müller, Richard Keiser, Eddy Boxerman, and Mark Carlson. State of the art reviews: Physically based deformable models in computer graphics. Computer Graphics Forum, 25(4):809–836, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nicoll:2005:TFF

[NMMK05] A. Nicoll, J. Meseth, G. Müller, and R. Klein. Texture: Fractional Fourier texture masks: Guiding near-regular texture synthesis. Computer Graphics Forum, 24(3):569–579, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:1998:ILM

[NMNP98] László Neumann, Krešimir Matković, Attila Neumann, and Werner Purgathofer. Incident light metering in computer graphics. Computer Graphics Forum, 17(4):??, December 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=287&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=17.

Nonato:2001:NAD

[NMOT01] L. Nonato, R. Minghim, M. C. F. Oliveira, and G. Tavares. A novel approach for Delaunay 3D reconstruction with a comparative analysis in the light of applications. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:1998:PBC

[NMP98] László Neumann, Kresimir Matkovic, and Werner Purgathofer. Perception based color image difference. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=270&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Neumann:1989:PIA

[NN89] Laszlo Neumann and Attila Neumann. Photosimulation: Interreflection with arbitrary reflectance models and illuminations. Computer Graphics Forum, 8(1):21–34, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishita:1993:NRA

[NN93] T. Nishita and E. Nakamae. A new radiosity approach using area sampling for parametric patches. Computer Graphics Forum, 12 (3):385–398, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishita:1994:MDM

[NN94] T. Nishita and E. Nakamae. A method for displaying metaballs by using Bézier clipping. Computer Graphics Forum, 13(3):C/271– C/280, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:1997:RWD

[NNB97] László Neumann, Attila Neumann, and Philippe Bekaert. Radiosity with well distributed ray sets. Computer Graphics Forum, 16 (3):C261–C269, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=163&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Neumann:1999:RMF

[NNSK99a] László Neumann, Attila Neumann, and László Szirmay-Kalos. Reflectance models with fast importance sampling. Computer Graphics Forum, 18(4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=378&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=18.

Neumann:1999:CMR

[NNSK99b] László Neumann, Attila Neumannn, and László Szirmay-Kalos. Compact metallic reflectance models. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=337&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Nagai:2009:RIS

[NOS09] Yukie Nagai, Yutaka Ohtake, and Hiromasa Suzuki. Reconstruction — irregular sampling: Smoothing of partition of unity implicit surfaces for noise robust surface reconstruction. Computer Graphics Forum, 28(5):1339–1348, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noe:2000:HRB

[NP00] N. Noé and B. Péroche. Hierarchical reconstruction of BRDFs using locally supported functions. Computer Graphics Forum, 19 (2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nehlig:1995:FQA

[NR95] P. W. Nehlig and J.-P. Reveilles. Fractals and quasi-affine transformations. Computer Graphics Forum, 14(2):147–158 (or 147– 157??), June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Navazo:2003:SCC

[NRJS03a] I. Navazo, J. Rossignac, J. Jou, and R. Shariff. ShieldTester: Cell-to-cell visibility test for surface occluders. Computer Graphics Forum, 22(3):291–302, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Navazo:2003:VSC

[NRJS03b] I. Navazo, J. Rossignac, J. Jou, and R. Shariff. Visibility: Shieldtester: Cell-to-cell visibility test for surface occluders. Computer Graphics Forum, 22(3):291–302, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ng:1993:MIR

[NS93] A. Ng and M. Slater. A multiprocessor implementation of radiosity. Computer Graphics Forum, 12(5):329–342, December 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2001:RHQ

[NS01] Ky Giang Nguyen and Dietmar Saupe. Rapid high quality compression of volume data for visualization. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nowrouzezahrai:2009:RTG

[NS09] Derek Nowrouzezahrai and John Snyder. Real time global illumination: Fast global illumination on dynamic height fields. Computer Graphics Forum, 28(4):1131–1139, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:2006:RCA

[NSGP06] László Neumann, Mateu Sbert, Bruce Gooch, and Werner Purgathofer. Reports: Computational Aesthetics 2005 Eurographics Workshop on Computational Aesthetics in Graphics, Visualization and Imaging Girona, Spain, 18–20 May 2005. Computer Graphics Forum, 25(1):145–146, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nichols:2009:RTG

[NSW09] Greg Nichols, Jeremy Shopf, and Chris Wyman. Real time global illumination: Hierarchical image-space radiosity for interactive global illumination. Computer Graphics Forum, 28(4):1141–1149, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nasri:2009:FSR

[NSY09] A. Nasri, M. Sabin, and Z. Yasseen. Filling N-sided regions by quad meshes for subdivision surfaces. Computer Graphics Forum, 28(6):1644–1658, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noser:1995:SVA

[NT95] H. Noser and D. Thalmann. Synthetic vision and audition for digital actors. Computer Graphics Forum, 14(3):C/325–C/336, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Norton-Wayne:1991:ILU

[NW91] Len Norton-Wayne. Inspection of lace using machine vision. Computer Graphics Forum, 10(2):113–119, June 1991. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nasri:2001:RSA

[NW01] Ahmad H. Nasri and Brian Wyvill. A recursive subdivision algorithm for piecewise circular spline. Computer Graphics Forum, 20 (1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Naruse:1987:SDC

[NYTN87] Tadashi Naruse, Masaharu Yoshida, Tokiichiro Takahashi, and Seiichiro Naito. Sight — a dedicated computer graphics machine. Computer Graphics Forum, 6(4):327–334, December 1987. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okabe:2009:TVA

[OAIS09] Makoto Okabe, Ken Anjyo, Takeo Igarashi, and Hans-Peter Seidel. Textures and video: Animating pictures of fluid using video examples. Computer Graphics Forum, 28(2):677–686, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohtake:2001:MOP

[OB01] Yutaka Ohtake and Alexander G. Belyaev. Mesh optimization for polygonized isosurfaces. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ouhyoung:1996:RRO

[OCL96] Ming Ouhyoung, Yung-Yu Chuang, and Rung-Huei Liang. Reusable radiosity objects. Computer Graphics Forum, 15(3): 347–356, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

OSullivan:2002:LDC

[OCV^{+}02] C. O’Sullivan, J. Cassell, H. Vilhjálmsson, J. Dingliana, S. Dobbyn, B. McNamee, C. Peters, and T. Giang. Levels of detail for crowds and groups. Computer Graphics Forum, 21(4):733–741, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Osland:1984:MS

[OF84] C. D. Osland and A. H. Francis. Metafile standards. Computer Graphics Forum, 3(4):301–303, December 1984. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oztireli:2009:MRF

[ÖGG09] A. C. Öztireli, G. Guennebaud, and M. Gross. Model reconstruction: Feature preserving point set surfaces based on nonlinear kernel regression. Computer Graphics Forum, 28(2):493– 501, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

OHara:2002:HIF

[O’H02] Noel O’Hara. Hierarchical impostors for the flocking algorithm in 3D. Computer Graphics Forum, 21(4):723–731, November 2002.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Obermaier:2009:FVV

[OHBKH09] Harald Obermaier, Martin Hering-Bertram, Jörg Kuhnert, and Hans Hagen. Flow visualization: Volume deformations in gridless flow simulations. Computer Graphics Forum, 28(3):879–886, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ojala:1991:DIW

[OIST91] Jari Ojala, Kenji Inoue, Ken Sasaki, and Masaharu Takano. Development of an intelligent wheelchair using computer graphics animation and simulation. Computer Graphics Forum, 10(4):285– 295, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Obert:2008:ARI

[OKP^{+}08] Juraj Obert, Jaroslav Křrivánek, Fabio Pellacini, Daniel Sýkora, and Sumanta Pattanaik. Artistic rendering: iCheat: a representation for artistic control of indirect cinematic lighting. Computer Graphics Forum, 27(4):1217–1223, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owens:2007:SAR

[OLG^{+}07] John D. Owens, David Luebke, Naga Govindaraju, Mark Harris, Jens Krüger, Aaron E. Lefohn, and Timothy J. Purcell. State of the art review: a survey of general-purpose computation on graphics hardware. Computer Graphics Forum, 26(1):80–113, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ollila:2004:RES

[Oll04] Mark Ollila. Reports: Eurographics Symposium on Rendering: Norrköping, Sweden, 21–23 June 2004. Computer Graphics Forum, 23(4):843, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oshita:2001:DMC

[OM01] Masaki Oshita and Akifumi Makinouchi. A dynamic motion control technique for human-like articulated figures. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohbuchi:2002:FDA

[OMT02] Ryutarou Ohbuchi, Akio Mukaiyama, and Shigeo Takahashi. A frequency-domain approach to watermarking 3D shapes. Computer Graphics Forum, 21(3):373–382, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Onoue:2005:IDS

[ON05] Koichi Onoue and Tomoyuki Nishita. An interactive deformation system for granular material. Computer Graphics Forum, 24(1): 51–60, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okabe:2005:MNI

[OOI05] Makoto Okabe, Shigeru Owada, and Takeo Igarash. Modelling nature: Interactive design of botanical trees using freehand sketches and example-based editing. Computer Graphics Forum, 24(3): 487–496, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oliva:1996:RCP

[OPC96] J. M. Oliva, M. Perrin, and S. Coquillart. 3D reconstruction of complex polyhedral shapes from contours using a simplified generalized Voronoi diagram. Computer Graphics Forum, 15(3): C397–C408, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Orti:1996:RDS

[ORDP96] Rachel Orti, Stephane Riviere, Fredo Durand, and Claude Puech. Radiosity for dynamic scenes in Flatland with the visibility complex. Computer Graphics Forum, 15(3):C237–C248, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ochotta:2008:IBS

[OS08] Tilo Ochotta and Dietmar Saupe. Image-based surface compression. Computer Graphics Forum, 27(6):1647–1663, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ovsjanikov:2008:SPL

[OSG08] Maks Ovsjanikov, Jian Sun, and Leonidas Guibas. Shape properties: Local and global: Global intrinsic symmetries of shapes.

Computer Graphics Forum, 27(5):1341–1348, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oshita:2008:ASM

[Osh08] Masaki Oshita. Animation: Smart motion synthesis. Computer Graphics Forum, 27(7):1909–1918, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Osland:1982:MA

[Osl82] C. D. Osland. Metafiles at Albuquerque. Computer Graphics Forum, 1(2):47–49, June 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oore:2002:LPM

[OTH02] Sageev Oore, Demetri Terzopoulos, and Geoffrey Hinton. Local physical models for interactive character animation. Computer Graphics Forum, 21(3):337–346, September 2002. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otaduy:2009:DIC

[OTSG09] Miguel A. Otaduy, Rasmus Tamstorf, Denis Steinemann, and Markus Gross. Deformations: Implicit contact handling for deformable objects. Computer Graphics Forum, 28(2):559–568, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otten:1990:RFE

[Ott90] Daan Otten. Report on the First Eurographics Workshop on Object-Oriented Graphics. Computer Graphics Forum, 9(3):285– 286, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oddy:1991:PBC

[OW91] Robert J. Oddy and Philip J. Willis. A physically based colour model. Computer Graphics Forum, 10(2):121–127, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1986:CGC

[Owe86] J. Owen. Computer graphics and CAD literature — a keywordindexed bibliography. Computer Graphics Forum, 5(3):233–263, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1987:CGC

[Owe87] Jon Owen. Computer graphics and CAD literature: a keywordindexed biography. Computer Graphics Forum, 6(3):235–254, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1988:CGC

[Owe88] Jon Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1987. Computer Graphics Forum, 7(3):215–226, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1989:BR

[Owe89a] J. Owen. Book reviews. Computer Graphics Forum, 8(4):359– 360, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1989:CGC

[Owe89b] Jon Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1988. Computer Graphics Forum, 8(3):249–261, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1990:BR

[Owe90a] J. Owen. Book reviews. Computer Graphics Forum, 9(4):363– 364, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1990:CGC

[Owe90b] Jon Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1989. Computer Graphics Forum, 9(3):257–267, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1992:CGC

[Owe92a] J. Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1991. Computer Graphics Forum, 11(4):247–259, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1992:CGL

[Owe92b] J. Owen. Computer graphics literature: a keyword-indexed bibliography for the year 1990. Computer Graphics Forum, 11(1): 65–75, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1993:CGC

[Owe93] J. Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1992. Computer Graphics Forum, 12(4):231–242, October 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1994:CGC

[Owe94] J. Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1993. Computer Graphics Forum, 13(4):253–266, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Owen:1995:CGC

[Owe95] Jon Owen. Computer graphics and computer-aided design literature: a keyword-indexed bibliography for the year 1994. Computer Graphics Forum, 14(5):295–306, December 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oshima:1992:CSC

[OYSO92] T. Oshima, S. Yuasa, K. Sakanoshita, and Y. Ogata. A CAD system for color design of a car. Computer Graphics Forum, 11 (3):C381–C390, C483, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Olson:2006:GSE

[OZ06] Matt Olson and Hao Zhang. Geometry: Silhouette extraction in Hough space. Computer Graphics Forum, 25(3):273–282, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Olson:2009:TDF

[OZ09] M. Olson and H. Zhang. Tangential distance fields for mesh silhouette problems. Computer Graphics Forum, 28(1):84–100, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Orvalho:2008:TRA

[OZS08] Verónica Costa Orvalho, Ernesto Zacur, and Antonio Susin. Transferring the rig and animations from a character to different face models. Computer Graphics Forum, 27(8):1997–2012, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Premoze:2001:RNW

[PA01] Simon Premoze and Michael Ashikhmin. Rendering natural waters. Computer Graphics Forum, 20(4):??, December 2001. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Papageorgiou:2006:TNV

[PA06] Stavros G. Papageorgiou and Nikos A. Aspragathos. Transformation and normal vector calculation of parametrically defined surfaces based on dual vectors and screw theory: Application to Phong’s shading model. Computer Graphics Forum, 25(2):183– 195, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pai:2002:SIS

[Pai02] Dinesh K. Pai. STRANDS: Interactive simulation of thin solids using Cosserat models. Computer Graphics Forum, 21(3):347–352, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parsons:1986:GLU

[Par86] Michael S. Parsons. Generating lines using quadgraph patterns. Computer Graphics Forum, 5(1):33–39, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parsons:1989:IRU

[Par89] Michael S. Parsons. Image representations using Miranda Laws. Computer Graphics Forum, 8(2):99–106, June 1989. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pascucci:2002:SGS

[Pas02] Valerio Pascucci. Slow growing subdivision (SGS) in any dimension: Towards removing the curse of dimensionality. Computer Graphics Forum, 21(3):451–460, September 2002. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patel:1989:MRE

[Pat89] Manjula Patel. A method for the representation, evaluation and display of CSG models in PHIGS and PHIGS+. Computer Graphics Forum, 8(4):337–345, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patterson:1993:FS

[Pat93] J. W. Patterson. Fast spheres. Computer Graphics Forum, 12 (3):C61–C71, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patel:1995:CIC

[Pat95] M. Patel. Colouration issues in computer generated facial animation. Computer Graphics Forum, 14(2):117–126, June 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pun:1990:RBI

[PB90] Thierry Pun and Edwin Blake. Relationships between image synthesis and analysis: Towards unification? Computer Graphics Forum, 9(2):149–163, June 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pattanaik:1994:FWR

[PB94] S. N. Pattanaik and K. Bouatouch. Fast wavelet radiosity method. Computer Graphics Forum, 13(3):C/407–C/420, ???? 1994. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palanque:1995:DSV

[PB95] P. Palanque and R. Bastide, editors. Design, specification, and verification of interactive systems ’95: proceedings of the Eurographics Workshop in Toulouse, France, June 7–9, 1995. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82739-0. LCCN T385 .E891 1995. This book is the final outcome of the Eurographics Workshop on Design, Specification and Verification of Interactive Systems, that was held in Bonas from June 7 to 9, 1995.

Pons:2007:LAD

[PB07] J.-P. Pons and J.-D. Boissonnat. A Lagrangian approach to dynamic interfaces through kinetic triangulation of the ambient space. Computer Graphics Forum, 26(2):227–239, June 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Promayon:1996:PBD

[PBP96] E. Promayon, P. Baconnier, and C. Puech. Physically-based deformations constrained in displacements and volume. Computer Graphics Forum, 15(3):C155–C164, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patterson:1992:CHS

[PC92] J. W. Patterson and G. Cockton. Composing hierarchically structured images. Computer Graphics Forum, 11(3):C311–C320, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palamidese:1994:ARM

[PC94] P. Palamidese and A. Crise. Animations to represent multivariate data. Computer Graphics Forum, 13(4):243–248, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paterson:2005:RIB

[PCF05] James A. Paterson, David Claus, and Andrew W. Fitzgibbon. Rendering II: BRDF and geometry capture from extended inhomogeneous samples using flash photography. Computer Graphics Forum, 24(3):383–391, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pavic:2009:GGI

[PCK09] D. Pavić, U. Ceumern, and L. Kobbelt. GIzMOs: Genuine image mosaics with adaptive tiling. Computer Graphics Forum, 28(8): 2244–2254, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perez-Cazares:1989:DAL

[PCR89] J.-R. Perez-Cazares and R. Rannou. Developing an animated local network simulator. Computer Graphics Forum, 8(2):139– 142, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paterno:1994:DSV

[PCS94] F. Paterno, I. Campari, and R. Scopigno. The design and specification of a visual language: An example for customising geographic information systems functionalities. Computer Graphics Forum, 13(4):199–210, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Platings:2004:CLS

[PD04] M. Platings and A. M. Day. Compression of large-scale terrain data for real-time visualization using a tiled quad tree. Computer Graphics Forum, 23(4):741–759, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Prime:1990:UIM

[PF90] Martin Prime and D. Faguy. User interface management systems. A current product review. Computer Graphics Forum, 9(1):53– 76, March 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pingi:2005:GIE

[PFC^{+}05] Paolo Pingi, Andrea Fasano, Paolo Cignoni, Claudio Montani, and Roberto Scopigno. Geometry III: Exploiting the scanning sequence for automatic registration of large sets of range maps. Computer Graphics Forum, 24(3):517–526, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pla-Garcia:1993:BOS

[PG93] N. Pla-Garcia. Boolean operations and spatial complexity of face octrees. Computer Graphics Forum, 12(3):C153–C164, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pla-Garcia:1994:RSB

[PG94] N. Pla-Garcia. Recovering a smooth boundary representation from an edge quadtree and from a Face Octree. Computer Graphics Forum, 13(4):189–198, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palmer:1995:CDA

[PG95] I. J. Palmer and R. L. Grimsdale. Collision detection for animation using sphere-trees. Computer Graphics Forum, 14(2): 105–116, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pekelny:2008:CAA

[PG08] Yuri Pekelny and Craig Gotsman. Character animation: Articulated object reconstruction and markerless motion capture from depth video. Computer Graphics Forum, 27(2):399–408, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Potter:2009:RIN

[PGG^{+}09] Kristin Potter, Amy Gooch, Bruce Gooch, Peter Willemsen, Joe Kniss, Richard Riesenfeld, and Peter Shirley. Resolution independent NPR-style 3D line textures. Computer Graphics Forum, 28(1):56–66, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peytavie:2009:ICT

[PGGM09a] A. Peytavie, E. Galin, J. Grosjean, and S. Merillou. Interactive cities and terrain: Arches: a framework for modeling complex terrains. Computer Graphics Forum, 28(2):457–467, April 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peytavie:2009:MDP

[PGGM09b] A. Peytavie, E. Galin, J. Grosjean, and S. Merillou. Modeling and deformation: Procedural generation of rock piles using aperiodic tiling. Computer Graphics Forum, 28(7):1801–1809, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palacios:1996:DCS

[PGR96] Antonio Palacios, Lee M. Gross, and Alyn P. Rockwood. Dynamics and chaos: The spherical pendulum. Computer Graphics Forum, 15(4):263–270, October 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popov:2007:RTS

[PGSS07] Stefan Popov, Johannes Günther, Hans-Peter Seidel, and Philipp Slusallek. Ray tracing: Stackless KD-tree traversal for high performance GPU ray tracing. Computer Graphics Forum, 26(3): 415–424, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patel:1987:SMS

[PH87] Manjula Patel and Roger J. Hubbold. A scanline method for solid model display. Computer Graphics Forum, 6(2):141–150, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Post:1991:RSE

[PH91] Frits H. Post and Andrea J. S. Hin. Report on the Second Eurographics Workshop on Visualization in Scientific Computing. Computer Graphics Forum, 10(3):261–264, September 1991. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Preston:1994:AUN

[PH94] M. Preston and W. T. Hewitt. Animation using NURBS. Computer Graphics Forum, 13(4):229–241, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Preston:1996:ICA

[PH96] Martin Preston and Terry Hewitt. Integrating computer animation and multimedia. Computer Graphics Forum, 15(3):C377– C386, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pritchard:2003:CSC

[PH03a] D. Pritchard and W. Heidrich. Capture and segmentation: Cloth motion capture. Computer Graphics Forum, 22(3):263–271, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pritchard:2003:CMC

[PH03b] D. Pritchard and W. Heidrich. Cloth motion capture. Computer Graphics Forum, 22(3):263–271, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patterson:1991:IDM

[PHL91] J. W. Patterson, S. G. Hoggar, and J. R. Logie. Inverse displacement mapping. Computer Graphics Forum, 10(2):129–139, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1986:BCD

[Pic86a] C. A. Pickover. Biomorphs: Computer displays of biological forms generated from mathematical feedback loops. Computer Graphics Forum, 5(4):313–316, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1986:MCA

[Pic86b] C. A. Pickover. A Monte Carlo approach for ϵ placement in fractaldimension calculations for waveform graphs. Computer Graphics Forum, 5(3):203–209, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1987:GBO

[Pic87] C. A. Pickover. Graphics, bifurcation, order and chaos. Computer Graphics Forum, 6(1):26–33, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1991:MCB

[Pic91a] C. Pickover. From math comes beauty: Monkey curves, spirals, and saddles. Computer Graphics Forum, 10(4):329–331, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1991:MPS

[Pic91b] Cliff Pickover. Million-point sculptures. Computer Graphics Forum, 10(4):333–336, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piller:1985:GPM

[Pil85] Ernst Piller. A general-purpose multi-microprocessor raster graphics display system with anti-aliasing. Computer Graphics Forum, 4(1):33–41, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pink:1985:EUC

[Pin85] J. Pink. EUROGRAPHICS UK chapter conference. Computer Graphics Forum, 4(3):172–175, September 1985. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pineda:1992:RSC

[Pin92] L. A. Pineda. Reference, synthesis and constraint satisfaction. Computer Graphics Forum, 11(3):C333–C344, ???? 1992. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poupyrev:1998:EOM

[PIWB98] I. Poupyrev, T. Ichikawa, S. Weghorst, and M. Billinghurst. Egocentric object manipulation in virtual environments: Empirical evaluation of interaction techniques. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=252&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Paulin:1994:AMG

[PJ94] M. Paulin and J.-P. Jessel. Adaptive mesh generation for progressive radiosity: a ray-tracing based algorithm. Computer Graphics Forum, 13(3):C/421–C/432, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pavic:2008:SRH

[PK08] Darko Pavić and Leif Kobbelt. Scanning and reconstruction: High-resolution volumetric computation of offset surfaces with feature preservation. Computer Graphics Forum, 27(2):165–174, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pauly:2003:CSM

[PKG03a] Mark Pauly, Richard Keiser, and Markus Gross. Capture and segmentation: Multi-scale feature extraction on point-sampled surfaces. Computer Graphics Forum, 22(3):281–289, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pauly:2003:MSF

[PKG03b] Mark Pauly, Richard Keiser, and Markus Gross. Multi-scale feature extraction on point-sampled surfaces. Computer Graphics Forum, 22(3):281–289, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pineda:1988:GUD

[PKL88] Luis A. Pineda, Ewan Klein, and John Lee. GRAFLOG: Understanding drawings through natural language. Computer Graphics Forum, 7(2):97–103, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Petz:2009:FVH

[PKPH09] Christoph Petz, Jens Kasten, Steffen Prohaska, and Hans-Christian Hege. Flow visualization: Hierarchical vortex regions in swirling flow. Computer Graphics Forum, 28(3):863–870, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paterno:1994:SBA

[PL94] F. Paterno and A. Leonardi. A semantics-based approach for the design and implementation of interaction objects. Computer Graphics Forum, 13(3):C/195–C/204, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poulet:1996:MMS

[PL96] François Poulet and Michel Lucas. Modelling megalithic sites. Computer Graphics Forum, 15(3):C279–C288, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pan:2007:GIP

[PLPB07] Minghao Pan, Rui Wang Xinguo Liu, Qunsheng Peng, and Hujun Bao. Global illumination (2): Precomputed radiance transfer field for rendering interreflections in dynamic scenes. Computer Graphics Forum, 26(3):485–493, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pandzic:1997:FAV

[PLT^{+}97] Igor Sunday Pandzic, Elwin Lee, Nadia Magnenat Thalmann, Tolga K. Capin, and Daniel Thalmann. A flexible architecture for virtual humans in networked collaborative virtual environments. Computer Graphics Forum, 16(3):C177– C188, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=154&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Pattanaik:1993:PEI

[PM93] S. N. Pattanaik and S. P. Mudur. The potential equation and importance in illumination computations. Computer Graphics Forum, 12(2):131–136, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popescu:2006:RRS

[PMDS06] Voicu Popescu, Chunhui Mei, Jordan Dauble, and Elisha Sacks. Rendering: Reflected-scene impostors for realistic reflections at interactive rates. Computer Graphics Forum, 25(3):313–322, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poet:1986:DRF

[PMW86] Ron Poet and E. James Milner-White. Displaying relevant features of protein molecules. Computer Graphics Forum, 5(3): 211–215, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peters:1997:CVS

[PN97] Jörg Peters and Ahmad Nasri. Computing volumes of solids enclosed by recursive subdivision surfaces. Computer Graphics Forum, 16(3):C89–C94, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=145&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Poller:1989:MGG

[PNR89] Jochem Poller, Stefan Noll, and Joachim Rix. Migration of GKS/GKS-3D and PHIGS discussed under the view of the computer graphics reference model. Computer Graphics Forum, 8(3): 239–248, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pitteway:1985:FEP

[PO85] M. L. V. Pitteway and P. M. Olive. Filtering edges by pixel integration. Computer Graphics Forum, 4(2):111–116, June 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peters:2002:SVM

[PO02] C. Peters and C. O’Sullivan. Synthetic vision and memory for autonomous virtual humans. Computer Graphics Forum, 21(4): 743–752, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pietroni:2007:ITI

[POB^{+}07] Nico Pietroni, Miguel A. Otaduy, Bernd Bickel, Fabio Ganovelli, and Markus Gross. Imaging (2): Texturing internal surfaces from a few cross sections. Computer Graphics Forum, 26(3):637–644, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Plummer:1989:MMA

[PP89] Michael Plummer and David Penna. Mass market applications for real time 3D graphics. Computer Graphics Forum, 8(2):143– 150, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patow:2005:SIS

[PP05] Gustavo Patow and Xavier Pueyo. A survey of inverse surface design from light transport behavior specification. Computer Graphics Forum, 24(4):773–789, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pegoraro:2009:CSA

[PP09a] Vincent Pegoraro and Steven G. Parker. Caustics and scattering: an analytical solution to single scattering in homogeneous participating media. Computer Graphics Forum, 28(2):329–335, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poelke:2009:FLD

[PP09b] Konstantin Poelke and Konrad Polthier. Functions: Lifted domain coloring. Computer Graphics Forum, 28(3):735–742, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paquette:1998:LHF

[PPD98] Eric Paquette, Pierre Poulin, and George Drettakis. A light hierarchy for fast rendering of scenes with many lights. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=254&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Paris:2007:CAP

[PPD07] Sébastien Paris, Julien Pettré, and Stéphane Donikian. Crowds and animation: Pedestrian reactive navigation for crowd simulation: a predictive approach. Computer Graphics Forum, 26(3): 665–674, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pickover:1993:EKM

[PR93] C. A. Pickover and L. Rudolph. Electronic kaleidoscopes for the mind. Computer Graphics Forum, 12(1):41–46, March 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Prince:1985:SER

[Pri85] M. D. Prince. A survey of electronics R and D in the USA. Computer Graphics Forum, 4(4):301–318, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paoluzzi:1989:PHS

[PRS89] A. Paoluzzi, M. Rosina, and R. Scopigno. A parallel HW/SW environment for image synthesis. Computer Graphics Forum, 8 (3):227–236, September 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pfeifle:1995:STB

[PS95] R. Pfeifle and H.-P. Seidel. Spherical triangular B-splines with application to data fitting. Computer Graphics Forum, 14(3): C/89–C/96, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pfeifle:1996:FTB

[PS96a] Ron Pfeifle and Hans-Peter Seidel. Fitting triangular B-splines to functional scattered data. Computer Graphics Forum, 15(1): 15–23, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pueyo:1996:RTP

[PS96b] Xavier Pueyo and Peter Schroder, editors. Rendering techniques ’96: proceedings of the Eurographics Workshop in Porto, Portugal, June 17–19, 1996, Springer Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-211-82883-4.

Patane:2004:GGB

[PSF04] Giuseppe Patanè, Michela Spagnuolo, and Bianca Falcidieno. Para-Graph: Graph-based parameterization of triangle meshes with arbitrary genus. Computer Graphics Forum, 23(4):783–797, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ponto:2009:WII

[PSK09] K. Ponto, M. Seracini, and F. Kuester. Wipe-Off: an intuitive interface for exploring ultra-large multi-spectral data sets for cultural heritage diagnostics. Computer Graphics Forum, 28(8): 2291–2301, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Park:1998:OFR

[PSL98] Tae-Joon Park, Sung Yong Shin, and Seungyong Lee. Optical flow rendering. Computer Graphics Forum, 17(3): ??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=255&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Pueyo:1988:SCA

[PT88] Xavier Pueyo and Daniela Tost. A survey of computer animation. Computer Graphics Forum, 7(4):281–300, December 1988. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Platis:2003:PHI

[PT03] Nikos Platis and Theoharis Theoharis. Progressive hulls for intersection applications. Computer Graphics Forum, 22(2):107–116, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Premzoe:2003:AFH

[PTB^{+}03a] Simon Premžoe, Tolga Tasdizen, James Bigler, Aaron Lefohn, and Ross T. Whitaker. Animation of fluids and hair: Particle-based simulation of fluids. Computer Graphics Forum, 22(3):401–410, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Premzoe:2003:PBS

[PTB^{+}03b] Simon Premžoe, Tolga Tasdizen, James Bigler, Aaron Lefohn, and Ross T. Whitaker. Particle-based simulation of fluids. Computer Graphics Forum, 22(3):401–410, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pudet:1994:RTF

[Pud94] T. Pudet. Real time fitting of hand-sketched pressure brushstrokes. Computer Graphics Forum, 13(3):C/205–C/220, ????

1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Purgathofer:2003:IPO

[Pur03a] Werner Purgathofer. Invited presentation: Open issues in photorealistic rendering. Computer Graphics Forum, 22(3):xix, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Purgathofer:2003:OIP

[Pur03b] Werner Purgathofer. Open issues in photo-realistic rendering. Computer Graphics Forum, 22(3):??, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Purgathofer:2007:RES

[Pur07] Werner Purgathofer. Reports: EUROGRAPHICS’2006 4 — 8 September, 2006 Vienna, Austria. Computer Graphics Forum, 26 (1):119–122, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pretorius:2008:GVV

[PV08] A. Johannes Pretorius and Jarke J. Van Wijk. Graph visualization: Visual inspection of multivariate graphs. Computer Graphics Forum, 27(3):967–974, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peeters:2009:IFS

[PVtHR09] T. H. J. M. Peeters, A. Vilanova, and B. M. ter Haar Romeny. Interactive fibre structure visualization of the heart. Computer Graphics Forum, 28(8):2140–2150, December 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pagendarm:1993:PCV

[PW93] H.-G. Pagendarm and B. Walter. A prototype of a cooperative visualization workplace for the aerodynamicist. Computer Graphics Forum, 12(3):C485–C496, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pan:2009:RAM

[PWC^{+}09] Minghao Pan, Rui Wang, Weifeng Chen, Kun Zhou, and Hujun Bao. Rendering and appearance modeling: Fast, sub-pixel antialiased shadow maps. Computer Graphics Forum, 28(7):1927– 1934, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poston:1998:MIE

[PWH98] Tim Poston, Tien-Tsin Wong, and Pheng-Ann Heng. Multiresolution isosurface extraction with adaptive skeleton climbing. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=261&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Pegoraro:2008:PMS

[PWP08] Vincent Pegoraro, Ingo Wald, and Steven G. Parker. Participating media: Sequential Monte Carlo adaptation in low-anisotropy participating media. Computer Graphics Forum, 27(4):1097–1104, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popa:2009:CMW

[PZB^{+}09] T. Popa, Q. Zhou, D. Bradley, V. Kraevoy, H. Fu, A. Sheffer, and W. Heidrich. Captured models: Wrinkling captured garments using space–time data-driven deformation. Computer Graphics Forum, 28(2):427–435, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pei:2008:SFK

[PZY08] Yuru Pei, Hongbin Zha, and Zhongbiao Yuan. Skull, face, and knits: The craniofacial reconstruction from the local structural diversity of skulls. Computer Graphics Forum, 27(7):1711–1718, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2003:FPR

[QNTN03a] Xueying Qin, Eihachiro Nakamae, Katsumi Tadamura, and Yasuo Nagai. Fast photo-realistic rendering of trees in daylight. Computer Graphics Forum, 22(3):243–252, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2003:RAN

[QNTN03b] Xueying Qin, Eihachiro Nakamae, Katsumi Tadamura, and Yasuo Nagai. Rendering and animating nature: Fast photo-realistic rendering of trees in daylight. Computer Graphics Forum, 22(3): 243–252, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:1992:RCU

[QSW92] Kaihuai Qin, Jiaguang Sun, and Xuefu Wang. Representing conics using NURBS of degree two. Computer Graphics Forum, 11(5): 285–291, December 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ris:1994:PRT

[RA94] P. Ris and D. Arques. Parallel ray tracing based upon a multilevel topological knowledge acquisition of the scene. Computer Graphics Forum, 13(3):C/221–C/232, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raffin:2005:RES

[Raf05] Bruno Raffin. Reports: 2004 Eurographics Symposium on Parallel Graphics and Visualization. Computer Graphics Forum, 24(1): 133, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rappoport:1992:EAA

[Rap92] Ari Rappoport. An efficient adaptive algorithm for constructing the convex differences tree of a simple polygon. Computer Graphics Forum, 11(4):235–240, October 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reitsma:2008:PRA

[RAP08] P. S. A. Reitsma, J. Andrews, and N. S. Pollard. Perception for rendering and animation: Effect of character animacy and preparatory motion on perceptual magnitude of errors in ballistic motion. Computer Graphics Forum, 27(2):201–210, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieger:2003:FSN

[RB03a] Thomas Rieger and Norbert Braun. Faces and speech: Narrative use of sign language by a virtual character for the hearing impaired. Computer Graphics Forum, 22(3):651–660, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieger:2003:NUS

[RB03b] Thomas Rieger and Norbert Braun. Narrative use of sign language by a virtual character for the hearing impaired. Computer Graphics Forum, 22(3):651–660, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2001:HMC

[RBB01] Holly Rushmeier, Laurent Balmelli, and Fausto Bernardini. Horizon map capture. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rautek:2008:IRV

[RBG08] Peter Rautek, Stefan Bruckner, and M. Eduard Gröller. Illustrative and realistic volume rendering: Interaction-dependent semantics for illustrative volume rendering. Computer Graphics Forum, 27(3):847–854, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rocchini:2001:LCS

[RCM^{+}01] C. Rocchini, P. Cignoni, C. Montani, P. Pingi, and R. Scopigno. A low cost 3D scanner based on structured light. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ryder:2005:SAR

[RD05] G. Ryder and A. M. Day. State of the art reviews: Survey of realtime rendering techniques for crowds. Computer Graphics Forum, 24(2):203–215, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roth:2001:RTT

[RDG01] S. H. Martin Roth, Patrick Diezi, and Markus H. Gross. Ray tracing triangular Bézier patches. Computer Graphics Forum, 20 (3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reddy:1996:SPD

[Red96] M. Reddy. SCROOGE: Perceptually-driven polygon reduction. Computer Graphics Forum, 15(4):191–203, October 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinhard:2003:FEW

[Rei03] Erik Reinhard. Fourth Eurographics Workshop on Parallel Graphics and Visualisation. Computer Graphics Forum, 22(2):218–219, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Requicha:1986:EPE

[Req86] Aristides A. G. Requicha, editor. EUROGRAPHICS ’86: proceedings of the European Computer Graphics Conference and Exhibition, the computer interface, Lisbon, Portugal, 25–19, August, 1986. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1986. ISBN 0-444-70065-X. LCCN T 385 E97 1986.

Reynolds:1986:TBG

[Rey86] Graham J. Reynolds. A token based graphics system. Computer Graphics Forum, 5(2):139–145, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ranjan:1996:MIS

[RF96] V. Ranjan and A. Fournier. Matching and interpolation of shapes using unions of circles. Computer Graphics Forum, 15(3):C129– C142, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reumann:1985:MMB

[RGM85] K. Reumann, G. Griebel, and K. Mildenstein. Manufacturing models of biomedical objects via CAD/CAM and GKS. Computer Graphics Forum, 4(4):375–382, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roth:1998:BBB

[RGTC98] S. H. Martin Roth, Markus H. Gross, Silvio Turello, and Friedrich R. Carls. A Bernstein–Bézier based approach to soft tissue simulation. Computer Graphics Forum, 17(3): ??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=275&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Rasche:2005:IVP

[RGW05] Karl Rasche, Robert Geist, and James Westall. Image and video processing: Re-coloring images for gamuts of lower dimension.

Computer Graphics Forum, 24(3):423–432, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roger:2006:RAS

[RH06] David Roger and Nicolas Holzschuch. Rendering: Accurate specular reflections in real-time. Computer Graphics Forum, 25(3): 293–302, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rohmer:2008:ALV

[RHC08] Damien Rohmer, Stefanie Hahmann, and Marie-Paule Cani. Animation: Local volume preservation for skinned characters. Computer Graphics Forum, 27(7):1919–1927, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roll:1995:FGL

[RHv95] Stefan Röll, Axel Haase, and Markus von Kienlin. Fast generation of leakproof surfaces from well-defined objects by a modified marching cubes algorithm. Computer Graphics Forum, 14(2): 127–138, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richard:1987:GAC

[Ric87a] F. Richard. Graphical analysis of complex O.D.E. solutions. Computer Graphics Forum, 6(4):335–341, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richards:1987:FCP

[Ric87b] J. E. Richards. Filling complex polygons by region-fill methods on raster graphics terminals. Computer Graphics Forum, 6(1):49– 54, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richards:1987:SAB

[Ric87c] J. E. Richards. Some aspects of binding GKS to C++. Computer Graphics Forum, 6(3):211–218, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritschel:2009:HPG

[RIF^{+}09] T. Ritschel, M. Ihrke, J. R. Frisvad, J. Coppens, K. Myszkowski, and H.-P. Seidel. Human perception and Green graphics: Temporal glare: Real-time dynamic simulation of the scattering in the human eye. Computer Graphics Forum, 28(2):183–192, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rossignac:1994:ABM

[RK94] J. Rossignac and A. Kaul. AGRELs and BIPs: Metamorphosis as a Bézier curve in the space of polyhedra. Computer Graphics Forum, 13(3):C/179–C/184, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiters:2009:AMR

[RK09a] Roland Ruiters and Reinhard Klein. Appearance measurement and rendering: BTF compression via sparse tensor decomposition. Computer Graphics Forum, 28(4):1181–1188, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiters:2009:MRH

[RK09b] Roland Ruiters and Reinhard Klein. Model reconstruction: Heightfield and spatially varying BRDF reconstruction for materials with interreflections. Computer Graphics Forum, 28(2): 513–522, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Redon:2002:FCC

[RKC02] Stéphane Redon, Abderrahmane Kheddary, and Sabine Coquillart. Fast continuous collision detection between rigid bodies. Computer Graphics Forum, 21(3):279–288, September 2002. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Russell:1984:CIA

[RL84] G. T. Russell and D. M. Lane. CAD for intelligent activity of an autonomous submersible system. Computer Graphics Forum, 3(1):13–27, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosenthal:2009:GRE

[RL09] Paul Rosenthal and Lars Linsen. Geometry & reconstruction: Enclosing surfaces for point clusters using 3D discrete Voronoi diagrams. Computer Graphics Forum, 28(3):999–1006, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rustamov:2009:MSI

[RLF09] R. M. Rustamov, Y. Lipman, and T. Funkhouser. Modeling and sculpting: Interior distance using barycentric coordinates. Computer Graphics Forum, 28(5):1279–1288, July 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rhee:2006:ART

[RLN06] Taehyun Rhee, J. P. Lewis, and Ulrich Neumann. Animation: Real-time weighted pose-space deformation on the GPU. Computer Graphics Forum, 25(3):439–448, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ray:2009:MST

[RLW^{+}09] N. Ray, B. Lévy, H. Wang, G. Turk, and B. Vallet. Material space texturing. Computer Graphics Forum, 28(6):1659–1669, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rush:1989:UTR

[RM89] Stephen Rush and Wendy J. Milne. Using a Transputer ring to transform raster images. Computer Graphics Forum, 8(2):151– 158, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rangel-Mondragon:1988:CGP

[RMA88] J. Rangel-Mondragon and S. J. Abas. Computer generation of Penrose tilings. Computer Graphics Forum, 7(1):29–37, March 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rudolf:2005:BDM

[RMN05] Dave Rudolf, David Mould, and Eric Neufeld. A bidirectional deposition model of wax crayons. Computer Graphics Forum, 24(1):27–39, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rump:2008:RSP

[RMS^{+}08] Martin Rump, Gero Müller, Ralf Sarlette, Dirk Koch, and Reinhard Klein. Rendering and simplification: Photo-realistic rendering of metallic car paint from image-based measurements. Computer Graphics Forum, 27(2):527–536, April 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ropinski:2008:RVE

[RMSD^{+}08] Timo Ropinski, Jennis Meyer-Spradow, Stefan Diepenbrock, Jörg Mensmann, and Klaus Hinrichs. Rendering volume effects: Interactive volume rendering with dynamic ambient occlusion and color bleeding. Computer Graphics Forum, 27(2):567–576, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruttkay:2003:EDE

[RNtH03] Zsófia Ruttkay, Han Noot, and Paul ten Hagen. Emotion disc and emotion squares: Tools to explore the facial expression space. Computer Graphics Forum, 22(1):49–53, March 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodriguez:2007:DDT

[RNV07] L. Rodríguez, I. Navazo, and À. Vinacua. Data-driven tetrahedral mesh subdivision. Computer Graphics Forum, 26(4):783–793, December 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Robinson:1987:ASD

[Rob87] K. Robinson. ACM SIGGRAPH ’86, Dallas, Texas, 18–22 August. Computer Graphics Forum, 6(1):55–62, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Robertson:1993:IVU

[Rob93] G. G. Robertson. Information visualization using 3D interactive animation. Computer Graphics Forum, 12(3):C512–??, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rokne:1997:ARI

[Rok97] Jon G. Rokne. Applying rounding-up integral linear interpolation to the scan-conversion of filled polygons. Computer Graphics Forum, 16(2):??, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=126&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Rosenthal:1982:RSW

[Ros82] D. S. H. Rosenthal. Report on: SIGGRAPH workshop on graphics input interaction techniques. Computer Graphics Forum, 1(3):91, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosenthal:1983:GIF

[Ros83] D. S. H. Rosenthal. The GKS input facilities and how to use them. Computer Graphics Forum, 2(2-3):97–103, August 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rossignac:1997:FEC

[Ros97] Jarek Rossignac. Finally everyone can work with highly complex 3D models. Computer Graphics Forum, 16(3):C1–??, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rougeron:1998:CFC

[RP98] Gilles Rougeron and Bernard Péroche. Color fidelity in computer graphics: a survey. Computer Graphics Forum, 17 (1):3–??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=212&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Reissell:2001:MSD

[RP01] L.-M. Reissell and Dinesh K. Pai. Modeling stochastic dynamical systems for interactive simulation. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roudier:1993:LSA

[RPP93] P. Roudier, B. Peroche, and M. Perrin. Landscapes synthesis achieved through erosion and deposition process simulation. Computer Graphics Forum, 12(3):375–383, ???? 1993. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2002:OSE

[RPZ02] Liu Ren, Hanspeter Pfister, and Matthias Zwicker. Object space EWA surface splatting: a hardware accelerated approach to high quality point rendering. Computer Graphics Forum, 21(3):461– 470, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ronfard:1994:TMC

[RR94] R. P. Ronfard and J. R. Rossignac. Triangulating multiplyconnected polygons: a simple, yet efficient algorithm. Computer Graphics Forum, 13(3):C/281–C/292, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ronfard:1996:FRA

[RR96] Remi Ronfard and Jarek Rossignac. Full-range approximation of triangulated polyhedra. Computer Graphics Forum, 15(3):C67– C76, C462, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rudolf:2000:MMD

[RR00] Marcin J. Rudolf and Jacek Raczkowski. Modeling the motion of dense smoke in the wind field. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=394&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Rieder:2008:VNI

[RRRP08] Christian Rieder, Felix Ritter, Matthias Raspe, and Heinz-Otto Peitgen. Visualization for neurology: Interactive visualization of multimodal volume data for neurosurgical tumor treatment. Computer Graphics Forum, 27(3):1055–1062, May 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roesch:1997:IVI

[RRS97a] Angela Roesch, Matthias Ruhl, and Dietmar Saupe. Interactive visualization of implicit surfaces with singularities. Computer Graphics Forum, 16(5):295–306, December 1997. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosch:1997:IVI

[RRS97b] Angela Rösch, Matthias Ruhl, and Dietmar Saupe. Interactive visualization of implicit surfaces with singularities. Computer Graphics Forum, 16(5):??, December 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=196&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=5&amp; src=
ard&amp; vid=16.

Reimers:2008:GRR

[RS08] Martin Reimers and Johan Seland. GPU rendering and raycasting: Ray casting algebraic surfaces using the frustum form. Computer Graphics Forum, 27(2):361–370, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rose:2001:ADI

[RSC01] Charles F. Rose III, Peter-Pike J. Sloan, and Michael F. Cohen. Artist-directed inverse-kinematics using radial basis function interpolation. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roberts:1988:FIN

[RSD^{+}88] W. Roberts, M. Slater, K. Drake, A. Simmins, A. Davison, and P. Williams. First impressions of NeWS. Computer Graphics Forum, 7(1):39–57, March 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rogers:1990:CGW

[RSK90] Greg Rogers, Kelvin Sung, and William Kubitz. Combining graphics and windowing standards in the XGKS system. Computer Graphics Forum, 9(3):229–237, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rezk-Salama:2006:VOP

[RSK06] Christof Rezk-Salama and Andreas Kolb. Visualization: Opacity peeling for direct volume rendering. Computer Graphics Forum, 25(3):597–606, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rungjiratananon:2008:ART

[RSKN08] Witawat Rungjiratananon, Zoltan Szego, Yoshihiro Kanamori, and Tomoyuki Nishita. Animation: Real-time animation of sandwater interaction. Computer Graphics Forum, 27(7):1887–1893, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rumpf:1996:FDA

[RSS96] Martin Rumpf, Alfred Schmidt, and Kunibert G. Siebert. Functions defining arbitrary meshes — a flexible interface between numerical data and visualization routines. Computer Graphics Forum, 15(2):129–141, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rezk-Salama:2008:IRV

[RSTK08] C. Rezk-Salama, S. Todt, and A. Kolb. Illustrative and realistic volume rendering: Raycasting of light field galleries from volumetric data. Computer Graphics Forum, 27(3):839–846, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinders:2002:VTV

[RSVP02] Freek Reinders, I. Ari Sadarjoen, Benjamin Vrolijk, and Frits H. Post. Vortex tracking and visualisation in a flow past a tapered cylinder. Computer Graphics Forum, 21(4):675–682, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Robinson:1997:FIP

[RSW^{+}97] Peter Robinson, Dan Sheppard, Richard Watts, Robert Harding, and Steve Lay. A framework for interacting with paper. Computer Graphics Forum, 16(3):C329–C334, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=170&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Reniers:2008:SSPb

[RT08a] Dennie Reniers and Alexandru Telea. Subdivision and segmentation: Part-type segmentation of articulated voxel-shapes using the junction rule. Computer Graphics Forum, 27(7):1845–1852, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reniers:2008:SSPa

[RT08b] Dennie Reniers and Alexandru Telea. Subdivision and segmentation: Patch-type segmentation of voxel shapes using simplified surface skeletons. Computer Graphics Forum, 27(7):1837–1844, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiz:2003:ICU

[RTN03a] David Gavilan Ruiz, Hiroki Takahashi, and Masayuki Nakajima. Image categorization using color blobs in a mobile environment.

Computer Graphics Forum, 22(3):427–432, September 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiz:2003:IQI

[RTN03b] David Gavilan Ruiz, Hiroki Takahashi, and Masayuki Nakajima. Image quantization: Image categorization using color blobs in a mobile environment. Computer Graphics Forum, 22(3):427–432, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2001:RIO

[Rus01] H. Rushmeier. Rendering: Input and output. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raskar:2004:ICM

[RvBWR04] Ramesh Raskar, Jeroen van Baar, Thomas Willwacher, and Srinivas Rao. Image calibration and matching: Quadric transfer for immersive curved screen displays. Computer Graphics Forum, 23 (3):451–460, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raman:2008:MVQ

[RW08] Sundaresan Raman and Rephael Wenger. Meshes and volumes: Quality isosurface mesh generation using an extended marching cubes lookup table. Computer Graphics Forum, 27(3):791–798, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roch:1988:IMP

[RWP88] Michel Roch, Jacques Weber, and Christian Pellegrini. 3D images of molecular properties by triangulation of contour lines. Computer Graphics Forum, 7(3):195–201, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2008:RGB

[RZLG08] Zhong Ren, Kun Zhou, Stephen Lin, and Baining Guo. Rendering: Gradient-based interpolation and sampling for real-time rendering of inhomogeneous, single-scattering media. Computer Graphics Forum, 27(7):1945–1953, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simonetto:2009:GVF

[SAA09] Paolo Simonetto, David Auber, and Daniel Archambault. Graph visualization: Fully automatic visualisation of overlapping sets. Computer Graphics Forum, 28(3):967–974, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sabin:1982:STS

[Sab82] M. A. Sabin. A short term step in the development of 3D graphics standards. Computer Graphics Forum, 1(3):129–132, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sabin:1986:SCM

[Sab86] Malcolm Sabin. A survey of contouring methods. Computer Graphics Forum, 5(4):325–339, December 1986. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Setlur:2005:PSV

[SABG^{+}05] Vidya Setlur, Conrad Albrecht-Buehler, Amy A. Gooch, Sam Rossoff, and Bruce Gooch. Perception: Semanticons: Visual metaphors as file icons. Computer Graphics Forum, 24(3):647– 656, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sadri:2009:RRS

[Sad09] Bardia Sadri. Reconstruction — regular sampling: Manifold homotopy via the flow complex. Computer Graphics Forum, 28 (5):1361–1370, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shirman:1993:CNT

[SAE93] L. A. Shirman and S. S. Abi-Ezzi. The cone of normals technique for fast processing of curved patches. Computer Graphics Forum, 12(3):C261–C272, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soede:1991:GIM

[SAHt91] Dirk Soede, Farhad Arbab, Ivan Herman, and Paul J. W. ten Hagen. The GKS input model in Manifold. Computer Graphics Forum, 10(3):209–224, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salesin:1996:IEI

[Sal96] D. Salesin. Image editing, image querying, and video clip art: Three multiresolution applications in computer graphics. Computer Graphics Forum, 15(3):452, September 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Samara:1993:DIG

[Sam93a] V. Samara. DaScript3D: integrating 3D graphics in a page description language. Computer Graphics Forum, 12(3):C315–C327, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Samet:1993:GIS

[Sam93b] H. Samet. Geographic information systems. Computer Graphics Forum, 12(3):C511–??, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santos:1992:CAH

[San92] A. Santos. A cooperative architecture for hypermedia editing — CoMEdiA. Computer Graphics Forum, 11(5):309–322, December 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santos:2006:RES

[San06] Luis Paulo Santos. Reports: 2006 Eurographics Symposium on Parallel Graphics and Visualization. Computer Graphics Forum, 25(4):838, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sara:2007:ITW

[Šár07] Radim Šára. Invited talks: What can computer graphics expect from 3D computer vision? Computer Graphics Forum, 26(3):xix, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sastry:1992:UIM

[Sas92] L. Sastry. User interface management systems for engineering applications. Computer Graphics Forum, 11(2):113–129, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sawyer:2007:ITS

[Saw07] Ben Sawyer. Invited talks: Serious games: Broadening games impact beyond entertainment. Computer Graphics Forum, 26 (3):xviii, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Streit:1998:IDH

[SB98] L. Streit and J. Buchanan. Importance driven halftoning. Computer Graphics Forum, 17(3):??, September 1998. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=268&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Samavati:1999:MCS

[SB99a] Faramarz F. Samavati and Richard H. Bartels. Multiresolution curve and surface representation: Reversing subdivision rules by least-squares data fitting. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=361&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Sousa:1999:CGG

[SB99b] Mario Costa Sousa and John W. Buchanan. Computer-generated graphite pencil rendering of 3D polygonal models. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=340&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Sousa:2000:OMG

[SB00] Mario Costa Sousa and John W. Buchanan. Observational models of graphite pencil materials. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=386&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Sbert:1993:IGB

[Sbe93] M. Sbert. An integral geometry based method for fast form-factor computation. Computer Graphics Forum, 12(3):C409–C420, ????

1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sbert:1997:OSS

[Sbe97] Mateu Sbert. Optimal source selection in shooting random walk Monte Carlo radiosity. Computer Graphics Forum, 16(3): C301–C308, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=167&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Suykens:2003:IRB

[SBLD03] Frank Suykens, Karl Berge, Ares Lagae, and Philip Dutré. Interactive rendering with bidirectional texture functions. Computer Graphics Forum, 22(3):463–472, September 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Staneker:2006:ODS

[SBW06] Dirk Staneker, Dirk Bartz, and Straßer Wolfgang. Occlusiondriven scene sorting for efficient culling. Computer Graphics Forum, 25(4):699–708, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schoenhut:1984:ICD

[SC84] J. Schoenhut and S. Contin. Implementation and comparison of different presentation layer codings for efficient transfer of GKS metafiles. Computer Graphics Forum, 3(1):71–78, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stolte:1995:DRT

[SC95] N. Stolte and R. Caubet. Discrete ray-tracing of huge voxel spaces. Computer Graphics Forum, 14(3):C/383–C/394, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seidel:1997:E

[SC97] Hans-Peter Seidel and Sabine Coquillart. Editorial. Computer Graphics Forum, 16(1):??, March 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shesh:2004:SMS

[SC04] Amit Shesh and Baoquan Chen. Shape modelling: SMARTPA-PER: an interactive and user friendly sketching system. Computer Graphics Forum, 23(3):301–310, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shesh:2008:PPF

[SC08a] Amit Shesh and Baoquan Chen. Peek-in-the-pic: Flying through architectural scenes from a single image. Computer Graphics Forum, 27(8):2143–2153, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shesh:2008:RSE

[SC08b] Amit Shesh and Baoquan Chen. Rendering and simplification: Efficient and dynamic simplification of line drawings. Computer Graphics Forum, 27(2):537–545, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

S2C:2004:RSA

[SCA04] SCA 2003 Chairs. Reports: Second Annual Symposium on Computer Animation (SCA 2003): San Diego, California, 26–27 July 2003. Computer Graphics Forum, 23(1):115–116, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scheidegger:2005:PCS

[SCD05] Carlos E. Scheidegger, João L. D. Comba, and Rudnei D. Da Cunha. Practical CFD simulations on programmable graphics hardware using SMAC. Computer Graphics Forum, 24(4):715– 728, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schoenhut:1984:MCG

[Sch84] J. Schoenhut. More on computer graphics metafile coding efficiency. Computer Graphics Forum, 3(4):287–288, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scheller:1985:RMW

[Sch85] A. Scheller. Report on the MICROGRAPHICS workshop. Computer Graphics Forum, 4(1):61–62, January 1985. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneider:1988:POO

[Sch88] Bengt-Olaf Schneider. A processor for an object-oriented rendering system. Computer Graphics Forum, 7(4):301–310, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schettini:1994:DSR

[Sch94a] R. Schettini. Deriving spectral reflectance functions of computersimulated object colours. Computer Graphics Forum, 13(4):211– 217, October 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlick:1994:IBM

[Sch94b] C. Schlick. An inexpensive BRDF model for physically-based rendering. Computer Graphics Forum, 13(3):C/233–C/246, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlick:1994:SSR

[Sch94c] C. Schlick. A survey of shading and reflectance models. Computer Graphics Forum, 13(2):121–131, June 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneider:1998:SEW

[Sch98] Bengt-Olaf Schneider. Siggraph/Eurographics Workshop on Graphics Hardware. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=227&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=1&amp; src=ard&amp; vid=17.

Schneider:2000:HAS

[Sch00] Uwe Schneider. A hybrid approach for stroke-based letterform composition including outline-based methods. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schilling:2001:AEM

[Sch01] Andreas Schilling. Antialiasing of environment maps. Computer Graphics Forum, 20(1):??, March 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2009:VRB

[SCM^{+}09] Yuyan Song, Wei Chen, Ross Maciejewski, Kelly P. Gaither, and David S. Ebertt. Volume rendering: Bivariate transfer functions on unstructured grids. Computer Graphics Forum, 28(3):783– 790, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scopigno:2002:STC

[Sco02] Roberto Scopigno. 3D scanning technology: Capabilities and issues. Computer Graphics Forum, 21(3):??, September 2002. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schleich:1994:BWD

[SD94a] R. Schleich and M. J. Dürst. Beyond WYSIWYG: Display of hidden information in graphics editors. Computer Graphics Forum, 13(3):C/185–C/194, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simiakakis:1994:FDA

[SD94b] G. Simiakakis and A. M. Day. Five-dimensional adaptive subdivision for ray tracing. Computer Graphics Forum, 13(2):133–140, June 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:2000:BVE

[SD00] Shamus P. Smith and David J. Duke. Binding virtual environments to toolkit capabilities. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=400&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Sen:2009:CPI

[SD09] Pradeep Sen and Soheil Darabi. Computational photography and image manipulation: Compressive dual photography. Computer Graphics Forum, 28(2):609–618, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sillion:1997:EIM

[SDB97] François Sillion, George Drettakis, and Benoit Bodelet. Efficient impostor manipulation for real-time visualization of urban scenery. Computer Graphics Forum, 16(3):C207–C218, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=158&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Soetebier:1999:SID

[SDB99] Ingo Soetebier, Ralf Dörner, and Norbert Braun. Seamless integration of databases in VR for constructing virtual environments. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=353&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Sykora:2009:JMC

[SDC09] Daniel Sýkora, John Dingliana, and Steven Collins. The joy of math and cartoon drawing: LazyBrush: Flexible painting tool for hand-drawn cartoons. Computer Graphics Forum, 28(2):599– 608, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slater:1992:SCT

[SDD^{+}92] M. Slater, K. Drake, A. Davison, E. Kordakis, A. Billyard, and E. Miranda. A statistical comparison of two hidden surface techniques: the scan-line and z-buffer algorithms. Computer Graphics Forum, 11(2):131–138, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Savransky:1999:MRE

[SDG99] Guillermo Savransky, Dan Dimerman, and Craig Gotsman. Modeling and rendering Escher-like impossible scenes. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=367&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Schnabel:2009:MRC

[SDK09] Ruwen Schnabel, Patrick Degener, and Reinhard Klein. Model reconstruction: Completion and reconstruction with primitive shapes. Computer Graphics Forum, 28(2):503–512, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:1999:HWV

[SDM99] Shamus Smith, David Duke, and Mieke Massink. The hybrid world of virtual environments. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=350&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Surazhsky:2002:ASR

[SE02a] Tatiana Surazhsky and Gershon Elber. Artistic surface rendering using layout of text. Computer Graphics Forum, 21(2):99–110, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See erratum [SE02b].

Surazhsky:2002:EAS

[SE02b] Tatiana Surazhsky and Gershon Elber. Erratum: Artistic surface rendering using layout of text. Computer Graphics Forum, 21(4): 801–802, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See [SE02a].

Schein:2004:PDO

[SE04] S. Schein and G. Elber. Placement of deformable objects. Computer Graphics Forum, 23(4):727–739, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sintorn:2008:SSL

[SEA08] Erik Sintorn, Elmar Eisemann, and Ulf Assarsson. Soft shadows and light fields: Sample based visibility for soft shadows using alias-free shadow maps. Computer Graphics Forum, 27(4):1285– 1292, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scholz:2009:EOB

[SEASM09] Volker Scholz, Sascha El-Abed, Hans-Peter Seidel, and Marcus Magnor. Editing object behaviour in video sequences. Computer Graphics Forum, 28(6):1632–1643, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seidel:1988:RIC

[Sei88] H.-P. Seidel. Report on the International Conference “Theory and Practice of Geometric Modelling”. Computer Graphics Forum, 7 (4):363, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seidel:1994:SCG

[Sei94] H.-P Seidel. Splines in computer graphics and CAD/CIM. Computer Graphics Forum, 13(3):C/514–??, ???? 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strothotte:1983:RDR

[SF83] T. Strothotte and B. Funt. Raster display of a rotating object using parallel processing. Computer Graphics Forum, 2(4):209– 217, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shinya:1992:SMM

[SF92] M. Shinya and A. Fournier. Stochastic motion-motion under the influence of wind. Computer Graphics Forum, 11(3):C119–C128, C469, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Streit:2005:MNM

[SFS05] L. Streit, P. Federl, and M. C. Sousa. Modelling nature: Modelling plant variation through growth. Computer Graphics Forum, 24 (3):497–506, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sousa:2003:NPRa

[SFWS03a] Mario Costa Sousa, Kevin Foster, Brian Wyvill, and Faramarz Samavati. Non-photorealistic rendering: Precise ink drawing of 3D models. Computer Graphics Forum, 22(3):369–379, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sousa:2003:PID

[SFWS03b] Mario Costa Sousa, Kevin Foster, Brian Wyvill, and Faramarz Samavati. Precise ink drawing of 3D models. Computer Graphics Forum, 22(3):369–379, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmalstieg:1996:DDG

[SG96a] Dieter Schmalstieg and Michael Gervautz. Demand-driven geometry transmission for distributed virtual environments. Computer Graphics Forum, 15(3):C421–C432, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sudarsky:1996:OSV

[SG96b] Oded Sudarsky and Craig Gotsman. Output-sensitive visibility algorithms for dynamic scenes with applications to virtual reality. Computer Graphics Forum, 15(3):C249–C258, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szalavari:1997:PIP

[SG97] Zsolt Szalavári and Michael Gervautz. The personal interaction panel — a two-handed interface for augmented reality. Computer Graphics Forum, 16(3):C335–C346, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=171&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Southern:2003:CCR

[SG03] Richard Southern and James Gain. Creation and control of realtime continuous level of detail on programmable graphics hardware. Computer Graphics Forum, 22(1):35–48, March 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strnad:2004:MTH

[SG04] D. Strnad and N. Guid. Modeling trees with hypertextures. Computer Graphics Forum, 23(2):173–187, June 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinbrecher:2008:RDS

[SG08] Tillmann Steinbrecher and Maik Gerth. Registering deforming surfaces I: Dental inlay and onlay construction by iterative Laplacian surface editing. Computer Graphics Forum, 27(5):1441–1447, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scopigno:2009:E

[SG09] Roberto Scopigno and Eduard Gröller. Editorial. Computer Graphics Forum, 28(1):i, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sung:2004:NCB

[SGC04] Mankyu Sung, Michael Gleicher, and Stephen Chenney. Navigation and crowd behaviour in virtual environments: Scalable behaviors for crowd simulation. Computer Graphics Forum, 23(3): 519–528, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schroder:1994:WPR

[SGCH94] P. Schröder, S. J. Gortler, M. F. Cohen, and P. Hanrahan. Wavelet projections for radiosity. Computer Graphics Forum, 13(2):141– 151, June 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sutcliffe:1993:ACS

[SGM^{+}93] D. C. Sutcliffe, J. R. Gallop, R. Maybury, R. A. Day, D. A. Duce, and G. J. Reynolds. The ARGOSI classification scheme for graphics and networking applications. Computer Graphics Forum, 12 (2):123–130, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlechtweg:2005:RMA

[SGS05] Stefan Schlechtweg, Tobias Germer, and Thomas Strothotte. RenderBots — multi-agent systems for direct image generation. Computer Graphics Forum, 24(2):137–148, June 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Streit:2002:BPF

[SH02] Lisa Streit and Wolfgang Heidrich. A biologically-parameterized feather model. Computer Graphics Forum, 21(3):665–674, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shaw:1997:HRD

[Sha97] Erin Shaw. Hierarchical radiosity for dynamic environments. Computer Graphics Forum, 16(2):107–118, June 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=127&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=2&amp; src=
ard&amp; vid=16.

Shamir:2008:SMS

[Sha08] Ariel Shamir. A survey on mesh segmentation techniques. Computer Graphics Forum, 27(6):1539–1556, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sauvage:2007:SMD

[SHB07] Basile Sauvage, Stefanie Hahmann, and Georges-Pierre Bonneau. Skinning and mesh deformation: Volume preservation of multiresolution meshes. Computer Graphics Forum, 26(3):275–283, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sherstyuk:1999:FRT

[She99] Andrei Sherstyuk. Fast ray tracing of implicit surfaces. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=364&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Scheib:2002:EFR

[SHLS02] Vincent Scheib, Jörg Haber, Ming C. Lin, and Hans-Peter Seidel. Efficient fitting and rendering of large scattered data sets using subdivision surfaces. Computer Graphics Forum, 21(3):353–362, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinke:2008:PTM

[SHPS08] Florian Steinke, Matthias Hein, Jan Peters, and Bernhard Schölkopf. Parameterization and transformations: Manifoldvalued thin-plate splines with applications in computer graphics. Computer Graphics Forum, 27(2):437–448, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schirmacher:1999:AAL

[SHS99] Hartmut Schirmacher, Wolfgang Heidrich, and Hans-Peter Seidel. Adaptive acquisition of lumigraphs from synthetic scenes. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=336&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Smyk:2005:RIT

[SiK– DM05] Miłosław Smyk, Shin ichi Kinuwaki, Roman – Durikovič, and Karol Myszkowski. Rendering II: Temporally coherent irradiance caching for high quality animation rendering. Computer Graphics Forum, 24(3):401–412, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sillion:1997:MPM

[Sil97] François X. Sillion. A million polygons, a million pixels: Which is heavier? Computer Graphics Forum, 16(3):C370– ??, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=175&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Segovia:2007:GIM

[SIP07] B. Segovia, J. C. Iehl, and B. Péroche. Global illumination (1): Metropolis instant radiosity. Computer Graphics Forum, 26(3): 425–434, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2009:CSB

[SJ09a] B. Spencer and M. W. Jones. Caustics and scattering: Into the blue: Better caustics through photon relaxation. Computer Graphics Forum, 28(2):319–328, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2009:RCC

[SJ09b] Ben Spencer and Mark W. Jones. Report: CGForum 2009 cover image. Computer Graphics Forum, 28(1):172, March 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sisneros:2008:MVC

[SJH08] Robert Sisneros, C. Ryan Johnson, and Jian Huang. Multivariate visualization: Concurrent viewing of multiple attribute-specific subspaces. Computer Graphics Forum, 27(3):783–790, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schumann:1986:MDT

[SK86] H. Schumann and A. Kotzauer. A method of displaying transformed picture rectangles using GKS raster functions. Computer Graphics Forum, 5(2):119–123, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:1999:SIN

[SK99] László Szirmay-Kalos. Stochastic iteration for non-diffuse global illumination. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=344&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=18.

Skala:1987:IMB

[Ska87] Vaclav Skala. An intersecting modification to the Bresenham algorithm for hidden-line solution. Computer Graphics Forum, 6(4): 343–347, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Skala:1996:EAL

[Ska96] Vaclav Skala. An efficient algorithm for line clipping by convex and non-convex polyhedra in E^{3}. Computer Graphics Forum, 15(1):61–68, March 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2005:GHA

[SKALP05] László Szirmay-Kalos, Barnabás Aszódi, István Lazányi, and Mátyás Premecz. GPU & hardware: Approximate ray-tracing on the GPU with distance impostors. Computer Graphics Forum, 24(3):695–704, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2001:CFP

[SKC01] Ming-Shing Su, Ming-Tat Ko, and Kuo-Young Cheng. Control of feature-point-driven facial animation using a hypothetical face. Computer Graphics Forum, 20(4):??, December 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2001:GIC

[SKCA01] László Szirmay-Kalos, Ferenc Csonka, and György Antal. Global illumination as a combination of continuous random walk and finite-element based iteration. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:1997:AQM

[SKFNC97] László Szirmay-Kalos, Tibor Fóris, László Neumann, and Balázs Csébfalvi. An analysis of quasi-Monte Carlo integration applied to the transillumination radiosity method. Computer Graphics Forum, 16(3):C271–C281, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0167-7055&iid=3&src=ard&vid=16.

Streit:2008:GVN

[SKKS08] M. Streit, M. Kalkusch, K. Kashofer, and D. Schmalstieg. Graph visualization: Navigation and exploration of interconnected pathways. Computer Graphics Forum, 27(3):951–958, May 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:2006:ITM

[SKMS06] Kaleigh Smith, Grzegorz Krawczyk, Karol Myszkowski, and Hans-Peter Seidel. Images and tone mapping: Beyond tone mapping: Enhanced depiction of tone mapped HDR images. Computer Graphics Forum, 25(3):427–438, September 2006. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2009:RDI

[SKS09] László Szirmay-Kalos and László Szécsi. Rendering I: Deterministic importance sampling with error diffusion. Computer Graphics Forum, 28(4):1055–1064, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlattmann:2007:FMC

[SKSK07] Markus Schlattmann, Ferenc Kahlesz, Ralf Sarlette, and Reinhard Klein. Fluid and motion capture: Markerless 4 gestures 6 DOF real-time visual tracking of the human hand with automatic initialization. Computer Graphics Forum, 26(3):467–476, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2008:DMG

[SKU08] László Szirmay-Kalos and Tamás Umenhoffer. Displacement mapping on the GPU — state of the art. Computer Graphics Forum, 27(6):1567–1592, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2009:SEG

[SKUP^{+}09] László Szirmay-Kalos, Tamás Umenhoffer, Gustavo Patow, László Szécsi, and Mateu Sbert. Specular effects on the GPU: State of the art. Computer Graphics Forum, 28(6):1586–1617, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:1989:IDS

[SL89] F. J. Smith and S. Leitch. The incremental display of a singlevalued curve. Computer Graphics Forum, 8(2):159–165, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shacked:2001:ALD

[SL01] Ram Shacked and Dani Lischinski. Automatic lighting design using a perceptual quality metric. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stam:2003:QTS

[SL03] Jos Stam and Charles Loop. Quad/triangle subdivision. Computer Graphics Forum, 22(1):79–85, March 2003. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Servin:2007:MCR

[SL07] M. Servin and C. Lacoursière. Massless cable for real-time simulation. Computer Graphics Forum, 26(2):172–184, June 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shor:2008:IWS

[SL08] Yael Shor and Dani Lischinski. Imaging and warping: The shadow meets the mask: Pyramid-based shadow removal. Computer Graphics Forum, 27(2):577–586, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shin:2009:AES

[SL09] Hyun Joon Shin and Yunjin Lee. Animation: Expression synthesis and transfer in parameter spaces. Computer Graphics Forum, 28 (7):1829–1835, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slater:1984:GP

[Sla84] M. Slater. GKS in Pascal. Computer Graphics Forum, 3(4):259– 267, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slater:1988:TMI

[Sla88] Mel Slater. A top down method for interactive drawing. Computer Graphics Forum, 7(4):323–329, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stich:2008:MIV

[SLAM08] Timo Stich, Christian Linz, Georgia Albuquerque, and Marcus Magnor. Moving images: View and time interpolation in image space. Computer Graphics Forum, 27(7):1781–1787, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2009:ESS

[SLCZ09] Benjamin Spencer, Robert S. Laramee, Guoning Chen, and Eugene Zhang. Evenly spaced streamlines for surfaces: an imagebased approach. Computer Graphics Forum, 28(6):1618–1631, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shi:2008:SSI

[SLLW08] Zhuo Shi, Shujin Lin, Xiaonan Luo, and Renhong Wang. Subdivision and segmentation: Interpolatory and mixed loop schemes. Computer Graphics Forum, 27(7):1829–1835, October 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schoner:2004:SHS

[SLS04] Jeffrey L. Schoner, Jochen Lang, and Hans-Peter Seidel. Simulation with haptics and sound: Measurement-based interactive simulation of viscoelastic solids. Computer Graphics Forum, 23(3): 547–556, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharf:2006:GRC

[SLS^{+}06] Andrei Sharf, Thomas Lewiner, Ariel Shamir, Leif Kobbelt, and Daniel Cohen-Or. Geometry reconstruction: Competing fronts for coarse–to–fine surface reconstruction. Computer Graphics Forum, 25(3):389–398, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharf:2007:MSF

[SLSK07] Andrei Sharf, Thomas Lewiner, Ariel Shamir, and Leif Kobbelt. Modeling and segmentation: On-the-fly curve-skeleton computation for 3D shapes. Computer Graphics Forum, 26(3):323–328, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:2008:PRA

[SLTM08] Kaleigh Smith, Pierre-Edouard Landes, Joëlle Thollot, and Karol Myszkowski. Perception for rendering and animation: Apparent greyscale: a simple and fast conversion to perceptually accurate images and video. Computer Graphics Forum, 27(2):193–200, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Straayer:1984:CGM

[SM84] D. Straayer and J. Maynard. Computer graphics, metafile coding efficiency. Computer Graphics Forum, 3(4):275–286, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shinde:1986:AHF

[SM86] Yogesh N. Shinde and S. P. Mudur. Algorithms for handling the fill area primitive of GKS. Computer Graphics Forum, 5(2):105– 117, June 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Samavati:2002:MSH

[SMAB02] Faramarz F. Samavati, Nezam Mahdavi-Amiri, and Richard H. Bartels. Multiresolution surfaces having arbitrary topologies by a reverse Doo subdivision method. Computer Graphics Forum, 21(2):121–136, June 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:1985:CA

[Smi85] B. Reffin Smith. Computers and art. Computer Graphics Forum, 4(2):141–147, June 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schirmacher:2001:FPG

[SMS01] Hartmut Schirmacher, Li Ming, and Hans-Peter Seidel. On-the-fly processing of generalized lumigraphs. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soulie:2007:MRH

[SMTG07] Romain Soulié, Stéphane Mérillou, Olivier Terraz, and Djamchid Ghazanfarpour. Modeling and rendering of heterogeneous granular materials: Granite application. Computer Graphics Forum, 26(1):66–79, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szilvasi-Nagy:1984:ADI

[SN84] M. Szilvási-Nagy. An algorithm for determining the intersection of two simple polyhedra. Computer Graphics Forum, 3(3):219–225, September 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szilvasi-Nagy:1986:TAD

[SN86] M. Szilvasi-Nagy. Two algorithms for decomposing a polyhedron into convex parts. Computer Graphics Forum, 5(3):197–201, September 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Snydre:2008:SSL

[SN08] John Snydre and Derek Nowrouzezahrai. Soft shadows and light fields: Fast soft self-shadowing on dynamic height fields. Computer Graphics Forum, 27(4):1275–1283, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Snibbe:1995:DMI

[Sni95] S. Sona Snibbe. A direct manipulation interface for 3D computer animation. Computer Graphics Forum, 14(3):C/271–C/283, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shiratori:2006:ADM

[SNI06] Takaaki Shiratori, Atsushi Nakazawa, and Katsushi Ikeuchi. Animation: Dancing-to-music character animation. Computer Graphics Forum, 25(3):449–458, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simari:2009:MNS

[SNKS09] P. Simari, D. Nowrouzezahrai, E. Kalogerakis, and K. Singh. Meshing — non-simplex: Multi-objective shape segmentation and labeling. Computer Graphics Forum, 28(5):1415–1425, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sips:2009:MDS

[SNLH09] Mike Sips, Boris Neubert, John P. Lewis, and Pat Hanrahan. Multidimensional data: Selecting good views of high-dimensional data using class consistency. Computer Graphics Forum, 28(3): 831–838, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Staadt:2001:JPS

[SNLW01] Oliver G. Staadt, Martin Näf, Edouard Lamboray, and Stephan Würmlin. JAPE: a prototyping system for collaborative virtual environments. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2009:SAC

[SOG09] Jian Sun, Maks Ovsjanikov, and Leonidas Guibas. Shape analysis: a concise and provably informative multi-scale signature based on heat diffusion. Computer Graphics Forum, 28(5):1383–1392, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sumner:1999:ASM

[SOH99] Robert W. Sumner, James F. O’Brien, and Jessica K. Hodgins. Animating sand, mud, and snow. Computer Graphics Forum, 18(1):??, March 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=299&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=18.

Sud:2004:DFC

[SOM04] Avneesh Sud, Miguel A. Otaduy, and Dinesh Manocha. Distance fields and collision detection: Difi: Fast 3D distance field computation using graphics hardware. Computer Graphics Forum, 23 (3):557–566, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sorkine:2006:SAR

[Sor06] Olga Sorkine. State of the art reviews: Differential representations for mesh processing. Computer Graphics Forum, 25(4):789–807, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sadarjoen:1997:DST

[SP97] I. Ari Sadarjoen and Frits H. Post. Deformable surface techniques for field visualization. Computer Graphics Forum, 16(3): C109–C116, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=147&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Serpaggi:2001:AMI

[SP01] Xavier Serpaggi and Bernard Péroche. An adaptive method for indirect illumination using light vectors. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sousa:2003:FGL

[SP03a] Mario Costa Sousa and Przemysław Prusinkiewicz. A few good lines: Suggestive drawing of 3D models. Computer Graphics Forum, 22(3):381–390, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sousa:2003:NPRb

[SP03b] Mario Costa Sousa and Przemysław Prusinkiewicz. Nonphotorealistic rendering: a few good lines: Suggestive drawing of 3D models. Computer Graphics Forum, 22(3):381–390, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sheffer:2006:RSG

[SP06] Alla Sheffer and Konrad Polthier. Reports: Symposium on Geometry Processing 2006: Cagliari, Italy, June 26–28 2006. Computer Graphics Forum, 25(4):839, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sparks:1985:GLB

[Spa85] Madeleine Sparks. Graphics language bindings — the what and why. Computer Graphics Forum, 4(4):387–392, December 1985.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sabha:2007:TSU

[SPD07] M. Sabha, P. Peers, and P. Dutré. Texture synthesis using exact neighborhood matching. Computer Graphics Forum, 26(2):131– 142, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Speer:1991:CIG

[Spe91] L. Richard Speer. A cross-indexed guide to the ray tracing literature. Computer Graphics Forum, 10(2):145–174, June 1991. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schott:2009:VRH

[SPH^{+}09] Mathias Schott, Vincent Pegoraro, Charles Hansen, Kévin Boulanger, and Kadi Bouatouch. Volume rendering & hardware acceleration: a directional occlusion shading model for interactive direct volume rendering. Computer Graphics Forum, 28(3): 855–862, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Savchenko:1995:FRS

[SPOK95] V. V. Savchenko, A. A. Pasko, O. G. Okunev, and T. L. Kunii. Function representation of solids reconstructed from scattered surface points and contours. Computer Graphics Forum, 14 (4):181–188, October 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strothotte:1994:HRF

[SPR^{+}94] T. Strothotte, B. Preim, A. Raab, J. Schumann, and D. R. Forsey. How to render frames and influence people. Computer Graphics Forum, 13(3):C/455–C/466, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slusallek:1994:IRP

[SPS94] Ph. Slusallek, Th. Pflaum, and H.-P. Seidel. Implementing RenderMan — practice, problems and enhancements. Computer Graphics Forum, 13(3):C/443–C/454, ???? 1994. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slusallek:1995:UPR

[SPS95] Ph. Slusallek, Th. Pflaum, and H.-P. Seidel. Using procedural RenderMan shaders for global illumination. Computer Graphics Forum, 14(3):C/311–C/324, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shamir:1996:ETE

[SR96] Ariel Shamir and Ari Rappoport. Extraction of typographic elements from outline representations of fonts. Computer Graphics Forum, 15(3):C259–C268, C473–C474, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Skrba:2009:AQM

[SRH^{+}09] Ljiljana Skrba, Lionel Reveret, Franck Hétroy, Marie-Paule Cani, and Carol O’Sullivan. Animating quadrupeds: Methods and applications. Computer Graphics Forum, 28(6):1541–1560, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seebold:1991:SWD

[SS91] Patrice Séébold and Karine Slowinski. The shortest way to draw a connected picture. Computer Graphics Forum, 10(4):319–327, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steed:1995:IDB

[SS95] Anthony Steed and Mel Slater. 3D interaction with the desktop bat. Computer Graphics Forum, 14(2):97–104, June 1995. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schaufler:1996:TDI

[SS96a] Germot Schaufler and Wolfgang Sturzlinger. A three-dimensional image cache for virtual reality. Computer Graphics Forum, 15(3): C227–C235, C471–C472, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smets-Solanes:1996:VFB

[SS96b] Jean-Paul P. Smets-Solanes. Vector field based texture mapping of animated implicit objects. Computer Graphics Forum, 15(3): C289–C300, C475–C476, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sadagic:2000:DPV

[SS00] Amela Sadagic and Mel Slater. Dynamic polygon visibility ordering for head-slaved viewing in virtual environments. Computer Graphics Forum, 19(2):??, June 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schwarz:2007:SVB

[SS07] Michael Schwarz and Marc Stamminger. Shadows and visibility: Bitmask soft shadows. Computer Graphics Forum, 26(3):515–524, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2008:ISBb

[SS08] Ryan Schmidt and Karan Singh. Interactive and sketch-based techniques: Sketch-based procedural surface modeling and compositing using surface trees. Computer Graphics Forum, 27(2): 321–330, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schwarz:2009:GFG

[SS09] Michael Schwarz and Marc Stamminger. GPU: Fast GPU-based adaptive tessellation with CUDA. Computer Graphics Forum, 28 (2):365–374, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schultz:2008:VNV

[SSA^{+}08] T. Schultz, N. Sauber, A. Anwander, H. Theisel, and H.-P. Seidel. Visualization for neurology: Virtual klingler dissection: Putting fibers into context. Computer Graphics Forum, 27(3):1063–1070, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinke:2005:GSV

[SSB05] Florian Steinke, Bernhard Schölkopf, and Volker Blanz. Geometry I: Support vector machines for 3D shape processing. Computer Graphics Forum, 24(3):285–294, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2008:ISBa

[SSB08] Ryan Schmidt, Karan Singh, and Ravin Balakrishnan. Interactive and sketch-based techniques: Sketching and composing widgets for 3D manipulation. Computer Graphics Forum, 27(2):301–310, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shapira:2009:CPI

[SSCO09] L. Shapira, A. Shamir, and D. Cohen-Or. Computational photography and image manipulation: Image appearance exploration by model-based navigation. Computer Graphics Forum, 28(2):629– 638, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schreiner:2006:MPD

[SSFS06] John Schreiner, Carlos E. Scheidegger, Shachar Fleishman, and Cláudio T. Silva. Mesh processing: Direct (re)meshing for efficient surface processing. Computer Graphics Forum, 25(3):527–536, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stamminger:2000:EGG

[SSG^{+}00] Marc Stamminger, Annette Scheel, Xavier Granier, Frederic Perez-Cazorla, George Drettakis, and François Sillion. Efficient glossy global illumination with interactive viewing. Computer Graphics Forum, 19(1):??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=385&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Saka:1993:PSF

[SSK93] G. Saka, F. Schroder, and H.-J. Koppert. Pseudo-satellite film using fractal clouds to enhance animated weather forecasting. Computer Graphics Forum, 12(3):C329–C338, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scholz:2005:IVP

[SSK^{+}05] Volker Scholz, Timo Stich, Michael Keckeisen, Markus Wacker, and Marcus Magnor. Image and video processing: Garment motion capture using color-coded patterns. Computer Graphics Forum, 24(3):439–447, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shevtsov:2007:RTH

[SSK07] Maxim Shevtsov, Alexei Soupikov, and Alexander Kapustin. Ray tracing: Highly parallel fast KD-tree construction for interactive ray tracing of dynamic scenes. Computer Graphics Forum, 26(3): 395–404, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schjoth:2008:DBP

[SSO08] L. Schjøth, J. Sporring, and O. Fogh Olsen. Diffusion based photon mapping. Computer Graphics Forum, 27(8):2114–2127, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stamminger:1997:BRI

[SSS97] Marc Stamminger, Philipp Slusallek, and Hans-Peter Seidel. Bounded radiosity — illumination on general surfaces and clusters. Computer Graphics Forum, 16(3):C309–C317, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=168&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Scheel:2000:TRI

[SSS00] A. Scheel, M. Stamminger, and H.-P. Seidel. Tone reproduction for interactive walkthroughs. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=422&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Scheel:2002:GBF

[SSS02] Annette Scheel, Marc Stamminger, and Hans-Peter Seidel. Grid based final gather for radiosity on complex clustered scenes. Computer Graphics Forum, 21(3):547–556, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherbaum:2007:FPI

[SSSB07] K. Scherbaum, M. Sunkel, H.-P. Seidel, and V. Blanz. Faces: Prediction of individual non-linear aging trajectories of faces. Computer Graphics Forum, 26(3):285–294, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sbert:2004:RTI

[SSSK04a] Mateu Sbert, László Szécsi, and László Szirmay-Kalos. Real-time illumination and shadows: Real-time light animation. Computer Graphics Forum, 23(3):291–299, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szecsi:2004:SGI

[SSSK04b] László Szécsi, Mateu Sbert, and László Szirmay-Kalos. Sampling for global illumination: Combined correlated and importance sampling in direct light source computation and environment mapping. Computer Graphics Forum, 23(3):585–593, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stamminger:1998:GRL

[SSSS98] M. Stamminger, H. Schirmacher, Ph. Slusallek, and H.-P. Seidel. Getting rid of links in hierarchical radiosity. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=264&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Schouten:1988:DCR

[St88] H. J. Schouten and P. J. W. ten Hagen. Dialogue cell resource model and basic dialogue cells. Computer Graphics Forum, 7(4): 311–322, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santos:1993:UMS

[ST93] A. Santos and B. Tritsch. Using multimedia to support cooperative editing. Computer Graphics Forum, 12(3):C73–C87, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sanso:1994:HCA

[ST94] R. M. Sanso and D. Thalmann. A hand control and automatic grasping system for synthetic actors. Computer Graphics Forum, 13(3):C/167–C/177, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spillmann:2008:FSA

[ST08] Jonas Spillmann and Matthias Teschner. Fluids and simulation: an adaptive contact model for the robust simulation of knots. Computer Graphics Forum, 27(2):497–506, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stam:1997:SDS

[Sta97] Jos Stam. Stochastic dynamics: Simulating the effects of turbulence on flexible structures. Computer Graphics Forum, 16(3): C159–C164, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=152&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Stam:2006:IPT

[Sta06] Jos Stam. Invited presentations: Towards a unified dynamical solver for computer graphics. Computer Graphics Forum, 25(3): xix, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Swindells:2009:USI

[STD09] Colin Swindells, Melanie Tory, and Rebecca Dreezer. User studies and interaction: Comparing parameter manipulation with mouse, pen, and slider user interfaces. Computer Graphics Forum, 28(3): 919–926, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinhart:1984:PGO

[Ste84] Jonathan E. Steinhart. Proposal for GKS output level 3 segment hierarchy and editing. Computer Graphics Forum, 3(4):289–297, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stenger:1985:HNH

[Ste85] Ludwig Stenger. HDTV — a new high definition television system. Computer Graphics Forum, 4(2):117–125, June 1985. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shlafman:2002:MPS

[STK02] Shymon Shlafman, Ayellet Tal, and Sagi Katz. Metamorphosis of polyhedral surfaces using decomposition. Computer Graphics Forum, 21(3):219–228, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmitz:2008:IIG

[STK08] Arne Schmitz, Markus Tavenrath, and Leif Kobbelt. Illumination: Interactive global illumination for deformable geometry in CUDA.

Computer Graphics Forum, 27(7):1979–1986, October 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seron:1993:QCI

[STM93] F. J. Seron, J. J. Torrens, and J. A. Magallon. Quality control of an interpolation method for discontinuous parametric surfaces. Computer Graphics Forum, 12(3):C461–C471, ???? 1993. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strasser:1982:GS

[Str82] W. Strasser. Graphics systems. Computer Graphics Forum, 1(1): 17–19, March 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Straayer:1983:GSP

[Str83] D. H. Straayer. Graphics standards: The pace quickens. Computer Graphics Forum, 2(1):67–75, March 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strubbe:1984:RMS

[Str84] H. J. Strubbe. Role, model, structure and construction of a UIMS — working group report. Computer Graphics Forum, 3(2):171– 174, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sugihara:1994:RCA

[Sug94] K. Sugihara. A robust and consistent algorithm for intersecting convex polyhedra. Computer Graphics Forum, 13(3):C/45–C/54, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sung:1992:ASB

[Sun92] K. Sung. Area sampling buffer: tracing rays with Z-buffer hardware. Computer Graphics Forum, 11(3):C299–C310, C480–C481, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Suzjevic:1989:RTI

[Suz89] V. Suzjevic. Report of the Third International Conference on Computer Graphics, Yugoslavia. Computer Graphics Forum, 8 (1):60, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schafhitzel:2008:FFE

[SVG^{+}08] T. Schafhitzel, J. E. Vollrath, J. P. Gois, D. Weiskopf, A. Castelo, and T. Ertl. Flow feature extraction: Topology-preserving λ_{2}- based vortex core line detection for flow visualization. Computer Graphics Forum, 27(3):1023–1030, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Suykens:2003:TBI

[SvLD03] Frank Suykens, Karl Berge vom, Ares Lagae, and Philip Dutré. Textures and BRDFs: Interactive rendering with bidirectional texture functions. Computer Graphics Forum, 22(3):463–472, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scateni:1995:VSC

[SvZ95] R. Scateni, J. van Wijk, and P. Zanarini, editors. Visualization in scientific computing ’95: proceedings of the Eurographics Workshop in Chia, Italy, May 3–5, 1995. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82729-3. LCCN T385 .E89 1995.

Sweeney:1983:REZ

[SW83] A. Sweeney and H. K. Watkins. Report of EUROGRAPHICS ’83, Zagreb, Yugoslavia. Computer Graphics Forum, 2(4):200– 202, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sakas:1992:FAV

[SW92a] G. Sakas and R. Westermann. A functional approach to the visual simulation of gaseous turbulence. Computer Graphics Forum, 11(3):C107–C117, C467–C468, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Siltanen:1992:NOM

[SW92b] P. Siltanen and C. Woodward. Normal orientation methods for 3D offset curves, sweep surfaces and skinning. Computer Graphics Forum, 11(3):C449–C457, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Suykens:1999:WMM

[SW99] Frank Suykens and Yves D. Willems. Weighted multipass methods for global illumination. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=341&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Schregle:2004:PVG

[SW04a] Roland Schregle and Jan Wienold. Physical validation of global illumination methods: Measurement and error analysis. Computer Graphics Forum, 23(4):761–781, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2004:IIP

[SW04b] Dan Su and Philip Willis. Image interpolation by pixel-level datadependent triangulation. Computer Graphics Forum, 23(2):189– 201, June 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schaefer:2005:DMC

[SW05] Scott Schaefer and Joe Warren. Dual marching cubes: Primal contouring of dual grids. Computer Graphics Forum, 24(2):195– 201, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherzer:2008:RTF

[SW08a] Daniel Scherzer and Michael Wimmer. Real time: Frame sequential interpolation for discrete level-of-detail rendering. Computer Graphics Forum, 27(4):1175–1181, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2008:IST

[SW08b] Yinlong Sun and Qiqi Wang. Interference shaders of thin films. Computer Graphics Forum, 27(6):1607–1631, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sanftmann:2009:SIS

[SW09] Harald Sanftmann and Daniel Weiskopf. Scatterplots: Illuminated 3D scatterplots. Computer Graphics Forum, 28(3):751– 758, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sloan:1998:IDT

[SWB98] Peter-Pike J. Sloan, David M. Weinstein, and J. Dean Brederson. Importance driven texture coordinate optimization. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=257&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Smith:2001:FCS

[SWB01] Jeffrey Smith, Andrew Witkin, and David Baraff. Fast and controllable simulation of the shattering of brittle objects. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sussmuth:2008:RDS

[SWG08] Jochen Süßmuth, Marco Winter, and Günther Greiner. Registering deforming surfaces II: Reconstructing animated meshes from time-varying point clouds. Computer Graphics Forum, 27 (5):1469–1476, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schnabel:2007:ERP

[SWK07] R. Schnabel, R. Wahl, and R. Klein. Efficient RANSAC for pointcloud shape detection. Computer Graphics Forum, 26(2):214–226, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahner:2008:FAG

[SWPL08] J. Sahner, B. Weber, S. Prohaska, and H. Lamecker. Flow advection and geometry processing: Extraction of feature lines on surface meshes based on discrete Morse theory. Computer Graphics Forum, 27(3):735–742, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneider:2009:FVS

[SWS09] Dominic Schneider, Alexander Wiebel, and Gerik Scheuermann. Flow visualization: Smooth stream surfaces of fourth order precision. Computer Graphics Forum, 28(3):871–878, June 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scopigno:2004:RSS

[SZAB04] Roberto Scopigno, Denis Zorin, Pierre Alliez, and Jean-Daniel Boissonnat. Reports: Second Symposium on Geometry Processing: Nice, France, 8–10 June 2004. Computer Graphics Forum, 23(4):842, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:1993:DFF

[SZG93] Jizhou Sun, L. Q. Zou, and R. L. Grimsdale. The determination of form-factors by lookup table. Computer Graphics Forum, 12 (4):191–198, October 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szyszkowicz:1991:BIC

[Szy91a] Mieczysław Szyszkowicz. Block iterations in the complex plane. Computer Graphics Forum, 10(1):67–70, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szyszkowicz:1991:CAN

[Szy91b] Mieczysław Szyszkowicz. Computer art from numerical methods. Computer Graphics Forum, 10(3):255–259, September 1991. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szyszkowicz:1991:SSR

[Szy91c] Mieczysław Szyszkowicz. A survey of several root-finding methods in the complex plane. Computer Graphics Forum, 10(2):141–144, June 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Telea:2008:SVC

[TA08] Alexandru Telea and David Auber. Software visualization: Code flows: Visualizing structural evolution of source code. Computer Graphics Forum, 27(3):831–838, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tamminen:1982:HLU

[Tam82] M. Tamminen. Hidden lines using the EXCELL method. Computer Graphics Forum, 1(3):96–105, September 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tucker:1984:EPE

[TB84] Hugh A. Tucker and Ketil Bo, editors. Eurographics ’84: proceedings of the European Graphics Conference and Exhibition, Copenhagen, Denmark 12–14 September. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1984. ISBN 0-444-87617-0. LCCN T 385 E97 1984. Sponsored by the European Association for Computer Graphics and the Nordic Computer Graphics Society.

Torres:1993:GOM

[TC93] J. C. Torres and B. Clares. Graphic objects: a mathematical abstract model for computer graphics. Computer Graphics Forum, 12(5):311–327, December 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Torres:1994:FAS

[TC94] J. C. Torres and B. Clares. A formal approach to the specification of graphic object functions. Computer Graphics Forum, 13(3): C/371–C/380, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarini:2005:TPT

[TC05] Marco Tarini and Paolo Cignoni. Texture: Pinchmaps: textures with customizable discontinuities. Computer Graphics Forum, 24 (3):557–568, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarrin:2003:PES

[TCH^{+}03a] N. Tarrin, S. Coquillart, S. Hasegawa, L. Bouguila, and M. Sato. Practice and experience: The stringed haptic workbench: a new haptic workbench solution. Computer Graphics Forum, 22(3): 583–589, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarrin:2003:SHW

[TCH^{+}03b] N. Tarrin, S. Coquillart, S. Hasegawa, L. Bouguila, and M. Sato. The Stringed Haptic Workbench: a new haptic workbench solution. Computer Graphics Forum, 22(3):583–589, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tariniy:2000:RTA

[TCRS00] M. Tariniy, P. Cignoniz, C. Rocchinix, and R. Scopigno. Real time, accurate, multi-featured rendering of bump mapped surfaces. Computer Graphics Forum, 19(3):??, August 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=404&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Thomas:2000:MVC

[TD00] Gwenola Thomas and Stéphane Donikian. Modelling virtual cities dedicated to behavioural animation. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=399&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=19.

Tal:1999:IMF

[TE99] Ayellet Tal and Gershon Elber. Image morphing with feature preserving texture. Computer Graphics Forum, 18(3): ??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=354&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

tenHagen:1982:AGS

[ten82a] P. J. W. ten Hagen. The area of graphics standardization. Computer Graphics Forum, 1(1):14–16, March 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

tenHagen:1982:RGV

[ten82b] P. J. W. ten Hagen. The review of GKS version 7.0, the finishing touch. Computer Graphics Forum, 1(4):204–212, December 1982. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Terzopoulos:2002:AAH

[Ter02] Demetri Terzopoulos. Artificial animals and humans: From physics to intelligence. Computer Graphics Forum, 21(3):??, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2003:AFH

[TFK^{+}03a] Tsunemi Takahashi, Hiroko Fujii, Atsushi Kunimatsu, Kazuhiro Hiwada, Takahiro Saito, Ken Tanaka, and Heihachi Ueki. Animation of fluids and hair: Realistic animation of fluid with splash and foam. Computer Graphics Forum, 22(3):391–400, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2003:RAF

[TFK^{+}03b] Tsunemi Takahashi, Hiroko Fujii, Atsushi Kunimatsu, Kazuhiro Hiwada, Takahiro Saito, Ken Tanaka, and Heihachi Ueki. Realistic animation of fluid with splash and foam. Computer Graphics Forum, 22(3):391–400, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Turner:1998:ICA

[TG98] Russell Turner and Enrico Gobbetti. Interactive construction and animation of layered elastically deformable characters. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=234&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

Turner:1996:HTS

[TGS96] Russell Turner, Enrico Gobbetti, and Ian Soboroff. Head-tracked stereo viewing with two-handed 3D interaction for animated character construction. Computer Graphics Forum, 15(3):C197–C206, C470, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

tenHagen:1983:EPI

[tH83a] P. J. W. ten Hagen, editor. Eurographics ’83: proceedings of the international conference and exhibition, University of Zagreb, Faculty of Electrical Engineering, Zagreb, Yugoslavia, 31 August-2 September, 1983. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1983. ISBN 0-444-86770-8. LCCN T385 .E97 1983 Sci-Eng. Fourth annual conference of the European Association for Computer Graphics.

tenHagen:1983:RMI

[tH83b] P. J. W. ten Hagen. Report on the 6th meeting of ISO/TC97/ SC5/WG2 Computer Graphics. Computer Graphics Forum, 2(4):

189–191, November 1983. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

tenHagen:1984:ET

[tH84] P. J. W. ten Hagen, editor. Eurographics tutorials ’83, Eurographic seminars. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-13644-4. LCCN T385 .E98 1984 Bar.

Theisel:2002:DVF

[The02a] Holger Theisel. Designing 2D vector fields of arbitrary topology. Computer Graphics Forum, 21(3):595–604, September 2002. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theisel:2002:EIM

[The02b] Holger Theisel. Exact isosurfaces for marching cubes. Computer Graphics Forum, 21(1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tokumasu:1993:SSG

[THN93] S. Tokumasu, I. Harashima, and N. Nakajima. Spanning a C^{1}- surface over a given wireframe. Computer Graphics Forum, 12 (4):211–218, October 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thomas:1986:OOR

[Tho86] A. L. Thomas. Overlap operations and raster graphics. Computer Graphics Forum, 5(1):13–32, March 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

tenHagen:1990:ECG

[tHS90] P. ten Hagen and J. Schönhut. EUROPE’92 — the computer graphics market: Challenge and opportunity. A R&D institute’s view of a workshop held in Darmstadt in December 1989. Computer Graphics Forum, 9(4):388, December 1990. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thurmer:2001:SNV

[Thü01] Grit Thürmer. Smoothing normal vectors on discrete surfaces while preserving slope discontinuities. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:1995:AEC

[TIS^{+}95] S. Takahashi, T. Ikeda, Y. Shinagawa, T. L. Kunii, and M. Ueda. Algorithms for extracting correct critical points and constructing topological graphs from discrete geographical elevation data. Computer Graphics Forum, 14(3):C/181–C/192, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theisel:1998:ESM

[TK98] Holger Theisel and Matthias Kreuseler. An enhanced spring model for information visualization. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=280&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Teschner:2005:SAR

[TKH^{+}05] M. Teschner, S. Kimmerle, B. Heidelberger, G. Zachmann, L. Raghupathi, A. Fuhrmann, M.-P. Cani, F. Faure, N. Magnenat-Thalmann, W. Strasser, and P. Volino. State of the art reviews: Collision detection for deformable objects. Computer Graphics Forum, 24(1):61–81, March 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Teler:2001:SCS

[TL01] Eyal Teler and Dani Lischinski. Streaming of complex 3D scenes for remote walkthroughs. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tecchia:2002:VCR

[TLC02] Franco Tecchia, Celine Loscos, and Yiorgos Chrysanthou. Visualizing crowds in real-time. Computer Graphics Forum, 21(4): 753–765, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Turner:1999:MOO

[TLG99] Russell Turner, Song Li, and Enrico Gobbetti. Metis — an object-oriented toolkit for constructing virtual reality applications. Computer Graphics Forum, 18(2):??, June 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=362&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=2&amp; src=
ard&amp; vid=18.

Tasaki:2004:SAR

[TMO04] Daisuke Tasaki, Shinji Mizuno, and Minoru Okada. Stylized animation and rendering: Virtual drypoint by a model-driven strategy. Computer Graphics Forum, 23(3):431–440, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thalmann:1986:AIT

[TMT86] D. Thalmann and N. Magnenat-Thalmann. Artificial intelligence in three-dimensional computer animation. Computer Graphics Forum, 5(4):341–348, December 1986. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thomas:1989:AAR

[TNF89] D. Thomas, Arun N. Netravali, and D. S. Fox. Anti-aliased ray tracing with covers. Computer Graphics Forum, 8(4):325–336, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tadamura:1993:MSR

[TNK^{+}93] K. Tadamura, E. Nakamae, K. Kaneda, M. Baba, H. Yamashita, and T. Nishita. Modeling of skylight and rendering of outdoor scenes. Computer Graphics Forum, 12(3):C189–C200, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tanaka:1997:PLI

[TO97] Toshimitsu Tanaka and Noboru Ohnishi. Painting-like image emphasis based on human vision systems. Computer Graphics Forum, 16(3):C253–C260, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0167-7055&iid=3&src=ard&vid=16.

Takahashi:2002:MSP

[TON^{+}02] Shigeo Takahashi, Naoya Ohta, Hiroko Nakamura, Yuriko Takeshima, and Issei Fujishiro. Modeling surperspective projection of landscapes for geographical guide-map generation. Computer Graphics Forum, 21(3):259–268, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theoharis:1988:IPR

[TP88] Theoharis Theoharis and Ian Page. Incremental polygon rendering on a SIMD processor array. Computer Graphics Forum, 7(4): 331–341, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theoharis:1989:TPM

[TP89] Theoharis Theoharis and Ian Page. Two parallel methods for polygon clipping. Computer Graphics Forum, 8(2):107–114, June 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Torchelsen:2009:DSC

[TPBC09] Rafael P. Torchelsen, Francisco Pinto, Rui Bastos, and João L. D. Comba. Data structure and collision detection: Approximate on-surface distance computation using quasi-developable charts. Computer Graphics Forum, 28(7):1781–1789, October 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thomaszewski:2009:DCB

[TPS09] Bernhard Thomaszewski, Simon Pabst, and Wolfgang Straßer. Deformations: Continuum-based strain limiting. Computer Graphics Forum, 28(2):569–576, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theisel:2003:CCV

[TRS03a] H. Theisel, Ch. Rössl, and H.-P. Seidel. Compression: Compression of 2D vector fields under guaranteed topology preservation. Computer Graphics Forum, 22(3):333–342, September 2003. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Theisel:2003:CVF

[TRS03b] H. Theisel, Ch. Rössl, and H.-P. Seidel. Compression of 2D vector fields under guaranteed topology preservation. Computer Graphics Forum, 22(3):333–342, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Todt:2008:GBS

[TRSKK08] S. Todt, C. Rezk-Salama, A. Kolb, and K.-D. Kuhnert. GPUbased spherical light field rendering with per-fragment depth correction. Computer Graphics Forum, 27(8):2081–2095, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Truckenbrod:1984:CA

[Tru84] J. Truckenbrod. Computer art. Computer Graphics Forum, 3(4): 315–318, December 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tricoche:2001:TTT

[TSH01] X. Tricoche, G. Scheuermann, and H. Hagen. Tensor topology tracking: a visualization method for time-dependent 2D symmetric tensor fields. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tabatabai:1994:VRN

[TSM94] Behnam Tabatabai, Emanuela A. Sessarego, and Harald F. Mayer. Volume rendering on non-regular grids. Computer Graphics Forum, 13(3):C/247–C/258, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tiddeman:2005:IVP

[TSP05] B. P. Tiddeman, M. R. Stirrat, and D. I. Perrett. Image and video processing: Towards realism in facial image transformation: Results of a wavelet MRF method. Computer Graphics Forum, 24(3):449–456, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Telleen:2007:ISS

[TSY^{+}07] J. Telleen, A. Sullivan, J. Yee, O. Wang, P. Gunawardane, I. Collins, and J. Davis. Imaging (1): Synthetic shutter speed imaging. Computer Graphics Forum, 26(3):591–598, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tanaka:2001:GSS

[TSYK01] Satoshi Tanaka, Akihiro Shibata, Hiroaki Yamamoto, and Hisakiyo Kotsuru. Generalized stochastic sampling method for visualization and investigation of implicit surfaces. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tanaka:1994:CSB

[TT94] T. Tanaka and T. Takahashi. Cross Scan Buffer and its applications. Computer Graphics Forum, 13(3):C/467–C/476, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tanaka:1995:FSA

[TT95a] T. Tanaka and T. Takahashi. Fast shadowing algorithm for linear light sources. Computer Graphics Forum, 14(3):C/205–C/216, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thalmann:1995:CAS

[TT95b] Daniel Thalmann and Demetri Terzopoulos, editors. Computer animation and simulation ’95: proceedings of the Eurographics Workshop in Maastricht, the Netherlands, September 2–3, 1995, Springer Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-211-82738-2. LCCN TR897.7.C66 1995.

Tanaka:1997:FAS

[TT97] Toshimitsu Tanaka and Tokiichiro Takahashi. Fast analytic shading and shadowing for area light sources. Computer Graphics Forum, 16(3):C231–C240, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=160&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Theoharis:1990:DSI

[TTW90] T. A. Theoharis, A. R. L. Travis, and N. E. Wiseman. 3D display: Synthetic image generation and visual effect simulation. Computer Graphics Forum, 9(4):337–348, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Turk:1984:CGE

[Tur84] S. Turk. Computer graphics education. Computer Graphics Forum, 3(2):139–145, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tierny:2009:PSR

[TVD09] Julien Tierny, Jean-Philippe Vandeborre, and Mohamed Daoudi. Partial 3D shape retrieval by Reeb pattern unfolding. Computer Graphics Forum, 28(1):41–55, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thalmann:1998:EWC

[TvdP98] Daniel Thalmann and Michiel van de Panne. 8th Eurographics Workshop on Computer Animation and Simulation. Computer Graphics Forum, 17(1):??, March 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=226&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=1&amp; src=ard&amp; vid=17.

Turner:1996:ELI

[TW96] Martin J. Turner and Neil E. Wiseman. Efficient lossless image contour coding. Computer Graphics Forum, 15(2):107–117, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thuermer:1997:NCD

[TW97a] Grit Thuermer and Charles A. Wüthrich. Normal computation for discrete surfaces in 3D space. Computer Graphics Forum, 16 (3):C15–C26, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thurmer:1997:NCD

[TW97b] Grit Thürmer and Charles A. Wüthrich. Normal computation for discrete surfaces in 3D space. Computer Graphics Forum, 16(3):C15–??, September 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=138&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Tournier:2009:MCM

[TWC^{+}09] M. Tournier, X. Wu, N. Courty, E. Arnaud, and L. Revéret. Motion compression: Motion compression using principal geodesics analysis. Computer Graphics Forum, 28(2):355–364, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tonietto:2006:RAP

[TWJ06] L. Tonietto, M. Walter, and C. R. Jung. A randomized approach for patch-based texture synthesis using wavelets. Computer Graphics Forum, 25(4):675–684, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2009:VHS

[TYK^{+}09] Shigeo Takahashi, Kenichi Yoshida, Taesoo Kwon, Kang Hoon Lee, Jehee Lee, and Sung Yong Shin. Virtual humans: Spectralbased group formation control. Computer Graphics Forum, 28(2): 639–648, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tak:2000:MBF

[TySK00] Seyoon Tak, Oh young Song, and Hyeong-Seok Ko. Motion balance filtering. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=436&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=19.

Tai:2004:PMS

[TZF04] Chiew-Lan Tai, Hongxin Zhang, and Jacky Chun-Kin Fong. Prototype modeling from sketched silhouettes based on convolution surfaces. Computer Graphics Forum, 23(1):71–83, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urtasun:2004:SBM

[UGB^{+}04] Raquel Urtasun, Pascal Glardon, Ronan Boulic, Daniel Thalmann, and Pascal Fua. Style-based motion synthesis. Computer Graphics Forum, 23(4):799–812, December 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Unger:2008:SSL

[UGLY08] J. Unger, S. Gustavson, P. Larsson, and A. Ynnerman. Soft shadows and light fields: Free form incident light fields. Computer Graphics Forum, 27(4):1293–1301, June 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urban:1992:RTG

[UH92] B. Urban and T. Haberstroh. Ray tracing and graphics standards. Computer Graphics Forum, 11(1):55–60, March 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urena:2000:CIT

[Ure00] C. Ureña. Computation of irradiance from triangles by adaptive sampling. Computer Graphics Forum, 19(2):??, June 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ulicny:2002:TIR

[UT02] Branislav Ulicny and Daniel Thalmann. Towards interactive realtime crowd behavior simulation. Computer Graphics Forum, 21 (4):767–775, November 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ulbricht:2006:SAR

[UWP06] Christiane Ulbricht, Alexander Wilkie, and Werner Purgathofer. State of the art review: Verification of physically based rendering algorithms. Computer Graphics Forum, 25(2):237–255, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vandoni:1980:EPI

[Van80] C. E. Vandoni, editor. Eurographics 80: proceedings of the international conference and exhibition, University of Geneva, Switzerland, 3–5 September 1980. Elsevier North-Holland, Inc., New York, NY, USA, 1980. ISBN 0-444-86107-6 (Elsevier North-Holland). LCCN T385 .E97 1980 Long Loan.

vandenBos:1982:IMC

[van82] J. van den Bos. Introduction to man-computer communication. Computer Graphics Forum, 1(4):194–203, December 1982. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vandoni:1985:EPE

[Van85] C. E. Vandoni, editor. EUROGRAPHICS ’85: proceedings of the European Graphics Conference and Exhibition, Nice, France, 9–13 September, 1985. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1985. ISBN 0-444-87831-9. LCCN T 385 E97 1985.

vandenBos:1987:RCO

[van87] Jan van den Bos. RasterCalc: Calculus for operations on graphics colour rasters. Computer Graphics Forum, 6(3):181–193, September 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanBerkel:1989:SSI

[van89a] Pierre van Berkel. SIAS: Strokes Interpreted Animated Sequences. Computer Graphics Forum, 8(1):35–47, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanEmmerik:1989:RIC

[van89b] M. van Emmerik. Report on the international conference on CAD and CG, Beijing, China. Computer Graphics Forum, 8(4):370, December 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanEmmerik:1990:DMT

[van90] Maarten J. G. M. van Emmerik. A direct manipulation technique for specifying 3D object transformations with a 2D input device. Computer Graphics Forum, 9(4):355–361, December 1990. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandePanne:1997:FAa

[van97] Michiel van de Panne. From footprints to animation. Computer Graphics Forum, 16(4):211–223, October 1997. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vassilev:1997:ISD

[Vas97] Tzvetomir Ivanov Vassilev. Interactive sculpting with deformable nonuniform B-splines. Computer Graphics Forum, 16 (4):191–199, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=179&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Vazquez:2007:ALS

[Váz07] P.-P. Vázquez. Automatic light source placement for maximum visual information recovery. Computer Graphics Forum, 26(2): 143–156, June 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veryovka:1999:CHS

[VB99] O. Veryovka and J. Buchanan. Comprehensive halftoning of 3D scenes. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=323&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=18.

Veryovka:2000:TBD

[VB00] Oleg Veryovka and John Buchanan. Texture-based dither matrices. Computer Graphics Forum, 19(1):??, March 2000. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=387&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=1&amp; src=
ard&amp; vid=19.

Vidal:2006:SAR

[VBB^{+}06] F. P. Vidal, F. Bello, K. W. Brodlie, N. W. John, D. Gould, R. Phillips, and N. J. Avis. State of the art review: Principles and applications of computer graphics in medicine. Computer Graphics Forum, 25(1):113–137, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vieira:2009:SDI

[VBP^{+}09] Thales Vieira, Alex Bordignon, Adelailson Peixoto, Geovan Tavares, Hélio Lopes, Luiz Velho, and Thomas Lewiner. Symmetry detection and intelligent galleries: Learning good views through intelligent galleries. Computer Graphics Forum, 28(2): 717–726, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Valette:2004:GPA

[VC04] Sébastien Valette and Jean-Marc Chassery. Geometry processing: Approximated centroidal Voronoi diagrams for uniform polygonal mesh coarsening. Computer Graphics Forum, 23(3):381–389, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vemuri:1998:FCD

[VCC98] B. C. Vemuri, Y. Cao, and L. Chen. Fast collision detection algorithms with applications to particle flow. Computer Graphics Forum, 17(2):??, June 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=233&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=17.

VanReeth:1995:RVP

[VCDF95] F. Van Reeth, K. Coninx, S. De Backer, and E. Flerackers. Realizing 3D visual programming environments within a virtual environment. Computer Graphics Forum, 14(3):C/361–C/370, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Valette:2009:HCP

[VCP09] Sébastien Valette, Raphaëlle Chaine, and Rémy Prost. Hierarchies and compression: Progressive lossless mesh compression via incremental parametric refinement. Computer Graphics Forum, 28(5):1301–1310, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vandoni:1990:REB

[VD90] C. E. Vandoni and D. A. Duce. Revisions to the EUROGRAPH-ICS bye-laws. Computer Graphics Forum, 9(3):295–296, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanDam:1998:FUC

[vD98] Andries van Dam. Frontiers in user-computer interaction. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=245&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

vandePanne:1997:FAb

[vdP97] Michiel van de Panne. From footprints to animation. Computer Graphics Forum, 16(4):??, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=181&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=4&amp; src=
ard&amp; vid=16.

Veltkamp:1991:RSE

[Vel91] R. Veltkamp. Report on the second EUROGRAPHICS workshop on object-oriented graphics. Computer Graphics Forum, 10(4): 357–358, December 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veltkamp:1992:SCC

[Vel92] R. C. Veltkamp. Survey of continuities of curves and surfaces. Computer Graphics Forum, 11(2):93–112, June 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veltkamp:1993:CM

[Vel93] R. C. Veltkamp. 3D computational morphology. Computer Graphics Forum, 12(3):C115–C127, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veltkamp:1999:GGP

[Vel99] Remco C. Veltkamp. Generic geometric programming in the computational geometry algorithms library. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=363&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Velho:1996:ACP

[VG96] Luiz Velho and Jonas Gomes. Approximate conversion of parametric to implicit surfaces. Computer Graphics Forum, 15(5): 327–337, December 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Velho:2000:VRM

[VG00] Luiz Velho and Jonas Gomes. Variable resolution 4-k meshes: Concepts and applications. Computer Graphics Forum, 19(4):??, December 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vinayagamoorthy:2004:EGM

[VGSS04] V. Vinayagamoorthy, M. Garau, A. Steed, and M. Slater. An eye gaze model for dyadic interaction in an immersive virtual environment: Practice and experience. Computer Graphics Forum, 23(1): 1–11, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanco:2008:SRU

[VHB08] Marek Vančo, Bernd Hamann, and Guido Brunnett. Surface reconstruction from unorganized point data with quadrics. Computer Graphics Forum, 27(6):1593–1606, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanWijk:1985:SID

[vJB85] J. J. van Wijk, F. W. Jansen, and W. F. Bronsvoort. Some issues in designing user interfaces to 3D raster graphics. Computer Graphics Forum, 4(1):5–10, January 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanNieuwenhuizen:1994:ILT

[vKB94] P. R. van Nieuwenhuizen, O. Kiewiet, and W. F. Bronsvoort. An integrated line tracking and vectorization algorithm. Computer Graphics Forum, 13(3):C/349–C/359, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKaick:2006:CEM

[vKP06] Oliver Matias van Kaick and Hélio Pedrini. A comparative evaluation of metrics for fast mesh simplification. Computer Graphics Forum, 25(2):197–210, June 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandenBos:1990:PDB

[vL90] Jan van den Bos and Chris Laffra. Project DIGIS — building interactive applications by direct manipulation. Computer Graphics Forum, 9(3):181–193, September 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanCampenhout:1993:PPR

[VL93] J. M. Van Campenhout and R. Lasure. PRIP-a parallel raster image processor. Computer Graphics Forum, 12(2):95–104, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vallet:2008:GPS

[VL08] B. Vallet and B. Lévy. Geometry processing: Spectral geometry processing with manifold harmonics. Computer Graphics Forum, 27(2):251–260, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vieira:2004:SMS

[VLV^{+}04] Antônio W. Vieira, Thomas Lewiner, Luiz Velho, Hélio Lopes, and Geovan Tavares. Stellar mesh simplification using probabilistic optimization. Computer Graphics Forum, 23(4):825–838, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vieira:2004:SSR

[VMA^{+}04] M. B. Vieira, P. P. Martins Jr., A. A. Araújo, M. Cord, and S. Philipp-Foliguet. Smooth surface reconstruction using tensor fields as structuring elements. Computer Graphics Forum, 23(4): 813–823, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vucini:2009:GRV

[VMG09] Erald Vuçini, Torsten Möller, and M. Eduard Gröller. Geometry & reconstruction: On visualization and reconstruction from nonuniform point sets using B-splines. Computer Graphics Forum, 28 (3):1007–1014, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vollmer:1999:ILS

[VMM99] J. Vollmer, R. Mencl, and H. Müller. Improved Laplacian smoothing of noisy surface meshes. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=334&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Voloboj:1993:MDP

[Vol93] A. G. Voloboj. The method of dynamic palette construction in realistic visualization systems. Computer Graphics Forum, 12(5): 289–296, December 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanWalsum:1994:SVV

[vP94] T. van Walsum and F. H. Post. Selective visualization of vector fields. Computer Graphics Forum, 13(3):C/339–C/347, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Valette:2008:DMC

[VPLL08] Gilles Valette, Stéphanie Prévost, Laurent Lucas, and Joël Léonard. A dynamic model of cracks development based on a 3D discrete shrinkage volume propagation. Computer Graphics Forum, 27(1):47–62, March 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vilbrandt:2004:CHP

[VPP^{+}04] C. Vilbrandt, G. Pasko, A. Pasko, P.-A. Fayolle, T. Vilbrandt, J. R. Goodwin, J. M. Goodwin, and T. L. Kunii. Cultural heritage preservation using constructive shape modeling. Computer Graphics Forum, 23(1):25–41, March 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanHintum:1995:MCS

[VR95] J. E. A. Van Hintum and G. J. Reynolds. A multimedia constraint system (or: do we have it MADE). Computer Graphics Forum, 14(3):C/135–C/148, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vorsatz:2001:FSR

[VRKS01] J. Vorsatz, C. Rössl, L. P. Kobbelt, and H.-P. Seidel. Feature sensitive remeshing. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2009:CCB

[VS09] L. Váša and V. Skala. COBRA: Compression of the basis for PCA represented animations. Computer Graphics Forum, 28(6): 1529–1540, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vassilev:2001:FCA

[VSC01] T. Vassilev, B. Spanlang, and Y. Chrysanthou. Fast cloth animation on walking avatars. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanDenHengely:2009:IBM

[VSD09] A. Van Den Hengely, D. Sale, and A. R. Dick. Image-based modeling and rendering: SecondSkin: An interactive method for appearance transfer. Computer Graphics Forum, 28(7):1735–1744, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanLiere:1987:LID

[vt87] R. van Liere and P. J. W. ten Hagen. Logical input devices and interaction. Computer Graphics Forum, 6(4):349–357, December 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Volino:1994:ESC

[VT94] P. Volino and N. M. Thalmann. Efficient self-collision detection on smoothly discretized surface animations using geometrical shape regularity. Computer Graphics Forum, 13(3):C/155–C/166, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandenBos:1984:IIP

[vvT84] J. van den Bos, M. van Naelten, and W. Teunissen. IDECAP: Interactive pictorial information system for demographic and environmental planning applications. Computer Graphics Forum, 3(1):91–102, March 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Visvalingam:1990:DPA

[VW90] M. Visvalingam and J. D. Whyatt. The Douglas–Peucker algorithm for line simplification: Re-evaluation through visualization. Computer Graphics Forum, 9(3):213–228, September 1990. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Visvalingam:1991:CAP

[VW91] M. Visvalingam and J. D. Whyatt. Cartographic algorithms: problems of implementation and evaluation and the impact of digitising errors. Computer Graphics Forum, 10(3):225–235, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veltkamp:1995:MCM

[VW95] R. C. Veltkamp and W. Wesselink. Modeling 3D curves of minimal energy. Computer Graphics Forum, 14(3):C/97–C/110, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanHam:2008:GVC

[VW08] F. Van Ham and M. Wattenberg. Graph visualization: Centrality based visualization of small world graphs. Computer Graphics Forum, 27(3):975–982, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2009:SGI

[WÅ09] Rui Wang and Oskar Åkerlund. Shadows and global illumination: Bidirectional importance sampling for unstructured direct illumination. Computer Graphics Forum, 28(2):269–278, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolter:2009:TMT

[WAH^{+}09] M. Wolter, I. Assenmacher, B. Hentschel, M. Schirski, and T. Kuhlen. A time model for time-varying visualization. Computer Graphics Forum, 28(6):1561–1571, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waite:1988:MNB

[Wai88] Kevin W. Waite. Modelling natural branching structures. Computer Graphics Forum, 7(2):105–115, June 1988. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waldschmidt:1987:ECG

[Wal87] H. Waldschmidt. An editor for constructing graphics with TEX. Computer Graphics Forum, 6(4):359–364, December 1987. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watkins:1982:IEE

[Wat82] H. K. Watkins. Impressions of the EUROGRAPHICS ’82 Exhibition. Computer Graphics Forum, 1(4):176, December 1982. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watkins:1987:GRS

[Wat87] Howard K. Watkins. Graphics in reservoir simulation. Computer Graphics Forum, 6(2):111–117, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watkins:1988:REC

[Wat88] H. K. Watkins. Report on the Eurographics ’88 conference. Computer Graphics Forum, 7(4):359–360, December 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watson:1996:ISE

[Wat96] V. Watson. 3D, interactive, scientific expeditions on the World Wide Web and remote, collaborative visualization over the Internet. Computer Graphics Forum, 15(3):325, September 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watanabe:2001:DSC

[WB01] Kouki Watanabe and Alexander G. Belyaev. Detection of salient curvature features on polygonal surfaces. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Winnemoller:2002:GAT

[WB02] Holger Winnemöller and Shaun Bangay. Geometric approximations towards free specular comic shading. Computer Graphics Forum, 21(3):309–316, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2009:JMC

[WBCG09a] Ofir Weber, Mirela Ben-Chen, and Craig Gotsman. The joy of math and cartoon drawing: Complex barycentric coordinates with applications to planar shape deformation. Computer Graphics Forum, 28(2):587–597, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wither:2009:AAT

[WBCG09b] J. Wither, F. Boudon, M.-P. Cani, and C. Godin. All about trees: Structure from silhouettes: a new paradigm for fast sketch-based design of trees. Computer Graphics Forum, 28(2):541–550, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:1997:MGP

[WBEF97] LiFeng Wang, David Botta, Chris Ellefson, and Alain Fournier. Modelling the garden of perfect brightness. Computer Graphics Forum, 16(3):C365–C369, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=174&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Wicke:2007:DFE

[WBG07] Martin Wicke, Mario Botsch, and Markus Gross. Deformation: a finite element method on convex polyhedra. Computer Graphics Forum, 26(3):355–364, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:1998:AWV

[WBP98] Yigang Wang, Hujun Bao, and Qunsheng Peng. Accelerated walkthroughs of virtual environments based on visibility preprocessing and simplification. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=266&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Winter:2002:ISV

[WC02] Andrew S. Winter and Min Chen. Image-swept volumes. Computer Graphics Forum, 21(3):441–450, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2005:GIE

[WC05] Chung-Ming Wang and Yu-Ming Cheng. Geometry IV: an efficient information hiding algorithm for polygon models. Computer Graphics Forum, 24(3):591–600, September 2005. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Whitaker:1995:OCA

[WCB^{+}95] R. T. Whitaker, C. Crampton, D. E. Breen, M. Tuceryan, and E. Rose. Object calibration for augmented reality. Computer Graphics Forum, 14(3):C/15–C/27, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2009:HCD

[WD09] Kenneth Weiss and Leila De Floriani. Hierarchies and compression: Diamond hierarchies of arbitrary dimension. Computer Graphics Forum, 28(5):1289–1300, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2008:VRV

[WDC^{+}08] Oliver Wang, James Davis, Erika Chuang, Ian Rickard, Krystle De Mesa, and Chirag Dave. Video and relighting: Video relighting using infrared illumination. Computer Graphics Forum, 27(2): 271–279, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2001:ANF

[WDGT01] Xunlei Wu, Michael S. Downes, Tolga Goktekin, and Frank Tendick. Adaptive nonlinear finite elements for deformable body simulation using dynamic progressive meshes. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2009:IPN

[WDR09] Hongzhi Wu, Julie Dorsey, and Holly Rushmeier. Important points and NPR: Characteristic point maps. Computer Graphics Forum, 28(4):1227–1236, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Westermann:1997:MAI

[WE97] Rüdiger Westermann and Thomas Ertl. A multiscale approach to integrated volume segmentation and rendering. Computer Graphics Forum, 16(3):C117–C127, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=148&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Weiskopf:2004:SVD

[Wei04] D. Weiskopf. Scientific visualization: Dye advection without the blur: a level-set approach for texture-based visualization of unsteady flow. Computer Graphics Forum, 23(3):479–488, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiskopf:2008:RES

[Wei08] Daniel Weiskopf. Report: 2008 Eurographics Symposium on Parallel Graphics and Visualization. Computer Graphics Forum, 27 (8):2246, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walter:1997:GAP

[WF97] Marcelo Walter and Alain Fournier. Growing and animating polygonal models of animals. Computer Graphics Forum, 16(3): C151–C158, September 4–8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=151&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Warman:1982:EPI

[WG82] Ernest Warman and D. S. Greenaway, editors. EUROGRAPH-ICS ’82: proceedings of the international conference and exhibition, University of Manchester, Institute of Science and Technology, U.M.I.S.T., Manchester, U.K., 8–10 September, 1982. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1982. ISBN 0-444-86480-6. LCCN T 385 E97 1982. Organized by Eurographics Association.

Wang:1993:SSR

[WG93] W. Wang and G. G. Grinstein. A survey of 3D solid reconstruction from 2D projection line drawings. Computer Graphics Forum, 12 (2):137–158, June 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinkauf:2009:ESP

[WG09] T. Weinkauf and D. Günther. Edges: Separatrix persistence: Extraction of salient edges on surfaces using topological methods. Computer Graphics Forum, 28(5):1519–1528, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wyvill:1999:ECT

[WGG99] Brian Wyvill, Andrew Guy, and Eric Galin. Extending the CSG tree. warping, blending and Boolean operations in an implicit surface modeling system. Computer Graphics Forum, 18(2):??, June 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=365&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=18.

Wood:1988:ECC

[WGK88] C. A. Wood, P. D. Gray, and A. C. Kilgour. Experience with Chisl, a configurable hierarchical interface specification language. Computer Graphics Forum, 7(2):117–127, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waldyz:2004:SGI

[WGS04] Ingo Waldyz, Johannes Günthery, and Philipp Slusalleky. Sampling for global illumination: Balancing considered harmful — faster photon mapping using the voxel volume heuristic —. Computer Graphics Forum, 23(3):595–603, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wyrwas:1989:SGP

[WH89] K. M. Wyrwas and W. T. Hewitt. A survey of GKS and PHIGS implementations. Computer Graphics Forum, 8(1):49–59, March 1989. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1991:PAB

[WH91] Philip Willis and Andrew Hunter. A picture archive browser. Computer Graphics Forum, 10(1):49–59, March 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wooten:1996:AHD

[WH96] Wayne L. Wooten and Jessica K. Hodgins. Animation of human diving. Computer Graphics Forum, 15(1):3–13, March 1996. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2004:ICM

[WH04] Shuzhen Wang and Wolfgang Heidrich. Image calibration and matching: The design of an inexpensive very high resolution scan camera system. Computer Graphics Forum, 23(3):441–450, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wen:2008:IEE

[WHCO08] Chung-Lin Wen, Chang-Hsi Hsieh, Bing-Yu Chen, and Ming Ouhyoung. Image editing: Example-based multiple local color transfer by strokes. Computer Graphics Forum, 27(7):1765–1772, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2004:RAF

[WHL^{+}04] Yang Wang, Xiaolei Huang, Chan-Su Lee, Song Zhang, Zhiguo Li, Dimitris Samaras, Dimitris Metaxas, Ahmed Elgammal, and Peisen Huang. Retargeting and animating faces: High resolution acquisition, learning and transfer of dynamic 3-D facial expressions. Computer Graphics Forum, 23(3):677–686, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:1997:OIT

[WHR97] D. Wang, I. Herman, and G. J. Reynolds. The Open Inventor Toolkit and the PREMO standard. Computer Graphics Forum, 16(4):159–175, October 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=177&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=16.

Wiegand:1996:IRC

[Wie96] T. F. Wiegand. Interactive rendering of CSG models. Computer Graphics Forum, 15(4):249–261, October 1996. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1984:CGF

[Wil84] P. J. Willis. Colour graphics and the fifth generation workstation. Computer Graphics Forum, 3(2):147–152, June 1984. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Williams:1987:PAR

[Wil87a] Richard Williams. Preserving the area of regions. Computer Graphics Forum, 6(1):43–48, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1987:VSA

[Wil87b] P. J. Willis. Visual simulation of atmospheric haze. Computer Graphics Forum, 6(1):35–41, January 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:2006:OMP

[Wil06a] Phil Willis. Obituaries: In memoriam: Professor Dick Grimsdale 1929–2005. Computer Graphics Forum, 25(1):1–2, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:2006:MCP

[Wil06b] Philip Willis. Matting and compositing: Projective Alpha colour. Computer Graphics Forum, 25(3):557–566, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2004:PBT

[WK04] Jainhua Wu and Leif Kobbelt. Point-based techniques: Optimized sub-sampling of point sets for surface splatting. Computer Graphics Forum, 23(3):643–652, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Westrup:1985:UIE

[WKG85] Terry D. Westrup, W. Kegel, and J. Gras. User interaction with an environment for image processing and graphics. Computer Graphics Forum, 4(3):187–202, September 1985. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Woo:2009:ASC

[WKM^{+}09] Insoo Woo, Sung Ye Kim, Ross Maciejewski, David S. Ebert, Timothy D. Ropp, and Krystal Thomas. Applications: SDViz: a context-preserving interactive visualization system for technical diagrams. Computer Graphics Forum, 28(3):943–950, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2008:SFK

[WL08] S. F. Wang and S. H. Lai. Skull, face, and knits: Reconstructing 3D shape, albedo and illumination from a single face image. Computer Graphics Forum, 27(7):1729–1736, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wilson:1999:PHM

[WLML99] A. Wilson, E. Larsen, D. Manocha, and M. C. Lin. Partitioning and handling massive models for interactive collision detection. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=352&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Wurmlin:2003:VRS

[WLSG03] Stephan Würmlin, Edouard Lamboray, Oliver G. Staadt, and Markus H. Gross. 3D video recorder: a system for recording and playing free-viewpoint video. Computer Graphics Forum, 22(2): 181–193, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1985:BHD

[WM85] Philip Willis and David Milford. Browsing high definition colour pictures. Computer Graphics Forum, 4(3):203–208, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2009:BWB

[WM09] H. Wang and W. Ma. Biorthogonal wavelets based on interpolatory subdivision. Computer Graphics Forum, 28(6):1572–1585, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wald:2009:SAR

[WMG^{+}09] Ingo Wald, William R. Mark, Johannes Günther, Solomon Boulos, Thiago Ize, Warren Hunt, Steven G. Parker, and Peter Shirley. State of the art in ray tracing animated scenes. Computer Graphics Forum, 28(6):1691–1722, September 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2008:VAI

[WMS^{+}08] Xiaoyu Wang, Erin Miller, Kathleen Smarick, William Ribarsky, and Remco Chang. Visual analytics: Investigative visual analysis of global terrorism. Computer Graphics Forum, 27(3):919–926, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Winnemoller:2005:IVP

[WMTG05] Holger Winnemöller, Ankit Mohan, Jack Tumblin, and Bruce Gooch. Image and video processing: Light waving: Estimating light positions from photographs alone. Computer Graphics Forum, 24(3):433–438, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2009:ICT

[WMWG09] Basil Weber, Pascal Müller, Peter Wonka, and Markus Gross. Interactive cities and terrain: Interactive geometric simulation of 4D cities. Computer Graphics Forum, 28(2):481–492, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wyman:2009:CSA

[WN09] Chris Wyman and Greg Nichols. Caustics and scattering: Adaptive caustic maps using deferred shading. Computer Graphics Forum, 28(2):309–318, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1992:RNR

[WO92] P. Willis and R. Oddy. Rendering NURB regions for 2D animation. Computer Graphics Forum, 11(3):C35–C44, C465, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:1994:RLH

[WO94] Jiann-Rong Wu and Ming Ouhyoung. Reducing the latency in head-mounted displays by a novel prediction method using Grey System theory. Computer Graphics Forum, 13(3):C/503–C/512, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2002:ISR

[WO02] Jianning Wang and Manuel M. Oliveira. Improved scene reconstruction from range images. Computer Graphics Forum, 21(3): 521–530, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Winnemoller:2009:ITT

[WOBT09] H. Winnemöller, A. Orzan, L. Boissieux, and J. Thollot. Images and textures: Texture design and draping in 2D images. Computer Graphics Forum, 28(4):1091–1099, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2004:DFC

[WP04] Xiaobin Wu and Jörg Peters. Distance fields and collision detection: Interference detection for subdivision surfaces. Computer Graphics Forum, 23(3):577–584, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walter:2002:UPT

[WPG02] Bruce Walter, Sumanta N. Pattanaik, and Donald P. Greenberg. Using perceptual texture masking for efficient image synthesis. Computer Graphics Forum, 21(3):393–399, September 2002. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wojdel:2005:PGF

[WR05] A. Wojdeł and L. J. M. Rothkrantz. Parametric generation of facial expressions based on FACS. Computer Graphics Forum, 24 (4):743–757, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Westermann:2001:RTV

[WRS01] Rüdiger Westermann and Christof Rezk-Salama. Real-time volume deformations. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wonka:1999:OSF

[WS99] Peter Wonka and Dieter Schmalstieg. Occluder shadows for fast walkthroughs of urban environments. Computer Graphics Forum, 18(3):??, September 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=327&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=18.

Way:2001:SRT

[WS01] Der-Lor Way and Zen-Chung Shih. The synthesis of rock textures in Chinese landscape painting. Computer Graphics Forum, 20(3): ??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wand:2002:MRR

[WS02] Michael Wand and Wolfgang Straßer. Multi-resolution rendering of complex animated scenes. Computer Graphics Forum, 21(3): 483–491, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wand:2003:GIR

[WS03a] M. Wand and W. Straßer. Global illumination: Real-time caustics. Computer Graphics Forum, 22(3):611–620, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wand:2003:RTC

[WS03b] M. Wand and W. Straßer. Real-time caustics. Computer Graphics Forum, 22(3):611–620, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wennersten:2009:TVT

[WS09a] Per Wennersten and Jacob Ström. Textures and video: Tablebased Alpha compression. Computer Graphics Forum, 28(2):687– 695, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Woodring:2009:VRS

[WS09b] Jonathan Woodring and Han-Wei Shen. Volume rendering: Semiautomatic time-series transfer functions via temporal clustering and sequencing. Computer Graphics Forum, 28(3):791–798, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wald:2001:IRC

[WSBW01] Ingo Wald, Philipp Slusallek, Carsten Benthin, and Markus Wagner. Interactive rendering with coherent ray tracing. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weng:2008:MIS

[WSBZ08] Yanlin Weng, Xiaohan Shi, Hujun Bao, and Jun Zhang. Moving images: Sketching MLS image deformations on the GPU. Computer Graphics Forum, 27(7):1789–1796, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walder:2006:MIS

[WSC06] C. Walder, B. Schölkopf, and O. Chapelle. Modelling I: Implicit surface modelling with a globally regularised basis of compact support. Computer Graphics Forum, 25(3):635–644, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiskopf:2004:GBR

[WSE04] Daniel Weiskopf, Tobias Schafhitzel, and Thomas Ertl. GPUbased rendering: GPU-based nonlinear ray tracing. Computer Graphics Forum, 23(3):625–633, September 2004. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wicke:2005:PEA

[WSG05] Martin Wicke, Denis Steinemann, and Markus Gross. Points: Efficient animation of point-sampled thin shells. Computer Graphics Forum, 24(3):667–676, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2007:SMD

[WSLG07] Ofir Weber, Olga Sorkine, Yaron Lipman, and Craig Gotsman. Skinning and mesh deformation: Context-aware skeletal shape deformation. Computer Graphics Forum, 26(3):265–274, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wainer:1993:LDP

[WT93] M. Wainer and Ting Tan. A library for developing PHIGS programming tools in a PEX environment. Computer Graphics Forum, 12(4):219–229, October 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2009:ADF

[WT09] Fernando J. Wong and Shigeo Takahashi. Abstraction and design: Flow-based automatic generation of hybrid picture mazes. Computer Graphics Forum, 28(7):1975–1984, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinkauf:2004:SVT

[WTHS04] Tino Weinkauf, Holger Theisel, Hans-Christian Hege, and Hans-Peter Seidel. Scientific visualization: Topological construction and visualization of higher order 3D vector fields. Computer Graphics Forum, 23(3):469–478, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinkauf:2006:VTS

[WTHS06] T. Weinkauf, H. Theisel, H.-C. Hege, and H.-P. Seidel. Visualization: Topological structures in two-parameter-dependent 2D vector fields. Computer Graphics Forum, 25(3):607–616, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:1990:SNL

[Wu90] Jintong Wu. Simulating natural light. Computer Graphics Forum, 9(4):349–353, December 1990. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Westenberg:2008:VAL

[WVKR08] M. A. Westenberg, S. A. F. T. Van Hijum, O. P. Kuipers, and J. B. T. M. Roerdink. Visualization applications in life sciences: Visualizing genome expression and regulatory network dynamics in genomic and metabolic context. Computer Graphics Forum, 27(3):887–894, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willems:2009:AVV

[WVV09] Niels Willems, Huub Van De Wetering, and Jarke J. Van Wijk. Applications: Visualization of vessel movements. Computer Graphics Forum, 28(3):959–966, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wijffelaars:2008:VTG

[WVVV08] Martijn Wijffelaars, Roel Vliegen, Jarke J. Van Wijk, and Erik-Jan Van Der Linden. Visualization tools: Generating color palettes using intuitive parameters. Computer Graphics Forum, 27(3):743–750, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Watters:1987:UNA

[WW87a] Geoff Watters and Philip Willis. UltraPaint: a new approach to a painting system. Computer Graphics Forum, 6(2):125–132, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1987:SCE

[WW87b] Philip Willis and Geoff Watters. Scan converting extruded lines at ultra high definition. Computer Graphics Forum, 6(2):133–140, May 1987. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willis:1988:CRO

[WW88] P. J. Willis and G. W. Watters. Colour raster operations. Computer Graphics Forum, 7(3):151–159, September 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weimer:1998:SST

[WW98] Henrik Weimer and Joe Warren. Subdivision schemes for thin plate splines. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=277&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-7055&
amp;   iid=3&amp; src=ard&amp; vid=17.

Wang:1999:ASI

[WW99] Wencheng Wang and Enhua Wu. Adaptable splatting for irregular volume rendering. Computer Graphics Forum, 18 (4):??, December 1999. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=375&amp;

http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=4&amp; src=ard&amp; vid=18.

Weidlich:2009:RAD

[WW09a] Andrea Weidlich and Alexander Wilkie. Rendering I: Anomalous dispersion in predictive rendering. Computer Graphics Forum, 28 (4):1065–1072, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wilkie:2009:ITR

[WW09b] Alexander Wilkie and Andrea Weidlich. Images and textures: a robust illumination estimate for chromatic adaptation in rendered images. Computer Graphics Forum, 28(4):1101–1109, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waschbusch:2007:MIB

[WWG07] Michael Waschbüsch, Stephan Würmlin, and Markus Gross. Material and image based rendering: 3D video billboard clouds. Computer Graphics Forum, 26(3):561–569, September 2007. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wonka:2001:IV

[WWS01] Peter Wonka, Michael Wimmer, and François X. Sillion. Instant visibility. Computer Graphics Forum, 20(3):??, September 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:1992:FNE

[WY92] Bennan Wang and Shi Yin. The fractal nature of an ecological model. Computer Graphics Forum, 11(1):61–64, March 1992. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2007:MIB

[XGL^{+}07] Kun Xu, Yue Gao, Yong Li, Tao Ju, and Shi-Min Hu. Material and image based rendering: Real-time homogeneous translucent material editing. Computer Graphics Forum, 26(3):545–552, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2009:IRO

[xHMC09] Qi xing Huang, Radomir Mech, and Nathan Carr. Image restructuring: Optimizing structure preserving embedded deformation for resizing images and vector art. Computer Graphics Forum, 28(7):1887–1896, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2008:NPR

[XJJ^{+}08] Songhua Xu, Hao Jiang, Tao Jin, Francis C. M. Lau, and Yunhe Pan. Non-photorealistic rendering: Automatic facsimile of Chinese calligraphic writings. Computer Graphics Forum, 27(7): 1879–1886, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2003:ADR

[XLTP03a] Songhua Xu, Francis C. M. Lau, Feng Tang, and Yunhe Pan. Advanced design for a realistic virtual brush. Computer Graphics Forum, 22(3):533–542, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2003:NPR

[XLTP03b] Songhua Xu, Francis C. M. Lau, Feng Tang, and Yunhe Pan. Non-photorealistic rendering and painting: Advanced design for a realistic virtual brush. Computer Graphics Forum, 22(3):533–542, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2009:IEG

[XM09] Xuezhong Xiao and Lizhuang Ma. Image editing: Gradientpreserving color transfer. Computer Graphics Forum, 28(7):1879– 1886, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2006:CRS

[XS06] Huiying Xu and Yinlong Sun. Compact representation of spectral BRDFs using Fourier transform and spherical harmonic expansion. Computer Graphics Forum, 25(4):759–775, December 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2007:ICG

[XTJ^{+}07] Songhua Xu, Haisheng Tan, Xiantao Jiao, Francis C. M. Lau, and Yunhe Pan. Image and color: a generic pigment model for digital painting. Computer Graphics Forum, 26(3):609–618, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2002:SMB

[XTLP02] Songhua Xu, Min Tang, Francis Lau, and Yunhe Pan. A solid model based virtual hairy brush. Computer Graphics Forum, 21 (3):299–308, September 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xuey:2008:IBT

[XWT^{+}08] Su Xuey, Jiaping Wang, Xin Tong, Qionghai Dai, and Baining Guo. Image-based techniques: Image-based material weathering. Computer Graphics Forum, 27(2):617–626, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2009:IEE

[XWT^{+}09] Kun Xu, Jiaping Wang, Xin Tong, Shi-Min Hu, and Baining Guo. Image editing: Edit propagation on bidirectional texture functions. Computer Graphics Forum, 28(7):1871–1877, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2009:AVB

[XYL09] Songhua Xu, Wenxia Yang, and Francis C. M. Lau. Applications: a visualization based approach for digital signature authentication. Computer Graphics Forum, 28(3):935–942, June 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoshizawa:2007:SMD

[YBS07] Shin Yoshizawa, Alexander Belyaev, and Hans-Peter Seidel. Skinning and mesh deformation: Skeleton-based variational mesh deformations. Computer Graphics Forum, 26(3):255–264, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2009:BVC

[YCLE09] Zhicheng Yan, Wei Chen, Aidong Lu, and David S. Ebert. Biomedical visualization: Context-aware volume modeling of skeletal muscles. Computer Graphics Forum, 28(3):887–894, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yau:1988:TMA

[YD88] John F. S. Yau and Neil D. Duffy. A texture mapping approach to 3-D facial image synthesis. Computer Graphics Forum, 7(2):

129–134, June 1988. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2008:ASI

[Ye08] Juntao Ye. Animation: Simulating inextensible cloth using impulses. Computer Graphics Forum, 27(7):1901–1907, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ying:1985:AAF

[YF85] Dao-Ning Ying and Xing Feng. Arbitrary area filling in a fast procedure. Computer Graphics Forum, 4(4):363–370, December 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2009:SPB

[YFGL09] Baoguang Yang, Jieqing Feng, Gaël Guennebaud, and Xinguo Liu. Shadows: Packet-based hierarchal soft shadow mapping. Computer Graphics Forum, 28(4):1121–1130, June 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2009:DGP

[YGL^{+}09] Yong-Liang Yang, Ren Guo, Feng Luo, Shi-Min Hu, and Xianfeng Gu. Digital geometry processing: Generalized discrete Ricci flow. Computer Graphics Forum, 28(7):2005–2014, October 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yue:2009:RAM

[YIC^{+}09] Yonghao Yue, Kei Iwasaki, Bing-Yu Chen, Yoshinori Dobashi, and Tomoyuki Nishita. Rendering and appearance modeling: Interactive rendering of interior scenes with dynamic environment illumination. Computer Graphics Forum, 28(7):1935–1944, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yagel:1992:TBV

[YK92] R. Yagel and A. Kaufman. Template-based volume viewing. Computer Graphics Forum, 11(3):C153–C167, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoon:2006:DSB

[YK06] Seung-Hyun Yoon and Myung-Soo Kim. Deformations: Sweepbased freeform deformations. Computer Graphics Forum, 25(3):

487–496, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuksel:2008:SLD

[YK08] Cem Yuksel and John Keyser. Shadows and lighting: Deep opacity maps. Computer Graphics Forum, 27(2):675–680, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoon:2009:APS

[YKH^{+}09] Jong-Chul Yoon, Hyeong Ryeol Kam, Jeong-Mo Hong, Shin Jin Kang, and Chang-Hun Kim. Animation: Procedural synthesis using vortex particle method for fluid simulation. Computer Graphics Forum, 28(7):1853–1859, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kwon:2008:ECM

[yKL08] Ji yong Kwon and In-Kwon Lee. Exaggerating character motions using sub-joint hierarchy. Computer Graphics Forum, 27 (6):1677–1686, September 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2007:FMC

[YLD07] Qian Yu, Qing Li, and Zhigang Deng. Fluid and motion capture: Online motion capture marker labeling for multiple interacting articulated targets. Computer Graphics Forum, 26(3):477–483, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoon:2008:NPR

[YLK08] Jong-Chul Yoon, In-Kwon Lee, and Henry Kang. Nonphotorealistic rendering: a hidden-picture puzzles generator. Computer Graphics Forum, 27(7):1869–1877, October 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2009:MSI

[YLL^{+}09] Dong-Ming Yan, Bruno Lévy, Yang Liu, Feng Sun, and Wenping Wang. Meshing — simplex: Isotropic remeshing with fast and exact computation of restricted Voronoi diagram. Computer Graphics Forum, 28(5):1445–1454, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoon:2006:ARC

[YM06] Sung-Eui Yoon and Dinesh Manocha. Animation rendering: Cache-efficient layouts of bounding volume hierarchies. Computer Graphics Forum, 25(3):507–516, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yeh:2009:SDI

[YM09] Yi-Ting Yeh and Radomír Měch. Symmetry detection and intelligent galleries: Detecting symmetries and curvilinear arrangements in vector art. Computer Graphics Forum, 28(2):707–716, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoshida:2006:ITM

[YMMS06] Akiko Yoshida, Rafał Mantiuk, Karol Myszkowski, and Hans-Peter Seidel. Images and tone mapping: Analysis of reproducing real-world appearance on displays of varying dynamic range. Computer Graphics Forum, 25(3):415–426, September 2006. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yamaguchi:1993:NPG

[YN93] F. Yamaguchi and M. Niizeki. A new paradigm for geometric processing. Computer Graphics Forum, 12(3):C177–C188, ???? 1993. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoon:2000:WBR

[YN00] I. Yoon and U. Neumann. Web-based remote rendering with IBRAC (image-based rendering acceleration and compression). Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=424&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Yu:2009:FBS

[YNBH09] Qizhi Yu, Fabrice Neyret, Eric Bruneton, and Nicolas Holzschuch. Fluids and beyond: Scalable real-time animation of rivers. Computer Graphics Forum, 28(2):239–248, April 2009. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:1995:MBS

[YP95] Yizhou Yu and Qunsheng Peng. Multiresolution B-spline radiosity. Computer Graphics Forum, 14(3):C/285–C/298, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:1991:FC

[YR91] C. Yao and J. Rokne. Fat curves. Computer Graphics Forum, 10(3):237–248, September 1991. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:1997:ARI

[YR97] Chengfu Yao and Jon G. Rokne. Applying rounding-up integral linear interpolation to the scan-conversion of filled polygons. Computer Graphics Forum, 16(2):101–106, June 1997. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2008:RTG

[YSL08] Lei Yang, Pedro V. Sander, and Jason Lawrence. Real time: Geometry-aware framebuffer level of detail. Computer Graphics Forum, 27(4):1183–1188, June 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoshida:1994:SMB

[YSY94] N. Yoshida, M. Shiokawa, and F. Yamaguchi. Solid modeling based on a new paradigm. Computer Graphics Forum, 13(3): C/55–C/64, ???? 1994. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:1997:RES

[YW97] Yizhou Yu and Hong Wu. A rendering equation for specular transfers and its integration into global illumination. Computer Graphics Forum, 16(3):C283–C292, September 4–8, 1997. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=165&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=16.

Yang:2003:RTC

[YWB03] Ruigang Yang, Greg Welch, and Gary Bishop. Real-time consensus-based scene reconstruction using commodity graphics hardware. Computer Graphics Forum, 22(2):207–216, June 2003.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2008:IIG

[YWY08] Xuan Yu, Rui Wang, and Jingyi Yu. Illumination: Interactive glossy reflections using GPU-based ray tracing with adaptive LOD. Computer Graphics Forum, 27(7):1987–1996, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

You:2007:MSB

[YYPZ07] L. H. You, X. S. Yang, M. Pachulski, and Jian J. Zhang. Modeling and segmentation: Boundary constrained swept surfaces for modelling and animation. Computer Graphics Forum, 26(3):313–322, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zachrisen:1989:YAR

[Zac89] M. Zachrisen. Yet another remark on the modelling clip problem. Computer Graphics Forum, 8(3):237–238, September 1989. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zara:2006:VRC

[Zar06] J. Zara. Virtual reality course — a natural enrichment of computer graphics classes. Computer Graphics Forum, 25(1):105–112, March 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ziegler:2007:GIB

[ZBA^{+}07] Remo Ziegler, Simon Bucheli, Lukas Ahrenberg, Marcus Magnor, and Markus Gross. Global illumination (1): a bidirectional light field — hologram transform. Computer Graphics Forum, 26(3): 435–446, September 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zaninetti:1999:AMA

[ZBP99] Jacques Zaninetti, Pierre Boy, and Bernard Peroche. An adaptive method for area light sources and daylight in ray tracing. Computer Graphics Forum, 18(3):??, September 1999. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=335&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=18.

Zemcik:1995:OCT

[ZC95] Pavel Zemcik and Alan Chalmers. Optimised CSG tree evaluation for space subdivision. Computer Graphics Forum, 14(2):139–146, June 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:1998:NAD

[ZCG98] J. M. Zheng, K. W. Chan, and I. Gibson. A new approach for direct manipulation of free-form curve. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=279&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=17.

Ziegler:2008:OVR

[ZCG08] Remo Ziegler, Simone Croci, and Markus Gross. Occlusion and volume rendering: Lighting and occlusion in a wave-based framework. Computer Graphics Forum, 27(2):211–220, April 2008. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2009:IRS

[ZCHM09] Guo-Xin Zhang, Ming-Ming Cheng, Shi-Min Hu, and Ralph R. Martin. Image restructuring: a shape-preserving approach to image resizing. Computer Graphics Forum, 28(7):1897–1906, October 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zalik:1997:ECB

[ŽCO97] Borut Žalik, Gordon Clapworthy, and Črtomir Oblonšek. An efficient code-based voxel-traversing algorithm. Computer Graphics Forum, 16(2):119–128, June 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=128&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.asp?
ref=0167-7055&amp;   iid=2&amp; src=ard&amp; vid=16.

Zhou:2007:ADF

[ZCP07] Tianshu Zhou, Jim X. Chen, and Mark Pullen. Accurate depth of field simulation in real time. Computer Graphics Forum, 26(1): 15–23, March 2007. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2009:SSL

[ZCQ^{+}09] Hong Zhou, Weiwei Cui, Huamin Qu, Yingcai Wu, Xiaoru Yuan, and Wei Zhuo. Scatterplots: Splatting the lines in parallel coordinates. Computer Graphics Forum, 28(3):759–766, June 2009. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeleznik:2000:SGB

[ZHC^{+}00] Bob Zeleznik, Loring Holden, Michael Capps, Howard Abrams, and Tim Miller. Scene-Graph-As-Bus: Collaboration between heterogeneous stand-alone 3-D graphical applications. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=401&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=19.

Zhang:2008:MIS

[ZHM08] Yi-Fei Zhang, Shi-Min Hu, and Ralph R. Martin. Moving images: Shrinkability maps for content-aware video resizing. Computer Graphics Forum, 27(7):1797–1804, October 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zander:2004:SAR

[ZISS04] Johannes Zander, Tobias Isenberg, Stefan Schlechtweg, and Thomas Strothotte. Stylized animation and rendering: High quality hatching. Computer Graphics Forum, 23(3):421–430, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:1992:OOV

[ZK92] T. Z.-Y. Zhou and W. J. Kubitz. An object-oriented view of the user interface. Computer Graphics Forum, 11(3):C81–C92, ???? 1992. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ziegler:2008:VAC

[ZK08] Hartmut Ziegler and Daniel A. Keim. Visualization applications: COPERNICUS: Context-preserving engine for route navigation with interactive user-modifiable scaling. Computer Graphics Forum, 27(3):927–934, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ziemkiewicz:2009:USI

[ZK09] Caroline Ziemkiewicz and Robert Kosara. User studies and interaction: Preconceptions and individual differences in understanding visual metaphors. Computer Graphics Forum, 28(3):911–918, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2005:PRH

[ZLK05] Jianhui Zhao, Ling Li, and Kwoh Chee Keong. 3D posture reconstruction and human animation from 2D feature points. Computer Graphics Forum, 24(4):759–771, December 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zotti:2008:SPE

[Zot08] G. Zotti. Selected papers from the EG 2007 cultural heritage track: Tangible heritage: Production of astrolabes on a laser engraver. Computer Graphics Forum, 27(8):2169–2177, December 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2003:EMA

[ZPS03] Yu Zhang, Edmond C. Prakash, and Eric Sung. Efficient modeling of an anatomy-based face and fast 3D facial expression synthesis. Computer Graphics Forum, 22(2):159–169, June 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2004:FHA

[ZQK04] Haitao Zhang, Feng Qiu, and Arie Kaufman. Fast hybrid approach for texturing point models. Computer Graphics Forum, 23(4):715–725, December 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2008:SSB

[ZQQS08] Hui Zhang, Guiming Qin, Kaihuai Qin, and Hanqiu Sun. Subdivision and segmentation: a biorthogonal wavelet approach based on dual subdivision. Computer Graphics Forum, 27(7):1815–1822, October 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zakai:1996:TDMa

[ZR96a] Y. Zakai and A. Rappoport. Three-dimensional modeling and effects on still images. Computer Graphics Forum, 15(3):C3–C10, C455–C456, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zakai:1996:TDMb

[ZR96b] Yaron Zakai and Ari Rappoport. Three-dimensional modeling and effects on still images. Computer Graphics Forum, 15(3):3– 10, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zayer:2005:GIH

[ZRKS05] Rhaleb Zayer, Christian Rössl, Zachi Karni, and Hans-Peter Seidel. Geometry IV: Harmonic guidance for surface deformation. Computer Graphics Forum, 24(3):601–609, September 2005. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2008:RDS

[ZSCO^{+}08] H. Zhang, A. Sheffer, D. Cohen-Or, Q. Zhou, O. Van Kaick, and A. Tagliasacchi. Registering deforming surfaces I: Deformationdriven shape correspondence. Computer Graphics Forum, 27 (5):1431–1439, July 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zaninetti:1998:VAG

[ZSP98] Jacques Zaninetti, Xavier Serpaggi, and Bernard Péroche. A vector approach for global illumination in ray tracing. Computer Graphics Forum, 17(3):??, September 1998. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

URL aid=262&amp;   http://www.blackwellpublishers.co.
uk/asp/journal.asp?ref=0167-7055&amp;   iid=3&amp; src=
ard&amp; vid=17.

Zair:1996:FMU

[ZT96] Chems Eddine Zair and Eric Tosan. Fractal modeling using free form techniques. Computer Graphics Forum, 15(3):C269–C278, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zickler:2009:TBB

[ZV09] Stefan Zickler and Manuela Veloso. Tactics-based behavioural planning for goal-driven rigid body control. Computer Graphics Forum, 28(8):2302–2314, December 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2002:CBC

[ZY02] Dongliang Zhang and Matthew M. F. Yuen. A coherence-based collision detection method for dressed human simulation. Computer Graphics Forum, 21(1):??, March 2002. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2004:SMF

[ZY04] Jian. J. Zhang and L. H. You. Shape modelling: Fast surface modelling using a 6th order PDE. Computer Graphics Forum, 23 (3):311–320, September 2004. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2008:PCV

[ZYQ^{+}08] Hong Zhou, Xiaoru Yuan, Huamin Qu, Weiwei Cui, and Baoquan Chen. Parallel coordinates: Visual clustering in parallel coordinates. Computer Graphics Forum, 27(3):1047–1054, May 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).